Gcal

Gcal

An Extended Calendar Program Edition 4 June 2000

by Thomas Esken

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994, 95, 96, 1997, 2000 Thomas Esken

Any suggestions, improvements, extensions, bug reports, donations, proposals for contract work, and so forth are welcome! Please send them directly to my eMail address esken@gmx.net. If you like my work, I’d appreciate a postcard from you!

      ------------------------oOO \\\_''/ OOo---------------------------
               Thomas Esken O (/o-o\) O eMail:  esken@gmx.net
               Im Hagenfeld 84 (( ^)) Phone:  +49 251 232585
      D-48147 Muenster; Germany \____) ~ (____/ MotD : 2old2live, 2young2die

This is Edition 4 of Gcal, an Extended Calendar Program, for the 4 (or later) version of the GNU implementation

of cal and calendar.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Free Software Foundation.

Short Contents

1 Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Gcal Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3 Invoking gcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4 Eternal Holidays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 5 Fixed Dates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 A Genesis of the Gregorian Calendar . . . . . . . . . . . . . . . . . . . . . 125 B Gcal Utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 C Aspects in Internationalization . . . . . . . . . . . . . . . . . . . . . . . . 131 D Metasymbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 E Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 F Summary of all Regular Expressions . . . . . . . . . . . . . . . . . . . . 143 G %?… Special Texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 H Summary of all %?… Special Texts . . . . . . . . . . . . . . . . . . . . 211 I Obsolete Special Texts and Format Elements . . . . . . . . . . . . . 245 J Format Instruction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 K Coding Scheme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 L Environment Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 M Error Codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 N Glossary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Command Line Argument Index . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 Environment Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 Table Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277 O GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . 279 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287

Table of Contents

1 Preface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Gcal Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3 Invoking gcal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3.1 Command line arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

3.1.1 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.1.1.1 Common options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.1.2 Global options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.1.1.3 Calendar options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.1.1.4 Fixed date options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

3.1.2 Response file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 3.1.3 Actual date modifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 3.1.4 Commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

3.1.4.1 Single commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 3.1.4.2 3-Month mode commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 3.1.4.3 Lists of commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 3.1.4.4 Ranges of commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

3.2 The GCAL environment variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

4 Eternal Holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 5 Fixed Dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

5.1 Resource file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 5.1.1 Structure of resource file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 5.1.2 Date part of a line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 5.1.3 Further date part attributes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

5.1.3.1 Lists of days . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 5.1.3.2 Ranges of days . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 5.1.3.3 Repetition factor of days . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 5.1.3.4 Appearance factor of days . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

5.1.4 Text part of a line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 5.1.5 %?… Special Texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 5.1.6 Comment line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 5.1.7 Include directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 5.1.8 Date variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 5.1.9 Text variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106

5.2 Resource file examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

Appendix A Genesis of the Gregorian Calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 Appendix B Gcal Utilities . . . . . . . . . . . . . . . . . . . . 127

B.1 Invoking tcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 B.2 Invoking txt2gcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 B.3 Invoking gcal2txt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

Appendix C Aspects in Internationalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

Appendix D Metasymbols . . . . . . . . . . . . . . . . . . . . 135

Appendix E Regular Expressions . . . . . . . . . . . . 137

E.1 How to Use Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 E.2 Regular Expression Operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

Appendix F Summary of all Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143

Appendix G %?… Special Texts . . . . . . . . . . . . . 145

G.1 %?… special texts for date exclusion . . . . . . . . . . . . . . . . . . . . . . . . 146

G.1.1 Exclusions with date argument %?[date] special texts … 146

G.1.1.1 Inclusive date period %i[date][#[date]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 G.1.1.2 Exclusive date period %e[date][#[date]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 G.1.2 Exclusions without any argument %? special texts . . . . . . . 148

G.1.2.1 Inclusive day period %? special texts . . . . . . . . . . . . . . . 149 G.1.2.2 Exclusive day period %? special texts. . . . . . . . . . . . . . . 149

G.2 %?… special texts for text replacement . . . . . . . . . . . . . . . . . . . . . 150

G.2.1 Replacements with date argument %?[date] special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 G.2.1.1 Weekday name %[format]K[date] special text . . . . . . . 150 G.2.1.2 Weekday number %[format]?[date] special texts …. 151 G.2.1.3 Day-of-year number %[format]N[date] special text . . 153 G.2.1.4 Day number %[format]D[date] special text. . . . . . . . . . 153 G.2.1.5 Week number %k[date] special text. . . . . . . . . . . . . . . . . 154 G.2.1.6 Month name %[format]U[date] special text . . . . . . . . . 154 G.2.1.7 Month number %[format]M[date] special text . . . . . . . 154 G.2.1.8 Year number %[format]Y[date] special text . . . . . . . . . 155 G.2.1.9 Moon phase %[format]?[date] special texts . . . . . . . . . 155 G.2.1.10 Biorhythm %?[date] special texts . . . . . . . . . . . . . . . . . 156 G.2.1.11 Difference value %[format]?[date] special texts …. 158

G.2.2 Replacements with other argument %[format]?[argument] special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 G.2.2.1 Actual clocktime %[format]?[argument] special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 G.2.2.2 Textual date %n[argument] special text . . . . . . . . . . . . . 162 G.2.2.3 Julian day number %[format]j[argument] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 G.2.2.4 Geographical distance and course angle

%[format]bargument special text . . . . . . . . . . . . . . . . . . . . . . . . 162 G.2.2.5 Sun data %[format]?argument special texts . . . . . . . . 165 G.2.2.6 Moon data %[format]?argument special texts. . . . . . . 188 G.2.2.7 Environment variable %[format]-[argument] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 G.2.3 Replacements without any argument %? special texts …. 206

G.2.3.1 Difference value %[format]? special texts . . . . . . . . . . . 206 G.2.3.2 Highlighting %? special texts . . . . . . . . . . . . . . . . . . . . . . . 208

G.3 Shell command %![argument] special text . . . . . . . . . . . . . . . . . . . . 210

Appendix H Summary of all %?… Special Texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

Appendix I Obsolete Special Texts and Format

Elements. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

I.1 Table of Obsolete Special Texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 I.2 Table of Obsolete Date Format Elements . . . . . . . . . . . . . . . . . . . . . 247 I.3 Table of Obsolete Date Formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247

Appendix J Format Instruction . . . . . . . . . . . . . . 249

Appendix K Coding Scheme . . . . . . . . . . . . . . . . . . 251

K.1 Coding Scheme Table 1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 K.2 Coding Scheme Table 2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253

Appendix L Environment Variables. . . . . . . . . . 255

Appendix M Error Codes . . . . . . . . . . . . . . . . . . . . . 257

Appendix N Glossary. . . . . . . . . . . . . . . . . . . . . . . . . . 259

Command Line Argument Index . . . . . . . . . . . . . . . 271

Environment Variable Index . . . . . . . . . . . . . . . . . . . . 275

Table Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 277

Appendix O GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279

Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287

1 Preface

The reason why I have written Gcal was mainly to have an off-line tool^{1} which is able to display Julian and Gregorian calendar sheets in a highly flexible and internationalized manner, fixed date lists for reminding purposes and country respectively territory specific holiday lists. At a later stage, Gcal was extended in the way to perform some basic astronomical calculations like Moon phase, moonrise, moonset and other Moon based data, sunset, sunrise, different twilight times and other Sun based data, zodiacal markers, solar and lunar eclipses, equinoxes and solstices. Other gimmicks like biorhythm calculation and the computation of a distance between two geographical point locations and their course angles (true tracks) are also done by it now, plus a lot more.

The accuracy of most of the astronomical calculations done by Gcal is low because it is really hard to perform accurate astronomical predictions more that 100 years ahead. Of course, this also has consequences for some country specific holidays which are based on such astronomical calculations. So it is really possible that some output may differ from the real astronomical date and time of the event calculated, and that by one or more minutes/hours/days. This can happen when the date of a real astronomical event occurs near around midnight. So don’t be surprised! Strictly speaking, take all such astronomically based output as a more or less good approximation of the true value, but do not rely on it!

Because Gcal is not primarily designed for the purpose of having a high precision astronomical calculation tool, the world currently has to live with these limitations. But if you have C source code functions, which

• perform the astronomical calculations more precise for the period that is respected by Gcal,

• have the similar or faster processing speed than Gcal’s actual functions,

• are either put into the public domain or are covered by the GPL^{2},

• have the same calling conventions and return types like Gcal’s actual functions,

• conform to the GNU coding standard and are therefore easy to maintain and extend,

• do not break Gcal internal limitations and concepts,

and you want to grant them to the public, I would be very pleased if you would contact me so I am able to implement your result to a next release of this software.

Some words to the build-in country specific holidays now…

As you certainly know, there are several other calendar systems actually existing around the world among the Gregorian calendar. The Gregorian calendar that Gcal mainly represents as a software is only one calendar of many others, but it is that one which is most respected around the world today. Even countries which use another calendar system officially mostly respect the Gregorian calendar for civil and administrative purposes.

But when Gcal offers the ability to include holidays from a definite country into its eternal holiday list, it is possible that not all holidays are displayed which are respected in that country. One reason is that my various sources simply don’t tell me about the existence of a holiday. A second reason is that the holiday is just designed or changed or removed by political or society reality during Gcal’s development phase. And another reason is that the method to calculate a holiday is not build-in Gcal yet. This case is marked by a ‘#’ character behind the country name in the list of countries, which are respected by Gcal. See [Calendar option --cc-holidays=cc[_tt][+…]], page 17, for the above mentioned list. The holidays of early years that are included into the eternal holiday list —i.e. those within the period AD 30 until the midth of the 20th century— are certainly not accomplishing the requirements of historians to such a list. Gcal cannot be understood as a complete and historically correct data base, and it is not designed for that purpose! This is caused by the fact that the territorial borders and names of the national structures and the settlement in earlier times was partly quite different as it is nowadays. Moreover, I am yet lacking reliable information that tells me the precise date of the introduction of a holiday in a region. Here I appeal to the historians to support me with the concrete country and culture based data if they are bothered by this fact and they want to have it removed.

1 I.e. no need to search for calendar information in the net so phone and Internet Service Provider (ISP)

costs are reduced, and furthermore, neither searching for such information in books or data files nor

the need to use several other software which altogether do what Gcal actually also does.

2 The “General Public License” of the Free Software Foundation.

Still unsupported alive and major calendar systems, respectively methods to derive country specific holidays from them, are:

• Religious Islamic lunar calendar

• Hindu solar and lunar calendars

• Buddhist calendar

• Sikh calendar

• Jain calendar

Even if the list of supported countries seems to be almost complete if you compare it with the actual list of all nations existing around the world, you should note that there are still some countries missing! If you can grant me support to complete Gcal’s list of respected countries, or if you would work out or have C functions which manage the missing calendar systems, I also would be very pleased about if you would contact me.

I have recorded the country specific holiday dates as good as I can, but I cannot guarantee that I did not made any mistakes while I hacked them in. So it is possible that some country specific holiday lists contain errors.

Well, if you think that I forgot to provide Gcal with the specific holidays of a/your country, or you think that a country specific holiday list is incomplete, or you think that some of the already implemented holidays are wrong, please feel free to send me a complete list of all the holidays which are celebrated in this country respectively to give me hints how to correct or to complete them!

Such a list should ideally include the following information:

• The English name of the holiday.

• The rule how it is built (like 10th October each year, or 47 days after Easter Sunday’s date, or three Fridays before the last Saturday in June…). If there are any exceptions from such a general rule, please state them in a detailed manner. And please tell me the date since what time the holiday is or was celebrated for the first time, respectively. By the way, if the holiday is tied to the Easter Sunday’s date, please tell me whether you mean the Easter Sunday of the Christian Western churches or the Easter Sunday of the Orthodox Christian Eastern churches.

• Whether it is a legal holiday which is respected in the whole country. (Use a ‘+’ character for marking this case.) In this sense, a legal holiday means a day people don’t have to work — as in most companies, and businesses are closed; like ‘New Year's Day’ on January 1st.

• Whether it is a legal holiday in some parts of your country only. Please tell me whether such a holiday is valid in major parts of the country or not, try to express that by using percent points. (If such a holiday is celebrated in more than 70% of the country, use a ‘#’ character for marking this case, otherwise use a ‘*’ character.)

• Whether it is a day only used for memorial or remembering purposes. (Use a ‘-’ character for marking this case.) Like ‘All Fool's Day’ on April 1st is not a holiday in the sense that anyone has a day off from school or work, just a play jokes day, pretty ordinary. Other holidays may be celebrated with cards or presents etc. — but they are working days such as:

         Valentine's day, Father's day, Mother's day.

The Islamic holidays inserted into the country specific holiday list are calculated arithmetically, not astronomically. This is what people call the civil Islamic lunar calendar. This type of calendar is just a rough estimate of the religious Islamic lunar calendar, which is determined quite astronomically. So don’t be surprised too much if some holiday dates differ from the religious Islamic calendar about ± 1… 2 days in some years. The main difficulty to implement the religious Islamic calendar correctly into Gcal are the several different criteria that are existing for the first sighting of the crescent Moon after a New Moon phase — that event, which defines the beginning of a new (lunar) month in the religious oriented Islamic calendar. So on the one hand, there are quite astronomically-based criteria, and on the other hand, there are also quite religiously-based criteria. Furthermore, the local predominant weather conditions, the geographical co-ordinate of the observer’s location and the kind of the observance method used also affect these criteria. Moreover, to render all more difficult, I as yet do not have any reliable information that tells me which kind of criteria is used in the countries (around 60 world-wide) which respect this religious Islamic calendar. So due to all the above mentioned facts of problematic nature, this religious Islamic calendar is not included into Gcal yet.

The Chinese (and Japanese) holidays also inserted into the country specific holiday list are calculated by using that algorithm as it is used by the Purple Mountain Observatory in Nanjing to determine the astronomical lunisolar Chinese calendar. This calendar is obligatory for (the Peoples Republic of) China since 1929 and is determined annually always new. Nevertheless and strictly speaking, in the historic reality and depending on the region, the Chinese calendar, especially in the 17th and 18th century, was either calculated using this modern algorithm, or by the use of a different method, so that a different calendar structure for one and the same year was possibly calculated as result.

And last…

This manual shows the possible use of the Gcal program only briefly and deliberately renounces to present the topics that are covered by this software arranged in a novel-like style. The use of Gcal is illustrated in an exemplary manner only! A detailed description of all possible use is inexpedient here because there are simply too many of such possibilities which are caused by the existing variety of combinations of all useable modes of operation.

Furthermore, this manual does not contain detailed explanations about the calendar systems supported by Gcal, nor it can be understood as a reference-book for the basics in celestial mechanics or spherical trigonometry! A detailed discussion of the above topics is completely beyond the scope of this manual. If you are interested to learn more about these subjects, you unfortunately have to read the specific pertinent literature.

Any further errors occurring in the output of the above mentioned holiday lists and in the astronomical calculations are my own fault, and are not intended to offend members of any culture, religion or profession!

2 Gcal Introduction

Apart from the usual and well known calendar functions like the output of a month or a year calendar sheet, or the output of an eternal holiday list, Gcal, the Gregorian calendar program of the Free Software Foundation, offers the facility to display fixed dates on the day of their occurrence and to remind or inform the user about them. So it is imaginable after booting the computer or starting the work session, that the user is informed on screen or by means of electronic mail about all holidays or appointments which are observed or scheduled for that day.

The period, for which Gcal respects occurring fixed dates, may be freely selected by the user. So it is possible that Gcal displays all fixed dates which occur on tomorrow’s date, the whole week, the whole month or in the whole year. Fixed dates which occur on a selected date of the year, and those that occur relative to another given date, are displayed either related to this single date only, or in listed manner starting on this date and ending on the actual date^{1}.

There are two methods to display a preview of fixed dates^{2} or retrospective view of fixed dates^{3}. On the one hand, Gcal can be started by using an option that sets the system date of the computer to the given date during the time of the program execution with the result, the program assumes the system date is set to this given date and the user can define any needed period that should be respected by an option. On the other hand, Gcal can be started with a command which forces the program to use a different year instead of the actual year, so Gcal will display all occurring fixed dates for this particular year. But this limits the user in that it disables defining any needed period by an option, because the period is always set to the whole year by default.

Gcal isn’t only able to display fixed dates which are stored for a concrete date, e.g. ‘Fixed date on 1st December 1995’, rather than fixed dates occurring periodically again and again. So it is possible to define repeated events like ‘This fixed date occurs every

day in May 1995’ or ‘Every 15th November in any years’. These fixed date definitions are

stored in resource files and whenever Gcal is started, an option to evaluate the necessary resource files can be given.

Once the user has set his/her preferred command line arguments for querying the fixed dates data base, it is possible to store them in a response file or shell script file. A response file contains all arguments delivered to Gcal, but unlike a shell script file, such a response file isn’t executable; it is only a pool of command line arguments which can be preloaded if needed. A shell script file can be started and calls Gcal directly with all arguments stored in it and all arguments which are given further in the command line.

A list of all usable command line arguments and their descriptions can be found in the next chapter, which helps one to use Gcal in the most efficient and productive way possible. After it follows the description how to use the eternal holiday list, and in the succeeding chapter, how to use the fixed date list. Comprehensive explanations respectively summaries and tables to definite details or themes from other disciplines can be found in the numerous appendices.

1 Inclusive or exclusive.

2 Future years.

3 Past years.

3 Invoking gcal

Gcal is a command line oriented program. It is usually called from the shell^{1} and processes given arguments that are options and commands. Options must be given before commands, i.e. you must call Gcal like this:

                   gcal [ [option. . .] [%date] [@file. . .] ] [command]

If Gcal is started without any options or commands, a calendar of the current month is displayed. If the calendar of a definite year is wanted, the year must be fully specified, e.g. ‘gcal 94’ displays a year calendar of the year 94, not of the year 1994.

If two arguments are given in the command part, the first argument denotes the month, and the second argument denotes the year. In case any incorrect commands are given running Gcal, the program will use internal defaults.

In the English program version, Gcal assumes the Gregorian Reformation has occurred in 1752 on the 3rd of September. See Appendix A [Genesis of the Gregorian Calendar], page 125, and Appendix C [Aspects in Internationalization], page 131, for further details.

1 This is an operating system program which interprets and runs given command lines.

3.1 Command line arguments

This section describes all command line arguments processed by Gcal. Four different types of command line arguments exists. One important type of arguments are the options which control how Gcal behaves. Other types of arguments are the %date and the @file options. The %date option sets the period Gcal shall work on to any starting date; the @file option preloads options and commands from a response file. The most important arguments are the commands which control the periods Gcal respects.

An option is defined by a leading switch character; either the ‘-’ (dash) or the ‘/’ (slash) character for traditional short-style options, or ‘--’ for mnemonic long-style options; a command may not have a leading switch character. Options^{1} must be given before commands!

Depending on operating system and used shell, some of the arguments and texts given in command line must be quoted by ‘"’ or ‘'’ characters respectively protected or depreciated by a ‘\’ character to avoid expansion by the shell.

Here is an incomplete list of characters which must potentially be protected:

‘(’, ‘)’, ‘<’, ‘>’, ‘[’, ‘]’, ‘{’, ‘}’, ‘\’, ‘|’, ‘$’, ‘@’, ‘!’, ‘&’, ‘~’, ‘"’, ‘'’, ‘`’, ‘;’

1 Inclusive %date and @file.

3.1.1 Options

The options processed by Gcal can be grouped into four major option classes. The options of the common option class are the standard options all GNU software should implement at least partially. The global option class contains options which affect the program output. The options of the calendar option class control the calendar layout, and the options of the fixed date option class control the fixed date layout and intensity.

Gcal supports both short-style options and GNU long-style options. Traditional shortstyle options are indicated by a single switch character, and trailed by the option character itself and perhaps a modifier or an argument. The most single character options^{1} can be composed into a single command line word: -Ax is equivalent to -A -x. GNU longstyle options are indicated with ‘--’, and trailed by the mnemonic option name itself and perhaps an argument. Long-style options and their arguments may be abbreviated if done unambiguously. When a long-style option takes an argument, connect the option name and the argument with ‘=’.

Brackets ([ and ]) indicate in the following tables, that an option takes an optional argument. The ‘|’ character is used to separate several arguments from each other.

Gcal processes the GNU long-style options in a special, non-standard way. There are five different types of long-style options:

 1. --foo

Enables option --foo.

 2. --foo=bar

Enables option --foo with the required argument bar.

 3. --foo[=bar[|. . .|bar]]

Option --foo may have one bar argument. If no argument list is given, any argument can be given to this option. If an argument list is given, exactly one bar argument may be selected from the given list. If there is no argument chosen in this case, the first bar argument of the argument list is preselected by default.

 4. --foo=bar|. . .|bar

Option --foo requires exactly one bar argument which must be selected from the given argument list.

 5. --foo=bar|. . .|bar|baz

Option --foo requires exactly one bar argument which must be selected from the given bar argument list, or the alternative baz argument.

Traditional short-style options differ as follows:

 1. -x

Enables option -x.

 2. -x bar

Enables option -x with the required argument bar. The bar argument may be separated by a leading whitespace character from the short-style option character x. This means, the following notations are valid for giving an argument, namely -x bar or -xbar.

 3. -x[bar|. . .|bar]

Option -x may have one or more bar modifier. In this sense, modifiers are one or more characters which define a special mode of operation enabled by the -x option. A modifier may not be separated by a leading whitespace character from the short-style option character.

1 Unless they take an argument.

3.1.1.1 Common options

-?
-h

--help Print a short usage message listing only some few of all available options, then

exit successfully.

-??
-hh
--usage[=argument]
--long-help[=argument]

Print an extended usage message listing all available options, then exit successfully. If an argument is given and it is a valid long-style option name, an extended help text related to the given long-style option name is displayed only, e.g.:

               --long-help=long-help

displays the extended help text for the long option --long-help.

If the argument only consists of the single ‘?’ character or is no valid long-style option name, a list of all valid long-style option names is displayed.

-L
--license
--copyleft
--copyright

Print the software license message, then exit successfully.

-V
--version

Print the version number and compilation options, then exit successfully.

--exit-status-help-non-zero

Set the exit state of program to 127 instead to 0, if one of the other options of the common option class is used.

3.1.1.2 Global options

-R name
--response-file=name

Write the contents of the environment variable GCAL (see [Environment Variable GCAL], page 255), and then the arguments of command line (in the given order) to file name, i.e. create response file name. See Section 3.1.2 [Response file], page 65, for more details.

-S name
--shell-script=name

Write the contents of the environment variable GCAL (see [Environment Variable GCAL], page 255), and then the arguments of command line (in the given order) to file name, i.e. create shell script file name. An automatically created shell script file is executable and calls Gcal directly with the arguments stored in it. You may start the shell script with other command line arguments which are directed to Gcal, too.

--debug[=internal|handled|unhandled|all|abort]

Display some debug information.

--debug=internal

Display informational texts if program internal maxima are reached or other conditions occurred, respectively.

--debug=handled

Like --debug=internal, additionally display the file names which can be processed respectively handled.

--debug=unhandled

Like --debug=internal, additionally display file names which cannot be processed respectively handled.

--debug=all
          Like --debug=handled and --debug=unhandled together.
--debug=abort

Like --debug=all and abort program with an error code if the file name cannot be handled or other unmanageable conditions occurred, respectively. See Appendix M [Error Codes], page 257.

-p

--pager Enables either an external pager or a simple, internal pager. If an environment variable PAGER is set, its contents will be used for detecting the external pager program. See [Environment Variable PAGER], page 256, for more information. If no PAGER environment variable is set or if its contents is invalid, Gcal tries to use the less pager; if this program cannot be found during scanning the PATH environment variable, Gcal tries to use the more pager, if this program cannot be found, the pg pager in the same way^{2}. See [Environment Variable PATH], page 256. If all these actions fail, Gcal will use its simple, built-in pager. If the internal pager is used, Gcal detects the number of lines shown before it prompts and waits for user input by using these methods:

1. Gcal respects the values set in the environment variables GCAL_LINES and GCAL_COLUMNS. See [Environment Variable GCAL_LINES], page 255, and [Environment Variable GCAL_COLUMNS], page 255, for further information.

2. If above action fails, Gcal respects the values set in the environment variables LINES and COLUMNS. See [Environment Variable LINES], page 255, and [Environment Variable COLUMNS], page 255, for more details.

3. If above action fails, Gcal respects the values set in the termcap^{3} file which refers to the terminal used (see [Environment Variable TERM], page 256).

2 See the standard manual pages for less, more and pg.

3 See the standard manual pages for Termcap.

This step is only done on systems which support the use of Termcap by default. On MS-DOS, OS/2 and some other operating systems, Gcal uses a system dependent low-level function and respects the reported values.

4. If all above actions have failed, Gcal uses default values^{4}.

-H yes
--force-highlighting
--highlighting=yes

If the output of the program is redirected^{5} or piped^{6}, the highlighting sequences are not automatically converted into the according marking characters, they remain unchanged. This option takes no effect if the output of the program is sent by means of electronic mail. See [Global option --mail[=address]], page 14.

-H no
--disable-highlighting
--highlighting=no

Disable highlighting sequence / marking character pairs of current day, holiday or text explicitly.

-H text
--highlighting=text

Set highlighting sequence / marking character pairs explicitly. In this sense, highlighting sequences are control character sequences which cause a color or intensity switch in output text. Typical control character sequences are the ANSI escape sequences which have a leading escape character, and trailing more characters that define the type of the ANSI escape sequence. In this sense, marking characters are single, printable characters which lead and trail the output text. The text argument must be a (‘:’) colon-separated text which is structured in this way: seq1 start:seq1 end:seq2 start:seq2 end. The first sequence is used for highlighting/marking an actual day, the second for a holiday. The sequences must be given in form of a sequence pair; seq? start enables the highlighting/marking, seq? end disables it. Only two sequence pairs will be processed, others are ignored. Either highlighting sequence pairs or marking character pairs may be defined, i.e. using them both in a mixed couple is not permitted! For example:

     -H \x20:\x20:\x1:# respectively
     --highlighting=\x20:\x20:\x1:#

marks the actual day like ‘\x20actual date\x20’^{7} and the holiday date like ‘\x1holiday date#’ using the given marking characters.

4 Either 23 or 24 lines, and 80 columns.

5 This means, sent to another device.

6 This means, used as an input data stream for another program.

7 This means with a leading and a trailing blank.

     -H \x1b[34;42m:\x1b[0;40m or
     -H \033[34;42m:\033[0;40m or
     -H \E[34;42m:\E[0;40m

defines a starting ANSI escape highlighting sequence ‘\x1b[34;42m’ used for actual day and ending ANSI escape highlighting sequence ‘\x1b[0;40m’ with no given highlighting sequence for holidays, so default highlighting sequences for holidays are used (non-given entries are always skipped). Please note the last abstract of this text part which informs you more detailed of this context. See [Environment Variable GCALANSI], page 255, too.

Control code definitions may contain any printable characters. Non-printable characters may be encoded in octal or hexadecimal notation. The abbreviation ‘\E’ directly encodes the escape character (octal \033 respectively hexadecimal

\x1B).

A character can be encoded octal by typing ‘\nnn’ (backslash-octal digit(s)), where n must be a valid octal digit (0…7). Normally, three octal digits must be given. If the octal character code consists of one or two octal digits, leading zeroes must be added, except the case, where the encoded octal character is given at last in the single sequence. A character can be encoded hexadecimal by typing ‘\xnn’ (backslash-x hexadecimal digit(s)), where n must be a valid hexadecimal digit (0…9A…Fa…f). Normally, two hexadecimal digits must be given. If the hexadecimal character code consists of one hexadecimal digit, a leading zero must be added, except the case, where the encoded hexadecimal character is given at last in the single sequence. If the sequence separator character, thus the ‘:’ (colon) character itself, is used as a marking character, it must be encoded either octal by \072 or hexadecimal

by \x3A.

If the C Preprocessor symbol USE_PAGER was defined and the output of the program is redirected or used in a pipeline, the highlighting sequences are automatically converted into the according marking characters; if USE_PAGER was not defined, they remain untouched. Incomplete or non-given highlighting sequences will be replaced by internal default ANSI escape highlighting sequences if a GCALANSI environment variable is defined; otherwise completely replaced by their according marking characters. See [Environment Variable GCALANSI], page 255.

--mail[=address]

Send Gcal’s output via mail^{1} program to the given address, e.g.:

     --mail=esken@gmx.net

If no address is given, Gcal tries to send the eMail by using the following methods:

1. If an environment variable MAILTO is defined and set, the eMail is send to the address which is listed in this environment variable. See [Environment Variable MAILTO], page 256, for more information.

2. If above action fails, and if an environment variable USER is defined and set, the eMail is send to the address which is listed in this environment variable. See [Environment Variable USER], page 256, for more information.

3. If above action fails, and if an environment variable LOGNAME is defined and set, the eMail is send to the address which is listed in this environment variable. See [Environment Variable LOGNAME], page 256, for more information.

4. If all above actions have failed, no eMail is send.

Generally, Gcal does not send electronic Mails whose message body is empty! An informational message will be shown on the standard error channel if this case occurs.

All highlighting sequences produced by Gcal itself are always disabled respectively automatically converted into the according marking characters if an eMail must be send; no matter if the --force-highlighting option was given or not. This behavior of Gcal is an imperative necessity, because it is possible that the mail program cannot perform the mailing correctly. Please pay attention in this context to the further explanations concerning the limitations of the text part of a resource file line (see Section 5.1.4 [Text part of a line], page 97).

If an environment variable MAILPROG is defined and set, its contents will be used as the program name of the mailer instead of the standard name mail. See [Environment Variable MAILPROG], page 256, for more information.

3.1.1.3 Calendar options

-n|N[-]
--holiday-list[=long|short]
--descending-holiday-list[=long|short]

Display the eternal holiday list. By default, there are no entries in the eternal holiday list. You have to choose country specific holidays or holidays from other calendar systems to provide the eternal holiday list with entries. See Chapter 4 [Eternal Holidays], page 77, and [Calendar option --cc-holidays=cc[+…]], page 17, for additional information.

1 See the standard manual pages for mail.

-n
--holiday-list=long

Display all holidays of eternal holiday list —this means, all legal holidays and all further memorial days— sorted in ascending order.

-n-
--descending-holiday-list=long

Display all holidays of eternal holiday list —this means, all legal holidays and all further memorial days— sorted in descending order.

-N
--holiday-list=short

Display legal holidays only of eternal holiday list, sorted in ascending order.

-N-
--descending-holiday-list=short

Display legal holidays only of eternal holiday list, sorted in descending order.

-G
--suppress-holiday-list-separator

Suppress displaying of the blank line which is always leading an eternal holiday list.

-X
--exclude-holiday-list-title

Suppress the title text line of the eternal holiday list.

--astronomical-holidays

Provide the eternal holiday list additionally with some astronomical data, and that the Full and New Moon phases, waning and waxing Half Moon phases, solar and lunar eclipses, and the solstices and equinoxes. See [Calendar option --time-offset=argument], page 33, how to change the timezone respectively base time for which the astronomical data is calculated.

--bahai-holidays

Provide the eternal holiday list additionally with Bahá’ì holidays (only for dates after AD 1843).

--celtic-holidays

Provide the eternal holiday list additionally with Celtic holidays. See [Calendar option --time-offset=argument], page 33, how to change the timezone for which the Celtic holidays are calculated.

--chinese-flexible-holidays

Provide the eternal holiday list additionally with Chinese holidays, that are determined in a flexible manner (only for dates after AD 1644). See [Calendar option --time-offset=argument], page 33, how to change the timezone for which the Chinese holidays are calculated in a flexible manner. See Chapter 1 [Preface], page 1, for further details.

--chinese-holidays

Provide the eternal holiday list additionally with Chinese holidays (only for dates after AD 1644). For dates until AD 1928 all computations done are depending fixed on Beijing local time, for later dates fixed on the timezone GMT-8. See Chapter 1 [Preface], page 1, for further details.

--christian-holidays

Provide the eternal holiday list additionally with Christian holidays.

--hebrew-holidays

Provide the eternal holiday list additionally with Hebrew holidays.

--islamic-civil-holidays

Provide the eternal holiday list additionally with Islamic holidays (only for dates after AD 621), that are based on the civil Islamic calendar. See Chapter 1 [Preface], page 1, for further details.

--japanese-flexible-holidays

Provide the eternal holiday list additionally with Japanese holidays, that are determined in a flexible manner (only for dates after AD 1644). See [Calendar option --time-offset=argument], page 33, how to change the timezone for which the Japanese holidays are calculated in a flexible manner.

--japanese-holidays

Provide the eternal holiday list additionally with Japanese holidays (only for dates after AD 1644). For dates until AD 1887 all computations done are depending fixed on Beijing local time, for later dates fixed on the timezone GMT-9.

--multicultural-new-year-holidays

Provide the eternal holiday list additionally with multicultural New Year holidays. See [Calendar option --time-offset=argument], page 33, how to change the timezone for which the multicultural New Year holidays are calculated.

--orthodox-new-holidays

Provide the eternal holiday list additionally with Orthodox new calendar holidays, and it is assumed that the Gregorian Reformation has occurred from 10th till 22nd March 1924. See [Calendar option --orthodox-calendar], page 30, for further details.

--orthodox-old-holidays

Provide the eternal holiday list additionally with Orthodox new calendar holidays, and it is assumed that the Gregorian Reformation has occurred from 10th till 22nd March 1924. See [Calendar option --orthodox-calendar], page 30, for further details.

--persian-jalaali-holidays

Provide the eternal holiday list additionally with Persian holidays (only for dates after AD 621), which are based on the Persian Jalaali calendar. All computations done are depending on the timezone GMT-3.5.

--zodiacal-marker-holidays

Provide the eternal holiday list additionally with zodiacal marker holidays, i.e. the dates when the Sun enters a zodiac sign or when the Sun reaches the turning-point in the zodiac sign. See [Calendar option --time-offset=argument], page 33, how to change the timezone for which the zodiacal marker holidays are calculated.

-q cc[_tt][+. . .]
--cc-holidays=cc[_tt][+. . .]

Provide the eternal holiday list additionally with country specific holidays. Furthermore, all additionally highlighted days of the eternal holiday list are highlighted in the calendar sheets, too. The cc argument is a two-letter country code as defined by the ISO-3166^{2} like ‘BE’ for Belgium or ‘ES’ for Spain. See the pertinent literature for more details. In some cases, such a country code may be trailed by a two-letter territory code tt for better specification, which is separated by a ‘_’ character from the country code. You can use more than one country code cc[ tt] by connecting them with a ‘+’ character, e.g.:

     --cc-holidays=be+Fr+IT resp.,
     -q be+Fr+IT

includes all the country specific holidays given in the preceding argument into the eternal holiday list, i.e. Belgian, French and Italian holidays. Actually, Gcal respects the following country codes, at which countries marked by a ‘#’ character have only an incomplete recording of holidays:

AD Andorra

AE United Arab Emirates

AF Afghanistan

AG Antigua and Barbuda

AI Anguilla

AL Albania

AM Armenia

AN_BO Netherlands Antilles/Bonaire

AN_CU Netherlands Antilles/Curaçao

AN_MA Netherlands Antilles/St Maarten

AN_SA Netherlands Antilles/Saba and Statia

AO Angola #

AR Argentina

AS American Samoa AT Austria AU_CT Australia/Canberra AU_NT Australia/Northern Territory AU_QU Australia/Queensland AU_SA Australia/Southern Australia AU_SW Australia/New South Wales AU_TA Australia/Tasmania AU_VI Australia/Victoria AU_WA Australia/Western Australia AW Aruba AZ Azerbaijan BA Bosnia-Herzegovina BB Barbados BD Bangladesh # BE Belgium BF Burkina Faso BG Bulgaria BH Bahrain BI Burundi # BJ Benin # BM Bermuda BN Brunei BO Bolivia BR Brazil BS Bahamas BT Bhutan # BV Bouvet Island BW Botswana BY Belarus BZ Belize CA_AL Canada/Alberta CA_BC Canada/British Columbia CA_MA Canada/Manitoba CA_NB Canada/New Brunswick CA_NF Canada/Newfoundland and Labrador CA_NS Canada/Nova Scotia CA_NW Canada/Northwest Territories CA_ON Canada/Ontario CA_PE Canada/Prince Edward Island CA_QU Canada/Québec CA_SA Canada/Saskatchewan CA_YU Canada/Yukon CC Cocos Islands (Keeling) CD Democratic Republic of Congo # CF Central African Republic # CG Republic of Congo # CH_AG Switzerland/Aargau CH_AI Switzerland/Appenzell Innerrhoden CH_AR Switzerland/Appenzell Ausserrhoden CH_BE Switzerland/Bern CH_BL Switzerland/Basel-Land CH_BS Switzerland/Basel-Stadt CH_FR Switzerland/Fribourg CH_GE Switzerland/Genève CH_GL Switzerland/Glarus CH_GR Switzerland/Graubünden CH_JU Switzerland/Jura CH_LU Switzerland/Luzern CH_NE Switzerland/Neuchâtel CH_NW Switzerland/Nidwalden CH_OW Switzerland/Obwalden CH_SG Switzerland/St Gallen CH_SH Switzerland/Schaffhausen CH_SO Switzerland/Solothurn CH_SZ Switzerland/Schwyz CH_TG Switzerland/Thurgau CH_TI Switzerland/Ticino CH_UR Switzerland/Uri CH_VD Switzerland/Vaud CH_VS Switzerland/Valais CH_ZG Switzerland/Zug CH_ZH Switzerland/Zürich CI Côte d’Ivoire CK Cook Islands CL Chile CM Cameroon CN China CO Colombia CR Costa Rica CU Cuba CV Cape Verde CX Christmas Islands CY Cyprus CZ Czech Republic DE_BB Germany/Brandenburg DE_BE Germany/Berlin DE_BW Germany/Baden-Württemberg DE_BY Germany/Bavaria DE_HB Germany/Bremen DE_HE Germany/Hesse DE_HH Germany/Hamburg DE_MV Germany/Mecklenburg-West Pomerania DE_NI Germany/Lower Saxony DE_NW Germany/North Rhine-Westphalia DE_RP Germany/Rhineland Palatinate DE_SH Germany/Schleswig-Holstein DE_SL Germany/Saar DE_SN Germany/Saxony DE_ST Germany/Saxony-Anhalt DE_TH Germany/Thuringia DJ Djibouti DK Denmark DM Dominica DO Dominican Republic DZ Algeria EC Ecuador EE Estonia EG Egypt EH Western Sahara ER Eritrea ES Spain ET Ethiopia FI Finland FJ Fiji # FK Falkland Islands (Malvinas) FM Federated States of Micronesia FO Faroes FR France GA Gabon # GB_EN Great Britain/England and Wales GB_NI Great Britain/Northern Ireland GB_SL Great Britain/Scotland GD Grenada GE Georgia GF French Guiana GH Ghana GI Gibraltar GL Greenland GM Gambia GN Guinea GP Guadeloupe GQ Equatorial Guinea GR Greece GS South Georgia and South Sandwich Islands GT Guatemala GU Guam GW Guinea-Bissau # GY Guyana # HK Hong Kong HM Heard and Mc Donald Islands HN Honduras HR Croatia HT Haiti HU Hungary ID Indonesia # IE Ireland IL Israel IN India # IQ Iraq IR Islamic Republic of Iran IS Iceland IT Italy JM Jamaica JO Jordan JP Japan KE Kenya # KG Kyrgyzstan KH Cambodia # KI Kiribati # KM Comoros KN St Kitts and Nevis KP Democratic People’s Republic of Korea # KR Republic of Korea KW Kuwait KY Cayman Islands KZ Kazakhstan LA Laos People’s Democratic Republic # LB Lebanon LC St Lucia LI Liechtenstein LK Sri Lanka # LR Liberia LS Lesotho LT Lithuania LU Luxembourg LV Latvia LY Libyan Arab Jamahiriya (Libya) MA Morocco MC Monaco MD Republic of Moldova MG Madagascar # MH Marshall Islands MK Former Yugoslav Republic of Macedonia ML Mali MN Mongolia # MO Macau MP Northern Marian Islands (Saipan) MQ Martinique MR Mauritania MS Montserrat MT Malta MU Mauritius MV Maldives MW Malawi MX Mexico MY Malaysia # MZ Mozambique NA Namibia NC New Caledonia NE Niger NF Norfolk Island NG Nigeria NI Nicaragua NL Netherlands NM Myanmar (Burma) # NO Norway NP Nepal # NR Nauru NU Niue NZ New Zealand OM Oman PA Panama PE Peru PF French Polynesia PG Papua New Guinea PH Philippines PK Pakistan PL Poland PM Saint-Pierre and Miquelon PN Pitcairn PR Puerto Rico PT Portugal PW Palau PY Paraguay QA Qatar RE Réunion RO Romania RU Russian Federation RW Rwanda SA Saudi Arabia SB Solomon Islands SC Seychellen SD Sudan SE Sweden SG Singapore # SH St Helena SI Slovenia SJ Svalbard and Jan Mayen Islands SK Slovakia SL Sierra Leone SM San Marino SN Senegal SO Somalia SR Suriname # ST Sao Tomé and Principe SV El Salvador SY Syrian Arab Republic (Syria) SZ Swaziland TC Turks and Caicos Islands TD Chad TG Togo # TH Thailand # TJ Tajikistan TK Tokelau TM Turkmenistan TN Tunisia TO Tonga TR Turkey TT Trinidad and Tobago # TV Tuvalu TW Taiwan TZ Tanzania UA Ukraine UG Uganda US_AK United States/Alaska US_AL United States/Alabama US_AR United States/Arkansas US_AZ United States/Arizona US_CA United States/California US_CO United States/Colorado US_CT United States/Connecticut US_DC United States/District of Columbia US_DE United States/Delaware US_FL United States/Florida US_GA United States/Georgia US_HI United States/Hawaii US_IA United States/Iowa US_ID United States/Idaho US_IL United States/Illinois US_IN United States/Indiana US_KS United States/Kansas US_KY United States/Kentucky US_LA United States/Louisiana US_MA United States/Massachusetts US_MD United States/Maryland US_ME United States/Maine US_MI United States/Michigan US_MN United States/Minnesota US_MO United States/Missouri US_MS United States/Mississippi US_MT United States/Montana US_NC United States/North Carolina US_ND United States/North Dakota US_NE United States/Nebraska US_NH United States/New Hampshire US_NJ United States/New Jersey US_NM United States/New Mexico US_NV United States/Nevada US_NY United States/New York US_OH United States/Ohio US_OK United States/Oklahoma US_OR United States/Oregon US_PA United States/Pennsylvania US_RI United States/Rhode Island US_SC United States/South Carolina US_SD United States/South Dakota US_TN United States/Tennessee US_TX United States/Texas US_UT United States/Utah US_VA United States/Virginia US_VT United States/Vermont US_WA United States/Washington US_WI United States/Wisconsin US_WV United States/West Virginia US_WY United States/Wyoming UY Uruguay UZ Uzbekistan VC St Vincent and Grenadines VE Venezuela VG British Virgin Islands VI U.S. Virgin Islands VN Viet Nam VU Vanuatu WF Wallis and Futuna Islands WS Samoa YE Yemen YT Mayotte YU Serbia and Montenegro ZA South Africa ZM Zambia

2 ISO is the abbreviation of International Organization for Standardization.

ZW Zimbabwe

--bahai-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Bahá’ì calendar. See [Calendar option --bahai-holidays], page 15, for further details.

--chinese-flexible-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Chinese calendar, that is determined in a flexible manner. See [Calendar option

--chinese-flexible-holidays], page 15, for further details.
--chinese-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Chinese calendar. See [Calendar option --chinese-holidays], page 16, for further details.

--coptic-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Coptic calendar (only for dates after AD 283).

--ethiopic-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Ethiopic calendar.

--french-revolutionary-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the French Revolutionary calendar (only for dates after AD 1791).

--hebrew-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Hebrew calendar.

--indian-civil-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the civil Indian calendar (only for dates after AD 1956).

--islamic-civil-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the civil Islamic calendar. See [Calendar option --islamic-civil-holidays], page 16, for further details.

--japanese-flexible-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Japanese calendar, that is determined in a flexible manner. See [Calendar option

--japanese-flexible-holidays], page 16, for further details.

--japanese-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Japanese calendar. See [Calendar option --japanese-holidays], page 16, for further details.

--old-armenic-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Old-Armenic calendar (only for dates after AD 551).

--old-egyptic-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Old-Egyptic calendar.

--persian-jalaali-months

Provide the eternal holiday list additionally with the starting dates of the common and leap months, as they result from the Persian Jalaali calendar. See [Calendar option --persian-jalaali-holidays], page 16, for further details.

-i[-]
--type=special|standard

To obtain the standard calendar format^{3}, either start Gcal omitting the -i[-] option because it is set by default, Appendix C [Aspects in Internationalization], page 131, or start Gcal with the -i- respectively --type=standard option:

      $ gcal -i-

⊣

      ⊣  September 1994
      ⊣ Su Mo Tu We Th Fr Sa
      ⊣  1  2  3
      ⊣ 4  5  6  7  8  9 10
      ⊣ 11 12 13 14 15 16 17
      ⊣ 18 19 20 21 22 23 24
      ⊣ 25 26 27 28 29 30

To obtain the special calendar format, start Gcal with the -i respectively

--type=special option:

3 Similar the BSD-cal program.

      $ gcal -i

⊣

      ⊣ September 1994

⊣

      ⊣ Sunday          4 11 18 25
      ⊣ Monday          5 12 19 26
      ⊣ Tuesday         6 13 20 27
      ⊣ Wednesday       7 14 21 28
      ⊣ Thursday     1  8 15 22 29
      ⊣ Friday       2  9 16 23 30
      ⊣ Saturday     3 10 17 24

-O
--orthodox-calendar

Use the leap year rule as used by the Eastern Orthodox churches.

Without specifying the --orthodox-calendar option, Gcal is unable to display Gregorian years later than 2799 in the correct way for the Eastern churches, because they use a different scheme for calculating the leap years. The method for computing leap years within the common Gregorian calendar, which Gcal uses by default, is as follows:

A leap year is any year which number can be divided by 4 without a remainder, and years ending in hundreds are no leap years unless they are divisible by 400.

But the Eastern Orthodox churches compute leap years within the Gregorian calendar by using another rule:

A leap year is any year which number can be divided by 4 without a remainder, and years ending in hundreds are leap years, if a remainder of 2 or 6 occurs when such a year is divided by 9.

The first difference therefore occurs in the year 2800 which is a leap year in the common Gregorian calendar, but an ordinary year only in the calendar as used by the Eastern Orthodox churches.

-K
--with-week-number

Provide the calendar sheet with week numbers. See [Calendar option --starting-day=argument], page 33, [Calendar option --iso-week-number=yes|no], page 30, and Appendix C [Aspects in Internationalization], page 131, for further details.

--iso-week-number=yes|no

Determine the type of week numbers which are used in the calendar sheet, in the fixed date list and by the %date actual date modifier. See [Calendar option --starting-day=argument], page 33, and Appendix C [Aspects in Internationalization], page 131, for more details.

--iso-week-number=yes

The methods of the ISO-8601:1988 are used for detecting week numbers; this means a week starts on Mondays, and the first week of a year is the one which includes the first Thursday; equivalently, the one which includes the 4th January. This method is called ISO week number in the further context. If the starting day of the week is not set to Monday, the week numbers are not represented correctly in most cases. If you use this option, you should take care of setting Monday as the starting day of the week!

--iso-week-number=no

Weeks start on the respective starting day of the week, and the days in a new year that are preceding the first starting day of the week are in the last week of the previous year, respectively in week zero of the new year. This method is called standard week number in the further context.

-u
--suppress-calendar

Suppress output of calendar sheet explicitly.

-b number
--blocks=number

Set number of calendar sheet blocks (valid arguments: 1|2|3|4|6|12). The default number for the standard calendar format is -b4 respectively --blocks=4, and for the special calendar format -b 3 respectively --blocks=3. If this option is found, the program sees that a year calendar output is desired!

-b 1
--blocks=1

Displays one block with twelve months at a time.

-b 2
--blocks=2

Displays two blocks with six months at a time.

-b 3
--blocks=3

Displays three blocks with four months at a time.

-b 4
--blocks=4

Displays four blocks with three months at a time.

-b 6
--blocks=6

Displays six blocks with two months at a time.

-b 12
--blocks=12

Displays twelve blocks with one month at a time.

-j[b]
--calendar-dates=special|both

Use alternative date format in calendar sheet instead of the default standard format which displays the days of month in consecutive manner.

-j
--calendar-dates=special

Display the calendar sheet by using the special date format. This means, the days of year are displayed in consecutive manner instead of the days of month.

-jb
--calendar-dates=both

Display the calendar sheet by using both the standard date format and special date format.

-jn[b]
--holiday-dates=special|both

Use alternative date format in eternal holiday list instead of the default standard format which displays the days of month in consecutive manner. See [Calendar

option --holiday-list[=long|short]], page 14.
-jn
--holiday-dates=special

Display the eternal holiday list by using the special date format. This means, the days of year are displayed in consecutive manner instead of the days of month.

-jnb
--holiday-dates=both

Display the eternal holiday list by using both the standard date format and special date format.

-jc[b]
--fixed-dates=special|both

Use alternative date format in fixed date list instead of the default standard format which displays the days of month in consecutive manner. See [Fixed

date option --list-of-fixed-dates[=short|long]], page 45.
-jc
--fixed-dates=special

Display the fixed date list by using the special date format. This means, the days of year are displayed in consecutive manner instead of the days of month.

-jcb
--fixed-dates=both

Display the fixed date list by using both the standard date format and special date format.

-s argument
--starting-day=argument

Set the starting day of the week (valid argument: 0, 1…7 | today | weekday name). For example:

     --starting-day=Sunday or
     --starting-day=7 or
     -s SUNDAY or
     -s sund or
     -sSu or
     -s 7

thus all specifies the Sunday (1≡Mon, 2≡Tue … 7≡Sun).

If the -s today option (or --starting-day=today) or the -s 0 option (or

--starting-day=0) is given, the starting day of the week is set to the actual weekday as it is delivered by the system date. See Appendix C [Aspects in Internationalization], page 131, for more details.

--time-offset=argument

Change the base time of the astronomical functions (valid argument:

t|@|[t|@][+|-]mmmm|hh:[mm]).

If no --time-offset=argument option is given, the astronomical data that is inserted into the eternal holiday list by the --astronomical-holidays option (see [Calendar option --astronomical-holidays], page 15), and all Sun and Moon related special texts are always calculated for 0 o’clock Universal time (UTC/GMT), thus civil midnight time. See Section G.2.2.5 [Sun data], page 165, and Section G.2.2.6 [Moon data], page 188, likewise Section G.2.1.9 [Moon phase], page 155, for further information. The argument is either the ‘t’ or the ‘@’ character —where ‘t’ means a relation to the actual local time^{4} and ‘@’ denotes a relation to the actual Universal time—, or one of these characters followed by a displacement value, or only a displacement value which has to be specified either by using the [+|-]mmmm format or the [+|-]hh:[mm] format. [+|-]mmmm adds respectively subtracts the specified amount of minutes mmmm from the base time value 0 o’clock Universal time (range 0…9999), while [+|-]hh:[mm] adds respectively subtracts the given amount of hours hh (range 0…99) and minutes mm (range 0…59) from the base time value 0 o’clock Universal time. The displacement value is always added to the base time value 0 o’clock Universal time in case it is specified without a +|- sign. For example, the --time-offset=+1: option causes that while displaying eternal holiday lists and fixed date lists, the time 0 o’clock of the timezone GMT-1 (≡ CET) is used as the base time by the astronomical functions instead of the base time 0 o’clock Universal time (≡ GMT). For example, the --time-offset=t-2: option causes that while displaying eternal holiday lists and fixed date lists, the time that is two hours earlier than the actual local time is used as the base time by the astronomical functions instead of the base time 0 o’clock Universal time (≡ GMT). For such a kind of relation, the term relative time offset value will be used in the further context.

4 The time which is set as system time in the computer’s clock.

--transform-year=argument

Change the base year of calendar (valid argument: -9999…[+]9999). For example, the --transform-year=-543 option causes that while displaying calendar sheets, eternal holiday lists and fixed date lists, the year 543 BC is used as the base year of the calendar instead of the year AD 1. This results in the case that —for example— for the year 1999 (Christian era), the year number 1999 is not used in the above mentioned outputs, but the year number 2542 which is used in the western oriented Thai calendar. Nevertheless, Gcal does not respect the --transform-year=argument option when using the actual date modifier %date (see Section 3.1.3 [Actual date modifier], page 67), the commands (see Section 3.1.4 [Commands], page 72), and the fixed date entries in a resource file (see Section 5.1.2 [Date part of a line], page 86). There, all references made to a definite year are always treated in the way that the year AD 1 is the base year of the calendar, so in fact it is only possible to use references based on the Christian era.

--gregorian-reform=1582|1700|1752|1753|argument

Set the period which was skipped during the Gregorian Reformation. By default, Gcal runs in the hybrid calendar mode, i.e. Gcal automatically changes from the Julian calendar system to the Gregorian calendar system if output is related to dates after the Gregorian Reformation has happened. See Appendix C [Aspects in Internationalization], page 131, for more details. Actually, four fixed default periods are supported, and that of the year 1582, of the year 1700, of the year 1752 and of the year 1753. If Gcal is called with the --gregorian-reform=1582 option, it assumes the Gregorian Reformation has occurred from 5th till 14th October 1582. If Gcal is called with the --gregorian-reform=1700 option, it assumes the Gregorian Reformation has occurred from 19th till 28th February 1700. If Gcal is called with the --gregorian-reform=1752 option, it assumes the Gregorian Reformation has occurred from 3rd till 13th September 1752. If Gcal is called with the --gregorian-reform=1753 option, it assumes the Gregorian Reformation has occurred from 18th till 28th February 1753. In case another period shall be respected, it can be arranged by the option argument like ‘yyyyy,mm,first-day,last-day’. If the Gregorian Reformation has occurred for example on the 7th till the 17th April 1802, this can be arranged as follows:

     --gregorian-reform=1802,4,7,17

Gcal is able to represent so-called proleptic calendars of a definite calendar system. This means, Gcal only uses a definite calendar system during a definite period, although there was a change to another calendar system in the historic reality during this definite period. The following proleptic calendar systems are actually supported by Gcal:

• Proleptic Gregorian calendars can be created from AD 302 until AD 9999 if --gregorian-reform=301,12,31,31 is used. Unfortunately, Gcal actually cannot represent years before AD 302 in a proleptic-Gregorian manner.

• Proleptic Julian calendars can be created from AD 1 until AD 9999 if

    --gregorian-reform=10002,12,31,31 is used. In this case, the actual

system date that is normally based on the Gregorian calendar and that is used by Gcal, is internally and automatically converted into the according Julian date^{5}. The same happens if the actual system date is modified by using the actual date modifier %date. In such a case, Gcal also assumes that the given date is a Gregorian date, and converts it automatically into the according Julian date internally. See Section 3.1.3 [Actual date modifier], page 67, for further details.

Please note that it is possible to corrupt the calendars likewise the fixed date feature logically (which works correctly now for the year in which the Gregorian Reformation has occurred) if the argument of the --gregorian-reform option is not used with care.

--date-format=de|us|gb|text

Set the date format which affects the ordering and representation of a displayed date. See Appendix C [Aspects in Internationalization], page 131, for more details. The date format text is respected by Gcal in the eternal holiday list, in the fixed date list and the calendar sheets. Moreover, Gcal internally tries to obtain the best representation of a displayed date in case the day-of-year numbers instead of the day-of-month numbers must be displayed, or both types of numbers are used in a combined manner. Actually, three fixed default date formats are supported, and that for German users, U.S. American users and for users in Great Britain. If Gcal is called with the --date-format=de option, the ‘%<2#K, %1%>2*D%2 %<3#U %>04*Y’ date format text is used. This result in that a date is displayed by using the ‘ww, dd mmm yyyy’ ordering, for example

‘Sa,  28 Aug 1999’.

If Gcal is called with the --date-format=us option, the ‘%<3#K, %<3#U %1%>2&*D%2 %>04*Y’ date format text is used. This result in that a date is displayed by using the ‘www, mmm dd yyyy’ ordering, for example

‘Sat, Aug  28th 1999’.

If Gcal is called with the --date-format=gb option, the ‘%<3#K, %1%>2&*D%2 %<3#U %>04*Y’ date format text is used. This result in that a date is displayed by using the ‘www, dd mmm yyyy’ ordering, for example

‘Sat,  28th Aug 1999’.

In case another format text shall be respected, this format text can either be set in the GCAL_DATE_FORMAT environment variable^{6}, or it can be arranged by the option argument text. For example,

5 The automatical conversion into the according Julian date is always done if the actual system date,

that is based on the Gregorian calendar, is before the date of the Gregorian Reformation.

6 But GCAL_DATE_FORMAT is only used in case no --date-format=de|us|gb|text option is given.

‘--date-format='%Y %D %>02*M ; %1(%>5u#K)%2'’  displays  a  date

by using the ‘[[[y]y]y]y [d]d mm ; (wwwww)’ ordering, thus for example

‘1999 28 08 ; (SATUR)’. See [Environment Variable GCAL_DATE_FORMAT],

page 255, for further information. The format text may contain on the one hand all characters which can be managed by Gcal, and on the other hand character replacement instructions and format elements which are transformed into their according values at runtime. Some format elements may have a format instruction which is called format in the further context. See Appendix J [Format Instruction], page 249, for the detailed description of the format instruction and its components. A minimum date format text must contain the following format elements minimum:

The day number, one component of the month group, the year number and both components of the highlighting group.

A weekday name format element may be included optionally into the date format text. The following format elements and character replacement instructions are currently supported:

%[format]D

Day number (must be defined)

%[format]Y

Year number (must be defined)

%[format]K

Weekday name (may be defined)

Month group (exactly one member must be defined):

%[format]M

Month number

%[format]U

Month name

Highlighting group (all members must be defined and %1 must be specified before %2):

%1 Start of highlighting sequence / marking character

%2 End of highlighting sequence / marking character

Character replacement instructions:

_Space/blank character ‘ ’

\_Underscore character ‘_’

\% Percent character ‘%’

\\ Backslash character ‘\’

See Section I.3 [Table of Obsolete Date Formats], page 247, and Section I.2 [Table of Obsolete Date Format Elements], page 247, for further information.

--translate-string=text

Define the country specific special character pairs which are respected or translated by a style format instruction component, respectively. See Appendix J [Format Instruction], page 249, for the detailed description of the format instruction and its components. The country specific special character pairs are arranged by the text option argument as a sequence of single character pairs, and that in any number and order. The upper-case representation of the country specific special character has to be specified at first, and after that its lowercase representation. Country specific special characters which do not have an upper-case resp., lower-case representation in the character set used, like e.g. the ‘ß’-character which is very usual in the German character set, are also specified as a special character pair (here: ‘ßß’), otherwise these special characters are not recognized and they are converted incorrectly by the style format instruction component. In case the lower-case representation of the country specific special character is specified at first, and after that its upper-case representation, this option does not cause any further affects to the style format instruction component; resulting, the country specific special characters specified are not recognized as such and they remain untranslated therefore.

For example, a ‘ÄäÖöÜüßß’ option argument causes the correct conversion of the preceding special characters in an individual date format, which has a style format instruction component, and that, how they are used by the character set used in Germany.

3.1.1.4 Fixed date options

-v argument
--date-variable=argument

Define global date variable which is visible in all referenced resource files (see Section 5.1.8 [Date variables], page 103). The option argument must either be a valid single date variable definition or a (‘:’) colon-connected series of definitions.

A date variable name dvar consists of a single, case-insensitive letter^{7}. In a definition, this name is trailed by the assignment operator character ‘=’, and the numerical values of month mm (or a short, three characters month name, e.g. ‘Jan’, ‘Feb’…) and day dd, this variable has to store^{8}.

For example:

     --date-variable=a=1127:c=a respectively
     -v a=1127:c=a

stores the 27th November into the global date variable a and into the global date variable c.

7 Except the date variable e which is internally reserved for the Easter Sunday’s date, so it cannot be

assigned or operated therefore. And except the date variable t which is internally reserved for today’s date, so it likewise cannot be assigned or operated.

8 Or another global date variable name which is already defined, but no date given in the

‘e|t|dvar[+|-]n[www]’, ‘mmwwwn’, or the ‘*d|wn[www]’ format.

--export-date-variables

Causes the export of local date variables from one resource file to another resource file. See Section 5.1.8 [Date variables], page 103, for more details.

-r argument
--text-variable=argument

Define global text variable which is visible in all referenced resource files (see Section 5.1.9 [Text variables], page 106). The option argument must either be a valid single text variable definition or a (‘:’) colon-connected series of definitions. A text variable name tvar consists of the ‘$’ prefix and a single, case-insensitive letter. A text variable is defined as follows:

• If the name of a text variable is trailed by the assignment operator character ‘=’, all the text which appears at the right side of this assignment operator is assigned to this text variable.

• If the name of a text variable is trailed by the assignment operator character ‘?’, all the text which appears at the right side of this assignment operator is assumed as an external command and executed by Gcal. The output, which is created on the standard output channel by the command run, is assigned to this text variable; and that in an interpreted manner. This means, Gcal interprets all special texts and references to other text variables which are contained in this output. See Section 5.1.9 [Text variables], page 106, and Section 5.1.5 [Special Texts processed], page 100, for further information.

• If the name of a text variable is trailed by the assignment operator character ‘:’, all the text which appears at the right side of this assignment operator is assumed as an external command and executed by Gcal. The output, which is created on the standard output channel by the command run, is assigned to this text variable; and that in an uninterpreted manner. This means, Gcal does not interpret any special texts and references to other text variables which are contained in this output. See also Section 5.1.9 [Text variables], page 106, and Section 5.1.5 [Special Texts processed], page 100, for further details.

Indeed, Gcal executes external commands only if the --execute-command option is given at program start-up. See [Fixed date option --execute-command], page 43, for more details. The text which appears at the right side of one of the valid assignment operator characters may reference names of already defined global text variables. These references are converted to the according textual values just before the assignment is performed by Gcal. A global text variable is defined with an empty contents in case no text appears at the right side of one of the valid assignment operator characters. For example:

     --text-variable=$a=foo:$c=$a respectively
     -r $a=foo:$c=$a

stores the ‘foo’ text into the global text variable $a and into the global text variable $c.

     --text-variable=$a\:foo:$c?bar respectively
     -r $a\:foo:$c?bar

stores the uninterpreted output of the foo command into the global text variable $a, and the interpreted output of the bar command into the global text variable $c.

You may depreciate the special meaning of the ‘:’ separator character —in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\:’. If you need the ‘\:’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\:’.

You may depreciate the special meaning of the ‘$’ prefix character —in case this character itself is needed in the text— by placing a ‘\’ (backslash) character before it, e.g. ‘\$’. If you need the ‘\$’ characters themselves in the text, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\$’.

Each time a ‘_’ (underscore) character is found in argument, this character is replaced by a real ‘ ’ (space) character.

You may depreciate the special meaning of the ‘_’ (underscore) character — in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\_’. If you need the ‘\_’ themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\_’.

--export-text-variables

Causes the export of local text variables from one resource file to another resource file. See Section 5.1.9 [Text variables], page 106, for more details.

--adjust-value=argument

Causes on the one hand, that the number given as argument to the --adjust-value option is used as reference altitude to calculate the rise and set times by the Sun and Moon oriented special texts. The option argument must be a rational number in range -90.0…+90.0, which declares the angle distance in decimal degrees from the center of astronomical object to the horizon. For example, you can detect the time when the center of the Sun or the Moon is 20.5 decimal degrees below the horizon if you use

--adjust-value=-20.5.

And causes on the other hand, that the number that is given as argument is used as factor for computing the shadow length of a vertical pole by the Sun oriented special texts. The option argument must be a rational number in range 0.0…+90.0, which declares the shadow length. For example, you can detect the time when the shadow length of vertical pole is three and a half times longer than the shadow length that appears at astronomical noon if you use

--adjust-value=3.5.

See Section G.2.2.5 [Sun data], page 165, and Section G.2.2.6 [Moon data], page 188, for further information.

--atmosphere=air-pressure[,temperature]

Set the base data of the Earth’s atmosphere that is necessary to calculate the amount of refraction^{9}, that is used by the Sun and Moon oriented special texts. The option argument consists of one or two rational numbers, which are separated by a ‘,’ character. The first number denotes the air pressure in millibar, and the second number specifies the air temperature in degree Celsius. The default value for the air pressure is 1013.25 millibar, and the default value for the temperature is 15.0 degree Celsius. For example, if you use --atmosphere=1010,10, you define atmospheric conditions as they are existing for an air pressure of 1010 millibar and a temperature of 10 degree Celsius. If you set an air pressure value less or equal zero, no atmospheric refraction will be respected by the Sun and Moon oriented special texts, as for example done

by --atmosphere=0.

See Section G.2.2.5 [Sun data], page 165, and Section G.2.2.6 [Moon data], page 188, for further information.

--limit Causes when calculating the rise and set times by the Sun oriented special texts for high latitudes, that a rise or set event which possibly occurs on a previous or next day, is not displayed as occurring on the actual day as it is done by default. At high latitudes, i.e. above the solar Arctic Circle, it is fact that no, or more than one sunrise or sunset happens during a day at definite dates of the year. As a special text basically leads in a single, atomical output, even it should result in several output in reality, Gcal puts such an event on a previous or next day. Then, data represented in this manner could be misinterpreted very easily.

The following example is used to explain this. It calculates the time the astronomical twilight ends for the location Münster, Germany, during some days in May 2000, and that with and without the --limit option in each case:

9 Refraction is an optical phenomenon caused by the Earth’s atmosphere, which leads to an apparent

raising of the location of a celestial body. The amount of refraction increases with growing air pressure and sinking temperature and vice-versa.

      $ gcal %20000506 -f/dev/null -cxl4+ \
      >      -#'0*d1#999_%s9+5158+00738,120'

⊣

      ⊣ Sun, May   7th 2000: 23:52
      ⊣ Mon, May   8th 2000: 23:57
      ⊣ Tue, May   9th 2000: 00:01
      ⊣ Wed, May  10th 2000: 00:06
      $ gcal %20000506 -f/dev/null -cxl4+ --limit \
      >      -#'0*d1#999_%s9+5158+00738,120'

⊣

      ⊣ Sun, May   7th 2000: 23:52
      ⊣ Mon, May   8th 2000: 23:57
      ⊣ Tue, May   9th 2000: **:**
      ⊣ Wed, May  10th 2000: **:**

As one can see well in the above example, the output for the days since 9th May differs in that the times shown without the --limit option obviously exceeded the civil midnight time and actually cannot happen during these days, whereas no times are shown for the days since 9th May if the --limit option is used because the time circle is exceeded. Astronomically seen, it is the case that the end of the astronomical twilight is at 23:57 o’clock local time on 8th May, no end of astronomical twilight happens on 9th May because the center of the Sun’s disk is always above 18 degrees (the reference altitude of the astronomical twilight) below a geometric horizon, but on 10th May at 00:01 o’clock local time, this reference altitude is passed again. Thus, the time that is detected for the 9th May has properly to be moved to the 10th May, et cetera, as it is expressed by the output that is created by using the --limit option. See Section G.2.2.5 [Sun data], page 165, for further information.

--cycle-end=argument

Calculates the Sun and Moon oriented special texts that are dynamical, i.e. depending on the respective clocktime, for a series of clocktimes for the current day, i.e. determination is done for any time interval that can be individually defined by the user. Gcal is always put into such a cycle mode if the --cycle-end option is given at program start-up. The argument of the --cycle-end option is syntactically according to the argument which may be given to the --time-offset option, thus t|@|[t|@][+|-]mmmm|hh:[mm]. See [Calendar option --time-offset=argument], page 33, for the detailed explanation of the preceding option argument. Thus, the --cycle-end option is used to fix the ending time of the time interval, whereas the --time-offset option is implicitly used to fix the starting time of the time interval, that is 0 o’clock by default^{10}. If values for the time interval are given to both above mentioned options that fall short of 0 o’clock, these values are automatically set to 0 o’clock. If values for the time interval are given that exceed 23:39 o’clock, these values are automatically set to 23:59 o’clock.

10 Correct, the --time-offset option has multiple modes of operation that are depending on the context

of its use!

The timestep value between the starting and the ending time of the time interval that is respected by Gcal is one minute by default. This value can at pleasure be changed by using the --cycle-step option. See [Fixed date option --cycle-step=argument], page 42, for more details. You should always keep in mind that the use of the cycle mode is efficient only if the current time of the time interval is also part of that resource file line that has to be evaluated in this manner. Otherwise, it is difficult to find out unobjectionably lateron when displaying the data and times, which Sun or Moon oriented value resulted in what time. The following example shows how the topocentric elevation angle of the Sun for the location Münster, Germany, is calculated during two days in April 2000, and that for the time interval 12:00 o’clock until 12:05 o’clock local time in each case:

      $ gcal %20000420 -Hno -f/dev/null -xdt -g::: \
      >      --here='0*d1#999_%ot+00+000_:_%oa+5158+00738,120' \
      >      --cycle-end=12:05 --time=12:00

⊣

      ⊣ Thu, Apr  20th 2000: +12h00': +045d58'
      ⊣ Thu, Apr  20th 2000: +12h01': +046d03'
      ⊣ Thu, Apr  20th 2000: +12h02': +046d07'
      ⊣ Thu, Apr  20th 2000: +12h03': +046d12'
      ⊣ Thu, Apr  20th 2000: +12h04': +046d17'
      ⊣ Thu, Apr  20th 2000: +12h05': +046d22'
      ⊣ :::
      ⊣ Fri, Apr  21st 2000: +12h00': +046d18'
      ⊣ Fri, Apr  21st 2000: +12h01': +046d23'
      ⊣ Fri, Apr  21st 2000: +12h02': +046d28'
      ⊣ Fri, Apr  21st 2000: +12h03': +046d32'
      ⊣ Fri, Apr  21st 2000: +12h04': +046d37'
      ⊣ Fri, Apr  21st 2000: +12h05': +046d42'

See Section G.2.2.5 [Sun data], page 165, and Section G.2.2.6 [Moon data], page 188, for further information.

--cycle-step=argument

Changes the timestep value that is used between the starting and ending clocktime of the time interval, in which the Sun and Moon oriented special texts are calculated, that are dynamical, i.e. depending on the respective clocktime. See [Fixed date option --cycle-end=argument], page 41, how to define such a time interval. Without the --cycle-end option, a --cycle-step option is completely ignored. It also does not enable Gcal’s cycle mode. The argument of the --cycle-step option is by change syntactically according to the argument which may be given to the --time-offset option, and that relative time offset value references based on the actual local time and Universal time cannot be made, respectively. Gcal always uses the amount of the given timestep value. Thus, the template for the argument that may be given to the --cycle-step

option is [+|-]mmmm|hh:[mm]. See [Calendar option --time-offset=argument],

page 33, for the detailed explanation of the preceding option argument. The timestep value between the starting and the ending time of the time interval that is respected by Gcal is one minute by default. The following example shows how the topocentric elevation angle of the Sun for the location Münster, Germany, is calculated during two days in April 2000, and that for the time interval 10:00 o’clock until 17:00 o’clock local time with a timestep value of one hour and 15 minutes in each case:

      $ gcal %20000420 -Hno -f/dev/null -xdt -g::: \
      >      --here='0*d1#999_%ot+00+000_:_%o*a+5158+00738,120' \
      >      --cycle-end=10:00 --time=17:00 --cycle-step=1:15

⊣

      ⊣ Thu, Apr  20th 2000: +10h45': +38.038
      ⊣ Thu, Apr  20th 2000: +12h00': +45.962
      ⊣ Thu, Apr  20th 2000: +13h15': +49.665
      ⊣ Thu, Apr  20th 2000: +14h30': +47.882
      ⊣ Thu, Apr  20th 2000: +15h45': +41.267
      ⊣ Thu, Apr  20th 2000: +17h00': +31.633
      ⊣ :::
      ⊣ Fri, Apr  21st 2000: +10h45': +38.355
      ⊣ Fri, Apr  21st 2000: +12h00': +46.299
      ⊣ Fri, Apr  21st 2000: +13h15': +50.006
      ⊣ Fri, Apr  21st 2000: +14h30': +48.198
      ⊣ Fri, Apr  21st 2000: +15h45': +41.545
      ⊣ Fri, Apr  21st 2000: +17h00': +31.881

See Section G.2.2.5 [Sun data], page 165, and Section G.2.2.6 [Moon data], page 188, for further information.

--precise

The astronomical data and times that are created by the Sun and Moon oriented special texts, and the geographical data that is created by the distance and course angle oriented special text, respectively, are displayed by using a higher precision, i.e. according to the mode or style of representation and the type of special text, decimal values are displayed with more decimal places, or output of the time seconds part or the arcseconds part. See Section G.2.2.5 [Sun data], page 165, and Section G.2.2.6 [Moon data], page 188, likewise Section G.2.2.4 [Geographical distance and course angle], page 162, for further information.

--execute-command

Causes on the one hand, that shell command %![argument] special texts are executed, and not displayed textually only. See Section G.3 [Shell command %![argument] special text], page 210, for more details. And causes on the other hand, that ‘tvar?[command]’ and ‘tvar:[command]’ text variable definitions is assigned the output of an external command, and not the text which contains the call of the external command. See Section 5.1.9 [Text variables], page 106, for further information.

-D argument
--filter-day=argument

Displays only those fixed dates, whose date is not excluded by the given argument. See Section 5.1.2 [Date part of a line], page 86, and likewise Section G.1.2 [Exclusions without any argument %? special texts], page 148. The argument consists of one or more characters as used in the exclusion without any argument %? special text; but without the leading ‘%’ character in each case. For example, it is possible to induce Gcal to display only those fixed dates from the resource file demo.rc for the whole year which are legals holidays, and moreover, Saturdays or Sundays:

     gcal -f demo.rc --year --filter-day=Rv
-P argument
--filter-period=argument

Displays only those fixed dates, whose date is not excluded by the given argument. See Section 5.1.2 [Date part of a line], page 86, and likewise Section G.1.1 [Exclusions with date argument %?[date] special texts], page 146. The argument consists of one or more expressions as used in the exclusion with date argument %?[date] special text; but without the leading ‘%’ character in each case. You can use more than one of these %?[date] expressions by connecting them with a ‘,’ character. For example, it is possible to induce Gcal to display only those fixed dates from the resource file demo.rc for the whole year which appear within the period of the 17th and the 20th of any month, and which appear on the 31st of any month.

     gcal -f demo.rc -y -P i00000017#00000020,i00000031#00000031
-I pattern
--filter-text=pattern

Displays only those fixed dates, whose completely expanded text^{11} is matched by the given pattern. See Section 5.1.4 [Text part of a line], page 97, and Section G.2 [%?… special texts for text replacement], page 150. The pattern is a regular expression, as recognized by the Unix ed line-editor. See Appendix E [Regular Expressions], page 137, for further information. For example, it is possible to induce Gcal to display all fixed dates from the resource file doctor.rc for the current month, which are defined for dentists or oculists:

     gcal -f doctor.rc --month -I "[dD]entist|[oO]culist"

Each time a ‘_’ (underscore) character is found in the pattern argument, this character is replaced by a real ‘ ’ (space) character.

11 Except Gcal’s line break-up characters ‘~’ and ‘^’, so you have to use ‘\~’ and ‘\\^’ in the pattern

argument if you want to search texts which contain these characters.

You may depreciate the special meaning of the ‘_’ (underscore) character — in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\_’. If you need the ‘\_’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g.

‘\\_’.
--ignore-case

Ignores all case distinctions in both the completely expanded text and the pattern of the --filter-text=pattern option. See [Fixed date option

--filter-text=pattern], page 44.
--revert-match

Displays all those fixed dates, whose completely expanded text does not match the pattern of the --filter-text=pattern option. See [Fixed date option

--filter-text=pattern], page 44.
-c|C[-]
--today
--list-of-fixed-dates[=short|long]
--descending-fixed-dates[=short|long]

Activate the fixed date function and display fixed date messages, if any.

-c
--today
--list-of-fixed-dates=short

Activate fixed date function (use standard resource file implicitly) and list all dates related to the actual system date (≡today), sorted in ascending order. If no fixed dates related to the current day are found, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 1], page 257.

-c-
--descending-fixed-dates=short

Activate fixed date function (use standard resource file implicitly) and list all dates related to the actual system date (≡today), sorted in descending order. If no fixed dates related to the current day are found, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 1], page 257.

-C
--list-of-fixed-dates=long

Activate fixed date function (use standard resource file implicitly) and list all dates related to the actual system date (≡today), sorted in ascending order. If no fixed dates related to the current day are found, an “empty” fixed date message is displayed which is only consisting of the date.

-C-
--descending-fixed-dates=long

Activate fixed date function (use standard resource file implicitly) and list all dates related to the actual system date (≡today), sorted in descending order. If no fixed dates related to the current day are found, an “empty” fixed date message is displayed which is only consisting of the date.

-f|F name[+. . .]
--resource-file=name[+. . .]

Activate fixed date function and use file name instead of the standard resource file. Then list all dates, sorted in ascending order which occur in the fixed date period. If no other fixed date period is specified, the actual system date (≡today) is used for the fixed date period. If the option letter f of the shortstyle option is used and no dates related to the fixed date period are found in name, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 1], page 257. If the option letter F of the shortstyle option is used and no dates related to the fixed date period are found in name, an “empty” fixed date message is displayed which is only consisting of the date. You can use more than one resource file name by connecting them with a ‘+’ character, e.g.:

     --resource-file=./foo+xyz+/data/bar+$HOME/.gcalrc resp.,
     -f ./foo+xyz+/data/bar+$HOME/.gcalrc

respects all files given in the preceding argument by using a special file search mechanism. See [File searching mechanism], page 82, for more details. Use /dev/null to avoid the use of the standard resource file, useful for creating empty fixed date lists. If a single ‘-’ character is given as file name, like

--resource-file=./foo+-+xyz or -F-, but not -F -, Gcal reads and processes

all input received from the standard input channel. You may depreciate the special meaning of the ‘+’ separator character —in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\+’. If you need the ‘\+’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\+’. Each time a ‘_’ (underscore) character is found in name, this character is replaced by a real ‘ ’ (space) character. You may depreciate the special meaning of the ‘_’ (underscore) character — in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\_’. If you need the ‘\_’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g.

‘\\_’.
-# line
--here=line

Activate fixed date function and use the line argument together with the standard resource file respectively additional resource files. The line argument has to be a valid line as it may occur in a Gcal resource file (see Section 5.1.1 [Structure of resource file], page 84), and is always evaluated last after the processing of all resource files, so dealing with references to exported local date or text variables is enabled. This option may be given multiple and it will be processed exactly in the given order. All dates are listed, sorted in ascending order which occur in the fixed date period. If no other fixed date period is specified, the actual system date (≡today) is used for the fixed date period. If no dates related to the fixed date period are found in the line, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 1], page 257.

For example, the call:

     gcal -# "0*d1su#99su.7 ^%Z" --here="#include <foo>" -y

causes the implicit processing of the standard resource file just before the further resource file line ‘0*d1su#99su.7 ^%Z’ and following ‘#include <foo>’ are processed, and that as if these lines were a physical part of the standard resource file.

Each time a ‘_’ (underscore) character is found in line, this character is replaced by a real ‘ ’ (space) character.

You may depreciate the special meaning of the ‘_’ (underscore) character — in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\_’. If you need the ‘\_’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g.

‘\\_’.

Please note: If the short-style option -c and/or its modifiers t|[w|m|y[+|-], or the short-style option -f is/are given in upper-case, e.g.:

-C
-F foo.bar
-CeT
-c-aW+
-CeaxY-
-c-M

Gcal displays fixed date messages for every day of the requested period.

The modifiers of a combined/complex/packed short-style fixed date option must be constructed according to following steps, because Gcal treats them as positional parameters:

First step: Representation-of-text-modifier

     a, A, e|E, J, k, o, Q, U, x, z, Z

Second step: Period-of-fixed-dates-modifier

     d, l, t|T, w|W|m|M|y|Y[+|-], n+|-, nd, nw, *dn[www], *wn[www],
     @e|t|dvar[[+|-]n], @e|t|dvar[+|-]nwww, mmdd, mmwwwn

If you want to operate with the preceding modifiers, an explicit -c|C[-] short-style option must be leading on the one hand, e.g.:

     -c-t
     -cw-
     -CZdekloQUzJaxA12+

and only one period defining fixed date period modifier may be given^{12}. Or on the other hand, the modifiers must be given as single character options or composed into a single command line word, and additionally either the short-style option -c|C[-], -f|F name[+…] or -# line^{13} must be given like:

     -c -t
     -w- -c
     -a -Ad -e -kloQUxz -C12+
     -F foo.bar -dJZA -l*d10
     --here="; Comment" -dZAa

If the modifiers are given as single character options or composed into a single command line word, and one of the t|T|[w|W|m|M|y|Y[+|-]] modifiers is used^{14}, it is unnecessary to give a -c|C[-] respectively -f|F name[+…] short-style option^{15} to activate the fixed date function of Gcal, because it is triggered implicitly by means of these modifiers / alias names.

See Appendix K [Coding Scheme], page 251, whose tables explain the relations between the date part (yyyy …) of a resource file line and the modifiers, the short-style option -c|C[] —which is used for enabling the fixed date feature— respectively the long-style option --period-of-fixed-dates=argument may have, i.e. they point out all possible correspondences. This is necessary because Gcal does not respect all entries found in a resource file if a fixed date argument is given, it respects only those entries which have a definite relation to the fixed date argument to avoid the output of redundant information!

12 This means, exclusive the --include-today option respectively d modifier, the --list-mode option or

the l modifier, and the --leap-day=february|march option.

13 Respectively their according long-style options.

14 Or their according alias names, like --today, --tomorrow…

15 Or their according long-style options.

Fixed date text modifiers and options:

--biorhythm-axis=number

The width of the biorhythm text graphics —which can be created by means of the biorhythm text graphics %;[date] special text— is changed to number characters for each axis of the bar. The number argument must be an integer value between 1…100, the default number of characters is 20. If the number divides 100 with a remainder, it is internally reduced to the nearest number that divides 100 without any remainder. This means, only the numbers 1, 2, 4, 5, 10, 20, 25, 50 and 100 are factually respected, so a number argument of 19 is internally reduced to 10. For example:

      $ cat bio.rc
      ⊣ 0*d1#999 %;19620921
      $ gcal -fbio.rc %19961212 -s0 -xw --biorhythm-axis=19

⊣

      ⊣ Thu, Dec <12th>1996: 0- I       P1 E        +0
      ⊣ Fri, Dec  13th 1996: 0-  I   P   1          +0
      ⊣ Sat, Dec  14th 1996: 0-    @   E 0          +0
      ⊣ Sun, Dec  15th 1996: 0-  P  IE   0          +0
      ⊣ Mon, Dec  16th 1996: 0- P  E  I  0          +0
      ⊣ Tue, Dec  17th 1996: 1-P  E     I0          +0
      ⊣ Wed, Dec  18th 1996: 0- @        1          +0

See Section G.2.1.10 [Biorhythm %?[date] special texts], page 156, for more details.

--moonimage-lines=number

The height of the Moon phase text graphics —which can be created by means of the Moon phase text graphics %Z[date] special text— is changed to number lines. The number argument must be an integer value between 6…30, the default number of lines is 12.

For example:

      $ gcal -f/dev/null -#0*d1#999_%Z__%O %19950407 -x --moon=8

⊣

      ⊣ Fri, Apr < 7th>1995:          (    @@@@@
      ⊣  (           @@@@@@@@
      ⊣  (              @@@@@@@@@@@
      ⊣  (               @@@@@@@@@@@@
      ⊣  (               @@@@@@@@@@@@
      ⊣  (              @@@@@@@@@@@
      ⊣  (             @@@@@@@@@@
      ⊣  (        @@@@@@@ 42%+

See Section G.2.1.9 [Moon phase %[format]?[date] special texts], page 155, for more details.

a
--include-resource-file-name

Extend fixed date text by the name of the resource file and the line number which the fixed date text comes from, e.g.:

     Mon, Jan   1st 1995: (`Eternal holiday list'#00003) New Year's day
     Mon, Jan   1st 1995: (.gcalrc#00987) No fixed dates
     ...
A
--alternative-format

Instead of using the standard list format, e.g.:

     Mon, Jan   1st 1995: New Year's day
     Mon, Jan   1st 1995: No fixed dates
     ...

Gcal uses the alternative list format, e.g.:

     Mon, Jan   1st 1995:
     New Year's day
     No fixed dates
     ...
e
--include-holidays=long

Enable inclusion of all built-in eternal holiday dates. This includes legal holidays and memorial days.

E
--include-holidays=short

Enable inclusion of those built-in eternal holiday dates which are legal holidays only.

g[text]
--grouping-text[=text]

Group fixed dates by day by using the text separator. If text is not given, the built-in text^{1} is used for grouping, e.g.:

      $ gcal -Cxw -cg

⊣

      ⊣ Sun, Nov   6th 1994: Text 1
      ⊣ Sun, Nov   6th 1994: Text 2

⊣

      ⊣ Mon, Nov   7th 1994: Text 3

⊣

      ⊣ Tue, Nov   8th 1994: Text 4
      ...

      $ gcal -Cxw -g'-A_user\_defined-~-grouping text-'

⊣

      ⊣ Sun, Nov   6th 1994: Text 1
      ⊣ Sun, Nov   6th 1994: Text 2
      ⊣ -A user_defined-
      ⊣ -grouping text-
      ⊣ Mon, Nov   7th 1994: Text 3
      ⊣ -A user_defined-
      ⊣ -grouping text-
      ⊣ Tue, Nov   8th 1994: Text 4
      ...

The text may contain references to global date and text variables (see [Fixed date option --date-variable=argument], page 37, and [Fixed date option --date-variable=argument], page 38). Furthermore, all special texts which cause a text replacement may be used in the text (see Section G.2 [%?… special texts for text replacement], page 150). Each time a ‘~’ (tilde) or ‘^’ (caret) character is found in text, this character is replaced by a real ‘\n’ (newline) character. You may depreciate the special meaning of the ‘~’ (tilde) character —in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\~’. If you need the ‘\~’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\_’. All above mentioned facts are also valid for the ‘^’ (caret) character. Each time a ‘_’ (underscore) character is found in text, this character is replaced by a real ‘ ’ (space) character. You may depreciate the special meaning of the ‘_’ (underscore) character — in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\_’. If you need the ‘\_’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g.

1 RC_GROUP_SEP “” ≡ empty line only.

‘\\_’.
k
--include-week-number

The leading date of a fixed date message is either extended by the ISO week number or the standard week number. See [Calendar option --starting-day=argument], page 33, [Calendar option --iso-week-number=yes|no], page 30, and Appendix C [Aspects in Internationalization], page 131, for further details.

o
--omit-multiple-date-part

Omit displaying the leading date of a fixed date message in case more than one message refers to that date^{2}, e.g.:

      $ gcal -cxo

⊣

      ⊣ Sun, Nov   6th 1994: Text 1
      ⊣  Text 2
      ...

Q
--suppress-fixed-dates-list-separator

Suppress displaying of the blank line which is always leading a fixed date list. For example, to cause Gcal to print only the number of days between 1st January 1970 and 1st January 1980, you can call the program like this:

      $ gcal -f /dev/null -QUx %19800101 -#0_%j-2440588
      ⊣ 3652

U
--suppress-date-part

Suppress displaying the leading date of a fixed date message, e.g.:

      $ gcal -cxU ⊣ ⊣ Text 1 ⊣ Text 2 ...

2 Only if the --alternative-format option or the A modifier is not given.

J
--suppress-text-part

Suppress displaying the text part of a fixed date message, e.g.:

      $ gcal -cxJ

⊣

      ⊣ Sun, Nov   6th 1994:
      ⊣ Sun, Nov   6th 1994:
      ...
x
--exclude-fixed-dates-list-title

Suppress the title text line of the fixed date list.

--heading-text=text

Changes the title text line of the fixed date list. Instead of displaying the default ‘Fixed date list:’ text, any other text can be used as heading text of the fixed date list, e.g.:

      $ gcal -cUJ --heading-text=Hello,_it_is_%K_,~%t__o'clock

⊣

      ⊣ Hello, it is Sunday,
      ⊣ 14:32 o'clock

⊣

      ⊣ Sun, Nov   6th 1994:
      ...

The text may contain references to global date and text variables (see [Fixed date option --date-variable=argument], page 37, and [Fixed date option --date-variable=argument], page 38). Furthermore, all special texts which cause a text replacement may be used in the text (see Section G.2 [%?… special texts for text replacement], page 150). Each time a ‘~’ (tilde) or ‘^’ (caret) character is found in text, this character is replaced by a real ‘\n’ (newline) character. You may depreciate the special meaning of the ‘~’ (tilde) character —in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\~’. If you need the ‘\~’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\_’. All above mentioned facts are also valid for the ‘^’ (caret) character. Each time a ‘_’ (underscore) character is found in text, this character is replaced by a real ‘ ’ (space) character. You may depreciate the special meaning of the ‘_’ (underscore) character — in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\_’. If you need the ‘\_’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g.

‘\\_’.

z
--include-consecutive-number

Display consecutive numbers of fixed date messages.

Z
--zero-dates-only

Display only those dates, for which fixed dates do not exist, i.e. all “empty” fixed dates only consisting of the date.

Fixed date period modifiers and options:

--leap-day=february|march

The leap day^{1} is observed in non-leap years.

          --leap-day=february

A fixed date that appears in every year and which is related to the leap day, is displayed on the 28th February in non-leap years.

          --leap-day=march

A fixed date that appears in every year and which is related to the leap day, is displayed on the 1st March in non-leap years.

d
--include-today

If lists of periods are generated, include the actual date into the list. See [Fixed date option --list-mode], page 55, how to create a list of periods.

l
--list-mode

Generate a list of periods instead of a single period.

Please note:

The following examples assumes the actual system date is the 17th February of any year (00000217), weeks start on Mondays and ISO week numbers are used!

--period-of-fixed-dates=argument

nd

*dn Single date of absolute day n ≡ ‘1…365|366|999’ of the actual year; the intensity level is the same as the simple -c option. If the value 999 for n is given, the last day of the year (31st December) is assumed.

            • E.g. ‘gcal -c10d’ respectively ‘gcal --period-of-fixed-dates=10d’
              or ‘gcal -c*d10’ respectively ‘gcal --period-of-fixed-dates=*d10’

displays all fixed dates which occur on the 10th day of the year.

lnd

l*dn List dates starting yesterday or tomorrow (depending on the given day number n) and ending on the n’th absolute day of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl10d’ respectively ‘gcal -cl*d10’ displays a list of all fixed dates which start on the 10th day of year and end on the 16th February. n+|- Single date of day actual day ± n days of the actual year; the intensity level is

1 The 29th February.

the same as the simple -c option.

• E.g. ‘gcal -c10+’ displays all fixed dates which occur 10 days after the 17th February (today).

• E.g. ‘gcal -c10-’ displays all fixed dates which occur 10 days before the 17th February (today).

ln+|- List dates starting yesterday ‘-’ or tomorrow ‘+’ and ending on actual day ± n days of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl10+’ displays a list of all fixed dates which start on the 18th February and end 10 days after.

• E.g. ‘gcal -cl10-’ displays a list of all fixed dates which start 10 days before the 16th February and end on the 16th February.

nw Single dates of week number n ≡ ‘0|1…52|53|99’ of the actual year; the intensity level is the same as the -cw option. See [Calendar option --starting-day=argument], page 33, [Calendar option --iso-week-number=yes|no], page 30, and Appendix C [Aspects in Internationalization], page 131, for more details.

• E.g. ‘gcal -c0w’ displays all fixed dates which occur in the 53rd week of the previous year, in case the previous year has a 53rd week. If the previous year has no 53rd week, all fixed dates occurring in the 1st week of the actual year are displayed.

• E.g. ‘gcal -c6w’ displays all fixed dates which occur in the 6th week of year.

• E.g. ‘gcal -c52w’ displays all fixed dates which occur in the 52nd week of year.

• E.g. ‘gcal -c53w’ displays all fixed dates which occur in the 53rd week of the actual year, in case the actual year has a 53rd week. If the actual year has no 53rd week, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -c99w’ displays all fixed dates which occur in the last week of the actual year, i.e. either the 52nd or the 53rd week.

lnw List dates starting yesterday or tomorrow (depending on the given week number n) and ending on the first respectively last day of given week number n ≡ ‘0|1…52|53|99’ of the actual year; the intensity level is the same as the -cy option. See [Calendar option --starting-day=argument], page 33, [Calendar option --iso-week-number=yes|no], page 30, and Appendix C [Aspects in Internationalization], page 131, for further details.

• E.g. ‘gcal -cl12w’ displays a list of all fixed dates which start on the 18th February and end on the last day of the 12th week.

• E.g. ‘gcal -cl3w’ displays a list of all fixed dates which start on the first day of the 3rd week and end on the 16th February. mmdd Single date of day dd in month mm of the actual year; the intensity level is the

same as the simple -c option.

• E.g. ‘gcal -c0225’ displays all fixed dates which occur on the 25th February.

lmmdd List dates starting yesterday or tomorrow (depending on the value given in mmdd) and ending on day dd of month mm of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl0225’ displays a list of all fixed dates which start on the 18th February and end on the 25th February.

• E.g. ‘gcal -cl0109’ displays a list of all fixed dates which start on the 9th January and end on the 16th February.

mmwwwn Single date of n’th ≡ ‘1…5|9’ weekday dd|www in month mm of the actual

year; the intensity level is the same as the simple -c option.

• E.g. ‘gcal -c01mon3’ displays all fixed dates which occur on the 3rd Monday in January.

• E.g. ‘gcal -c02fri9’ displays all fixed dates which occur on the last Friday in February.

lmmwwwn List dates starting yesterday or tomorrow (depending on the value given in mmwwwn) and ending on n’th ≡ ‘1…5|9’ weekday dd of month mm of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl01mon3’ displays a list of all fixed dates which start on the 3rd Monday in January and end on the 16th February.

• E.g. ‘gcal -cl02fri9’ displays a list of all fixed dates which start on the 18th February and end on the last Friday in February.

*dnwww Single date of n’th ≡ ‘1…51|52|53|99’ weekday www of the actual year; the

intensity level is the same as the simple -c option.

• E.g. ‘gcal -c*d16sun’ displays all fixed dates which occur on the 16th Sunday of the actual year.

• E.g. ‘gcal -c*d52sun’ displays all fixed dates which occur on the 52nd Sunday of the actual year. If the actual year has no 52nd Sunday, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -c*d53sun’ displays all fixed dates which occur on the 53rd Sunday of the actual year. If the actual year has no 53rd Sunday, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -c*d99sun’ displays all fixed dates which occur on the last Sunday^{1} of the actual year. l*dnwww List dates starting yesterday or tomorrow (depending on the value given in n)

1 That is either the 51st, 52nd or 53rd Sunday.

and ending on n’th ≡ ‘1…51|52|53|99’ weekday www of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl*d16sun’ displays a list of all fixed dates which start on the 18th February and end on the 16th Sunday of the actual year.

• E.g. ‘gcal -cl*d52sun’ displays a list of all fixed dates which start on the 18th February and end on the 52nd Sunday of the actual year. If the actual year has no 52nd Sunday, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -cl*d53sun’ displays a list of all fixed dates which start on the 18th February and end on the 53nd Sunday of the actual year. If the actual year has no 53nd Sunday, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -cl*d99sun’ displays a list of all fixed dates which start on the 18th February and end on the last Sunday^{1} of the actual year.

*wn[www] Single date of weekday www of n’th ≡ ‘0|1…52|53|99’ week of the actual year; the intensity level is the same as the simple -c option. If no weekday www is given, the starting day of the week is assumed for the weekday www. See [Calendar option --starting-day=argument], page 33, [Calendar option --iso-week-number=yes|no], page 30, and Appendix C [Aspects in Internationalization], page 131, for more details.

• E.g. ‘gcal -c*w0sun’ displays all fixed dates which occur on the Sunday of the 52nd respectively 53rd week of the previous year. In case such a Sunday does not exist, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -c*w16sun’ displays all fixed dates which occur on the Sunday of the 16th week of the actual year.

• E.g. ‘gcal -c*w52sun’ displays all fixed dates which occur on the Sunday of the 52nd week of the actual year. If the actual year has no Sunday of the 52nd week, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -c*w53sun’ displays all fixed dates which occur on the Sunday of the 53rd week of the actual year. If the actual year has no Sunday of the 53rd week, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -c*w99sun’ displays all fixed dates which occur on the last Sunday^{2} of the actual year.

l*wn[www] List dates starting yesterday or tomorrow (depending on the value given in n) and ending on weekday www of n’th ≡ ‘0|1…52|53|99’ week of the actual year; the intensity level is the same as the -cy option. If no weekday www is given, the starting day of the week is assumed for the weekday www. See [Calendar option --starting-day=argument], page 33, [Calendar option --iso-week-number=yes|no], page 30, and Appendix C [Aspects in Internationalization], page 131, for further details.

1 That is either the 51st, 52nd or 53rd Sunday.

2 That is either the Sunday of the 51st, 52nd or 53rd week.

• E.g. ‘gcal -cl*w0sun’ displays a list of all fixed dates which start on the Sunday of the 52nd respectively 53rd week of the previous year and end on the 16th February. In case such a Sunday does not exist, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -cl*w16sun’ displays a list of all fixed dates which start on the 18th February and end on the Sunday of the 16th week of the actual year.

• E.g. ‘gcal -cl*w52sun’ displays a list of all fixed dates which start on the 18th February and end on the Sunday of the 52nd week of the actual year. If the actual year has no Sunday of the 52nd week, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -cl*w53sun’ displays a list of all fixed dates which start on the 18th February and end on the Sunday of the 53rd week of the actual year. If the actual year has no Sunday of the 53rd week, no fixed date messages are displayed and the program is terminated with an error code. See [Error Code 126], page 257.

• E.g. ‘gcal -cl*w99sun’ displays a list of all fixed dates which start on the 18th February and end on the last Sunday^{3} of the actual year.

@e[[+|-]n] Single date of day n relative to the Easter Sunday’s date of the actual year; the

intensity level is the same as the simple -c option.

• E.g. ‘gcal -c@e’ displays all fixed dates which occur on the Easter Sunday’s date.

• E.g. ‘gcal -c@e10’ respectively ‘gcal -c@e+10’ displays all fixed dates which occur 10 days after the Easter Sunday’s date.

• E.g. ‘gcal -c@e-10’ displays all fixed dates which occur 10 days before the Easter Sunday’s date.

3 That is either the Sunday of the 51st, 52nd or 53rd week.

l@e[[+|-]n]

List dates starting yesterday or tomorrow (depending on the value given in n) and ending on n’th day relative to the Easter Sunday’s date of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl@e’ displays a list of all fixed dates which start on the 18th February and end on the Easter Sunday’s date.

• E.g. ‘gcal -cl@e10’ respectively ‘gcal -cl@e+10’ displays a list of all fixed dates which start on the 18th February and end 10 days after the Easter Sunday’s date.

• E.g. ‘gcal -cl@e-10’ displays a list of all fixed dates which start on the 18th February and end 10 days before the Easter Sunday’s date.

@e[+|-]nwww

Single date of n’th weekday www relative to the Easter Sunday’s date of the actual year; the intensity level is the same as the simple -c option.

• E.g. ‘gcal -c@e3fri’ respectively ‘gcal -c@e+3fri’ displays all fixed dates which occur on the 3rd Friday after the Easter Sunday’s date.

• E.g. ‘gcal -c@e-3fri’ displays all fixed dates which occur on the 3rd Friday before the Easter Sunday’s date.

l@e[+|-]nwww

List dates starting yesterday or tomorrow (depending on the value given in n) and ending on n’th weekday www relative to the Easter Sunday’s date of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl@e3fri’ respectively ‘gcal -cl@e+3fri’ displays a list of all fixed dates which start on the 18th February and end on the 3rd Friday after the Easter Sunday’s date.

• E.g. ‘gcal -cl@e-3fri’ displays a list of all fixed dates which start on the 18th February and end on the 3rd Friday before the Easter Sunday’s date.

@t[[+|-]n] Single date of day n relative to today’s date of the actual year; the intensity

level is the same as the simple -c option.

• E.g. ‘gcal -c@t’ displays all fixed dates which occur on today’s date (≡

‘gcal -c’).

• E.g. ‘gcal -c@t10’ respectively ‘gcal -c@t+10’ displays all fixed dates which occur 10 days after today’s date (≡ ‘gcal -c10+’).

• E.g. ‘gcal -c@t-10’ displays all fixed dates which occur 10 days before today’s date (≡ ‘gcal -c10-’).

l@t[[+|-]n]

List dates starting yesterday or tomorrow (depending on the value given in n) and ending on n’th day relative to today’s date of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl@t’ displays nothing.

• E.g. ‘gcal -cl@t10’ respectively ‘gcal -cl@t+10’ displays a list of all fixed dates which start on the 18th February and end 10 days after today’s date

    (≡ ‘gcal -cl10+’).

• E.g. ‘gcal -cl@t-10’ displays a list of all fixed dates which start on the 18th February and end 10 days before today’s date (≡ ‘gcal -cl10-’).

@t[+|-]nwww

Single date of n’th weekday www relative to today’s date of the actual year; the intensity level is the same as the simple -c option.

• E.g. ‘gcal -c@t3fri’ respectively ‘gcal -c@t+3fri’ displays all fixed dates which occur on the 3rd Friday after today’s date.

• E.g. ‘gcal -c@t-3fri’ displays all fixed dates which occur on the 3rd Friday before today’s date.

l@t[+|-]nwww

List dates starting yesterday or tomorrow (depending on the value given in n) and ending on n’th weekday www relative to today’s date of the actual year; the intensity level is the same as the -cy option.

• E.g. ‘gcal -cl@t3fri’ respectively ‘gcal -cl@t+3fri’ displays a list of all fixed dates which start on the 18th February and end on the 3rd Friday after today’s date.

• E.g. ‘gcal -cl@t-3fri’ displays a list of all fixed dates which start on the 18th February and end on the 3rd Friday before today’s date.

@dvar[[+|-]n]

Single date of the day n relative to that date of the actual year, which is referenced by the date variable dvar; the intensity level is the same as the simple -c option.

• E.g. ‘gcal -v a=0422 -c@a’ displays all fixed dates which occur on the 22nd April.

  • E.g. ‘gcal -v a=0422 -c@a10’ respectively ‘gcal -v a=0422 -c@a+10’ dis-

plays all fixed dates which occur 10 days after the 22nd April.

• E.g. ‘gcal -v a=0422 -c@a-10’ displays all fixed dates which occur 10 days before the 22nd April.

l@dvar[[+|-]n]

List dates starting yesterday or tomorrow (depending on the value given in n) and ending on the n’th day relative to that date of the actual year, which is referenced by the date variable dvar; the intensity level is the same as the -cy option.

• E.g. ‘gcal -v a=0422 -cl@a’ displays a list of all fixed dates which start on the 18th February and end on the 22nd April.

• E.g. ‘gcal -v a=0422 -cl@a10’ respectively ‘gcal -v a=0422 -cl@a+10’

displays a list of all fixed dates which start on the 18th February and end 10 days after the 22nd April.

• E.g. ‘gcal -v a=0422 -cl@a-10’ displays a list of all fixed dates which start on the 18th February and end 10 days before the 22nd April.

@dvar[+|-]nwww

Single date of the n’th weekday www relative to that date of the actual year, which is referenced by the date variable dvar; the intensity level is the same as the simple -c option.

• E.g. ‘gcal -v a=0122 -c@a2sat’ respectively ‘gcal -v a=0122 -c@a+2sat’

displays all fixed dates which occur on the 2nd Saturday after the 22nd January.

• E.g. ‘gcal -v a=0122 -c@a-2sat’ displays all fixed dates which occur on the 2nd Saturday before the 22nd January.

l@dvar[+|-]nwww

List dates starting yesterday or tomorrow (depending on the value given in n) and ending on the n’th weekday www relative to that date of the actual year, which is referenced by the date variable dvar; the intensity level is the same as the -cy option.

• E.g. ‘gcal -v a=0122 -cl@a2sat’ respectively ‘gcal -v a=0122

-cl@a+2sat’ displays a list of all fixed dates which start on the 2nd Saturday after the 22nd January and end on the 16th February.

• E.g. ‘gcal -v a=0122 -cl@a-2sat’ displays a list of all fixed dates which start on the 2nd Saturday before the 22nd January and end on the 16th February.

t|T

--tomorrow long-style option

List dates related to tomorrow.

• E.g. ‘gcal -ct’ displays all fixed dates which occur on the 18th February (tomorrow).

• E.g. ‘gcal -cdt’ displays all fixed dates which occur on the 17th February (today) and on the 18th February (tomorrow).

w|W[+|-]

--week long-style option --end-of-week long-style option --start-of-week long-style option

Display dates related to the week. See [Calendar option --starting-day=argument], page 33.

• w or --week long-style option List dates between the starting day of the current week and the last day of the current week

− E.g. ‘gcal -cw’ displays all fixed dates which occur in the whole week the 17th February is part of.

• w+ or --end-of-week long-style option List dates between the day after the current day of the current week and the last day of the current week.

− E.g. ‘gcal -cw+’ displays all fixed dates which start on the 18th February (tomorrow) and end on the last day of the week.

− E.g. ‘gcal -cdw+’ displays all fixed dates which start on the 17th February (today) and end on the last day of the week.

• w- or --start-of-week long-style option List dates between the starting day of the current week and the day before the current day of the current week

− E.g. ‘gcal -cw-’ displays all fixed dates which start on the first day of the week and end on the 16th February (yesterday).

− E.g. ‘gcal -cdw-’ displays all fixed dates which start on the first day of the week and end on the 17th February (today).

m|M[+|-]

--month long-style option --end-of-month long-style option

--start-of-month long-style option

Display dates related to the month.

• m or --month long-style option List dates between the first day of the current month and the last day of the current month.

− E.g. ‘gcal -cm’ displays all fixed dates which occur in the whole month of February.

• m+ or --end-of-month long-style option List dates between the day after the current day of the current month and the last day of the current month.

− E.g. ‘gcal -cm+’ displays all fixed dates which start on the 18th February (tomorrow) and end on the last day of the month February.

− E.g. ‘gcal -cdm+’ displays all fixed dates which start on the 17th February (today) and end on the last day of the month February.

• m- or --start-of-month long-style option List dates between the first day of the current month and the day before the current day of the current month. − E.g. ‘gcal -cm-’ displays all fixed dates which start on the first day of

the month February and end on the 16th February (yesterday). − E.g. ‘gcal -cdm-’ displays all fixed dates which start on the first day

of the month February and end on the 17th February (today).

y|Y[+|-]

--year long-style option --end-of-year long-style option --start-of-year long-style option

Display dates related to the year.

• y or --year long-style option List dates between the first day of the current year and the last day of the current year. − E.g. ‘gcal -cy’ displays all fixed dates which occur in the whole year.

• y+ or --end-of-year long-style option List dates between the day after the current day of the current year and the last day of the current year. − E.g. ‘gcal -cy+’ displays all fixed dates which start on the 18th Febru-

ary (tomorrow) and end on the last day of the year. − E.g. ‘gcal -cdy+’ displays all fixed dates which start on the 17th

February (today) and end on the last day of the year.

• y- or --start-of-year long-style option List dates between the first day of the current year and the day before the current day of the current year. − E.g. ‘gcal -cy-’ displays all fixed dates which start on the first day of

the year and end on the 16th February (yesterday). − E.g. ‘gcal -cdy-’ displays all fixed dates which start on the first day

of the year and end on the 17th February (today).

3.1.2 Response file

If a @file command line argument is given, a response file by the name of file will be used, i.e. options and commands are preloaded from that file. A response file may contain options and commands —which are preloaded by Gcal—, but no references to further response files. If such references to further response files are found, they are ignored. See [Global option --response-file=name], page 10, for hints how to generate a response file automatically.

The searching scheme for a response file is the same as that for resource files, except no standard response file is respected. See [File searching mechanism], page 82, for more details. Multiple response files may be given in the command line when Gcal is started, and they are processed exactly in that order as given, i.e. strictly sequential.

Each option and command must be on a single line, i.e. separated by a real ‘\n’ (newline) character. A line beginning with a ‘;’ (semicolon) character in the response file is treated as a remark and will not be used by Gcal (see Section 5.1.6 [Comment line], page 101).

Options must be defined before commands. If a response file contains any commands, then all further arguments after the @file option of the command line are ignored. For example:

1. Contents of response file file:

         Line     Text
         1        -j
         2        -Cw
         EOF

         $ gcal -i -b 3 @file -s 3 1994
         ↦ gcal -i -b 3 -j -Cw -s 3 1994

2. Contents of response file file:

         Line     Text
         1        -j
         2        -Cw
         3        may 1992
         EOF

         $ gcal -i -b 3 @file -s 3 1994

3. Contents of response file file:

Line     Text
1        -j
2        -Cw
3        may
4        1992
EOF

 $ gcal -i -b 3 @file -s 3 1994

3.1.3 Actual date modifier

It is allowed to use that date —instead of the actual system date— which is defined by the %date command line argument. This means, fixed dates can be checked for any year and are respected in the same way, as if they would be fixed dates of the actual year.

The date must be denoted in one of these formats:

• yyyy[mm[dd|wwwn]]
• yyyy*d|wn[www]
• yyyy@e|t|dvar[[+|-]n[www]]
• month name[dd]
• weekday name[n]

• dd

Some examples to this:

Please note: The following examples assumes the actual system date is Wednesday, the 17th February 1993 (19930217), weeks start on Mondays and ISO week numbers are used!

The yyyy[mm[dd|wwwn]] format:

• ‘gcal %1994’ (≡ yyyy)

The actual date is set to the date 17th February 1994 ⇒ 19940217.

• ‘gcal %1994Mar’
• ‘gcal %19943’

• ‘gcal %199403’ (≡ yyyymm)

The actual date is set to the date 17th March 1994 ⇒ 19940317.

• ‘gcal %1994Mar3’
• ‘gcal %1994033’

• ‘gcal %19940303’ (≡ yyyymmdd) The actual date is set to the date 3rd March 1994 ⇒ 19940303.

• ‘gcal %1994MarWed3’

• ‘gcal %199403wed3’ (≡ yyyymmwwwn) The actual date is set to the date 3rd Wednesday in March 1994

(n ≡ ‘1…5|9’) ⇒ 19940316.

• ‘gcal %0’
• ‘gcal %00’
• ‘gcal %000’

• ‘gcal %0000’ (≡ yyyy)

The actual date remains unchanged ⇒ 19930217.

• ‘gcal %0000Mar’
• ‘gcal %00003’

• ‘gcal %000003’ (≡ yyyymm)

The actual date is set to the date 17th March 1993 ⇒ 19930317.

• ‘gcal %0000Mar3’
• ‘gcal %0000033’

• ‘gcal %00000303’ (≡ yyyymmdd) The actual date is set to the date 3rd March 1993 ⇒ 19930303.

• ‘gcal %0000MarWed3’

• ‘gcal %000003wed3’ (≡ yyyymmwwwn) The actual date is set to the date 3rd Wednesday in March 1993

(n ≡ ‘1…5|9’) ⇒ 19930317.

The yyyy*d|wn[www] format:
• ‘gcal %1994*d10’ (≡ yyyy*dn)

The actual date is set to the date 10th absolute day in 1994

(n ≡ ‘1…365|366|999’) ⇒ 19940110.

• ‘gcal %1994*d10sun’ (≡ yyyy*dnwww)

The actual date is set to the date 10th Sunday in 1994

(n ≡ ‘1…51|52|53|99’) ⇒ 19940306.

• ‘gcal %0*d10’ (≡ yyyy*dn)

The actual date is set to the date 10th absolute day in 1993 ⇒ 19930110.

• ‘gcal %0*d10sun’ (≡ yyyy*dnwww)

The actual date is set to the date 10th Sunday in 1993 ⇒ 19930307.

• ‘gcal %1994*w10’ (≡ yyyy*wn)

The actual date is set to the date Monday of the 10th week in 1994

(n ≡ ‘0|1…52|53|99’) ⇒ 19940307.

• ‘gcal %1994*w10sun’ (≡ yyyy*wnwww)

The actual date is set to the date Sunday of the 10th week in 1994 ⇒ 19940313.

• ‘gcal %0*w10’ (≡ yyyy*wn)

The actual date is set to the date Monday of the 10th week in 1993 ⇒ 19930308.

• ‘gcal %0*w10sun’ (≡ yyyy*wnwww)

The actual date is set to the date Sunday of the 10th week in 1993 ⇒ 19930314.

The yyyy@e|t|dvar[[+|-]n[www]] format:
• ‘gcal %1994@e’ (≡ yyyy@e)

The actual date is set to the date Easter Sunday 1994 ⇒ 19940403.

• ‘gcal %1994@e10’
• ‘gcal %1994@e+10’ (≡ yyyy@e[+]n)

The actual date is set to the date 10 days after Easter Sunday 1994 ⇒ 19940413.

• ‘gcal %1994@e-10’ (≡ yyyy@e-n)

The actual date is set to the date 10 days before Easter Sunday 1994 ⇒ 19940324.

• ‘gcal %0@e’
• ‘gcal %00@e’
• ‘gcal %000@e’
• ‘gcal %0000@e’ (≡ yyyy@e)

The actual date is set to the date Easter Sunday 1993 ⇒ 19930411.

• ‘gcal %0@e10’
• ‘gcal %0@e+10’ (≡ yyyy@e[|]n)

The actual date is set to the date 10 days after Easter Sunday 1993 ⇒ 19930421.

• ‘gcal %0@e-10’ (≡ yyyy@e-n)

The actual date is set to the date 10 days before Easter Sunday 1993 ⇒ 19930401.

• ‘gcal %1994@e3Wed’
• ‘gcal %1994@e+3Wed’ (≡ yyyy@e[+]nwww)

The actual date is set to the date 3rd Wednesday after Easter Sunday 1994 ⇒ 19940420.

• ‘gcal %1994@e-3wed’ (≡ yyyy@e-nwww)

The actual date is set to the date 3rd Wednesday before Easter Sunday 1994 ⇒ 19940316.

• ‘gcal %0@e3wed’
• ‘gcal %0@e+3wed’ (≡ yyyy@e[+]nwww)

The actual date is set to the date 3rd Wednesday after Easter Sunday 1993 ⇒ 19930428.

• ‘gcal %0@e-3we’ (≡ yyyy@e-nwww)

The actual date is set to the date 3rd Wednesday before Easter Sunday 1993 ⇒ 19930324.

• ‘gcal %1994@t’ (≡ yyyy@t)

The actual date is set to today’s date in 1994 ⇒ 199400217.

• ‘gcal %1994@t10’
• ‘gcal %1994@t+10’ (≡ yyyy@t[+]n)

The actual date is set to the date 10 days after today’s date in 1994 ⇒ 19940227.

• ‘gcal %1994@t-10’ (≡ yyyy@e-n)

The actual date is set to the date 10 days before today’s date in 1994 ⇒ 19940207.

• ‘gcal %0@t’
• ‘gcal %00@t’
• ‘gcal %000@t’
• ‘gcal %0000@t’ (≡ yyyy@t)

The actual date is set to today’s date ⇒ 19930217.

• ‘gcal %0@t10’
• ‘gcal %0@t+10’ (≡ yyyy@t[|]n)

The actual date is set to the date 10 days after today’s date ⇒ 19930227.

• ‘gcal %0@t-10’ (≡ yyyy@t-n)

The actual date is set to the date 10 days before today’s date ⇒ 19930207.

• ‘gcal %1994@t3Wed’
• ‘gcal %1994@t+3Wed’ (≡ yyyy@t[+]nwww)

The actual date is set to the date 3rd Wednesday after today’s date in 1994 ⇒ 19940309.

• ‘gcal %1994@t-3wed’ (≡ yyyy@t-nwww)

The actual date is set to the date 3rd Wednesday before today’s date in 1994 ⇒ 19940302.

• ‘gcal %0@t3wed’
• ‘gcal %0@t+3wed’ (≡ yyyy@t[+]nwww)

The actual date is set to the date 3rd Wednesday after today’s date ⇒ 19930310.

• ‘gcal %0@t-3we’ (≡ yyyy@t-nwww)

The actual date is set to the date 3rd Wednesday before today’s date ⇒ 19930127.

• ‘gcal -v a=0303 %1994@a’ (≡ yyyy@dvar)

The actual date is set to the date 3rd March 1994 ⇒ 19940303.

• ‘gcal -v a=0303 %1994@a10’
• ‘gcal -v a=0303 %1994@a+10’ (≡ yyyy@dvar[+]n)

The actual date is set to the date 10 days after 3rd March 1994 ⇒ 19940313.

• ‘gcal -v a=0303 %1994@a-10’ (≡ yyyy@dvar-n)

The actual date is set to the date 10 days before 3rd March 1994 ⇒ 19940221.

• ‘gcal -v a=0303 %0@a’ (≡ yyyy@dvar)

The actual date is set to the date 3rd March 1993 ⇒ 19930303.

• ‘gcal -v a=0303 %0@a10’
• ‘gcal -v a=0303 %0@a+10’ (≡ yyyy@dvar[+]n)

The actual date is set to the date 10 days after 3rd March 1993 ⇒ 19930313.

• ‘gcal -v a=0303 %0@a-10’ (≡ yyyy@dvar-n)

The actual date is set to the date 10 days before 3rd March 1993 ⇒ 19930221.

• ‘gcal -v a=0303 %1994@a3wed’
• ‘gcal -v a=0303 %1994@a+3wed’ (≡ yyyy@dvar[+]nwww)

The actual date is set to the date 3rd Wednesday after 3rd March 1994 ⇒ 19940323.

• ‘gcal -v a=0303 %1994@a-3wed’ (≡ yyyy@dvar-nwww)

The actual date is set to the date 3rd Wednesday before 3rd March 1994 ⇒ 19940216.

• ‘gcal -v a=0303 %0@a3wed’
• ‘gcal -v a=0303 %0@a+3wed’ (≡ yyyy@dvar[+]nwww)

The actual date is set to the date 3rd Wednesday after 3rd March 1993 ⇒ 19930324.

• ‘gcal -v a=0303 %0@a-3wed’ (≡ yyyy@dvar-nwww)

The actual date is set to the date 3rd Wednesday before 3rd March 1993 ⇒ 19930210.

The month name[dd] format:

• ‘gcal %April’ (≡ month name)

The actual date is set to the date 17th April 1993 ⇒ 19930417.

• ‘gcal %april3’

• ‘gcal %april03’ (≡ month name dd)

The actual date is set to the date 3rd April 1993 ⇒ 19930403.

The weekday name[n] format:

• ‘gcal %Friday’ (≡ weekday name)

The actual date is set to the date Friday of current week in February 1993 ⇒ 19930219.

• ‘gcal %Friday2’

• ‘gcal %Friday02’ (≡ weekday name n)

The actual date is set to the date 2nd Friday in February 1993

(n ≡ ‘1…5|9’) ⇒ 19930212.

The dd format:

• ‘gcal %3’

• ‘gcal %03’ (≡ dd)

The actual date is set to the date 3rd February 1993 ⇒ 19930203.

3.1.4 Commands

The commands control the periods Gcal respects. They can be partitioned into four major classes, namely single commands, special 3-Month mode commands, lists of commands and ranges of commands. Single commands and special 3-Month mode commands only create one calendar sheet, ranges of commands or list of commands create more than one calendar sheet per program run. If a list of commands is given, Gcal works sequentially on each single command given in the list, one by one. A range of commands is expanded first by Gcal and after expansion is done, Gcal works sequentially on each single command produced during the internal expansion step.

For understanding the notation used in the text below, mm is either a number, a month name or one of the special 3-Month mode commands ., .., .+ and .-. No lists or ranges of months or years can be created in case a special 3-Month mode command is given.

The range of mm is valid from 1 to 12 or valid from January to December. Both notations may be mixed. Month names may be abbreviated up to their first three characters. yyyy is a number. The range of yyyy is valid from 1 to 9999. A range (mm-mm or yyyy+yyyy or …) must consist of two elements. A list (mm,…,mm or yyyy;…;yyyy or …) must contain two elements minimum and may have any number of elements.

Here is a list of all ranges of commands, lists of commands and other command separator characters:

• The ‘-’ character is the range of months separator.

• The ‘,’ character is the list of months separator.

• The ‘+’ character is the range of years separator.

• The ‘;’ character is the list of years separator.

• The ‘/’ character is the month of year separator.

• The ‘:’ character is the fiscal year separator.

3.1.4.1 Single commands

NOTHING Current month of current year, called single month mode in the further context.

month name

Selected month of current year, e.g.:

               gcal may
               gcal OCTOBER
               gcal ja

yyyy Selected year, called single year mode in the further context, e.g.:

gcal 1992

But there is an exception of this general rule. If the specified year number is less or equal twelve, Gcal assumes that a selected month of the current year is wanted. If you want to have a selected year less or equal twelve, call Gcal like this

gcal 1:6

or like this

gcal 6+6

to obtain the year AD 6.

mm yyyy

mm/yyyy Single month of selected year, e.g.:

gcal 9 1992
gcal 9/1992
gcal Nov 1777
gcal Nov/1777

: Single fiscal year, starting on the actual month of the actual year and ending on month actual month-1 of the actual year+1, called implicit fiscal year in the further context, e.g.:

gcal :

mm: Single fiscal year, starting on month mm of the actual year and ending on month

mm-1 of the actual year+1, called explicit fiscal year in the further context, e.g.:

gcal 6:
gcal feb:
gcal NOVEMBER:

:yyyy Single fiscal year, starting on the actual month of year yyyy and ending on month actual month-1 of year yyyy+1, also called explicit fiscal year in the further context, e.g.:

gcal :1999

mm:yyyy Single fiscal year, starting on month mm of year yyyy and ending at month

mm-1 of year yyyy+1, called explicit fiscal year too in the further context, e.g.:

gcal 7:1999
gcal oct:1992
gcal February:1777

3.1.4.2 3-Month mode commands

. Previous, actual and next month of the current year, e.g.:

gcal .

. yyyy Previous, actual and next month of selected year, e.g.:

gcal . 1992

.. Actual quarter of the current year, e.g.:

gcal ..

.. yyyy Actual quarter of selected year, e.g.:

gcal .. 1992

.+ Actual and next two months of the current year, e.g.:

gcal .+

.+ yyyy Actual and next two months of selected year, e.g.:

gcal .+ 1992

.- Actual and previous two months of the current year, e.g.:

gcal .-

.- yyyy Actual and previous two months of selected year, e.g.:

gcal .- 1992

3.1.4.3 Lists of commands

mm,. . .,mm

List of specified months of the current year, e.g.:

gcal 1,5,12
gcal 3,1,5,3
gcal june,9,jan
mm/yyyy,. . .,mm/yyyy

List of months in specified year, e.g.:

gcal 1/1992,5,12/2001
gcal june/1991,9/1801,jan
mm,. . .,mm yyyy

List of specified months of selected year, e.g.:

gcal 3,7 1999
gcal 1,dec,july 1777
yyyy;. . .;yyyy

List of specified years, e.g.:

gcal 1992;1777;1899
mm:yyyy;. . .;mm:yyyy

List of specified fiscal years, e.g.:

gcal 7:1999;8:1992;April:3

Other valid lists:

mm,mm/yyyy,. . .

Mixed list of months consisting of: mm mm/yyyy

gcal 6,8,3/1999,feb/3,january

mm:;mm:yyyy;. . .

Mixed list of fiscal years of the given / actual year consisting of: mm: :yyyy

mm:yyyy yyyy
gcal 3:;1994;february:1999;:1777;JAN:
mm:;mm:;. . .

List of fiscal years of the actual year consisting of: mm:

gcal 3:;august:;DEC:

3.1.4.4 Ranges of commands

mm-mm Range of current months, e.g.:

gcal 3-7
gcal 11-8
gcal MAY-dec
gcal 12-aug
mm/yyyy-mm/yyyy

Range of months of year, e.g.:

gcal 3/1991-july/1992
gcal dec/1994-3
gcal sep-dec/2000
mm-mm yyyy

Range of specified months of selected year, e.g.:

gcal 3-7 1999
gcal nov-3 1777
gcal aug-dec 1992
mm-mm yyyy+yyyy

Range of specified months of selected years, e.g.:

gcal 3-7 1999+2001
gcal nov-3 1777+1600
gcal aug-dec 1992+1994
mm yyyy-yyyy
mm/yyyy-yyyy

Range of specified month of selected years, e.g.:

gcal AUGUST 1494-1497
gcal 3/1993-1999
gcal nov/3-1777
yyyy+yyyy

Range of specified years, e.g.:

gcal 1992+1994
mm:yyyy+mm:yyyy

Range of specified fiscal years, e.g.:

gcal 7:1999+8:1992
gcal aug:1992+july:1999

3.2 The GCAL environment variable

The environment variable GCAL can hold a set of default options for Gcal (see [Environment Variable GCAL], page 255). These options are interpreted first at program start-up, and some can be overwritten by explicit command line arguments or values found in a @file response file. See Section 3.1 [Command line arguments], page 8, and Section 3.1.2 [Response file], page 65, for further information.

For example:

• for sh:

           GCAL="-i @color -b 3"; export GCAL

• for csh:

           setenv GCAL "-i @color -b 3"

• for MS-DOS:

           set GCAL=-i @color -b 3

On Vax/VMS, the name of the environment variable is GCAL_OPT, to avoid a conflict with the symbol GCAL set for invocation of the program.

For example:

• for VMS:

           define GCAL_OPT "-i @color -b 3"

4 Eternal Holidays

The eternal holiday list is created only for dates after AD 29. It is assumed that the Gregorian Reformation has occurred from 10th till 22nd March 1924 for all Christian-Orthodox holidays, which are inserted into the eternal holiday list.

The eternal holiday list can be displayed in different ways:

If you start Gcal without an explicit (fiscal) date and the eternal holiday list argument (see [Calendar option --holiday-list[=long|short]], page 14), e.g.:

     gcal -qfr --holiday-list
     gcal -qfr -n
     gcal -qfr -n :

Gcal displays all holidays of the eternal holiday list that refer to the actual / fiscal year, without a leading calendar sheet.

In case you start Gcal with the eternal holiday list argument and a month (plus an additional year), e.g.:

     gcal -qbe --holiday-list july
     gcal -qbe -n july
     gcal -qbe -n 7 1993
     gcal -qbe -n 7/1993

Gcal displays a month calendar sheet of the specified month (of year), and trailing those holidays of the eternal holiday list that refer to the given month (of the given year).

If you start Gcal with the eternal holiday list argument and a simple year (or an explicit fiscal year), e.g.:

     gcal -qbe --holiday-list 1993
     gcal -qbe -n 1993
     gcal -qbe -n july:
     gcal -qbe -n :1993
     gcal -qbe -n 7:1993

Gcal displays a year calendar sheet of the specified year or fiscal year, and trailing all holidays of the eternal holiday list that refer to the given year or fiscal year.

In case you start Gcal with the eternal holiday list argument and a special 3-Month mode command (see Section 3.1.4.2 [3-Month mode commands], page 73), e.g.:

     gcal -qbe --holiday-list .
     gcal -qbe -n .
     gcal -qbe -n ..
     gcal -qbe -n .+
     gcal -qbe -n .-

Gcal displays the according three months by using a fixed-style year calendar sheet, and trailing all holidays of the eternal holiday list that refer to these months.

In case you start Gcal with the eternal holiday list argument and a list or range of commands, it produces the according series of eternal holiday lists. See Section 3.1.4.3 [Lists of commands], page 74, and Section 3.1.4.4 [Ranges of commands], page 75, for further details.

Some annotations:

The following table lists all those characters which are used for marking an entry in the eternal holiday list, i.e. directly lead its date:

Character  Description

+

Legal holiday which is valid in the whole country. Is automatically provided with highlighting sequences respectively marking characters.

#

Legal holiday which is valid in major parts of the whole country. Is automatically provided with highlighting sequences respectively marking characters.

*

Legal holiday which is valid in minor parts of the whole country. Is not provided with highlighting sequences respectively marking characters.

-

Other holiday which serves for memorial or remarking purposes only. Is not provided with highlighting sequences respectively marking characters.

By using the --cc-holidays=cc[+…] option, it is possible to provide the eternal holiday list with additional country specific holidays. Thereafter, those country specific holidays are part of the list. See [Calendar option --cc-holidays=cc[+…]], page 17. For example:

     --cc-holidays=be+fr

causes the inclusion of Belgian and French holidays into the eternal holiday list. The inclusion of country specific holidays is not coupled to Gcal’s internationalization (see Appendix C [Aspects in Internationalization], page 131). This means, no country specific holidays are automatically respected at program start-up for a definite territory or country. So it is up to the user to decide which country specific holidays are included into the eternal holiday list. See Section 3.2 [The GCAL environment variable], page 76, for information how Gcal can be induced to include the country specific holidays automatically into the eternal holiday list at program start-up.

The free selection of the different country specific holidays can be used very ingenious in case information of a definite territory or country is required. Assuming a business man from Germany proposes to travel to Italy in November and needs the information which legal holidays are celebrated there on working days so he is able to plan his dates. This can be determined as follows:

     gcal -f /dev/null -q IT -D r -cE nov

5 Fixed Dates

This chapter describes how to use the fixed date feature of Gcal. Normally, the fixed dates are stored line by line in a so-called resource file (see Section 5.1 [Resource file], page 82). On request, Gcal searches any of such resource files for those fixed dates, that are happening in the period for which a fixed date list is wanted. If Gcal has found any fixed dates that are happening in the requested period, the program displays them collectively as a fixed date list.

Thus, each time you execute Gcal in simple single month mode, i.e. no explicit month, year or fiscal year is given in the command line, but the fixed date argument, e.g.:

     gcal --list-of-fixed-dates
     gcal -c

it checks the resource file for dates and processes those that refer to the actual system date (≡today). Then Gcal displays all fixed dates found in the resource file which refer to the actual day (week/month/year) of the actual year, without a leading calendar sheet.

Some more examples:

     gcal --ast -ce
     gcal -cw-
     gcal -c-m
     gcal -c-y-
     gcal -qbe -jcb -c-et
     gcal --zod -cey-
     gcal -qfr -jc -cey

The fixed date list can be embraced by different other leading and trailing output, respectively, that is also created by Gcal:

In case you start Gcal with the fixed date argument and/or a month and/or an explicit year or fiscal year; respectively with the special 3-Month mode command, e.g.:

     gcal -c july
     gcal -qfr -ce 7 1994
     gcal -c-y 7/1994
     gcal -c 7:1993
     gcal -c 1993
     gcal -c .
     gcal -c .+

Gcal displays a (three) month / year calendar sheet of the specified respectively actual month / year, and trailing those fixed dates found in the resource file that refer to the given period; the intensity level is the same as the -cy option if the period refers to a year or fiscal year or to a three month period, respectively, is the same as the -cm option if the period refers to a month.

If you start Gcal without an explicit date and the fixed date argument in connection with the eternal holiday list argument, e.g.:

     gcal -qbe -n -c
     gcal -qbe -n -Cm
     gcal -qbe -n -Cl1224
     gcal -qbe -n -cl3+
     gcal -qbe -n -4-

then Gcal displays all fixed dates found in the resource file that refer to the requested period, and after that all holidays of the eternal holiday list that refer to the whole actual year, without a leading calendar sheet. See Chapter 4 [Eternal Holidays], page 77, for more details how to use an eternal holiday list.

Annotation to previous examples:

Each time you execute Gcal in simple single year mode, i.e. no explicit month, year or fiscal year is given in the command line, but the fixed date argument (with optional modifiers) and the number of blocks argument (see [Calendar option --blocks=number], page 31), e.g.:

     gcal -c --blocks=3
     gcal -C-em -b 3 --cc-holidays=be
     gcal -c-t -b 6 -jc

it checks the resource file for dates and processes those that refer to the actual system date (≡today). Gcal displays a leading year calendar sheet, and all fixed dates found in the resource file that refer to the actual day (week/month/year) of the actual year.

You can list fixed dates of past, present or future month/years, too. This can be done on the one hand, if you use the %date option in the way you require (see Section 3.1.3 [Actual date modifier], page 67), and on the other hand, if a command respectively a list or range of commands is used.

For example:

‘gcal -qbe -c -n- 1993’

Lists all fixed dates of entire 1993 in ascending order, and the eternal holiday list of entire 1993 in descending order; the intensity level is the same as the -cy option.

‘gcal -c- july:1993’

Lists all fixed dates of fiscal year —starting in July 1993 and ending in June 1994— in descending order; the intensity level is the same as the -cy option.

‘gcal -qbe -ce july 1993’

Lists all fixed dates of July 1993 inclusive all eternal holidays merged into this list in ascending order; the intensity level is the same as the -cm option.

‘gcal -qbe -c-e -n 1993+1996’

Lists all fixed dates of entire 1993, 1994, 1995 and 1996 (series of years) inclusive all eternal holidays merged into this list in descending order, and the eternal holiday list related to all above years in ascending order; the intensity level is the same as the -cy option.

‘gcal -cm %19920317’

Lists all fixed dates of March 1992 in ascending order.

‘gcal -qbe %1994@e -n -c’

Lists all fixed dates of Easter Sunday 1994 in ascending order, and the eternal holiday list of entire 1994 in ascending order.

5.1 Resource file

This section describes how to use a resource file that contains fixed dates and appointments. The name of the standard resource file is HOME/.gcalrc^{1}.

Gcal uses a special file searching mechanism:

1. The Gcal program scans the file directories which are referenced in the environment variable GCALPATH (see [Environment Variable GCALPATH], page 255) to find the standard resource file. The GCALPATH environment variable contains a (‘:’) colon-separated^{2} list of file directories. If no such environment variable is defined and set, Gcal omits this step.

2. If the above action fails, Gcal looks for the standard resource file in the file directory which is referenced in the environment variable HOME. If no HOME environment variable is defined and set, Gcal also omits this step. See [Environment Variable HOME], page 255.

3. If the above action fails, Gcal inspects the user data file directory^{3} to find the standard resource file. This file directory is a file directory relative to the HOME file directory. If an environment variable GCAL_USR_DATADIR is set, Gcal appends the contents of this environment variable to the contents of the HOME environment variable, and tries to use this file directory instead of using the burned-in default name of this file directory (see [Environment Variable GCAL_USR_DATADIR], page 255). If no HOME environment variable is defined and set, Gcal of course also omits this step.

4. If all above actions fail, Gcal inspects the system data file directory^{4} to find the standard resource file. This file directory is a file directory absolute to the root/$prefix file directory. If an environment variable GCAL_SYS_DATADIR is set, Gcal tries to use this file directory instead of using the burned-in default name of this file directory (see [Environment Variable GCAL_SYS_DATADIR], page 255).

If the standard resource file is not found during the steps 1…4 of the file searching mechanism, no fixed date messages will be created!

In case the --debug=abort option is given and no standard resource file is found during the steps 1…4 of the file searching mechanism, the Gcal program will be aborted with an error code. See [Error Code 118], page 257, and [Global option --debug=abort], page 11, for more information.

If a -f|F name[+…] respectively --resource-file=name command line argument is given, a file name will be used for the resource file instead of the standard resource file. See [Fixed date option --resource-file=name], page 46.

1 On MS-DOS, OS/2 and some other operating systems, the name of the standard resource file is

HOME/gcalrc (without a leading ‘.’).

2 On MS-DOS, OS/2 and some other operating systems, another character will be used, e.g. ‘;’ on

MS-DOS and OS/2.

3 Normally HOME/share/gcal.

4 Normally /usr/local/share/gcal or $prefix/share/gcal.

Annotation:

An extended file name contains one ore more ‘/’ (slash) characters^{5} and denotes a file, whose fixed access path either starts from the root file directory, like ‘/foo/bar/file’, or from the actual file directory, like ‘./bar/file’ or ‘../bar/file’. If name is an extended file name:

• In case file name is found, Gcal will use it.

• In case file name is not found, Gcal will not use it.

A simple file name denotes a file, whose access-path either does not start from the root file directory or from the actual file directory, like ‘file’ or ‘bar/file’. If name is a simple file name:

• Gcal tries to find this file name by using the previously explained file searching mechanism.

5 Operating systems like MS-DOS or OS/2 use the ‘\’ (Backslash) character for separating a file directory

from a file name.

5.1.1 Structure of resource file

First of all, a Gcal resource file is a plain ascii text file. This text file may be created by any text editor or by redirecting the standard output channel to a file, e.g.:

      $ echo '19930217 Text'>> resource-file RET

A special —but simple— line structure is required so Gcal is able to interpret its contents. Each fixed date entry in a resource file must be split into two parts, namely a date part and an optional text part which must be separated by one whitespace^{1} character minimum. It is unnecessary to give a whitespace separator character if no text part is specified.

A line must always end with a ‘\n’ (newline) character, except it is the last line of a resource file. The maximum length of a line is limited to INT_MAX^{2} characters. A newline character is automatically appended to the line if the standard output channel is directed to a file. A newline character is appended to the typed line in a text editor window if it is completed by pressing the RET key. In case the used text editor does not generate the newline character in this way, it should be set to this mode of operation, otherwise this text editor is useless for creating Gcal resource files.

The line structure of fixed date entries is:

     date part [ whitespace text part ] newline

or more concrete, e.g.:

     yyyy[mm[dd|wwwn]] [ whitespace text ] newline

or much more concrete, e.g.:

     19940217 Hi, I'm the text!

1 For example Tab, Space, Form-feed and similar characters.

2 Normally, this is for example equivalent to a value of 32,767 ≡ (2^15)-1 on a 16-Bit machine archi-

tecture; to a value of 2,147,483,648 ≡ (2^31)-1 on a 32-Bit machine architecture etc.

Besides fixed date entries, a resource file may contain further entries like:

Comments…

; A remarked line
;     A formatted and multi-line \
      remark

Include directives…

#include <file name>
#include "file name"

Date variable assignments respectively operations…

dvar=NOTHING
dvar=mmdd
dvar=mmwwwn
dvar=*dn[www]
dvar=*wn[www]
dvar=dvar[[+|-]n[www]]
dvar++
dvar--
dvar+=[+|-]n
dvar-=[+|-]n
dvar+=nwww
dvar-=nwww

Text variable assignments respectively operations…

tvar=[text]
tvar?[command]
tvar:[command]
tvar++
tvar--
tvar+=[+|-]n
tvar-=[+|-]n

Text variable references…

tvar

Text variable references at the beginning of a Gcal resource file line may only be used if it is ensured that they are expanded to a valid Gcal resource file line.

5.1.2 Date part of a line

The structure of a date part —which gives Gcal the information at which date a fixed date happens— of a line in the resource file is

either: yyyy[mm[dd|wwwn]]

yyyy (4 digits), is the year including the century (range 0000…9999). Leading zeroes are required in case the defined year is less than 1000 and other components of the date part are trailing the year.

mm (2 digits or 3 characters), is the month (range 00…12 or 99, respectively Jan, Feb…). A given 99 for the month means the last month of the year (≡ December). Leading zeroes are required in case the defined month is less than 10 and other components of the date part are trailing the month.

dd (2 digits), is the day (range 00…31 or 99). A given 99 for the day means the last day of the month. Leading zeroes are required in case the defined day is less than 10 and other components of the date part are trailing the day.

www (2…3 characters), is a short weekday name (range Mon…Sun).

n (1 digit), is the n’th weekday www of month (range 1…5 or 9).

• n ≡ 1…5 n’th weekday www of month.

• n ≡ 9 Last weekday www of month.

or: yyyy*d|wn[www]

yyyy (4 digits), is the year including the century (range 0000…9999). Leading zeroes are required in case the defined year is less than 1000 and other components of the date part are trailing the year.

*d (2 characters), is the reference to an ordinary date.

*w (2 characters), is either the reference to a date of an ISO week or the reference to a date of a standard week.

n (1…3 digits), is the value in days or in weeks, the fixed date occurs. A given 99 connected with a short weekday name means the last weekday www of the year. A given 999 connected with *d means the last day of a year, i.e. the 31st December. If the computed date does not occur in the year, i.e. exceeds the year bounds, the fixed date message is suppressed.

www (2…3 characters), is a short weekday name (range Mon…Sun).

or: yyyy@e|t|dvar[[+|-]n]

yyyy (4 digits), is the year including the century (range 0000…9999). Leading zeroes are required in case the defined year is less than 1000 and other components of the date part are trailing the year.

@e (2 characters), is the reference to the Easter Sunday’s date.

@t (2 characters), is the reference to today’s date.

@dvar (2 characters), is the reference to a date variable.

[+|-]n (1…4 alphanumeric characters), is the optional displacement value in days, the fixed date occurs relative to the Easter Sunday’s date, relative to today’s date, or relative to a date variables date. A given -999 means the first day of a year, i.e. the 1st January. A given +999 or 999 means the last day of a year, i.e. the 31st December. If the computed date does not occur in the year, i.e. exceeds the year bounds, the fixed date message is suppressed.

or: yyyy@e|t|dvar[+|-]nwww

yyyy (4 digits), is the year including the century (range 0000…9999). Leading zeroes are required in case the defined year is less than 1000 and other components of the date part are trailing the year.

@e (2 characters), is the reference to the Easter Sunday’s date.

@t (2 characters), is the reference to today’s date.

@dvar (2 characters), is the reference to a date variable.

[+|-]nwww (3…7 alphanumeric characters), is the optional displacement value in weekdays, the fixed date occurs relative to the Easter Sunday’s date, relative to today’s date, or relative to a date variables date. A given -99 means the first weekday www of the year. A given +99 or 99 means the last weekday www of the year. If the computed date does not occur in the year, i.e. exceeds the year bounds, the fixed date message is suppressed.

Please note:

• If yyyy is specified as 0000, 000, 00 or 0:

The month and day are assumed to be annual events and the text part will be displayed for any year.

• If mm is not defined or specified as 00 or 0:

The day is assumed to be a monthly event for the specified year and the text part will be displayed for any month.

• If dd is not defined or specified as 00 or 0:

Each day is assumed to be a daily event for the specified year and month and the text part will be displayed for any day.

• If www is specified as a short (2…3 characters) weekday name: The given weekday name is assumed to be a weekly event for the specified year and month and the text part will be displayed for any week.

Hint: Preceding table ease misleading assumption that a fixed date, which shall be valid for any days of the year, can be defined with a date part only consisting of a 0. But that is only correct in that the so-defined fixed date is respected in some few fixed date periods only (see Appendix K [Coding Scheme], page 251).

If a fixed date is defined which shall be respected for any days of the year, it must be designed with a 0*d1#999 date part. For example, the same is likewise valid for fixed dates which represent a weekly event, and that shall be valid during all the year. Instead of creating such a fixed date with a 000000www date part (that is likewise respected in some few fixed date periods only), it should be defined with a date part like 0*d1www#99www.7. See Section 5.1.3.2 [Ranges of days], page 90, and Section 5.1.3.4 [Appearance factor of days], page 94, for further information.

5.1.3 Further date part attributes

The date part —which is leading a fixed date text of a Gcal resource file line— may be provided with further attributes, that are either specifications of lists of days or ranges of days, which provide the information at which date a fixed date happens. More further attributes are the repetition factor as well as the appearance factor.

All of those date parts in a Gcal resource file line, which are structured as follows, may generally be provided with further attributes:

  • yyyymm. . .
  • yyyy*d|w. . .
  • yyyy@e|t|dvar. . .

Attention: If the date part is supplied with further attributes and the year of the fixed date is not given in a concrete manner, i.e. the year yyyy is set to zero, such kinds of fixed dates are only computed correctly in simple year bounds —which means during a current year— and not in fixed dates shown after the current year has changed, e.g. in fiscal years.

5.1.3.1 Lists of days

A list of days is used to define recurrent fixed dates and to use only one line in the resource file for them, instead of using several lines in a resource file needed for defining these recurrent fixed dates separately.

A list of days is specified by a ‘,’ separator character; it must contain two elements minimum and may have any number of elements. A single element of the list may not be set to a zero value and be not concrete therefore. (This would not make any sense in this context.)

Assuming a fixed date shall always occur on the 1st and the 15th day in every month of the year 1996, e.g. ‘Tax returns’, one solution would be on the one hand a fixed date entry in the resource file for the 1st of the month and another entry for the 15th of the month, which would be a total of two entries in the resource file for such a fixed date.

On the other hand, this expense can be reduced to a total of only one entry in the resource file by using a list of days, which is likewise valid for the 1st and the 15th day in every month, that is

     19960001,15 Tax returns

The use of lists of days in the date part is permitted as follows:

  • yyyymmdd|www|wwwn,[mm]dd|[mm]www|[mm]wwwn . . .
  • yyyy*d|wn|nwww,n|nwww . . .
  • yyyy@e|t|dvar,|[+|-]n[www],[+|-]n[www] . . .

Some examples to this:

000001fr3,12,99,mon,apr14,993,julfri3,08fr

In every year in January: on the 3rd Friday, on the 12th, on the last day of the month and on every Monday. And in every year: on the 14th April, on the 3rd December, on the 3rd Friday in July and on every Friday in August.

199600mon,fr,12

In the year 1996 in every month: on every Monday, on every Friday and on the 12th.

0*d1,1fr,999,17mo

In every year: on the first day of the year, on the 1st Friday of the year, on the last day of the year and on the 17th Monday of the year.

1996*w1fr,1,17mo,99fr

In the year 1996: on Friday of the 1st week, on Monday of the first week, on Monday of the 17th week and on Friday of the last week.

0@a,+1,-3,5,+4fr,2fr,-10sa

In every year: on the date of the date variable a, one day after the date of a, three days before the date of a, five days after the date of a, four Fridays after the date of a, two Fridays after the date of a and ten Saturdays before the date of a.

1996@e+1,-3,5,+4fr,2fr,-10sa

In the year 1996: one day after the Easter Sunday’s date, three days before the Easter Sunday’s date, five days after the Easter Sunday’s date, four Fridays after the Easter Sunday’s date, two Fridays after the Easter Sunday’s date and ten Saturdays before the Easter Sunday’s date.

1996@t+1,-3,5,+4fr,2fr,-10sa

In the year 1996: one day after today’s date, three days before today’s date, five days after today’s date, four Fridays after the today’s date, two Fridays after today’s date and ten Saturdays before today’s date.

5.1.3.2 Ranges of days

A range of days just as a list of days is used to define recurrent fixed dates and to use only one line in the resource file for them, instead of using several lines in a resource file needed for defining these recurrent fixed dates separately.

A range of days is specified by a ‘#’ separator character and must consist of two elements, namely the starting day and the final day. The starting day and likewise the final day of the range may not be set to a zero value and be not concrete therefore. (This would not make any sense in this context.)

Assuming a fixed date shall always occur during the 1st and the 15th day (inclusive) in every month of the year 1996, e.g. ‘Inside work’, one solution would be on the one hand a fixed date entry in the resource file for the 1st of the month, another one for the 2nd of the month until the 15th of the month, which would be a total of 15 entries in the resource file for such a fixed date^{1}.

On the other hand, this expense can be reduced to a total of only one entry in the resource file by using a range of days, which is likewise valid for the 1st until the 15th day in every month, that is

     19960001#15 Inside work

1 Apart from the possibility, that a list of days can be used in such a case.

The use of ranges of days in the date part is permitted as follows:

  • yyyymmwww#www
  • yyyymmdd|wwwn#[mm]dd|[mm]wwwn
  • yyyy*d|wn[www]#n[www]
  • yyyy@e|t|dvar[[+|-]n[www]]#[+|-]n[www]

Preceding table shows that the starting and ending day can only be specified by using date formats of the same type left and right the ‘#’ separator character. Thus, it is not explicitly possible to define a range of days by using different types of date formats, means, fixed date entries like:

     1996@e+3fr#1012 In the year 1996: every day that appears within the\
                     period of the date 3 Fridays after Easter Sunday's\
                     date until the 12th October.

cannot be defined like this. But this problem can be solved by using an inclusive date period %i[date][#[date]] special text for defining the ending day of the range^{1}, and the fixed dates are simply produced until the last day of the year, e.g.:

     1996@e+3fr#+999 In the year 1996: every day that appears within the\
                     period of the date 3 Fridays after Easter Sunday's\
                     date until the 12th October.%i0#19961012

See Section G.1.1.1 [Inclusive date period %i[date][#[date]] special text], page 146, for more details. Some examples to this:

199600mon#fr

In the year 1996 in every month: every day that appears within the day sequence

          ‘mon, tue, wed, thu, fri’.
000000fr#mon

In every year in every month: every day that appears within the day sequence

          ‘fri, sat, sun, mon’.
000001fr3#12

In every year in January: every day that appears within the period of the 3rd Friday of the month until the 12th of the month.

00000112#fr3

In every year in January: every day that appears within the period of the 12th of the month until the 3rd Friday of the month.

00000112#augfri3

In every year: every day that appears within the period of the 12th January until the 3rd Friday of August.

0000sep13#99fr9

In every year: every day that appears within the period of the 13th September until the last Friday of December.

1 Implicit solution.

0*d1#1fr In every year: every day that appears within the period of the 1st day of the year until the 1rd Friday of the year.

0*d99fr#333

In every year: every day that appears within the period of the last Friday of the year until the 333rd day of the year.

1996*w1fr#17mo

In the year 1996: every day that appears within the period of the Friday of the 1st week until the Monday of the 17th week.

0@a#+4fr In every year: every day that appears within the period of the date of the date variable a until the 4th Friday after the date of the date variable a.

1996@e-3#+9fr

In the year 1996: every day that appears within the period of the date three days before the Easter Sunday’s date until the 9th Friday after the date of the Easter Sunday’s date.

1996@t-3#+99fr

In the year 1996: every day that appears within the period of the date three days before today’s date until the last Friday of the year.

5.1.3.3 Repetition factor of days

A repetition factor of days is used to define recurrent fixed dates and to use only one line in the resource file for them, instead of using several lines in a resource file needed for defining these recurrent fixed dates separately.

A repetition factor of days (‘:n’) is specified by a ‘:’ character and must trail the day field —which must have a concrete value in a date part of a Gcal resource file and has not been set to a zero value— respectively either lead or trail an appearance factor of days. Except ranges of days or fixed dates which occur only on a definite weekday like ‘199600mon Every Monday 1996’, such a repetition factor may be specified in all possible codings of date parts of a Gcal resource file.

This factor may have values in range 1…999. Values greater than the real difference between the date of the first occurrence of the fixed date and the last day of the year are always reduced to this real difference. In this sense, any value greater 365^{2} means a repetition factor until the last day of the year.

Assuming a fixed date shall always occur on the 15th day in every month of the year 1996 and covers a period of four days (inclusive the 15th itself), e.g. ‘Co-workers training’, one solution would be on the one hand a fixed date entry in the resource file for the 15th of the month and the succeeding three days, which would be a total of four entries in the resource file for such a fixed date^{3}.

2 Or 366 in leap years.

3 Apart from the possibility, that a list or a range of days can be used in such a case.

On the other hand, this expense can be reduced to a total of only one entry in the resource file by using a repetition factor of days, which is likewise valid for the 15th and the three days which succeed the 15th in every month, namely

19960015:4 Co-workers training

The use of repetition factors of days in the date part is permitted as follows:

• yyyymmdd:n|wwwn:n
• yyyy*d|wn:n|nwww:n
• yyyy@e|t|dvar[[+|-]n[www]]:n
• yyyymmdd:n|wwwn:n,[mm]dd:n|[mm]wwwn:n . . .
• yyyy*d|wn:n|nwww:n,n:n|nwww:n . . .
• yyyy@e|t|dvar:n,|[+|-]n[www]:n,[+|-]n[www]:n . . .

Some examples to this:

000001fr3:11

In every year in January: every day that appears within the period of the 3rd Friday of the month and the succeeding ten days.

00000112:3

In every year in January: every day that appears within the period of the 12th of the month and the succeeding two days.

00000112:3,fr3:5

In every year in January: every day that appears within the period of the 12th of the month and the succeeding two days, and that appears within the period of the 3rd Friday of the month and the succeeding four days.

0*d1:1 In every year: every day that appears within the period of the 1st day of the year and no succeeding day.

0*d1:2 In every year: every day that appears within the period of the 1st day of the year and the succeeding day.

0*d99fr:333

In every year: every day that appears within the period of the last Friday of the year and the succeeding 332 days. Well, in cases a fixed date exceeds the year bounds, it will only be produced until the last day of the year.

1996*w1fr:17

In the year 1996: every day that appears within the period of the Friday of the 1st week and the succeeding 16 days.

0@a:4 In every year: every day that appears within the period of the date of the date variable a and the succeeding three days.

1996@e-3:9

In the year 1996: every day that appears within the period of the date three days before the Easter Sunday’s date and the succeeding eight days.

1996@t-3:9

In the year 1996: every day that appears within the period of the date three days before today’s date and the succeeding eight days.

5.1.3.4 Appearance factor of days

An appearance factor of days is used to define a concrete displacement of recurrent fixed dates.

An appearance factor of days (‘.n’) is specified by a ‘.’ character and must trail the day field —which must have a concrete value in a date part of a Gcal resource file and has not been set to a zero value— respectively either lead or trail a repetition factor of days. Except fixed dates which occur only on a definite weekday and are not specified by using a range of days, like ‘199600mon Every Monday 1996’, such an appearance factor may be specified in all possible codings of date parts of a Gcal resource file. This factor may be specified with each single element of lists of days, but in a range of days, this factor may trail only the final day of the range. Well, the use of such an appearance factor is only helpful if it is either given in a range of days, or if it is given together with a repetition factor.

This factor may have values in range 1…999. Fixed dates will be ignored if the factor takes values greater than the real difference between the date of the occurrence of the fixed date and the last day of the year, respectively the end of the period, for which the fixed dates shall either be produced or respected.

Assuming a fixed date shall always occur on the 15th day in every month of the year 1996 and covers a period of seven days (inclusive the 15th itself), but shall only be respected every third day (i.e. two days have to be skipped at a time) within this period, e.g. ‘Training-college’, one solution would be on the one hand a fixed date entry in the resource file for the 15th of the month, for the 18th and for the 21st of the month, which would be a total of three entries in the resource file for such a fixed date^{4}.

On the other hand, this expense can be reduced to a total of only one entry in the resource file by using an appearance factor of days, which is likewise valid for the 15th, the 18th and the 21st in every month, namely on the one hand by the use of a repetition factor

     19960015:7.3 Training-college

or on the other hand by the use of a range of days

     19960015#21.3 Training-college

The use of appearance factors of days in the date part is permitted as follows:

  • yyyymmdd:n.n|dd.n:n
  • yyyymmwwwn:n.n|wwwn.n:n
  • yyyy*d|wn:n.n|n.n:n
  • yyyy*d|wnwww:n.n|nwww.n:n
  • yyyy@e|t|dvar[[+|-]n[www]]:n.n
  • yyyy@e|t|dvar[[+|-]n[www]].n:n
  • yyyymmdd:n.n|wwwn:n.n,[mm]dd:n.n|[mm]wwwn:n.n . . .
  • yyyymmdd.n:n|wwwn.n:n,[mm]dd:n.n|[mm]wwwn.n:n . . .
  • yyyy*d|wn:n.n|nwww:n.n,n:n.n|nwww:n.n . . .
  • yyyy*d|wn.n:n|nwww.n:n,n.n:n|nwww.n:n . . .
  • yyyy@e|t|dvar:n.n,|[+|-]n[www]:n.n,[+|-]n[www]:n.n . . .

4 Apart from the possibility, that a list of days can be used in such a case.

• yyyy@e|t|dvar.n:n,|[+|.]n[www].n:n,[+|.]n[www].n:n . . .
• yyyymmwww#www.n
• yyyymmdd|wwwn#[mm]dd|[mm]wwwn.n
• yyyy*d|wn[www]#n[www].n
• yyyy@e|t|dvar[[+|-]n[www]]#[+|-]n[www].n

Some examples to this:

000001fr3:11.3

In every year in January: every day that appears within the period of the 3rd Friday of the month and the succeeding ten days, but only every 3rd day within this period (skip two days at a time).

00000112:3.2

In every year in January: every day that appears within the period of the 12th of the month and the succeeding two days, but only every 2nd day within this period (skip one day at a time).

00000112:3.2,fr3:5.3

In every year in January: every day that appears within the period of the 12th of the month and the succeeding two days, but only every 2nd day within this period (skip one day at a time), and that appears within the period of the 3rd Friday of the month and the succeeding four days, but only every 3rd day within this period (skip two days at a time).

0*d1:10.1

In every year: every day that appears within the period of the 1st day of the year and the succeeding nine days, and this for any day within this period (skip zero days at a time).

0*d1:2.5 In every year: every day that appears within the period of the 1st day of the year and the succeeding day, but only every 5th day within this period (skip four days at a time). Well, the succeeding days (only one in this example) of the starting day (1st day of year) are not respected, because the next day resulted by the appearance factor exceeds the final day (resulted by the repetition factor) of the period.

0*d99fr:333.61

In every year: every day that appears within the period of the last Friday of the year and the succeeding 332 days, but only every 61st day within this period (skip 60 days at a time). Well, in cases a fixed date exceeds the year bounds, it will only be produced until the last day of the year. No succeeding day (332 in this example) of the starting day (last Friday of the year) is respected by reason of the displacement value of 60 days, because the next day resulted by the appearance factor exceeds the final day (resulted by the repetition factor) of the period (the last day of the year).

1996*w1fr:17.8

In the year 1996: every day that appears within the period of the Friday of the 1st week and the succeeding 16 days, but only every 8th day within this period (skip seven days at a time).

0@a:4.3 In every year: every day that appears within the period of the date of the date variable a and the succeeding three days, but only every 3rd day within this period (skip two days at a time).

1996@e-3:9.4

In the year 1996: every day that appears within the period of the date three days before the Easter Sunday’s date and the succeeding eight days, but only every 4th day within this period (skip three days at a time).

1996@e3#-20sun.15

In the year 1996: every day that appears within the period of the date three days after the Easter Sunday’s date until the 20th Sunday before the Easter Sunday’s date, but only every 15th day within this period (skip 14 days at a time).

1996@t3#-20sun.15

In the year 1996: every day that appears within the period of the date three days after today’s date until the 20th Sunday before today’s date, but only every 15th day within this period (skip 14 days at a time).

5.1.4 Text part of a line

The text part of a line in a resource file can be any text you like. Indeed, some characters have a special meaning (‘%’, ‘$’, ‘~’, ‘^’ and ‘\’) and must be protected should the occasion arise that special character combinations are formed with them^{1} which might be used only textually. If the text part contains characters that are used for highlighting the text or format it for a printer, or characters with decimal values above 127 in the code table of the used character set^{2} not produced by Gcal itself, such characters respectively sequences are displayed by Gcal in an uninterpreted manner! For that very reason, it can happen that the output of such characters can potentially create problems with the used screen device driver software and/or the external pager program, likewise mailing of such texts by means of electronic mail.

So-called NUL characters^{3} may also occur in the text part, but they only cause the suppression of all succeeding characters in the line. In my opinion, it makes no perceptible sense to output these NUL characters uninterpreted, so they are used for remarking purposes only; besides, the NUL characters would lead to the same problems as already mentioned above.

A line of the resource file is continued on the next line if a ‘\-\n’ (backslash-newline) character sequence is found, e.g.: The line:

     000000Mo Every Monday

and the lines:

     000000Mo \ Every \ Monday

produce the same output and are essentially equivalent.

Furthermore, you can break-up the text of a long text part at any place you like. The term long means in this context, that the text displayed by Gcal would override the right text margin of the screen respectively break-up at that margin.

Each time a ‘~’ (tilde) character is found in the text part, this character is replaced by a real ‘\n’ (newline) character. Such texts will be displayed by Gcal in a formatted manner at a left margin, this means, they are lead by a definite number of space characters.

You may depreciate the special meaning of the ‘~’ (tilde) character —in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\~’. If you need the ‘\~’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\~’.

Each time a ‘^’ (caret) character is found in the text part, this character is also replaced by a real ‘\n’ (newline) character. Such texts will be displayed by Gcal at column 1, this means, they are not lead by space characters. The rules for protecting the ‘^’ character are the same as the rules used for protecting the ‘~’ character.

1 Like ‘%i’, ‘$a’…

2 These are potentially non-printable.

3 That is the character with the value zero in the code table of the used character set.

The resource file lbrk-1.rc

$ cat lbrk-1.rc
⊣ ; Hi, I'm `lbrk-1.rc' and alive now
⊣ ;
⊣ 0 I know I'm a short text
⊣ 0 I hope I'm long enough~here, a line break-up~\
⊣ and again~and now for the last time...
⊣ 0 I hope I'm also long enough^here, a line break-up\
⊣ ~and again^and now for the last time...
⊣ 0 Am I another short text? Dunno...

is displayed as follows:

$ gcal %00000101 --resource-file=./lbrk-1.rc --disable-highlighting

⊣

⊣ Fixed date list:

⊣

⊣ Sat, Jan   1st 2000: Am I another short text? Dunno...
⊣ Sat, Jan   1st 2000: I hope I'm also long enough
⊣ here, a line break-up
⊣  and again
⊣ and now for the last time...
⊣ Sat, Jan   1st 2000: I hope I'm long enough
⊣  here, a line break-up
⊣  and again
⊣  and now for the last time...
⊣ Sat, Jan   1st 2000: I know I'm a short text

Because whitespace characters are used to separate the date part from the text part^{4}, it is not possible to supply the text part with leading whitespace characters without further ado. If one or more whitespace characters shall lead the text part, this can be arranged by protecting the first of these whitespace characters (and that by placing a ‘\’ (backslash) character before it), e.g. ‘\ ’ if it is a space character. By means of such a character sequence, Gcal notices that the text trailing the ‘\’ character is no longer used for separating purposes, but is member of the text part. Gcal skips this specific, marking backslash character (avoids its output) and processes the rest of the line as usual.

4 These are internally skipped and remain undisplayed therefore.

The following example should be enough to elucidate the above facts. The resource file

wspc-1.rc
$ cat wspc-1.rc
⊣ 0 BANG BANG
⊣ 0 bogus
⊣ 0         bogomips
⊣ 0    \hello world
⊣ 0    \ main(){for(;;)fork();}
⊣ 0 \     sh $0 & $0 &
⊣ 0    \  a  \  b  \\  c  \\\  d
⊣ 0    What happens now?\
⊣ ~0  \  This!
⊣ 0    What happens now?\\
⊣ ~0  \  That!

is displayed as follows:

$ gcal --resource-file=./wspc-1.rc --disable-highlighting

⊣

⊣ Fixed date list:

⊣

⊣ Wed, Jun  14th 2000:      sh $0 & $0 &
⊣ Wed, Jun  14th 2000:   a  \  b  \\  c  \\\  d
⊣ Wed, Jun  14th 2000:  main(){for(;;)fork();}
⊣ Wed, Jun  14th 2000: BANG BANG
⊣ Wed, Jun  14th 2000: What happens now?~0  \  That!
⊣ Wed, Jun  14th 2000: What happens now?
⊣  0  \  This!
⊣ Wed, Jun  14th 2000: \hello world
⊣ Wed, Jun  14th 2000: bogomips
⊣ Wed, Jun  14th 2000: bogus

5.1.5 %?… Special Texts

The text part of a resource file line may be provided with different special texts which may occur in any^{1} number and order. All these texts start with a ‘%’ (percent) character which may be protected by a leading ‘\’ (backslash) character if the ‘%’ (percent) character itself is wanted to be used in the text part, e.g. ‘\%’. The ‘%’ character is mostly trailed by an optional format instruction, (see Appendix J [Format Instruction], page 249, for the detailed description of the format instruction and its components), and then trailed by a distinguishing mark —which defines the kind of the special text to be used— and if it is a letter, Gcal accepts it case-sensitive.

The purpose of these special texts is mainly to suppress output of fixed dates in definite cases, either, or to provide them with particular texts respectively to start external commands. See Appendix G [Description of all %?… Special Texts], page 145, for a complete description of all usable special texts, and Appendix H [Summary of all %?… Special Texts], page 211, for an according short-list of them.

For example, the resource file speci-1.rc

      $ cat speci-1.rc
      ⊣ ; I'm veracious very special :)
      ⊣ ;
      ⊣ 0 1. Today is %>1u*K , %>02&*D  %U  %Y !
      ⊣ 0 2. It's the %>03&*N  day of the year.
      ⊣ 0 3. The actual week number is: %k .
      ⊣ 0 4. Currently, it's %t*  o'clock, Mr. %-USER .
      ⊣ 0 5. Hurry up with your work~\
      ⊣  by reason sunrise is at %o+5158+00738+61,2: .

is displayed as follows (in case today’s date is the 4th October 1996):

      $ gcal --resource-file=speci-1.rc -H no -ox

⊣

      ⊣ Mon, Oct   4th 1999: 1. Today is MONDAY, 04th October 1999!
      ⊣  2. It's the 277th day of the year.
      ⊣  3. The actual week number is: 40.
      ⊣  4. Currently, it's 06:53pm o'clock, Mr. esken.
      ⊣  5. Hurry up with your work
      ⊣  by reason sunrise is at 07:33.

1 Rational!

5.1.6 Comment line

A line^{1} beginning with a ‘;’ (semicolon) character in the resource file is treated as a remark and will not be used by Gcal! Any whitespace characters may lead the remark character. See Section 5.1.4 [Text part of a line], page 97, for information how a NUL character can be used for remarking purposes in a text part.

The following example of the remark.rc resource file —which only consists of remarks— should be enough to elucidate the above facts:

$ cat remark.rc
⊣ ;00001031 ^\
⊣  .......^\
⊣  ;::;::,^\
⊣  ;::;::;,^\
⊣  ;;::;;:;;,^\
⊣  .vmnv\%vnv\%,.;;;:::;;:;;,  .,vnmv\%vnn,^\
⊣  vmmmn\%vmmnv\%vnmmnv\%;;;;;;\%nmmmnv\%vnmv\%vmmv^\
⊣  vmmnv\%vmmmnv\%vnmmmmmnv\%;:\%nmmmmmmnv\%vnm\%vnmmmv^\
⊣  vmmnv\%vmmmnv\%vnmmmmmmmmnv\%nmmmmmmmmnv\%vnmm\%vnmmmv^\
⊣  vmmnv\%vmmmnv\%vnmmmmmmmmnv\%vmmmmmmmmmmnv\%vnmv\%vnmmmv^\
⊣ vmmnv\%vmmmnv\%vnmm;mmmmmmnv\%vmmmmmmmm;mmnv\%vnmv\%vnmmmv,^\
⊣ vmmnv\%vmmmnv\%vnmm;' mmmmmnv\%vmmmmmmm;' mmnv\%vnmv\%vnmmmv,^\
⊣ vmmnv\%vmmmnv\%vn;;    mmmmnv\%vmmmmmm;;    nv\%vnmmv\%vnmmmv^\
⊣ vmmnv\%vmmmmnv\%v;;      mmmnv\%vmmmmm;;      v\%vnmmmv\%vnmmmv^\
⊣ vmmnv\%vmmmmnv\%vnmmmmmmmmm;;      mmmmmmmmmnv\%vnmmmmv\%vnmmmv^\
⊣ vmmnv\%vmmmmnv\%vnmmmmmmmmmm;;    mmmmmmmmmmnv\%vnmmmmv\%vnmmmv^\
⊣ vmmnv\%vmmmmnv\%vnmmmmmmmmmmnv;,mmmmmmmmmmmmnv\%vn;mmmv\%vnmmmv^\
⊣ vmmnv\%vmmm. nv\%vnmmmmmmmmmnv\%;mmmmmmmmmmmnv\%vn; mmmv\%vnmmmv^\
⊣ `vmmnv\%vmm,  v\%vnmmmmmmmmmmnv\%nmmmmmmmmmmnv\%v;  mmv\%vnnmmv'^\
⊣ vmmnv\%vmm;,   \%vnmmmmmmmmmnv\%nmmmmmmmmmnv\%;'   mv\%vnmmmmv^\
⊣ vmmnv\%vmm;;,   nmmm;'             mmmm;;'    mv\%vnmmmmv'^\
⊣ `vmmnv\%vmmm;;,.         mmnvv;,            mmv\%vnmmmmv'^\
⊣  `vmmnv\%vmmmmnv\%vnmmmmmmmmnvvnmmmmmmnv\%vnmmmv\%vnmmmmv'^\
⊣  `vmvn\%vmmmmnv\%vnmmmmmmmnvvnmmmmmnv\%vnmmmm\%vnmmmv'^\
⊣  `v\%mmmmmn%:\%vnmnmmmmn\%vnmmmnv%:\%vnmv\%nmv'

⊣

⊣ ; A remarked line
⊣  ; Also a remarked line
⊣ ;     A                          \
⊣  formatted and multi-line   \
⊣  remark

1 Physically and logically.

5.1.7 Include directives

You may add #include directives in a resource file for searching and loading further resource files^{1}. An #include directive tells Gcal to suspend reading the current resource file and read one other resource file before continuing. The directive is a line in the resource file that looks like this template:

     #include whitespace argument [whitespace] newline

The argument can either be

     "file name"

or^{2}

     <file name>

One whitespace character minimum is required between #include and the argument. The argument may be trailed by extra whitespace characters and the line must always end with a ‘\n’ (newline) character, except it is the last line of a resource file. Let us inspect a example which is a bit more concrete:

1. #include "foo/bar"

2. #include <bar>

The first #include directive tells Gcal to load the file bar in the file directory foo from the actual file directory. If this fails, Gcal tries to load this file by using steps 1…4 of the previously explained mechanism used for searching files (see [File searching mechanism], page 82).

The second #include directive tells Gcal to load the file bar from the user respectively system data file directory by using steps 3…4 of the previously explained mechanism used for searching files. It is not allowed to include files which have a fixed access path starting from the root file directory by such an #include directive like ‘#include </file>’ or ‘#include </foo/bar/file>’; just as the specification of a single or an extended file name containing a disk/drive specifier^{3}, because the resulting path name^{4} would not be a valid file name any longer.

Included files may include other files again. But the nesting level of included files is limited on the one hand by the amount of usable working storage of the computer, and on the other hand by the value which is given by the operating system respectively the compiler for the number of files which can be opened simultaneously. Recursive includes^{5} or cyclic includes^{6} are not permitted, because such backward references would produce an infinite loop in the program internally! Gcal recognizes such situations and terminates immediately with a fatal error. See [Error Code 119], page 257, for more information.

If an included resource file cannot be found and no --debug=abort option is given, processing of the resource file containing the #include continues.

1 A mechanism similar the method which is used by the C Preprocessor.

2 Thus, a file name which is embraced by ‘""’ or ‘<>’.

3 Even if operating systems like MS-DOS or OS/2 permits such a disk/drive specification on principle.

4 A path name lead by the user respectively system data file directory name.

5 The included file tries to include itself again.

6 The included file is included again by another file that occurs at a deeper place in such a chain.

5.1.8 Date variables

Gcal respects global and/or local date variables which may be used either in the date part of a Gcal resource file line, or which may be part of a special text in the text part of a Gcal resource file line. This means, up to 24 user-defined date variables can be set and referenced during program execution. See [Fixed date option --date-variable=argument], page 37, for more information how global date variables can be used.

A date variable name dvar consists of a single, case-insensitive letter^{1}. In a definition, this name is trailed by the assignment operator character ‘=’, and the numerical values of month mm (or a short, three characters month name, e.g. ‘Jan’, ‘Feb’…) and day dd, this variable has to store^{2}, e.g.:

     a=1127 c=a

stores the 27th November into the local date variable a and into the local date variable c. You should always remember that date variables can only be referenced in a line if they are already defined, because Gcal processes a resource file line by line.

Only local date variables can store dynamical dates given in the mmwwwn, dvar[+|-]n[www] or *d|wn[www] format, e.g.:

     a=03su2
     b=JunMon9
     c=b-3
     d=c+4Sa
     x=*d10
     y=*d10fri
     z=*w3Su

which means, date variable a stores the date of second Sunday in March and b stores date of last Monday in June. The date variable c stores the date which is three days before the date of the last Monday in June and this date is set in the date variable d to that date which is four Saturdays after the date of c. The date variable x stores the date of the 10th absolute day of the year (≡ 10th January). y stores the date of the 10th Friday of the year, and the assignment to z is the date of the Sunday which occurs in the 3rd week of the year. Caution: These kinds of assignments to date variables are only computed correctly in simple year bounds —which means during a current year— and not in fixed dates shown after the current year has changed, e.g. in fiscal years.

Date variables which are defined in a resource file are always local and only respected in this specific file, and that file which is included by it. This means for the included file, that all local variables of the caller are visible. The included file itself may define its own local variables in its own name-space, which are not exported to caller on return. All other definitions of date variables^{3} are always global. If a resource file line contains a reference to a date variable which is not locally defined in that file, Gcal will try to use the global values held by this variable. If this fails because no global definition of this date variable is given, Gcal will not process this line. If the simple --debug respectively --debug=internal option is given, informational messages for each undefined date variable will be shown on the standard error channel (see [Global option --debug=internal], page 11). If the --debug=abort option is given, the Gcal program will be aborted with an error code when the first undefined global date variable reference occurs. See [Error Code 113], page 257, and [Global option --debug=abort], page 11, for further information.

1 Except the date variable e which is internally reserved for the Easter Sunday’s date, so it cannot be

assigned or operated therefore. And except the date variable t which is internally reserved for today’s

date, so it likewise cannot be assigned or operated.

2 Or the name of another date variable, which is already defined.

A local date variable can be deleted. This means, that this date variable is no longer valid in its local scope and be undefined therefore. So it is possible to use its potentially defined global value in succeeding lines if this date variable is referenced there. The following example of the resource file dvar-1.rc elucidates these facts:

$ cat dvar-1.rc
⊣ ; dvar-1.rc
⊣ ;
⊣ z=0202
⊣ 0@z The local assigned date to `z'
⊣ z=
⊣ 0@z The global assigned date to `z'
⊣ z=0404
⊣ 0@z The local assigned date to `z'

So Gcal creates the following output:

$ gcal %1777 -H no -x -v z=1212 -f ./dvar-1.rc -y

⊣

⊣ Sun, Feb   2nd 1777: The local assigned date to `z'
⊣ Fri, Apr   4th 1777: The local assigned date to `z'
⊣ Fri, Dec  12th 1777: The global assigned date to `z'
$ gcal %1777 -H no -x -f ./dvar-1.rc -y

⊣

⊣ Sun, Feb   2nd 1777: The local assigned date to `z'
⊣ Fri, Apr   4th 1777: The local assigned date to `z'

3 In the command line, environment variable GCAL or response file.

Only advanced users should apply the --export-date-variables option which causes that the actual incarnation of a local date variable —which was previously defined in a resource file and not in an included file— is being exported to further resource files instead of using its global value, in case that date variable is not locally defined in that further resource file. See Section 5.1.7 [Include directives], page 102, for more details.

But be aware, the use of this option could create unwanted results, because the order of processing the resource files is an important entity in managing the --export-date-variables option, so it is not recommended to use it. You, the user, must exactly know what you are doing when applying this option; you are expressively warned now!

Some basic operations can be performed on date variables. These are:

Operation  Description
dvar++  Simple increment by one day.
dvar--  Simple decrement by one day.
dvar+=[+|-]n  Addition of a constant numerical day factor [+|-]n.
dvar-=[+|-]n  Subtraction of a constant numerical day factor [+|-]n.
dvar+=nwww  Addition of n weekdays www.
dvar-=nwww  Subtraction of n weekdays www.

The scope of the operations which are done on a local date variable, is that resource or include file, where the local date variable is defined.

If operations on a global date variable are performed in a resource file or that file which is included by it, these operations are only valid in that specific file^{4}, not in further resource files processed. This means, the initial values of global date variables are always restored if the resource file changes.

Please finally note, that each date variable assignment/operation must be given separately on a single line in the resource file.

4 They are only done locally.

5.1.9 Text variables

Gcal respects global and/or local text variables which may be used anywhere in a line of a Gcal resource file. This means, up to 26 user-defined text variables can be set and referenced during program execution. See [Fixed date option --text-variable=argument], page 38, for more information how global text variables can be used.

A text variable name tvar consists of the ‘$’ prefix and a single, case-insensitive letter.

A text variable is defined as follows:

• If the name of a text variable is trailed by the assignment operator character ‘=’, all the text which appears at the right side of this assignment operator is assigned to this text variable.

• If the name of a text variable is trailed by the assignment operator character ‘?’, all the text which appears at the right side of this assignment operator is assumed as an external command and executed by Gcal. The output which is created on the standard output channel by the command run is assigned to this text variable; and that in an interpreted manner. This means, Gcal interprets all special texts and references to other text variables that are contained in this output. See Section 5.1.5 [Special Texts processed], page 100, for further information.

• If the name of a text variable is trailed by the assignment operator character ‘:’, all the text which appears at the right side of this assignment operator is assumed as an external command and executed by Gcal. The output which is created on the standard output channel by the command run is assigned to this text variable; and that in an uninterpreted manner. This means, Gcal does not interpret any special texts and references to other text variables that are contained in this output. See also Section 5.1.5 [Special Texts processed], page 100, for further details.

Indeed, Gcal executes external commands only if the --execute-command option is given at program start-up. See [Fixed date option --execute-command], page 43, for more details.

The text which appears at the right side of one of the valid assignment operator characters may reference names of already defined text variables. These references are converted to the according textual values just before the assignment is performed by Gcal. You should always remember that text variables can only be referenced in a line if they are already defined, because Gcal processes a resource file line by line. Gcal always expands text variable references recursively until all sub-references to other text variables are resolved.

Local text variables are set to an empty value and thus defined in a special mode in case no text appears at the right side of one of the valid assignment operator characters.

If a text variable is referenced, an optional format instruction may be specified between the ‘$’ prefix and the single letter (here: the variable name), which makes it possible to modify the representation of the text (here: the contents) at which the text variable points to. Let us assume a resource file by the name of tvarf-1.rc exists with the following contents:

      $ cat tvarf-1.rc
      ⊣ ; tvarf-1.rc
      ⊣ ;
      ⊣ $a=123
      ⊣ $b=$:010*a
      ⊣ $c=$b
      ⊣ $b=
      ⊣ 0 1. \$c=.$c.
      ⊣ 0 2. \$c=.$:20*c.
      ⊣ 0 3. \$b=.$>1w*b.
      ⊣ 0 4. \$a=.$>5#a.

So Gcal creates the following output:

      $ gcal -Ux --text-variable='$b=XXX YY' --resource-file=./tvarf-1.rc

⊣

      ⊣ 1. $c=.0000000123.
      ⊣ 2. $c=.      0000000123       .
      ⊣ 3. $b=.Xxx Yy.
      ⊣ 4. $a=.  123.

See Appendix J [Format Instruction], page 249, for the detailed description of the format instruction and its components.

You may depreciate the special meaning of the ‘$’ prefix character —in case this character itself is needed in the text— by placing a ‘\’ (backslash) character before it, e.g. ‘\$’. If you need the ‘\$’ characters themselves in the text, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\$’.

External commands are not directly executed by Gcal, they are executed by means of the command line interpreter of the operating system, the shell. As consequence to this, of course the commands have to be specified according to the syntax conventions of the used shell, concerning the grouping conventions, list processing conventions, redirection conventions et cetera! For example, it is possible that you have to write ‘$a?(cat text)’ instead of the simple ‘$a?cat text’. See the pertinent literature for more details.

If the output of external commands, which appears on the standard output channel, is assigned to text variables in an interpreted manner, this output is internally directed by Gcal into a temporary file. Thereafter, the contents of the temporary file is transfered to the text variable and the temporary file is removed.

If the output of external commands, which appears on the standard output channel, is assigned to text variables in an uninterpreted manner, this output is internally processed by the Txt2gcal program first (see Section B.2 [Invoking txt2gcal], page 127), and then this output is directed by Gcal into a temporary file. Thereafter, the contents of the temporary file is transfered to the text variable and the temporary file is removed. In case a TXT2GCALPROG environment variable (see [Environment Variable TXT2GCALPROG], page 256) is defined and set with the file name of the executable Txt2gcal program, Gcal will use this file name for calling Txt2gcal. Otherwise, the file name txt2gcal —which is burned-in during the compilation step of the Gcal program— is used for calling the Txt2gcal program. The TXT2GCALPROG environment variable must always be set if the Txt2gcal program is installed under another name than the standard name txt2gcal, otherwise Gcal is unable to execute the Txt2gcal program automatically!

If the simple --debug respectively --debug=internal option is given at program startup, informational messages about the executed command and its exit code will be shown on the standard error channel (see [Global option --debug=internal], page 11).

If the --debug=abort option is given, the Gcal program will be aborted with an error code in case an exit code not equal zero occurred during the execution of the command. See [Error Code 2], page 257, and [Global option --debug=abort], page 11, for further information.

Here is an example which explains how the output of external commands can be assigned to text variables. Supposing there is a resource file by the name of tvarc-1.rc with the following contents:

$ cat tvarc-1.rc
⊣ %t  $a $>+06*b $>1u*c

This file is used in the resource file tvarc-1a.rc and processed by Gcal as follows:

$ cat tvarc-1a.rc
⊣ ; tvarc-1a.rc
⊣ ;
⊣ $a=bonjour
⊣ $b=123
⊣ $c=bonsoir
⊣ $x?cat $f
⊣ 0 \$x=---$:30*x---
⊣ $y:cat $f
⊣ 0 \$y=---$y---
⊣ 0 \$y=:::$>1u*y:::
$ gcal -f tvarc-1a.rc -QUx --exe -r'$f=./tvarc-1.rc'
⊣ $x=---    18:54 bonjour +00123 BONSOIR     ---
⊣ $y=---%t  $a $>+06*b $>1u*c---
⊣ $y=:::%T  $A $>+06*B $>1U*C:::

Here are some more examples showing how Gcal processes text variables. Let us assume a resource file by the name of tvar-1.rc exists with the following contents:

$ cat tvar-1.rc
⊣ ; tvar-1.rc
⊣ ;
⊣ $a=foo
⊣ 0 \$a:1=$a
⊣ $c=$a
⊣ 0 \$c=$c
⊣ $b=$c bar $a
⊣ 0 \$b=$b
⊣ $a=bar $b baz $a booz
⊣ 0 \$a:2=$a
⊣ $a=
⊣ $b=0 $b $c frozz $a frozz
⊣ $b ->That's \$b
⊣ 0 \$x='$x' and \$d is undefined: $d...
⊣ $a= 0 ~ 1~2~\$3~%n~$c~\
⊣ now it's enough!
⊣ 0 \$a=---$a---
⊣ $a=0 \   \ \\ And this...
⊣ $a works too!

So Gcal creates the following output:

$ gcal %19960101 -H no -x -r '$A=FROBOZZ:$x=' -f ./tvar-1.rc

⊣

⊣ Mon, Jan   1st 1996:    \ \\ And this... works too!
⊣ Mon, Jan   1st 1996: $a:1=foo
⊣ Mon, Jan   1st 1996: $a:2=bar foo bar foo baz foo booz
⊣ Mon, Jan   1st 1996: $a=--- 0
⊣  1
⊣  2
⊣  $3
⊣  01-Jan-1996
⊣  foo
⊣  now it's enough!---
⊣ Mon, Jan   1st 1996: $b=foo bar foo
⊣ Mon, Jan   1st 1996: $c=foo
⊣ Mon, Jan   1st 1996: $x='' and $d is undefined: $d...
⊣ Mon, Jan   1st 1996: foo bar foo foo frozz FROBOZZ frozz ->That's $b

Or a resource file by the name of tvar-2.rc exists with the following contents:

      $ cat tvar-2.rc
      ⊣ ; tvar-2.rc
      ⊣ ;
      ⊣ $a=$b foo
      ⊣ $b=0@e
      ⊣ $a bar
      ⊣ $b \$b
      ⊣ 0 \$a:$a
      ⊣ 0 \$b:$b

So Gcal creates the following output:

      $ gcal %19960101 -H no -x -f ./tvar-2.rc

⊣

      ⊣ Mon, Jan   1st 1996: $a:0@e foo
      ⊣ Mon, Jan   1st 1996: $b:0@e
      $ gcal %19960101 -H no -x -f ./tvar-2.rc -y

⊣

      ⊣ Sun, Apr   7th 1996: $b
      ⊣ Sun, Apr   7th 1996: foo bar

As seen before, it is allowed to store complete (or partial) date parts (see Section 5.1.2 [Date part of a line], page 86), likewise special texts (see Section 5.1.5 [%?… Special Texts], page 100) into text variables or references to other text variables, which are processed by Gcal after their expansion. See Appendix G [Description of all %?… Special Texts], page 145, for limitations concerning the assignment of special texts to text variables.

Text variables which are defined in a resource file are always local and only respected in this specific file, and that file which is included by it. This means for the included file, that all local variables of the caller are visible. The included file itself may define its own local variables in its own name-space, which are not exported to caller on return. All other definitions of text variables^{1} are always global. If a resource file line contains a reference to a text variable which is not locally defined in that file, Gcal will try to use the global values held by this variable. If this fails because no global definition of this text variable is given, Gcal keeps the name of this text variable untouched in this line, except this text variable held an empty value. In such a case, Gcal completely ignores this text variable while processing and displaying instead of keeping its name untouched in this line.

An example to this. Supposing a resource file by the name of tvar-3.rc exists with the following contents:

1 In the command line, environment variable GCAL or response file.

      $ cat tvar-3.rc
      ⊣ ; tvar-3.rc
      ⊣ ;
      ⊣ $c=+00+000
      ⊣ $h=
      ⊣ $d=$c$h
      ⊣ 0 %o$d  %s$d
      ⊣ 0 \$c=$c  \$h=$h  \$d=$d  \$x=$x

So Gcal creates the following output:

      $ gcal -QUx -f ./tvar-3.rc
      ⊣ $c=+00+000  $h=  $d=+00+000  $x=$x
      ⊣ 05:57 18:04
      $ gcal -QUx -r '$h=+1000:$x=' -f ./tvar-3.rc
      ⊣ $c=+00+000  $h=+1000  $d=+00+000+1000  $x=
      ⊣ 05:52 18:09

Only advanced users should apply the --export-text-variables option which causes that the actual incarnation of a local text variable —which was previously defined in a resource file and not in an included file— is being exported to further resource files instead of using its global value, in case that text variable is not locally defined in that further resource file. See Section 5.1.7 [Include directives], page 102, for more details.

But be aware, the use of this option could create unwanted results, because the order of processing the resource files is an important entity in managing the --export-text-variables option, so it is not recommended to use it. You, the user, must exactly know what you are doing when applying this option; you are expressively warned now!

Some basic operations can be performed on text variables in case they contain integer values. These are:

Operation  Description
tvar++  Simple increment by one.
tvar--  Simple decrement by one.
tvar+=[+|-]n  Addition of a constant numerical factor [+|-]n.
tvar-=[+|-]n  Subtraction of a constant numerical factor [+|-]n.

Here is an example showing how Gcal processes text variable operations. Let us assume a resource file by the name of tvaro-1.rc exists with the following contents:

$ cat tvaro-1.rc
⊣ ; tvaro-1.rc
⊣ ;
⊣ $a=130
⊣ $b=2
⊣ 0 1. \$b=$b
⊣ $b++
⊣ 0 2. \$b=$b
⊣ $b+=7
⊣ 0 3. \$b=$b
⊣ $b--
⊣ 0 4. \$b=$b
⊣ $b-=-5
⊣ 0 5. \$b=$b
⊣ $b+=123
⊣ 0 6. \$b=$b
⊣ $b-=$a
⊣ 0 7. \$b=$b
⊣ $b-=10000
⊣ 0 8. \$b=$b
⊣ $b+=10000
⊣ 0 9. \$b=$b
⊣ $b=02
⊣ 0 a. \$b=$b
⊣ $b++
⊣ 0 b. \$b=$b

So Gcal creates the following output:

$ gcal -QUx -f ./tvaro-1.rc
⊣ 1. $b=2
⊣ 2. $b=3
⊣ 3. $b=10
⊣ 4. $b=09
⊣ 5. $b=14
⊣ 6. $b=137
⊣ 7. $b=007
⊣ 8. $b=-9993
⊣ 9. $b=00007
⊣ a. $b=02
⊣ b. $b=03

As you can see in the former example, if operations are made like these, the default behavior of Gcal is to fill with leading zeroes to preserve a former length of a text variable, in case a carry into one of the next decimal places has happened and is taken back by an operation at a later place in the resource file. But this default behavior can be disabled respectively changed by using the already above mentioned format instruction.

The scope of the definitions which are done on a local text variable, is that resource or include file, where the local text variable is defined.

If a global text variable is redefined in a resource file or that file which is included by it, these redefinitions are only valid in that specific file^{2}, not in further resource files processed. This means, the initial values of global text variables are always restored if the resource file changes.

Please finally note, that each text variable assignment/operation must be given separately on a single line in the resource file.

2 They are only done locally.

5.2 Resource file examples

Here are some examples of problem oriented resource files which might be useful to get a deeper impression about the attributes of Gcal’s fixed date feature.

Let us begin with a common example, the resource file common.rc:

$ cat common.rc
⊣ ; common.rc, a common example of a resource file.
⊣ ;
⊣ 19940715 Local time: %t  --> Dentist at 10:00
⊣ 00000921 My %B1962  birthday%i1952#2062
⊣ 0000093 Gregorian~Reformation\~is %y1752  years ago
⊣ 0000Sep03 Gregorian~Reformation\~is %y1752 years ago
⊣ 0 Every day in every month in every year
⊣ 0000 Every day in every month in every year
⊣ 00000000 Every day in every month in every year
⊣ 199400fri        Every Friday in 1994
⊣ 000007mo3 Every third Monday~in July every year less 1980%i#1979
⊣ 0000@e-10 Easter Sunday-10 days every year
⊣ 0@e-10 Easter Sunday-10 days every year
⊣ 0000@e+10 Easter Sunday+10 days every year
⊣ a=1127
⊣ 0@a+20 20 days after date of date variable `a'
⊣ a=*W10FR
⊣ 0@a-1 The Thursday of the 10th week every year
⊣ 0@a   The Friday of the 10th week every year
⊣ 0@a1  The Saturday of the 10th week every year
⊣ 1995*d10    The 10th absolute day in 1995
⊣ 1995*d99tue The last Tuesday in 1995
⊣ 0*w99su The Sunday of the last week every year
⊣ ;
⊣ ; Next entry is respected by ALL fixed date period modifiers.
⊣ ;
⊣ 1994*d1sun#99SUN.7 Each Sunday in 1994
⊣ ;
⊣ ; Day of Prayer & Repentance is Wednesday before the Sunday,
⊣ ;   which is before the 1st Advent's Sunday.
⊣ ;
⊣ a=1225
⊣ a=a-5sun
⊣ 0@a-4 Day of Prayer & Repentance
⊣ ; or
⊣ 0@a-1wed Day of Prayer & Repentance

And here is a resource file by the name of demo-1.rc, used to detect all Fridays the 13th of a year:

$ cat demo-1.rc
⊣ ; demo-1.rc
⊣ ;
⊣ $a=Friday the 13th%i0000
⊣ $b=#0000
⊣ $c=fri2
⊣ $d=fri3
⊣ 00000113 $a01$c$b01$c
⊣ 00000113 $a01$d$b01$d
⊣ 00000213 $a02$c$b02$c
⊣ 00000213 $a02$d$b02$d
⊣ 00000313 $a03$c$b03$c
⊣ 00000313 $a03$d$b03$d
⊣ 00000413 $a04$c$b04$c
⊣ 00000413 $a04$d$b04$d
⊣ 00000513 $a05$c$b05$c
⊣ 00000513 $a05$d$b05$d
⊣ 00000613 $a06$c$b06$c
⊣ 00000613 $a06$d$b06$d
⊣ 00000713 $a07$c$b07$c
⊣ 00000713 $a07$d$b07$d
⊣ 00000813 $a08$c$b08$c
⊣ 00000813 $a08$d$b08$d
⊣ 00000913 $a09$c$b09$c
⊣ 00000913 $a09$d$b09$d
⊣ 00001013 $a10$c$b10$c
⊣ 00001013 $a10$d$b10$d
⊣ 00001113 $a11$c$b11$c
⊣ 00001113 $a11$d$b11$d
⊣ 00001213 $a12$c$b12$c
⊣ 00001213 $a12$d$b12$d

The result for the years 1996…2002 is as follows:

$ gcal -H no -u -x -f ./demo-1.rc 1996+2002

⊣

⊣ Fri, Sep  13th 1996: Friday the 13th
⊣ Fri, Dec  13th 1996: Friday the 13th

⊣

⊣ Fri, Jun  13th 1997: Friday the 13th

⊣

⊣ Fri, Feb  13th 1998: Friday the 13th
⊣ Fri, Mar  13th 1998: Friday the 13th
⊣ Fri, Nov  13th 1998: Friday the 13th

⊣

⊣ Fri, Aug  13th 1999: Friday the 13th

⊣

⊣ Fri, Oct  13th 2000: Friday the 13th

⊣

⊣ Fri, Apr  13th 2001: Friday the 13th
⊣ Fri, Jul  13th 2001: Friday the 13th

⊣

⊣ Fri, Sep  13th 2002: Friday the 13th
⊣ Fri, Dec  13th 2002: Friday the 13th

Or a bit shorter version of the resource file demo-1.rc by the name of demo-2.rc, likewise used to detect all Fridays the 13th of a year, and that by using the fact, that a Friday the 13th only occurs if the first day of a month is a Sunday:

$ cat demo-2.rc
⊣ ; demo-2.rc
⊣ ;
⊣ $a=%n+12 is a Friday%i0000
⊣ $b=#0000
⊣ $c=sun1
⊣ 00000101 $a01$c$b01$c
⊣ 00000201 $a02$c$b02$c
⊣ 00000301 $a03$c$b03$c
⊣ 00000401 $a04$c$b04$c
⊣ 00000501 $a05$c$b05$c
⊣ 00000601 $a06$c$b06$c
⊣ 00000701 $a07$c$b07$c
⊣ 00000801 $a08$c$b08$c
⊣ 00000901 $a09$c$b09$c
⊣ 00001001 $a10$c$b10$c
⊣ 00001101 $a11$c$b11$c
⊣ 00001201 $a12$c$b12$c

The result for the years 1996…2002 is as follows:

$ gcal -H no -u -x -f ./demo-2.rc 1996+2002

⊣

⊣ Sun, Sep   1st 1996: 13-Sep-1996 is a Friday
⊣ Sun, Dec   1st 1996: 13-Dec-1996 is a Friday

⊣

⊣ Sun, Jun   1st 1997: 13-Jun-1997 is a Friday

⊣

⊣ Sun, Feb   1st 1998: 13-Feb-1998 is a Friday
⊣ Sun, Mar   1st 1998: 13-Mar-1998 is a Friday
⊣ Sun, Nov   1st 1998: 13-Nov-1998 is a Friday

⊣

⊣ Sun, Aug   1st 1999: 13-Aug-1999 is a Friday

⊣

⊣ Sun, Oct   1st 2000: 13-Oct-2000 is a Friday

⊣

⊣ Sun, Apr   1st 2001: 13-Apr-2001 is a Friday
⊣ Sun, Jul   1st 2001: 13-Jul-2001 is a Friday

⊣

⊣ Sun, Sep   1st 2002: 13-Sep-2002 is a Friday
⊣ Sun, Dec   1st 2002: 13-Dec-2002 is a Friday

Of course, preceding textual expense for the purpose of detecting all Fridays the 13th of a year can be dramatically reduced by simply using of

0*d1fri#99fri.7 Friday the 13th%i00000013#00000013

in the resource file demo-1.rc, respectively,

0*d1sun#99sun.7 %n+12 is a Friday%i00000001#00000001

in the demo-2.rc resource file.

On the one hand, preceding one-liners reduce the coding expense quite considerably —from 28 respectively 15 lines to 1 line in each case—, but on the other hand, this partly increases the checking expense of these one-liners considerably and result in a longer program runtime therefore. Both one-liners produce about 52 internal productions respectively effective fixed date entries of themselves for each year checked by Gcal, at which the demo-1.rc and demo-2.rc resource files cause 24 respectively 12 effective fixed date entries only.

And here is a resource file by the name of demo-3.rc, used to detect all working-days of any month:

$ cat demo-3.rc
⊣ ; demo-3.rc
⊣ ;
⊣ $a=%VWork day
⊣ $c=%i#0000$b99
⊣ a=$bmon1
⊣ 0@a#+4 $a
⊣ a=$btue1
⊣ 0@a#+3 $a
⊣ a=$bwed1
⊣ 0@a#+2 $a
⊣ a=$bthu1
⊣ 0@a#+1 $a
⊣ a=$bfri1
⊣ 0@a $a
⊣ a=$bmon2
⊣ 0@a#+4 $a
⊣ a=$bmon3
⊣ 0@a#+4 $a
⊣ a=$bmon4
⊣ 0@a#+4 $a$c
⊣ a=$bmon9
⊣ 0@a#+4 $a$c

Respectively a bit shorter version of the resource file demo-3.rc:

; demo-3.rc
;
$a=%VWork day
$c=%i#0000$b99
0000$bmon1:5 $a
0000$btue1:4 $a
0000$bwed1:3 $a
0000$bthu1:2 $a
0000$bfri1:1 $a
0000$bmon2:5 $a
0000$bmon3:5 $a
0000$bmon4:5 $a$c
0000$bmon9:5 $a$c

And a resource file by the name of include.rc exists which is only used to include one given resource file:

$ cat include.rc
⊣ #include "$f"

Then the result for the month August 1996 is as follows:

$ gcal -H no -u -x -r '$b=08:$f=demo-3.rc' -f ./include.rc Aug 1996

⊣

⊣ Thu, Aug   1st 1996: Work day
⊣ Fri, Aug   2nd 1996: Work day
⊣ Mon, Aug   5th 1996: Work day
⊣ Tue, Aug   6th 1996: Work day
⊣ Wed, Aug   7th 1996: Work day
⊣ Thu, Aug   8th 1996: Work day
⊣ Fri, Aug   9th 1996: Work day
⊣ Mon, Aug  12th 1996: Work day
⊣ Tue, Aug  13th 1996: Work day
⊣ Wed, Aug  14th 1996: Work day
⊣ Thu, Aug  15th 1996: Work day
⊣ Fri, Aug  16th 1996: Work day
⊣ Mon, Aug  19th 1996: Work day
⊣ Tue, Aug  20th 1996: Work day
⊣ Wed, Aug  21st 1996: Work day
⊣ Thu, Aug  22nd 1996: Work day
⊣ Fri, Aug  23rd 1996: Work day
⊣ Mon, Aug  26th 1996: Work day
⊣ Tue, Aug  27th 1996: Work day
⊣ Wed, Aug  28th 1996: Work day
⊣ Thu, Aug  29th 1996: Work day
⊣ Fri, Aug  30th 1996: Work day

And of course, the preceding textual expense for the purpose of detecting all work days of any month can be dramatically reduced by simply using of

00000001#0099 %V%rWork day

in the resource file demo-3.rc.

And next the resource file by the name of sun-1.rc which detects the actual local time, the approximate time of sunrise and sunset, day and night length of two geographic locations:

120 Gcal 4

$ cat sun-1.rc
⊣ ; sun-1.rc, comparison of sunrise / sunset, day length and night
⊣ ; length of two locations, namely:  Muenster.DE / Bangkok.TH
⊣ ; All references are based on Muenster.DE !
⊣ ;
⊣ ; Common definitions
⊣ $p=0*d1#999
⊣ $v=%2%4 for that places local time
⊣ $z=actual local time in
⊣ ; Definitions for winter time
⊣ $a=%e#1980 %e0@a#0@b-1
⊣ $x=Sun:~%o$c,$w rise, %s$c,$w set, %u$c,$w day, %z$c,$w night~$l
⊣ ; Winter time in Germany == CEWT/GMT-1 (+60 minutes)
⊣ $w=+60
⊣ ; Starting date of winter time in Germany
⊣ b=10sun9
⊣ ; Definitions for summer time (daylight-saving time)
⊣ $b=%e#1980 %i0@a#0@b-1
⊣ $y=Sun:~%o$c,$s rise, %s$c,$s set, %u$c,$s day, %z$c,$s night~$l
⊣ ; Summer time in Germany == CEST/GMT-2 (+120 minutes)
⊣ $s=+120
⊣ ; Starting date of summer time in Germany
⊣ a=03sun9
⊣ ; Geographic co-ordinate and name of the location Muenster, Germany
⊣ $c=+5158+00738+61
⊣ $l=%1Muenster
⊣ ; Muenster's local time sunrise etc. for that places local time
⊣ $p $a $x$v
⊣ $p $b $y$v
⊣ ; Muenster's actual local time
⊣ 0 $a %1%@$w %2 $z $l
⊣ 0 $b %1%@$s %2 $z $l
⊣ ; Geographic co-ordinate and name of the location Bangkok, Thailand
⊣ $c=+1345+10031+10
⊣ $l=%3Bangkok%4-time converted to $l's%2 local time
⊣ ; Bangkok's sunrise etc. based on Muenster's local time
⊣ $p $a $x
⊣ $p $b $y
⊣ ; Bangkok's local time == WAST/GMT-7 (+420 minutes)
⊣ $t=+420
⊣ $w=$t
⊣ ; Bangkok's actual local time
⊣ $l=%3Bangkok
⊣ 0 $a %3%@$t %4 $z $l
⊣ 0 $b %3%@$t %4 $z $l
⊣ ; Bangkok's sunrise etc. for that places local time
⊣ $l=$l$v
⊣ $p $x

The result for the 12th September 1999 is as follows:

$ gcal -fsun-1.rc -Ux -Hno

⊣

⊣ 18:53/12-Sep-1999 actual local time in Muenster
⊣ 23:53/12-Sep-1999 actual local time in Bangkok
⊣ Sun:
⊣ 01:06rise, 13:22set, 12h16'day, 11h44'night
⊣ Bangkok-time converted to Muenster's local time
⊣ Sun:
⊣ 06:06rise, 18:22set, 12h16'day, 11h44'night
⊣ Bangkok for that places local time
⊣ Sun:
⊣ 06:56rise, 19:54set, 12h58'day, 11h02'night
⊣ Muenster for that places local time

The next example is the resource file redir-1.rc which demonstrates the use of the shell command %![argument] special text:

$ cat redir-1.rc
⊣ ; redir.rc, run shell command $c and redirect it to $f
⊣ ;
⊣ $p=0*d1#999
⊣ $c=$p %!$c
⊣ $g=|txt2gcal - $p
⊣ ;
⊣ $c $g>$f

The result of succeeding call of Gcal, to which you can also add the --debug=all option for a better understanding of the execution methods, is as follows:

$ gcal --exe -QUx -f./redir-1.rc+XX -r'$f=XX:$c=echo $f' -#'0 %!rm $f'
⊣ XX

And finally a resource file by the name of swtim-1.rc which produces fixed date messages reminding Daylight Saving that are valid since 1981 for the Federal Republic of Germany^{1}.

$ cat swtim-1.rc
⊣ ; swtim-1.rc, daylight-saving time BRD
⊣ ;
⊣ $a=daylight-saving time starts Sunday
⊣ ;
⊣ a=03sun9
⊣ 0@a-2#+2 Summer $a~+1 hrs. (02:00 --> 03:00 h.)%i1981
⊣ ;
⊣ a=09sun9
⊣ 0@a-2#+2 Winter $a~-1 hrs. (03:00 --> 02:00 h.)%i1981#1995
⊣ ;
⊣ a=10sun9
⊣ 0@a-2#+2 Winter $a~-1 hrs. (03:00 --> 02:00 h.)%i1996

The result for the year 1998 is as follows:

$ gcal %1998 -Hno -xy -f ./swtim-1.rc

⊣

⊣ Fri, Mar  27th 1998: Summer daylight-saving time starts Sunday
⊣  +1 hrs. (02:00 --> 03:00 h.)
⊣ Sat, Mar  28th 1998: Summer daylight-saving time starts Sunday
⊣  +1 hrs. (02:00 --> 03:00 h.)
⊣ Sun, Mar  29th 1998: Summer daylight-saving time starts Sunday
⊣  +1 hrs. (02:00 --> 03:00 h.)
⊣ Mon, Mar  30th 1998: Summer daylight-saving time starts Sunday
⊣  +1 hrs. (02:00 --> 03:00 h.)
⊣ Tue, Mar  31st 1998: Summer daylight-saving time starts Sunday
⊣  +1 hrs. (02:00 --> 03:00 h.)
⊣ Fri, Oct  23rd 1998: Winter daylight-saving time starts Sunday
⊣  -1 hrs. (03:00 --> 02:00 h.)
⊣ Sat, Oct  24th 1998: Winter daylight-saving time starts Sunday
⊣  -1 hrs. (03:00 --> 02:00 h.)
⊣ Sun, Oct  25th 1998: Winter daylight-saving time starts Sunday
⊣  -1 hrs. (03:00 --> 02:00 h.)
⊣ Mon, Oct  26th 1998: Winter daylight-saving time starts Sunday
⊣  -1 hrs. (03:00 --> 02:00 h.)
⊣ Tue, Oct  27th 1998: Winter daylight-saving time starts Sunday
⊣  -1 hrs. (03:00 --> 02:00 h.)

1 Two days before adaptation till two days after adaptation.

Appendix A Genesis of the Gregorian Calendar

The nowadays calendar was first formulated in several inaccurate variations by the Romans based on methods developed by the Babylonians and Egyptians. The aim of all these calendars was to harmonize the cycles of the Moon and the Sun. During Julius Caesar’s reign, January was falling in autumn so he ordered Sosigenes to make changes to the calendar. He added 90 days to the year 46 BC to make up for the seasonal drift and adjusted the lengths of the months similarly as we know them to be today. He introduced the leap year by adding one day to February every four years. For the present, the leap year regulation was made in the way that all four years after the 23rd day in February a leap day was laid in, so the 24th February occurred twice. The use of the leap year was an improvement but not entirely accurate.

But in the later years, the leap rule was used in the wrong way so that the errors are corrected by emperor Augustus in the year 8 BC. A curious sequel happened on this occasion. Because Augustus reacted with great jealousy to all things previously made and propagated by Julius Caesar, he did not like Caesar’s name in the calendar, namely the today’s month of July. Offhandedly he ordered to name another month to himself and so the month name August arose. Furthermore, Augustus did not tolerate the fact that his month of birth (the August) was shorter than Caesar’s month in the sense of the periodical sequence of months with 30 and 31 days. Consequently, the month of August got 31 days, too. Due to this modification, the number of days in February were reduced to 28 and 29 days, respectively, so the 29th February was designed to be the leap day now.

This calendar is well known under the term Julian calendar and is based on a plain solar year. The nominal length of a solar year (respectively a so-called tropical year) is 365 days, 5 hours, 48 minutes, and 46 seconds. One 366-day year every four years equates to an average Julian year of 365 days and 6 hours, consequently to 365.25 days. This means, every four years, an error of 44 minutes, 56 seconds was accumulated by this kind of calendar calculation. Because of this counting method, the length of the years becomes a bit too long, by more than 11 minutes.

By the 16th century, the vernal equinox occurred around March 11, rather than March 21, by reason of an accumulated error of ten days. The feast of Easter shifted away more and more from the habitual vernal date, which must have always been celebrated on an earlier date. So Pope Gregory XIII introduced the new style calendar in 1582. Thursday, the 4th October 1582, was followed by Friday, the 15th October, by suppressing the ten days between both dates. Moreover, he ordained that years ending in hundreds should not be leap years unless they are divisible by 400. Incidentally, the Gregorian reform compensates by 72 hours (3 days) every 400 years. The actual excess accumulated is 74 hours, 53 minutes and 20 seconds. The error of 2 hours, 53 minutes and 20 seconds every 400 years accumulates to one day in about 3300 years. Thereby, the Gregorian year has an average length of 365.2425 days.

But this Gregorian calendar was accepted very slowly by others. Catholic countries adopted the Gregorian calendar almost immediately. Most Protestant countries on the Continent adopted the new calendar around 1700. England and the American colonies changed in 1752, by act of Parliament. Orthodox Christian countries adopted the Gregorian calendar later. Russia was the last European country to do so, after the communist revolution of 1917. As a result, the former U.S.S.R. celebrated the October Revolution (happened on October 25th, 1917) in the old style calendar on November 7th.

The era of a world wide uniform calendar is already part of history today. The Iran returned to the traditional Mohammedan lunar calendar in 1979 after removal of the Shah reign. There are some efforts to improve our currently valid Gregorian calendar. Its disadvantages are the reason why an appointed day is not always on the same week day. Besides, the month lengths are not equal and the holidays, which have relations to the feast of Easter, are moved within the calendar from one year to another. A very sophisticated suggestion was proposed by the United Nations, but the international establishment of this suggestions has failed, since it was resisted by some countries as well as the churches.

Appendix B Gcal Utilities

Three small utility programs are part of the Gcal 4 package. The Tcal program runs the Gcal program with the date set one day ahead. The Txt2gcal program creates a verbatim Gcal resource file from a text file, and in contrast to this, the Gcal2txt program creates a verbatim text file from a Gcal resource file.

B.1 Invoking tcal

The Tcal program runs the Gcal program with the date of tomorrow’s day^{1}. Its arguments are as follows:

     tcal [--help | --version] | [--shift=[+|-]number] [argument. . .]

All given arguments are passed unmodified to the Gcal program. If the Gcal program shall be called with another date than tomorrow’s date, this desired date can be selected by using the --shift=[+|-]number option, in which [+|-]number is the distance of days the desired date is distant from the actual system date (≡today). (Works only for Gregorian years.) The --shift option must be given before all other arguments that are passed to the Gcal program.

In case a GCALPROG environment variable (see [Environment Variable GCALPROG], page 255) is defined and set with the file name of the executable Gcal program, Tcal will use this file name for calling Gcal. Otherwise, the file name gcal —which is burned-in during the compilation step of the Tcal program— is used for calling the Gcal program. The GCALPROG environment variable must always be set if the Gcal program is installed under another name than the standard name gcal, otherwise Tcal is unable to execute the Gcal program automatically! An exit status of 0 means all processing is successfully done, any other value means an error has occurred. The program accepts the following options:

--help Print a usage message listing all available options, then exit successfully.

--version

Print the version number, then exit successfully.

--shift=[+|-]number

Define the displacement in ‘[+|-]number’ days the desired date is distant from the actual date.

B.2 Invoking txt2gcal

The Txt2gcal program creates a verbatim Gcal resource file from a text file. Its arguments are as follows:

     txt2gcal [--help | --version] | [text-file|-] [date-part]

If no text-file argument is given or/but a single ‘-’ character, the program reads and processes all input received from the standard input channel. If no date-part argument is given, Txt2gcal creates a ‘0’ for the date part. All results are always shown on the standard output channel. An exit status of 0 means all processing is successfully done, any other value means an error has occurred.

1 This is equivalent the --shift=1 option.

The program accepts the following options:

--help Print a usage message listing all available options, then exit successfully.

--version

Print the version number, then exit successfully.

Here comes an example how to use Txt2gcal. Let us suppose there is a text file by the name of tdemo-1.txt with the following contents:

$ cat tdemo-1.txt
⊣  Hi friends,

⊣

⊣ I'm the demo text containing funny characters and character
⊣ sequences like ~~ \~ % %% $a $1 %%%\
⊣ %\ %s %%foo %bar \%Baz \\~ \~\ and so on...
⊣ I'm be anxious how I'll be transformed by `txt2gcal'.

⊣

and Txt2gcal processes this file, redirecting the output to tdemo-1.rc:

$ txt2gcal tdemo-1.txt 0*d1#999 > tdemo-1.rc
$ cat tdemo-1.rc
⊣ 0*d1#999 \        Hi friends,~\
⊣ ~\
⊣ I'm the demo text containing funny characters and character~\
⊣ sequences like \~\~ \\~ % %% \$a $1 %%%\ ~\
⊣ %\ \%s %\%foo \%bar \\%Baz \\\~ \\~\ and so on...~\
⊣ I'm be anxious how I'll be transformed by `txt2gcal'.~\

then Txt2gcal has now created a new Gcal resource file tdemo-1.rc from the text file tdemo-1.txt. Let us finally see, how Gcal will interpret this new resource file:

$ gcal %19960101 -H no -Ax -f ./tdemo-1.rc

⊣

⊣ Mon, Jan   1st 1996:
⊣  Hi friends,

⊣

⊣ I'm the demo text containing funny characters and character
⊣ sequences like ~~ \~ % %% $a $1 %%%\
⊣ %\ %s %%foo %bar \%Baz \\~ \~\ and so on...
⊣ I'm be anxious how I'll be transformed by `txt2gcal'.

⊣

B.3 Invoking gcal2txt

The Gcal2txt program creates a verbatim text file from a Gcal resource file. Its arguments are as follows:

     gcal2txt [--help | --version] | [resource-file|-]

If no resource-file argument is given or/but a single ‘-’ character, the program reads and processes all input received from the standard input channel. All results are always shown on the standard output channel. An exit status of 0 means all processing is successfully done, any other value means an error has occurred.

The sense and purpose of Gcal2txt is to retrieve all additional texts, which are put into the output by means of the Txt2gcal program. The program accepts the following options:

--help Print a usage message listing all available options, then exit successfully.

--version

Print the version number, then exit successfully.

Here comes an example how to use Gcal2txt. For that purpose, let us use the resource file tdemo-1.rc of the previous section (see Section B.2 [Invoking txt2gcal], page 127), which was created by means of the Txt2gcal program from a text file and which looks like this:

      $ txt2gcal tdemo-1.txt 0*d1#999 > tdemo-1.rc
      $ cat tdemo-1.rc
      ⊣ 0*d1#999 \        Hi friends,~\
      ⊣ ~\
      ⊣ I'm the demo text containing funny characters and character~\
      ⊣ sequences like \~\~ \\~ % %% \$a $1 %%%\ ~\
      ⊣ %\ \%s %\%foo \%bar \\%Baz \\\~ \\~\ and so on...~\
      ⊣ I'm be anxious how I'll be transformed by `txt2gcal'.~\

Now let us see, how this resource file will be processed by Gcal2txt:

      $ gcal2txt tdemo-1.rc
      ⊣  Hi friends,

⊣

      ⊣ I'm the demo text containing funny characters and character
      ⊣ sequences like ~~ \~ % %% $a $1 %%%\
      ⊣ %\ %s %%foo %bar \%Baz \\~ \~\ and so on...
      ⊣ I'm be anxious how I'll be transformed by `txt2gcal'.

Appendix C Aspects in Internationalization

Starting with version 2.00, Gcal is able to display message texts using any native languages instead of using the English language only, because parts of the GNU gettext package are integrated into the Software. See Section “Introduction” in GNU gettext tools, for more details.

By default, Gcal displays all message texts using the English native language in case no other native language is wanted. A so-called message catalog is read by Gcal at run-time if message texts from another native language are required. Gcal 4 supports the following native languages:

Native Language  Language Code
English  en
German  de
French  fr
Dutch  nl
Polish  pl
Russian  ru
Swedish  sv

It is only necessary to set one of the environment variables^{1}:

1. LANGUAGE
2. LC_ALL
3. LC_MESSAGES
4. LANG

with a language code to select another native language instead of the English native language.

Normally, users only have to set the LANG environment variable to tell Gcal the native language to use at run-time level. Presuming users want to run Gcal using the German native language for displaying message texts, they merely have to execute ‘setenv LANG de’ (in csh) or ‘export LANG; LANG=de’ (in sh) at the shell prompt. Of course they could even do this from their .login or .profile file. See Section “The User’s View” in GNU gettext tools, for more details.

As shown above, a simple setting of de in the environment variable LANG is sufficient to enable German message texts. de is the two-letter language code for the German language defined in the ISO-639:1988, and is called simple language code information in the further context. Other language codes can be taken from this ISO-document^{2}.

1 Listed in decreasing priority as they are respected.

2 For example fr for French, es for Spanish…

Because Gcal as calendar program must also comply the specifics of a used native language concerning the ordering of day, month and year (and further things) of a displayed date, the period of Gregorian Reformation, the type of week number and the representation of calendar sheets, these criteria are likewise bound to the language code^{3}.

A en language code causes the following internal defaults of above criteria:

• U.S. American date format

• Gregorian Reformation has occurred 1752

• Week starts on Sunday

• Standard week numbers

• Representation of calendar sheet

   September 1994
Su Mo Tu We Th Fr Sa
             1  2  3
 4  5  6  7  8  9 10
11 12 13 14 15 16 17
18 19 20 21 22 23 24
25 26 27 28 29 30

And a de language code^{4} causes the following internal defaults:

• German date format

• Gregorian Reformation has occurred 1582

• Week starts on Monday

• ISO-8601:1988 week numbers

• Representation of calendar sheet

September 1994
Monday          5 12 19 26
Tuesday         6 13 20 27
Wednesday       7 14 21 28
Thursday     1  8 15 22 29
Friday       2  9 16 23 30
Saturday     3 10 17 24
Sunday       4 11 18 25

Remember, all these internal defaults are modifiable by the options --date-format,

--gregorian-reform, --starting-day, --iso-week-number and --type.

3 Strictly speaking, an extended language code information.

4 Or other language codes, for which a message catalog will be created and distributed in future.

If no language code is detected, Gcal takes the internal defaults of the en language code^{5}.

If a language code is specified for which no message catalog is installed, Gcal takes the internal defaults of the de language code, but displays the message texts using the English native language. Actually, this behavior seems to me the most proper solution in such a case. The English native language is spoken all over the world unlike the German or other native languages, so it is wise to use it here. But the other criteria bound to the English native language are so special for users of other native languages, that it is wise to use the criteria taken for internal defaults of the de language code, because most European countries (taken as standard) essentially use them.

Now British users will certainly ask whether they could use their date format as an internal default^{6}. The answer to this is a simple yes, nevertheless, these users have to set the environment variable LANG with an extended language code information instead of a simple language code information.

The usual template of an extended language code information is as follows:

• For XPG syntax (X/Open Portability Guide):

           language[_territory][.codeset][@modifier]

• For CEN syntax (Comité Europén de Normalisation):

           language[_territory][+audience][+special][,sponsor][_revision]

Both syntaxes contain the language and territory components, which are used by Gcal to select the native language and the other criteria. The language component is equivalent to the simple language code information and the territory component is a two-letter territory or country code as defined by the ISO-3166 like ‘GB’ for Great Britain or ‘US’ for the U.S.A. See the pertinent literature for more details. So British users only have to set the LANG environment variable with a en_GB contents, and after that, they can use the British date format as an internal default.

5 Or to be more precise, of the extended language code information en_US.

6 All other internal defaults of the simple en language code information just meet their criteria.

Appendix D Metasymbols

Within this document, the following metasyntactic conventions will be used when discussing command line options, commands, arguments and templates:

n Argument is not optional and n must be a valid number. e|t|dvar Argument is not optional and must be a single alphabetic character. $tvar Argument is not optional and must be a single alphabetic character, that is

lead by a ‘$’ character.

[text] Argument is optional and text is a valid string of text.

[ab] Arguments are optional, any number may be used in any order, e.g.:

               NOTHING, a, b, ab, ba ...

{ab} Arguments are not optional, minimum one up to any number may be used in

any order, e.g.:

               a, b, ab, ba ...

[a|b] Arguments are optional, select either a or b, but not both. a|b Arguments are not optional, select either a or b, but not both. [a[b]] Arguments are optional, select either a or ab, but not b. [a|b[c]] Arguments are optional, select a or b or ac or bc, but not c. [a|bc] Arguments are optional, select either ac or bc, but not a or b or c. [a]|[b] Argument lists are optional, select either list a or list b, but not both. […] Additional arguments/files may be specified and are optional.

Appendix E Regular Expressions

This appendix is derived from the gawk-3.0.0 and regex-0.12 manuals.

A regular expression, or regexp, is a way of describing a set of strings. The simplest regular expression is a sequence of letters, numbers, or both. Such a regexp matches any string that contains that sequence. Thus, the regexp ‘foo’ matches any string containing ‘foo’. Other kinds of regular expressions let you specify more complicated classes of strings.

Gcal uses exactly one of the following regular expression programming libraries respectively methods, and that in the decreasing priority how it is chosen at configuration time, or better, compile time of the software:

No. Function  Symbol

1.  GNU re_compile_pattern() and regex.h                    GNU-REGEX
2.  POSIX regcomp() and regex.h                              POSIX-REGEX
3.  BSD re_comp()                                           BSD-REGEX
4.  System V regcmp()                                        SysV-REGEX
5.  Henry Spencer V8 regcomp() and regexp.h                   V8-REGEX

6.  Pattern matching is supported, but without metacharacters  NO-REGEX

Try ‘gcal --version’ for detecting the kind of regular expression programming library respectively method that is burned in your Gcal program!

E.1 How to Use Regular Expressions

A regular expression can be used as a pattern if Gcal’s --filter-text=pattern option is specified at program start-up.

See [Fixed date option --filter-text=pattern], page 44, for more details.

Then the regular expression is tested against the entire, completely expanded text of each valid fixed date, and the fixed date is only displayed in case the pattern matches the text^{1}.

E.2 Regular Expression Operators

You can combine regular expressions with the following characters, called regular expression operators, or metacharacters, to increase the power and versatility of regular expressions.

1 Normally, it only needs to match some part of the text in order to succeed.

Here is a table of these metacharacters. All characters that are not listed in the table stand for themselves. \ This is used to suppress the special meaning of a character when matching. For

example:

     \$

matches the character ‘$’.

^This matches the beginning of a string. For example:

     ^@chapter

matches the ‘@chapter’ at the beginning of a string, and can be used to identify chapter beginnings in Texinfo source files. The ‘^’ is known as an anchor, since it anchors the pattern to matching only at the beginning of the string.

$ This is similar to ‘^’, but it matches only at the end of a string. For example:

     p$

matches a string that ends with a ‘p’. The ‘$’ is also an anchor.

. The period, or dot, matches any single character. For example:

     .P

matches any single character followed by a ‘P’ in a string. Using concatenation we can make a regular expression like ‘U.A’, which matches any three-character sequence that begins with ‘U’ and ends with ‘A’.

[…] This is called a character list. It matches any one of the characters that are

enclosed in the square brackets. For example:

     [MVX]

matches any one of the characters ‘M’, ‘V’, or ‘X’ in a string. Ranges of characters are indicated by using a hyphen between the beginning and ending characters, and enclosing the whole thing in brackets. For example:

     [0-9]

matches any digit. Multiple ranges are allowed. E.g., the list [A-Za-z0-9] is a common way to express the idea of “all alphanumeric characters.” To include one of the characters ‘\’, ‘]’, ‘-’ or ‘^’ in a character list, put a ‘\’ in front of it. For example:

     [d\]]

matches either ‘d’, or ‘]’. Character classes are a new feature introduced in the POSIX standard. A character class is a special notation for describing lists of characters that have a specific attribute, but where the actual characters themselves can vary from country to country and/or from character set to character set. For example, the notion of what is an alphabetic character differs in the U.S.A. and in France. A character class is only valid in a regexp inside the brackets of a character list. Character classes consist of ‘[:’, a keyword denoting the class, and ‘:]’. Here are the character classes defined by the POSIX standard:

[:alnum:]

Alphanumeric characters.

[:alpha:]

Alphabetic characters.

[:blank:]

Space and tab characters.

[:cntrl:]

Control characters.

[:digit:]

Numeric characters.

[:graph:]

Characters that are printable and are also visible^{1}.

[:lower:]

Lower-case alphabetic characters.

[:print:]

Printable characters^{2}.

[:punct:]

Punctuation characters^{3}.

[:space:]

Space characters^{4}.

[:upper:]

Upper-case alphabetic characters.

[:xdigit:]

Characters that are hexadecimal digits.

For example, before the POSIX standard, to match alphanumeric characters, you had to write [A-Za-z0-9]. If your character set had other alphabetic characters in it, this would not match them. With the POSIX character classes, you can write [[:alnum:]], and this will match all the alphabetic and numeric characters in your character set.

Two additional special sequences can appear in character lists. These apply to non-ascii character sets, which can have single symbols (called collating elements) that are represented with more than one character, as well as several characters that are equivalent for collating, or sorting, purposes. (E.g., in French, a plain ‘e’ and a grave-accented ‘è’ are equivalent.)

1 A space is printable, but not visible, while an ‘a’ is both.

2 Characters that are not control characters.

3 Characters that are not letter, digits, control characters, or space characters.

4 Such as Space, Tab, and Form-feed, to name a few.

Collating Symbols A collating symbol is a multi-character collating element enclosed in ‘[.’ and ‘.]’. For example, if ‘ch’ is a collating element, then [[.ch.]] is a regexp that matches this collating element, while [ch] is a regexp that matches either ‘c’ or ‘h’.

Equivalence Classes An equivalence class is a list of equivalent characters enclosed in ‘[=’ and ‘=]’. Thus, [[=eè=]] is a regexp that matches either ‘e’ or ‘è’.

These features are very valuable in non-English speaking locales. Caution: The library functions that Gcal uses for regular expression matching currently only recognize POSIX character classes (possibly); they do not recognize collating symbols or equivalence classes.

[^…] This is a negated character list respectively complemented character list. The first character after the ‘[’ must be a ‘^’. It matches any characters except those in the square brackets. For example:

     [^0-9]

matches any character that is not a digit.

| This is the alternation operator, and it is used to specify alternatives. For

example:

     ^P|[0-9]

matches any string that matches either ‘^P’ or ‘[0-9]’. This means it matches any string that starts with ‘P’ or contains a digit. The alternation applies to the largest possible regexps on either side. In other words, ‘|’ has the lowest precedence of all the regular expression operators.

(…) Parentheses are used for grouping in regular expressions as in arithmetic. They can be used to concatenate regular expressions containing the alternation operator, ‘|’. For example, ‘@(samp|code)\{[^}]+\}’ matches both ‘@code{foo}’ and ‘@samp{bar}’. (These are Texinfo formatting control sequences.)

* This symbol means that the preceding regular expression is to be repeated as

many times as necessary to find a match. For example:

     ph*

applies the ‘*’ symbol to the preceding ‘h’ and looks for matches of one ‘p’ followed by any number of ‘h’s. This will also match just ‘p’ if no ‘h’s are present. The ‘*’ repeats the smallest possible preceding expression. (Use parentheses if you wish to repeat a larger expression.) It finds as many repetitions as possible. For example:

     gcal --filter-text='\(c[ad][ad]*r x\)' -f sample.rc -y

prints every fixed date in sample.rc containing a fixed date text of the form ‘(car x)’, ‘(cdr x)’, ‘(cadr x)’, and so on. Notice the escaping of the parentheses by preceding them with backslashes.

+ This symbol is similar to ‘*’, but the preceding expression must be matched at

least once. This means that:

     wh+y

would match ‘why’ and ‘whhy’ but not ‘wy’, whereas ‘wh*y’ would match all three of these strings. This is a simpler way of writing the last ‘*’ example:

     gcal --filter-text='\(c[ad]+r x\)' -f sample.rc -y

? This symbol is similar to ‘*’, but the preceding expression can be matched

either once or not at all. For example:

     fe?d

will match ‘fed’ and ‘fd’, but nothing else.

{n}
{n,}

{n,m} One or two numbers inside braces denote an interval expression which is available in the POSIX standard. If there is one number in the braces, the preceding regexp is repeated n times. If there are two numbers separated by a comma, the preceding regexp is repeated n to m times. If there is one number followed by a comma, then the preceding regexp is repeated at least n times.

wh{3}y matches ‘whhhy’ but not ‘why’ or ‘whhhhy’.

wh{3,5}y matches ‘whhhy’ or ‘whhhhy’ or ‘whhhhhy’, only.

wh{2,}y matches ‘whhy’ or ‘whhhy’, and so on.

GNU software that deals with regular expressions provides a number of additional regexp operators. These operators are described here.

Most of the additional operators are for dealing with word matching. For our purposes, a word is a sequence of one or more letters, digits, or underscores (‘_’).

\w This operator matches any word-constituent character, i.e. any letter, digit, or underscore. Think of it as a short-hand for [A-Za-z0-9_] or [[:alnum:]_].

\W This operator matches any character that is not word-constituent. Think of it

as a short-hand for [^A-Za-z0-9_] or [^[:alnum:]_].

\< This operator matches the empty string at the beginning of a word. For example, \<away matches ‘away’, but not ‘stowaway’.

\> This operator matches the empty string at the end of a word. For example, stow\> matches ‘stow’, but not ‘stowaway’.

\b This operator matches the empty string at either the beginning or the end of a word (the word boundary). For example, ‘\bballs?\b’ matches either ‘ball’ or ‘balls’ as a separate word.

\B This operator matches the empty string within a word. In other words, ‘\B’ matches the empty string that occurs between two word-constituent characters. For example, \Brat\B matches ‘crate’, but it does not match ‘dirty rat’. ‘\B’ is essentially the opposite of ‘\b’.

There are two other operators that work on buffers. In Emacs, a buffer is, naturally, an Emacs buffer. For other programs, the regexp library routines that Gcal uses consider the entire string to be matched as the buffer^{5}.

For Gcal, since ‘^’ and ‘$’ always work in terms of the beginning and end of strings, these operators do not add any new capabilities. They are provided for compatibility with other GNU software.

\` This operator matches the empty string at the beginning of the buffer.

\' This operator matches the empty string at the end of the buffer.

In regular expressions, the ‘*’, ‘+’, and ‘?’ operators, as well as the braces ‘{’ and ‘}’, have the highest precedence, followed by concatenation, and finally by ‘|’. As in arithmetic, parentheses can change how operators are grouped.

Case is normally significant in regular expressions, both when matching ordinary characters (i.e. not metacharacters), and inside character sets. Thus a ‘w’ in a regular expression matches only a lower-case ‘w’ and not an upper-case ‘W’.

The simplest way to do a case-independent match is to use a character list: ‘[Ww]’. However, this can be cumbersome if you need to use it often; and unfortunately, it can make the regular expressions harder to read. Supplying a want, Gcal offers the --ignore-case option which ignores all case distinctions in both the regular expression and the completely expanded text of each valid fixed date. See [Fixed date option --ignore-case], page 45.

5 That is the text part of a Gcal resource file line.

Appendix F Summary of all Regular Expressions

Regular expressions are based on POSIX EREs (Extended Regular Expressions). Regexps are composed of characters. Here is a short-list of them all:

c matches the character c (assuming c is none of the characters listed below). \c matches the literal character c. . matches any character. ^matches the beginning of a string. $ matches the end of a string. [abc…] matches any of the characters abc… (character list).

[[:keyword:]]

matches any character in the character class keyword. Allowable classes are

          alnum, alpha, blank, cntrl, digit, graph, lower, print, punct, space,
          upper, and xdigit.
[[.element.]]

matches the multi-character collating element. Gcal does not currently support collating symbols.

[[=list=]]

matches any of the equivalent characters in list. Gcal does not currently support equivalence classes.

[^abc. . .]

matches any character except abc… (negated resp. complemented character list).

r1|r2 matches either r1 or r2 (alternation). r1r2 matches r1, and then r2 (concatenation). r+ matches one or more r’s. r* matches zero or more r’s. r? matches zero or one r’s.

(r) matches r (grouping).

r{n}
r{n,}

r{n,m} matches at least n, n to any number, or n to m occurrences of r (interval

expressions).

\b matches the empty string at either the beginning or the end of a word. \B matches the empty string within a word. \< matches the empty string at the beginning of a word.

\> matches the empty string at the end of a word. \w matches any word-constituent character (alphanumeric characters and the un-

derscore).

\W matches any character that is not word-constituent. \` matches the empty string at the beginning of a buffer^{1}. \' matches the empty string at the end of a buffer.

1 The same as a text part of a resource file line in Gcal.

Appendix G %?… Special Texts

The %?… special texts which can be used in the text part of a Gcal resource file line can coarsely be separated into three categories. So there are special texts used for suppression output of fixed dates in definite cases. Others are replaced by particular texts or cause the shell to start external commands. See Section 5.1.5 [%?… Special Texts], page 100, for a general description of the special texts which may be used in the text part of a Gcal resource file line.

Some of these special texts may have an optional format instruction (called format in the further context), which affects the representation of an expanded special text. See Appendix J [Format Instruction], page 249, for the detailed description of the format instruction and its components. A possibly specified format within special texts which may not have a format instruction is ignored by Gcal and not displayed in output.

But Gcal is also able to represent such special texts in a special way by using a format instruction, and that by using an indirect method. For displaying, you simply assign the special text exclusively to a text variable which contains a format instruction. Nevertheless, it is important to note that the field width component of the format instruction is chosen in the way that it refers to the assigned contents of the text variable, if the format instruction has a fixed format component. The contents assigned to the text variable is only the special text, and not its expanded result. Gcal has to proceed like that, because the special text which is hidden in the text variable could not be used (here: displayed) only for a definite date, but possibly also for several dates, and that adjusted to each date it is referenced. Thus exists a special text in the text variable which has a definite length, and a format instruction also exists, whose field width component has to affect the expanded result of the special text. Indeed, Gcal does not know at the moment when the assignment of the special text to the text variable is made, of which length the expanded result will be, because this expansion is made at a much later phase in the program execution. So, in such a case, the field width component of a format instruction should have the length of the special text minimum, otherwise it is crippled and expanded respectively displayed incorrectly by the internal methods which effect the expansion.

Of course, not all possibilities concerning the representation —which are offered by the format instruction— are listed here for those special texts those expanded representation may be affected by an optional format instruction, but only some few exemplary.

And some of these special texts may also have an optional simple or special date argument (called date in the further context) or another argument (called argument in the further context). The date argument must be given in one of the following date formats in case a special text is directly trailed by it:

  • yyyy[mm[dd|wwwn]]
  • yyyy*d|wn[www]
  • yyyy@e|t|dvar[[+|-]n[www]]

Special texts that may have a date argument must always be trailed by a whitespace character which is removed in output, no matter whether date was given or not!

See Appendix H [Summary of all %?… Special Texts], page 211, for a short-list of all applicable special texts.

G.1 %?… special texts for date exclusion

%?… special texts used for exclusion of points of time or periods of dates are very useful if they are used together with ranges of days (see Section 5.1.3.2 [Ranges of days], page 90). The complete special text is always suppressed in output in case the fixed date text must be displayed.

G.1.1 Exclusions with date argument %?[date] special texts

First of all, an inclusive date period %?[date] special text can be specified with the effect, that only those fixed dates are respected and displayed which are part of this period. And the specification of an exclusive date period %?[date] special text has the effect, that only those fixed dates are respected and displayed which are not part of this period.

Cleverly combined, these special texts offer very effective filtration capabilities used for the definition of special periods in a highly flexible manner, for example:

     0*d1#999 Every day of year: from January...August except April\
              %i00001#00008 %e00004#00004
     ;
     0*d1#999 Every day of year: from 1991...1993 and from 1996\
              except February 1992 and all September\
              %i1991#1993 %i1996# %e19922#19922 %e00009#00009

G.1.1.1 Inclusive date period %i[date][#[date]] special text

%i[date][#[date]] references an inclusive date period, i.e. all dates which are part of the specified period are valid and are displayed. Of course this means, that all fixed dates which are not part of this period (not covered by it) are invalid and remain undisplayed therefore.

The first specified date defines the starting date of the fixed date, i.e. the first date the fixed date may occur. For the starting date, the following rules concerning definite omission values are valid, elucidated by using the yyyy[mm[dd|wwwn]] date format:

• If yyyy is set to 0000, the current year is assumed for the year.

• If mm is not given, then 01 (≡January) is assumed for the month.

• If mm is set to 00, the current month is assumed for the month, and the December is assumed in case mm is set to 99.

• If dd|www is not given, then 01 is assumed for the day.

• If dd is set to 00, the current day is assumed for the day, and the last day of the month mm is assumed in case dd is set to 99.

• If the n field is given (range 1…5 or 9), then a www field must also be specified.

The second specified date defines the ending date of the fixed date, i.e. the last date the fixed date may occur. For the ending date, the following rules concerning definite omission values are valid, likewise elucidated by using the yyyy[mm[dd|wwwn]] date format:

• If yyyy is set to 0000, the current year is assumed for the year.

• If mm is not given, then 12 (≡December) is assumed for the month.

• If mm is set to 00, the current month is assumed for the month, and the December is assumed in case mm is set to 99.

• If dd|www is not given, the last day of the month mm is assumed for the day.

• If dd is set to 00, the current day is assumed for the day, and the last day of the month mm is assumed in case dd is set to 99.

• If the n field is given (range 1…5 or 9), then a www field must also be specified.

The preceding rules are analogously valid for the other permitted date formats.

If the starting date, which is encoded in the %i… special text, is later the ending date, the %i… special text will be ignored.

In case neither the starting date nor the ending date is specified, means only ‘%i’ is given, the implicit period 1st January 1…31st December 9999 (00010101…99991231) will be used.

If only the starting date is specified, means either ‘%idate’ or ‘%idate#’ is given, the implicit ending date 31st December 9999 (99991231) will be used.

In case only the ending date is specified, means ‘%i#date’ is given, the implicit starting date 1st January 1 (00010101) will be used.

For example:

     0d*1#999 Every day of year: only April and July\
              %i00004#00004 %i00007#00007
     ;
     0*d1#999 Every day of year: from January...October\
              %i#000010
     ;
     0*d1#999 Every day of year: from August 1990\
              %i19908
     ;
     0*d1#999 Every day of year: from August...December\
              %i00008
     ;
     0*d1#999 Every day of year: only today%i0@t#0@t

G.1.1.2 Exclusive date period %e[date][#[date]] special text

%e[date][#[date]] references an exclusive date period, i.e. all dates which are part of the specified period are invalid and remain undisplayed therefore. Of course this means, that all fixed dates which are not part of this period (not covered by it) are valid and are displayed.

The first specified date defines the starting date of the fixed date, i.e. the first date the fixed date may occur. For the starting date, the following rules concerning definite omission values are valid, elucidated by using the yyyy[mm[dd|wwwn]] date format:

• If yyyy is set to 0000, the current year is assumed for the year.

• If mm is not given, then 01 (≡January) is assumed for the month.

• If mm is set to 00, the current month is assumed for the month, and the December is assumed in case mm is set to 99.

• If dd|www is not given, then 01 is assumed for the day.

• If dd is set to 00, the current day is assumed for the day, and the last day of the month mm is assumed in case dd is set to 99.

• If the n field is given (range 1…5 or 9), then a www field must also be specified.

The second specified date defines the ending date of the fixed date, i.e. the last date the fixed date may occur. For the ending date, the following rules concerning definite omission values are valid, likewise elucidated by using the yyyy[mm[dd|wwwn]] date format:

• If yyyy is set to 0000, the current year is assumed for the year.

• If mm is not given, then 12 (≡December) is assumed for the month.

• If mm is set to 00, the current month is assumed for the month, and the December is assumed in case mm is set to 99.

• If dd|www is not given, the last day of the month mm is assumed for the day.

• If dd is set to 00, the current day is assumed for the day, and the last day of the month mm is assumed in case dd is set to 99.

• If the n field is given (range 1…5 or 9), then a www field must also be specified.

The preceding rules are analogously valid for the other permitted date formats.

If the starting date, which is encoded in the %e… special text, is later the ending date, the %e… special text will be ignored.

In case neither the starting date nor the ending date is specified, means only ‘%e’ is given, the implicit period 1st January 1…31st December 9999 (00010101…99991231) will be used.

If only the starting date is specified, means either ‘%edate’ or ‘%edate#’ is given, the implicit ending date 31st December 9999 (99991231) will be used.

In case only the ending date is specified, means ‘%e#date’ is given, the implicit starting date 1st January 1 (00010101) will be used.

For example:

     0d*1#999 Every day of year: except April and July\
             %e00004#00004 %e00007#00007
     ;
     0*d1#999 Every day of year: except January...October\
              %e#000010
     ;
     0*d1#999 Every day of year: until July 1990\
              %e19908
     ;
     0*d1#999 Every day of year: except August...December\
              %e00008
     ;
     0*d1#999 Every day of year: except today%e0@t#0@t

G.1.2 Exclusions without any argument %? special texts

First of all, an inclusive day period %? special text can be specified with the effect, that only those fixed dates are respected and displayed which are part of this period. And the specification of an exclusive day period %? special text has the effect, that only those fixed dates are respected and displayed which are not part of this period.

Cleverly combined, these special texts also offer very effective filtration capabilities used for the definition of very special periods.

For example:

     0*d1#999 Every day of year: all Mondays...Fridays\
              except Wednesdays and all legal holidays\
              %r%F%V
     ;
     00001001#9999 Every day in last quarter of year: all holidays\
                   which are not on a Sunday%x %P

G.1.2.1 Inclusive day period %? special texts

The following inclusive day periods %? special texts are respected, at which inclusive is meant for the period which is not excluded. %v Excludes fixed date in case it is not listed as legal holiday in the eternal holiday

list.

%x Excludes fixed date in case it is neither listed as legal holiday nor memorial day

in the eternal holiday list.

%a Excludes fixed date in case it is no Monday. %c Excludes fixed date in case it is no Tuesday. %f Excludes fixed date in case it is no Wednesday. %g Excludes fixed date in case it is no Thursday. %h Excludes fixed date in case it is no Friday. %l Excludes fixed date in case it is no Saturday. %p Excludes fixed date in case it is no Sunday. %q Excludes fixed date in case it is no Monday…Thursday. %r Excludes fixed date in case it is no Monday…Friday.

G.1.2.2 Exclusive day period %? special texts

The following exclusive day periods %? special texts are respected, at which exclusive is meant for the period which is excluded. %V Excludes fixed date in case it is listed as legal holiday in the eternal holiday list. %X Excludes fixed date in case it is either listed as legal holiday or memorial day

in the eternal holiday list.

%A Excludes fixed date in case it is a Monday. %C Excludes fixed date in case it is a Tuesday. %F Excludes fixed date in case it is a Wednesday. %G Excludes fixed date in case it is a Thursday. %H Excludes fixed date in case it is a Friday.

%L Excludes fixed date in case it is a Saturday.

%P Excludes fixed date in case it is a Sunday.

%Q Excludes fixed date in case it is a Monday…Thursday.

%R Excludes fixed date in case it is a Monday…Friday.

G.2 %?… special texts for text replacement

%?… special texts which are replaced by particular texts at program run-time causes the interspersing of data into the fixed date text, which quality is essentially depending on the command line arguments given for running the program. The complete special text is always replaced accordingly in output if the fixed date text must be displayed.

Cleverly combined, these special texts also offer very effective mechanisms used for the definition of particular texts in a highly flexible manner, for example:

     0*d1#999 Every day of year: Today is %K, the %n (%N)

G.2.1 Replacements with date argument %?[date] special texts

Apart from different representations of a weekday name, it is possible to create its weekday number. Likewise, the current day-of-year number, the day number, the week number, the month name, the month number, the year number, the Moon phase, the biorhythm, and some useful difference values can be produced.

For the date argument these %?[date] special text may have, the following rules concerning definite omission values are valid, elucidated by using the yyyy[mm[dd|wwwn]] date format^{1}:

• If yyyy … is not given (≡no date encoded), the current year is assumed for the date.

• If yyyy is set to 0000, the current year is assumed for the year.

• If mm is not given or set to 00, the current month is assumed for the month, and the December is assumed in case mm is set to 99.

• If dd|www is not given or set to 00, the current day is assumed for the day, and the last day of the month is assumed in case dd is set to 99.

• If the n field is given (range 1…5 or 9), then a www field must also be specified.

G.2.1.1 Weekday name %[format]K[date] special text

%K[date] Is replaced by the complete weekday name of the current day, e.g.: The text ‘Today is %K , the %n’ will be expanded to

               ↦ ‘Today is Monday, the 10-Jun-1996’, in case the actual system

date is the 10th June 1996.

%>3#K[date]

Is replaced by the 3-letter weekday name of the current day, e.g.:

               The text ‘Today is %>3#K0@t+1 , the %n+1’ will be expanded to
               ↦ ‘Today is Tue, the 11-Jun-1996’, in case the actual system

date is the 10th June 1996.

1 The succeeding rules are analogously valid for the other permitted date formats.

%>2#K[date]

Is replaced by the 2-letter weekday name of the current day, e.g.:

The text ‘Today is %>2#K , the %n’ will be expanded to ↦ ‘Today is Mo, the 10-Jun-1996’, in case the actual system date is the 10th June 1996.

G.2.1.2 Weekday number %[format]?[date] special texts

The following weekday number %[format]?[date] special texts are respected, e.g.:

%W[date] Is replaced by the weekday number of the current day, and which is determined

by the Monday≡1…Sunday≡7 counting method, e.g.:

The text ‘Today is weekday %W , the %n’ will be expanded to

↦ ‘Today is weekday 1, the 10-Jun-1996’, in case the actual sys-

tem date is the 10th June 1996.

%>1&*W[date]

Is replaced by the weekday number with trailing Ordinal Number suffix of the current day, and which is determined by the Monday≡1…Sunday≡7 counting method, e.g.:

The text ‘Today is weekday %>1&*W0@t+1 , the %n+1’ will be ex-

panded to

↦ ‘Today is weekday 2nd, the 11-Jun-1996’, in case the actual

system date is the 10th June 1996.

%E[date] Is replaced by the weekday number of the current day, and which is determined

by the Monday≡0…Sunday≡6 counting method, e.g.:

The text ‘Today is weekday %E , the %n’ will be expanded to

↦ ‘Today is weekday 0, the 10-Jun-1996’, in case the actual sys-

tem date is the 10th June 1996.

%>1&*E[date]

Is replaced by the weekday number with trailing Ordinal Number suffix of the current day, and which is determined by the Monday≡0…Sunday≡6 counting method, e.g.:

The text ‘Today is weekday %>1&*E , the %n’ will be expanded to
↦ ‘Today is weekday 0th, the 10-Jun-1996’, in case the actual

system date is the 10th June 1996.

%I[date] Is replaced by the weekday number of the current day, and which is determined

by the Sunday≡1…Saturday≡7 counting method, e.g.:

The text ‘Today is weekday %I , the %n’ will be expanded to

↦ ‘Today is weekday 2, the 10-Jun-1996’, in case the actual sys-

tem date is the 10th June 1996.

%>1&*I[date]

Is replaced by the weekday number with trailing Ordinal Number suffix of the current day, and which is determined by the Sunday≡1…Saturday≡7 counting method, e.g.:

The text ‘Today is weekday %>1&*I , the %n’ will be expanded to
↦ ‘Today is weekday 2nd, the 10-Jun-1996’, in case the actual

system date is the 10th June 1996.

%J[date] Is replaced by the weekday number of the current day, and which is determined by the Sunday≡0…Saturday≡6 counting method, e.g.:

The text ‘Today is weekday %J , the %n’ will be expanded to

↦ ‘Today is weekday 1, the 10-Jun-1996’, in case the actual sys-

tem date is the 10th June 1996.

%>1&*J[date]

Is replaced by the weekday number with trailing Ordinal Number suffix of the current day, and which is determined by the Sunday≡0…Saturday≡6 counting method, e.g.:

The text ‘Today is weekday %>1&*J , the %n’ will be expanded to
↦ ‘Today is weekday 1st, the 10-Jun-1996’, in case the actual

system date is the 10th June 1996.

%S[date] Is replaced by the weekday number of the current day, and which is determined by the starting day of week ≡1…ending day of week ≡7 counting method (see [Calendar option --starting-day=argument], page 33), e.g.:

The text ‘Today is weekday %S , the %n’ will be expanded to

↦ ‘Today is weekday 4, the 10-Jun-1996’, in case the actual sys-

tem date is the 10th June 1996 and Friday (≡5) is the starting day of week.

%>1&*S[date]

Is replaced by the weekday number with trailing Ordinal Number suffix of the current day, and which is determined by the starting day of week ≡1…ending day of week ≡7 counting method (see [Calendar option

--starting-day=argument], page 33), e.g.:
The text ‘Today is weekday %>1&*S , the %n’ will be expanded to
↦ ‘Today is weekday 4th, the 10-Jun-1996’, in case the actual

system date is the 10th June 1996 and Friday (≡5) is the starting day of week.

%T[date] Is replaced by the weekday number of the current day, and which is determined by the starting day of week ≡0…ending day of week ≡6 counting method (see [Calendar option --starting-day=argument], page 33), e.g.:

The text ‘Today is weekday %T , the %n’ will be expanded to

↦ ‘Today is weekday 3, the 10-Jun-1996’, in case the actual sys-

tem date is the 10th June 1996 and Friday (≡5) is the starting day of week.

%>1&*T[date]

Is replaced by the weekday number with trailing Ordinal Number suffix of the current day, and which is determined by the starting day of week ≡0…ending day of week ≡6 counting method (see [Calendar option

--starting-day=argument], page 33), e.g.:
The text ‘Today is weekday %>1&*T , the %n’ will be expanded to
↦ ‘Today is weekday 3rd, the 10-Jun-1996’, in case the actual

system date is the 10th June 1996 and Friday (≡5) is the starting day of week.

G.2.1.3 Day-of-year number %[format]N[date] special text

%N[date] Is replaced by the day-of-year number of the current day of year, e.g.:

The text ‘Day %N of year, the %n’ will be expanded to

↦ ‘Day 53 of year, the 22-Feb-1996’, in case the actual system

date is the 22nd February 1996.

%>03*N[date]

Is replaced by the 3-digit day-of-year number with leading zeroes of the current day of year, e.g.:

The text ‘Day %>03*N0@t-1  of year, the %n-1’ will be expanded

to

↦ ‘Day 052 of year, the 21-Feb-1996’, in case the actual system

date is the 22nd February 1996.

%>1&*N[date]

Is replaced by the day-of-year number with trailing Ordinal Number suffix of the current day of year, e.g.:

The text ‘Day %>1&*N of year, the %n’ will be expanded to

↦ ‘Day 53rd of year, the 22-Feb-1996’, in case the actual sys-

tem date is the 22nd February 1996.

%>03&*N[date]

Is replaced by the 3-digit day-of-year number with leading zeroes and trailing Ordinal Number suffix of the current day of year, e.g.:

The text ‘Day %>03&*N of year, the %n’ will be expanded to

↦ ‘Day 053rd of year, the 22-Feb-1996’, in case the actual sys-

tem date is the 22nd February 1996.

G.2.1.4 Day number %[format]D[date] special text

%D[date] Is replaced by the day number of the current day of month, e.g.:

The text ‘Day %D , %n’ will be expanded to ↦ ‘Day 2, 02-Feb-1996’, in case the actual system date is the 2nd February 1996.

%>02*D[date]

Is replaced by the 2-digit day number with leading zero of the current day of month, e.g.:

The text ‘Day %>02*D0@t-1 , %n-1’ will be expanded to ↦ ‘Day 01, 01-Feb-1996’, in case the actual system date is the 2nd February 1996.

%>1&*D[date]

Is replaced by the day number with trailing Ordinal Number suffix of the current day of month, e.g.:

The text ‘Day %>1&*D , %n’ will be expanded to ↦ ‘Day 2nd, 02-Feb-1996’, in case the actual system date is the 2nd February 1996.

%>02&*D[date]

Is replaced by the 2-digit day number with leading zero and trailing Ordinal Number suffix of the current day of month, e.g.:

The text ‘Day %>02&*D , %n’ will be expanded to ↦ ‘Day 02nd, 02-Feb-1996’, in case the actual system date is the 2nd February 1996.

G.2.1.5 Week number %k[date] special text

%k[date] Is replaced by either the 2-digit ISO week number or the standard week number with leading zero, or a 4-alphanumeric character text of the current week of year (see [Calendar option --starting-day=argument], page 33, [Calendar option --iso-week-number=yes|no], page 30, and Appendix C [Aspects in Internationalization], page 131, for further details), e.g.:

The text ‘Today is %K of week %k’ will be expanded to ↦ ‘Today is Monday of week 24’, in case the actual system date is the 10th June 1996.

G.2.1.6 Month name %[format]U[date] special text

%U[date] Is replaced by the complete month name of the current month of year, e.g.: The text ‘Today is %U , %n’ will be expanded to ↦ ‘Today is June, 10-Jun-1996’, in case the actual system date is the 10th June 1996.

%>3#U[date]

Is replaced by the 3-letter month name of the current month of year, e.g.: The text ‘Today is %>3#U19961010 , %n’ will be expanded to ↦ ‘Today is Oct, 10-Jun-1996’, in case the actual system date is the 10th June 1996.

G.2.1.7 Month number %[format]M[date] special text

%M[date] Is replaced by the month number of the current month of year, e.g.:

The text ‘Month %M , %n’ will be expanded to ↦ ‘Month 2, 22-Feb-1996’, in case the actual system date is the 22nd February 1996.

%>02*M[date]

Is replaced by the 2-digit month number with leading zero of the current month of year, e.g.:

The text ‘Month %>02*M000001 , %n’ will be expanded to ↦ ‘Month 01, 22-Feb-1996’, in case the actual system date is the 22nd February 1996.

%>1&*M[date]

Is replaced by the month number with trailing Ordinal Number suffix of the current month of year, e.g.:

The text ‘Month %>1&*M , %n’ will be expanded to ↦ ‘Month 2nd, 22-Feb-1996’, in case the actual system date is the 22nd February 1996.

%>02&*M[date]

Is replaced by the 2-digit month number with leading zero and trailing Ordinal Number suffix of the current month of year, e.g.:

The text ‘Month %>02&*M , %n’ will be expanded to ↦ ‘Month 02nd, 22-Feb-1996’, in case the actual system date is the 22nd February 1996.

G.2.1.8 Year number %[format]Y[date] special text

%Y[date] Is replaced by the year number of the current year, e.g.:

The text ‘Year %Y , %n’ will be expanded to ↦ ‘Year 933, 22-Feb-0933’, in case the actual system date is the 22nd February 933.

%>04*Y[date]

Is replaced by the 4-digit complete year number with leading zeroes of the current year, e.g.:

The text ‘Year %>04*Y0015 , %n’ will be expanded to ↦ ‘Year 0015, 22-Feb-0933’, in case the actual system date is the 22nd February 933.

G.2.1.9 Moon phase %[format]?[date] special texts

Gcal uses a very simple algorithm for detecting the Moon phase, which computes approximate values only. The Moon phase is always calculated for 0 o’clock Universal time (UTC/GMT). See [Calendar option --time-offset=argument], page 33, how to change the base time for which the astronomical functions are calculated. The following Moon phase %[format]?[date] special texts are respected: %O[date] Is replaced by the Moon phase text of the current day of year, e.g.:

The text ‘Moon phase %O , %n’ will be expanded to

↦ ‘Moon phase 94%+, 22-Dec-1996’, in case the actual system

date is the 22nd December 1996.

The constructed Moon phase text consists of a percent value, which informs about the illuminated fraction of the Moon’s disk, and a trailing sign that tells something about the state of the Moon. A trailing ‘+’ sign indicates a waxing Moon, a ‘-’ sign a waning Moon, a ‘@’ sign indicates the Full Moon phase, and a ‘!’ sign the New Moon phase.

%>03*O[date]

Is replaced by the 3-digit Moon phase text with leading zeroes of the current day of year, e.g.:

The text ‘Moon phase %>03*O0@t+1 , %n+1’ will be expanded to
↦ ‘Moon phase 098%+, 23-Dec-1996’, in case the actual system

date is the 22nd December 1996. The constructed Moon phase text consists of a percent value, which informs about the illuminated fraction of the Moon’s disk, and a trailing sign that tells something about the state of the Moon. A trailing ‘+’ sign indicates a waxing Moon, a ‘-’ sign a waning Moon, a ‘@’ sign indicates the Full Moon phase, and a ‘!’ sign the New Moon phase.

%Z[date] Is replaced by the Moon phase text graphics of the current day of year, e.g.:

The text ‘Date %n %Z  --- Moon phase %>03*O ~Text’ will be ex-

panded to

↦ ‘Date 08-Mar-1995 ’
↦ ‘  (       @@@@@@’
↦ ‘  (              @@@@@@@@@’
↦ ‘ (                @@@@@@@@@@@’
↦ ‘ (                   @@@@@@@@@@@@’
↦ ‘ (                    @@@@@@@@@@@@@’
↦ ‘(                      @@@@@@@@@@@@@’
↦ ‘(                      @@@@@@@@@@@@@’
↦ ‘ (                    @@@@@@@@@@@@@’
↦ ‘ (                   @@@@@@@@@@@@’
↦ ‘ (                  @@@@@@@@@@@’
↦ ‘  (             @@@@@@@@@’
↦ ‘  (        @@@@@@@ --- Moon phase 041%+’

↦ ‘Text’, in case the actual system date is the 8th March 1995. See [Fixed date option --moonimage-lines=number], page 49, how to change the size of a Moon phase text graphics. A possibly specified format within this special text is ignored by Gcal.

G.2.1.10 Biorhythm %?[date] special texts

Gcal is able to create a biorhythm for any specified date of birth. The created text shows the three standard biorhythm cycles, and that are the 28 day emotional cycle, the 33 day intellectual cycle, and the 23 day physical cycle. The emotional cycle governs sensibility, nerves, moodiness, and creative ability. The intellectual cycle reflects intelligence, memory, mental alertness, and reasoning power. The physical cycle represents physical strength, endurance, energy and resistance. All cycles start in the zero point at the date of birth and swing like sine curves between their positive and negative maximum values. The periods above the zero point show the days of full vitality and efficiency while the periods below the zero point indicate days of reduced efficiency. The biorhythm is implemented for entertaining purposes only! It is up to the user to interpret the biorhythm texts.

Gcal counts critical days, positive likewise negative days. Critical days are those days in which one or more of the biological cycles crosses the zero point. At that time, one’s system is said to be in a state of flux and it may be desirable to exhibit caution. Positive days are those days in which one or more of the biological cycles have a positive maximum value. At that time, one’s system is said to be in a raised state. Negative days are those days in which one or more of the biological cycles have a negative maximum value. At that time, one’s system is said to be in a lessen state.

The following biorhythm %?[date] special texts are respected:

%,[date] Is replaced by the biorhythm text of the current day of year, e.g.: The text ‘%,19620921’ will be expanded to

     ↦ ‘1! 0+ 0- , Emo=+022%- Int=-091%+ Phy=-014%-’, in case the

actual system date is the 12th December 1996. The preceding example shows the calculated biorhythm expressed as a series of values for a person born in 1962, September 21st. The constructed text consists of two parts, and that is on the one hand a triplet of total values which indicate the critical, positive and negative day. And on the other hand a triplet of values which indicate the emotional, intellectual and physical cycle. In the first triplet of total values, a ‘!’ suffix means a critical day, a ‘+’ indicates a positive day, and a ‘-’ suffix marks a negative day. In the second triplet of values, each single cycle percent value consists of a leading positive or negative sign that tells something about the distance of this value to the zero point, and a trailing character that tells something about the state of each cycle. A trailing ‘+’ character marks a waxing phase, a ‘-’ suffix means a waning phase, and a ‘@’ indicates the maximum value of a phase.

%;[date] Is replaced by the biorhythm text graphics of the current day of year, e.g.: The text ‘%;19620921’ will be expanded to ↦ ‘0- I P 1 E +0’, in case the actual system date is the 12th December 1996.

The preceding example shows the calculated biorhythm expressed as a text graphics line for a person born in 1962, September 21st. The constructed line is a bar with a negative and positive axis of adjustable length (see [Fixed date option --biorhythm-axis=number], page 49), which left margin is represented by the total value of the negative days (the ‘0-’ in this case), and which right margin is represented by the total value of the positive days (the ‘+0’ in this case). The zero point of the co-ordinate is represented by the total value of the critical days, which is the ‘1’ in this case. The emotional, intellectual and the physical phase value is accordingly placed in scaled manner on this bar, and that by using the initial letter of the cycle in each case (‘E’, ‘I’ and ‘P’). Coincidental phase values are marked by a ‘@’ character.

G.2.1.11 Difference value %[format]?[date] special texts

The following difference value %[format]?[date] special texts are respected:

%ydate Is replaced by a year difference value, e.g.:

               The text ‘Sylvester 1912 is %y1912  years ago’ will be

expanded to

               ↦ ‘Sylvester 1912 is -82 years ago’, in case the year of the

actual system date is 1994.

%B[date] Is replaced by an age value, e.g.:

               The text ‘My %B1962  birthday%i1952#2062’ will be expanded to

↦ ‘My 32 birthday’, in case the year of the actual system date is 1994.

Those age values are only displayed, if the computation of an age value is greater zero. The fixed date warning in preceding example is displayed only in case the current year is greater than 1952 and less than 2062.

%>1&*B[date]

Is replaced by an age value with trailing Ordinal Number suffix, e.g.:

               The text ‘My %>1&*B1962  birthday%i1952#2062’ will be

expanded to

↦ ‘My 32nd birthday’, in case the year of the actual system date is 1994.

Those age values are only displayed, if the computation of an age value is greater zero. The fixed date warning in preceding example is displayed only in case the current year is greater than 1952 and less than 2062.

G.2.2 Replacements with other argument %[format]?[argument] special texts

Apart from different representations of the actual system time, a fixed format date text can be created. Moreover, it is possible to produce a day number, which bears the Julian date as base date and which can be deferred if needed. Furthermore, the approximate distance and the course angle between two geographic point locations, and different Sun and Moon oriented data and times for at pleasure any geographical location can be created. It is also possible to display the contents of environment variables in the fixed date text.

For the argument these %?[argument] special texts may have, no special rules concerning definite omission values are valid.

G.2.2.1 Actual clocktime %[format]?[argument] special texts

Gcal is able to represent the actual clocktime as local/zone time and as Universal time (UTC/GMT), and additionally the RFC-822^{2} style numerical Universal time timezone offset value and the actual numerical local time timezone offset value. See [Actual local time in hh:mm format %t[argument] special text], page 159, for the detailed description of all components of the argument, which may trail the actual clocktime %[format]?[argument] special texts. All actual clocktime %[format]?[argument] special texts must always be trailed by a whitespace character which is removed in output! The following actual clocktime %[format]?[argument] special texts are respected:

2 RFC is the abbreviation of Request for Comments.

%t[argument]

%t[*][[+|-]mmmm|hh:[mm]] references the actual local time, which is displayed by using the hh:mm output format, e.g.:

     The text ‘Local time: %t  --> Dentist at 10:00’ will be

expanded to

     ↦ ‘Local time: 07:32 --> Dentist at 10:00’, in case the actual

local time value is 07:32 o’clock (hh:mm). If only %t is specified, the local time value will implicitly be displayed by using the 24-hours format. For displaying using the 12-hours format, add a ‘*’ character directly behind %t, e.g. ‘%t*’. A displacement value may trail the %t special text, which has to be specified either by using the [+|-]mmmm format or the [+|-]hh:[mm] format. [+|-]mmmm adds respectively subtracts the specified amount of minutes mmmm from the local time value (range 0…9999), while [+|-]hh:[mm] adds respectively subtracts the given amount of hours hh (range 0…99) and minutes mm (range 0…59) from the local time value. The displacement value is always added to the local time value in case it is specified without a +|- sign. In case Gcal is unable to compute the actual local time in hh:mm format by reason of a misspecified argument, a ??:?? text will be created instead of the actual local time in hh:mm format.

     The text ‘Local time: %t-3:  --> Dentist at 10:00’ will be ex-

panded to

     ↦ ‘Local time: 04:32 --> Dentist at 10:00’, in case the actual

local time value is 07:32 o’clock (hh:mm).

%[format]'[argument]

%[format]'[[+|-]mmmm|hh:[mm]] references the actual local time in minutes since midnight, which is displayed by using the m[m…] output format, e.g.:

     The text ‘Local time: %'  --> Dentist at 10:00’ will be

expanded to

     ↦ ‘Local time: 452 --> Dentist at 10:00’, in case the actual

local time value is 07:32 o’clock (hh:mm). In case Gcal is unable to compute the actual local time in m[m…] format by reason of a misspecified argument, a ?? text will be created instead of the actual local time in m[m…] format.

%[format]_[argument]

%[format]_[*][[+|-]mmmm|hh:[mm]] references the actual local time hour, which is displayed by using the h[h…] output format, e.g.:

The text ‘Local time hour: %_-1:’ will be expanded to

↦ ‘Local time hour: 6’ in case the actual local time value is 07:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual local time hour by reason of a misspecified argument, a ?? text will be created instead of the actual local time hour h[h…].

%[format]?[argument]

%[format]?[[+|-]mmmm|hh:[mm]] references the actual local time minute, which is displayed by using the m[m…] output format, e.g.:

     The text ‘Local time minute: %>04*?+:2’ will be expanded to

↦ ‘Local time minute: 0034’ in case the actual local time value is 07:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual local time minute by reason of a misspecified argument, a ?? text will be created instead of the actual local time minute m[m…].

%[format]{[argument]

%[format]{[[+|-]mmmm|hh:[mm]] references the actual local time 12-hour format suffix, which is displayed as am|pm, e.g.:

The text ‘Local time suffix: %{’ will be expanded to

↦ ‘Local time suffix: am’ in case the actual local time value is 07:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual local time 12-hour format suffix by reason of a misspecified argument, a ?? text will be created instead of the actual local time 12-hour format suffix am|pm.

%@[argument]

%@[*][[+|-]mmmm|hh:[mm]] references the actual Universal time (UTC/GMT), which is displayed by using the hh:mm/date output format, e.g.:

The text ‘Universal time: %@*-3.’ will be expanded to

     ↦ ‘Universal time: 03:32am/15-Feb-1999’, in case the actual

date is the 15th February 1999 and the actual Universal time value is 06:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual Universal time in hh:mm/date format by reason of a misspecified argument, a ??:??/??-???-???? text will be created instead of the actual Universal time in hh:mm/date format.

%[format]`[argument]

%[format]`[[+|-]mmmm|hh:[mm]] references the actual Universal time in minutes since midnight, which is displayed by using the m[m…]/date output format, e.g.:

The text ‘Universal time: %>06*`-3:’ will be expanded to

     ↦ ‘Universal time: 000212/15-Feb-1999’ in case the actual date

is the 15th February 1999 and the actual Universal time value is 06:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual Universal time in m[m…]/date format by reason of a misspecified argument, a ??/??-???-???? text will be created instead of the actual Universal time in m[m…]/date format.

%[format].[argument]

%[format].[*][[+|-]mmmm|hh:[mm]] references the actual Universal time hour, which is displayed by using the h[h…] output format, e.g.:

The text ‘Universal time hour: %.-1:’ will be expanded to ↦ ‘Universal time hour: 5’ in case the actual Universal time value is 06:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual Universal time hour by reason of a misspecified argument, a ?? text will be created instead of the actual Universal time hour h[h…].

%[format]/[argument]

%[format]/[[+|-]mmmm|hh:[mm]] references the actual Universal time minute, which is displayed by using the m[m…] output format, e.g.:

     The text ‘Universal time minute: %>04*/+:2’ will be expanded

to

↦ ‘Universal time minute: 0034’ in case the actual Universal time value is 06:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual Universal time minute by reason of a misspecified argument, a ?? text will be created instead of the actual Universal time minute m[m…].

%[format]}[argument]

%[format]}[[+|-]mmmm|hh:[mm]] references the actual Universal time 12-hour format suffix, which is displayed as am|pm, e.g.:

     The text ‘Universal time suffix: %>1w*}+10:’ will be expanded

to

↦ ‘Universal time suffix: Pm’ in case the actual Universal time value is 06:32 o’clock (hh:mm).

In case Gcal is unable to compute the actual Universal time 12-hour format suffix by reason of a misspecified argument, a ?? text will be created instead of the actual Universal time 12-hour format suffix am|pm.

%"[argument]

%"[[+|-]mmmm|hh:[mm]] references the RFC-822 style numerical Universal time timezone offset value, which is displayed by using the +|-hhmm output format, e.g.:

     The text ‘Universal time timezone offset value: %"-90’ will

be expanded to

     ↦ ‘Universal time timezone offset value: -0130’, at which

calculations are always based on the timezone of the Universal time UTC/GMT.

In case Gcal is unable to compute the numerical Universal time timezone offset value by reason of a misspecified argument, a +???? text will be created instead of the numerical Universal time timezone offset value +|-hhmm.

%=[argument]

%=[[+|-]mmmm|hh:[mm]] references the RFC-822 style actual numerical local time timezone offset value, which is displayed by using the +|-hhmm output format, e.g.:

     The text ‘Local time timezone offset value CET: %=-10’ will

be expanded to

               ↦ ‘Local time timezone offset value CET: +0050’, in case

the actual timezone is equal the Central European (winter)time CEWT/CET (+hhmm). Calculations are always based on the timezone of the local time.

In case Gcal is unable to compute the actual numerical local time timezone offset value by reason of a misspecified argument, a +???? text will be created instead of the actual numerical local time timezone offset value +|-hhmm.

G.2.2.2 Textual date %n[argument] special text

%n[[+|-]n] references the current respectively queried day number relative to the current date ± n days and is replaced by a date text by using the fixed ‘%>02*D-%>3#U-%>04*Y’ format, e.g.:

The resource file line ‘1962Sep21 10000 days old: %n+10000’ will be expanded to ↦ ‘10000 days old: 06-Feb-1990’, in case you call Gcal with the simple -c option and the command sep 1962 (see [Single command mm yyyy], page 73).

G.2.2.3 Julian day number %[format]j[argument] special text

%[format]j[[+|-]n] references the current respectively queried Julian day number relative to the actual system date (≡today). This day number is based on the date 1st January 4713 BCE —which is the starting day zero of a consecutive day counting used in astronomical computations— and is known as the Julian Date (J.D.). The real zero of this date is at 12 o’clock Universal time (UTC/GMT); the day does not change at midnight, but at noon Universal time. Here, Gcal does not evaluate the timezone returned by the systems date function. For that reason, this day number is represented without a time fraction on the supposition that the day has already changed at noon. If you do not like the feature that the day displayed has already changed at noon, you can decrease the resulting Julian day number of that special text always by one, e.g. ‘%j-1’.

For example:

     The resource file line ‘0 Julian day %j since 01-Jan-4713 BCE’ will be ex-

panded to

     ↦ ‘Julian day 2437929 since 01-Jan-4713 BCE’, in case you call Gcal with

the -c %19620921 option and no command.

If the %[format]j text is directly trailed by an unsigned number, this number is always subtracted from the real Julian day number. So you are able to work with any quantities referenced, e.g.:

     The resource file line ‘0 Julian day %j2415021 since 01-Jan-1900’ will be

expanded to

↦ ‘Julian day 4 since 01-Jan-1900’, in case you call Gcal with the -c %19000105 option and no command.

G.2.2.4 Geographical distance and course angle %[format]bargument special text

%[format]b[*][mode]ISO-6709:1983-co-ordinate-1/ISO-6709:1983-co-ordinate-2

references either the approximate air line distance or the approximate course angle (true track)^{3} between any of two geographic point locations. The selection, which value has to be calculated by using this special text is done by specifying the mode part of the preceding argument. Actually, exactly three different modes can be used that are represented by the ‘0’…‘2’ characters:

Mode  Description

0

Calculates the air line distance between ISO-6709:1983-co-ordinate-1 and ISO-6709:1983-co-ordinate-2. The calculated air line distance value is displayed in kilometers by default. A ‘*’ character directly before this mode character causes Gcal to display the distance value for another quantity, and that in statute miles; where one statute mile is equal to 1.609344 kilometer. If Gcal is unable to compute the approximate distance between the two geographic point locations by reason of a misspecified argument, a ?? text will be created instead of the distance value.

1

Calculates  the  course  angle  (true  track)  between
ISO-6709:1983-co-ordinate-1  and  ISO-6709:1983-co-ordinate-2.

The calculated course angle value is displayed in degrees and arcminutes by default. The course angle is measured clockwise relative to the geographic, true North (not the magnetic North as shown by a compass), where angle values for the North direction are both denoted as 0 degree and 360 degree. A ‘*’ character directly before this mode character causes Gcal to display the course angle value using another style; and that in decimal degrees. If Gcal is unable to compute the approximate course angle between the two geographic point locations by reason of a misspecified argument, a ???d??' text will be created instead of the course angle value.

2

Like mode 1, but the course angle (true track) between ISO-6709:1983-co-ordinate-2 and ISO-6709:1983-co-ordinate-1 is calculated.

If no mode is given, Gcal automatically uses that mode, which is enabled by the mode character ‘0’. If a mode character is given that is not according to one of the ‘0’…‘2’ characters, Gcal also automatically uses that mode, which is enabled by the mode character ‘0’.

After the optional style and mode characters, the latitude and longitude of the geographic co-ordinates follows, for which the calculations must be made. They must be conform the ISO-6709:1983 standard representation of latitude and longitude for geographic point locations. The two co-ordinates have to be separated by a ‘/’ termination character from each other.

3 The course angle gives the direction, which is the vertex of a great circles arc (Orthodrome) that is

casted between two points on a surface of a sphere, at a meridian (Gcal uses a geoid (rotation ellipsoid)

that is calculated taking pattern from the World Geodetic System, short W.G.S, that was designed in 1961). Such an arc is the shortest geodetic connecting line between two points on a sphere. Indeed, the course angles on such an arc are changing continually, because the circular line that is casted intersects each meridian with another angle — except the case, that both co-ordinates are referring to exactly the same meridian (longitude), or to exactly the same latitude.

See [Arguments of the Sun oriented special texts], page 182, for the detailed description of the components of the ISO-6709:1983 standard representation of latitude and longitude for geographic point locations.

For example:

The text ‘Distance Paris-Tokyo: %b+4852+00220/+3542+13946 km’ will be

expanded to

↦ ‘Distance Paris-Tokyo: 9746km’.
The text ‘Distance Paris-Tokyo: %b*0+4852+00220/+3542+13946 ms’ will be

expanded to

↦ ‘Distance Paris-Tokyo: 6056ms’.
The text ‘Course angle Paris-Tokyo: %b1+4852+00220/+3542+13946’ will be

expanded to

↦ ‘Course angle Paris-Tokyo: 033d22'’.
The text ‘Course angle Tokyo-Paris: %b*2+4852+00220/+3542+13946’ will

be expanded to

↦ ‘Course angle Tokyo-Paris: 333.548’.

While praying, people of Islamic faith always turn their heads into the direction of Makkah, Saudi-Arabia. Now by means of Gcal, these people can easily find out for their respective location, where they have to turn to, and that by:

%b1ISO-6709:1983-co-ordinate-1/+212516+0394929

where ISO-6709:1983-co-ordinate-1 is simply replaced by the co-ordinate of the respective location.

See [Fixed date option --precise], page 43, how to obtain a more precise representation of the values that are caused by this special text.

This special text must always be trailed by a whitespace character which is removed in output!

G.2.2.5 Sun data %[format]?argument special texts

%[format]o[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate time of sunrise by default,

%[format]s[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate time of sunset by default,

%[format]u[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate period of visibility of the Sun (solar day length) by default,

%[format]z[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate period of non-visibility of the Sun (solar night length) by default.

All these special texts can be used for at pleasure any geographic point location, i.e. it is possible to determine different astronomical values for any location on the globe, and that for at pleasure any clocktime with a resolution of a single minute within the period of the years AD 1 until AD 9999, that is respected by Gcal.

The selection which value has to be calculated by these special texts is done by specifying the mode part of the preceding argument. Actually, exactly 54 different modes can be used that are represented by the ‘0’…‘9’, ‘a’…‘z’ and ‘A’…‘R’ characters, and which create different kind of results that are depending on the special text used. First of all, here is a table in which all usable modes are described and explained sufficiently. You can also see from this table, which Sun oriented special text or texts are corresponding to which mode, i.e. cause the determination of an astronomical value as it is described in the table:

Mode  Special text Description

0

          o, s

Calculates the approximate midnight time of the Sun. The astronomical midnight time of the Sun is at that clocktime, when the Sun holds an azimuth (horizontal angular distance between the vertical circle, that passes the Sun, and the North point) of either precisely 0 degrees of precisely 180 degrees (noon line), which depends on the season and the geographical location. At that clocktime, the Sun is close its lowest culmination point, i.e. close the lowest point below or above the horizontal plane the Sun transits during this day.

1

o, s

Calculates the approximate noon time of the Sun. The astronomical noon time of the Sun is at that clocktime, when the Sun holds an azimuth of either precisely 180 degrees of precisely 0 degrees (noon line), which depends on the season and the geographical location. At that clocktime, the Sun is close its highest culmination point, i.e. close the highest point above or below the horizontal plane the Sun transits during this day. People of Islamic faith normally pray for the second time on the day during the period, which is between the astronomical noon time of the Sun (or some minutes later) and the Islamic prayer time by the name of Asr. These people commonly use the term Zuhr for this prayer time. The timing of Asr depends on the length of the shadow cast by a vertical pole (gnomon). According to the Shafi school of jurisprudence, Asr begins when the length of the shadow of a vertical pole exceeds the length of the pole. According to the Hanafi school of jurisprudence, Asr begins when the length of the shadow exceeds twice the length of the vertical pole. In both cases, the minimum length of the shadow at astronomical noon time of the Sun is subtracted from the length of the shadow before comparing it with the length of the pole. See [Islamic Asr-1 prayer time], page 179, and [Islamic Asr-2 prayer time], page 180, for further details.

2

o

Calculates the approximate time when the center of the Sun passes a reference altitude of 0 degrees on a mathematicalgeocentric horizon in the morning; thus rising. A mathematical horizon is a purely geometrically-built horizon which disregards the phenomenon of refraction as it arises in reality by the influence of the Earth’s atmosphere. A geocentrical horizon is the horizontal plane that passes through the Earth’s center, orthogonal to the observer’s local vertical. In the further context, the shorter term mathematical horizon is used which actually means the mathematical-geocentric horizon.

2

s

Calculates the approximate time when the center of the Sun passes a reference altitude of 0 degrees on a mathematical horizon in the evening; thus setting.

2

u

Calculates the approximate period while the center of the Sun is above a reference altitude of 0 degrees on a mathematical horizon; thus is visible.

2

z

Calculates the approximate period while the center of the Sun is below a reference altitude of 0 degrees on a mathematical horizon; thus is non-visible.

3

o

Calculates the approximate time when the upper limb of the Sun passes a reference altitude of 0 degrees on a mathematical horizon in the morning; thus rising. The above mentioned reference altitude is computed from the value of the Sun’s semidiameter as it appear at that clocktime. If the reference altitude that is referring to the Sun’s upper limb is converted to a reference altitude that is referring to the Sun’s center, this results in a value of about 16 arcminutes below the geocentric horizon.

3

s

Calculates the approximate time when the upper limb of the Sun passes a reference altitude of 0 degrees on a mathematical horizon in the evening; thus setting. The above mentioned reference altitude is computed from the value of the Sun’s semidiameter as it appear at that clocktime.

3

u

Calculates the approximate period while the upper limb of the Sun is above a reference altitude of 0 degrees on a mathematical horizon; thus is visible.

3

z

Calculates the approximate period while the upper limb of the Sun is below a reference altitude of 0 degrees on a mathematical horizon; thus is non-visible.

4

o

Calculates the approximate time when the center of the Sun passes a reference altitude of 34 arcminutes below the geocentric horizon in the morning; thus rising. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the --atmosphere option. [Fixed dates option

--atmosphere=air-pressure[,temperature]], page 40, how to

change the base data of the atmosphere, so that the atmospheric conditions as defined by it are used to calculate the amount of refraction.

4

s

Calculates the approximate time when the center of the Sun passes a reference altitude of 34 arcminutes below the geocentric horizon in the evening; thus setting. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the --atmosphere option.

4

u

Calculates the approximate period while the center of the Sun is above a reference altitude of 34 arcminutes below the geocentric horizon; thus is visible.

4

z

Calculates the approximate period while the center of the Sun is below a reference altitude of 34 arcminutes below the geocentric horizon; thus is non-visible.

5

o

Calculates the approximate time when the upper limb of the Sun passes a reference altitude of 34 arcminutes below the geocentric horizon in the morning; thus rising. This kind of rise time calculation is done according to the standard calculation method as it is commonly used internationally. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the --atmosphere option. [Fixed dates

option --atmosphere=air-pressure[,temperature]], page 40,

how to change the base data of the atmosphere, so that the atmospheric conditions as defined by it are used to calculate the amount of refraction. The above mentioned reference altitude is computed from the respective values of the Sun’s semidiameter and (standard) refraction as they appear at that clocktime. If the reference altitude that is referring to the Sun’s upper limb is converted to a reference altitude that is referring to the Sun’s center, this results in a value of about 50 arcminutes below the geocentric horizon.

5

s

Calculates the approximate time when the upper limb of the Sun passes a reference altitude of 34 arcminutes below the geocentric horizon in the evening; thus setting. This kind of set time calculation is done according to the standard calculation method as it is commonly used internationally. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the --atmosphere option. The above mentioned reference altitude is computed from the respective values of the Sun’s semidiameter and (standard) refraction as they appear at that clocktime. People of Islamic faith normally pray for the second-last time on the day at this clocktime, or some minutes later. These people commonly use the term Maghrib for this prayer time.

5

u

Calculates the approximate period while the upper limb of the Sun is above a reference altitude of 34 arcminutes below the geocentric horizon; thus is visible. This kind of visibility period calculation is done according to the standard calculation method as it is commonly used internationally.

5

z

Calculates the approximate period while the upper limb of the Sun is above a reference altitude of 34 arcminutes below the geocentric horizon; thus is non-visible. This kind of non-visibility period calculation is done according to the standard calculation method as it is commonly used internationally.

6

o

Calculates the approximate time when the center of the Sun passes a reference altitude of 6 degrees below a mathematical horizon in the morning; thus the beginning of civil twilight. The scattered light of the Sun that is remaining at the beginning of the civil twilight phase is in general not yet sufficient for reading outside without artificial illumination.

6

s

Calculates the approximate time when the center of the Sun passes a reference altitude of 6 degrees below a mathematical horizon in the evening; thus the ending of civil twilight.

6

u

Calculates the approximate period while the center of the Sun is above a reference altitude of 6 degrees below a mathematical horizon; thus the period, while the center of the Sun is always above -6 degrees.

6

z

Calculates the approximate period while the center of the Sun is below a reference altitude of 6 degrees below a mathematical horizon; thus the period, while the center of the Sun is always below -6 degrees.

7

o

Calculates the approximate time when the center of the Sun passes a reference altitude of 12 degrees below a mathematical horizon in the morning; thus the beginning of nautical twilight. The scattered light of the Sun that is remaining at the beginning of the nautical twilight phase is in general not yet sufficient for navigation using a sea horizon.

7

s

Calculates the approximate time when the center of the Sun passes a reference altitude of 12 degrees below a mathematical horizon in the evening; thus the ending of nautical twilight.

7

u

Calculates the approximate period while the center of the Sun is above a reference altitude of 12 degrees below a mathematical horizon; thus the period, while the center of the Sun is always above -12 degrees.

7

z

Calculates the approximate period while the center of the Sun is below a reference altitude of 12 degrees below a mathematical horizon; thus the period, while the center of the Sun is always below -12 degrees.

8

o

Calculates the approximate time when the center of the Sun passes a reference altitude of 15 degrees below a mathematical horizon in the morning; thus the beginning of amateurastronomical twilight. The scattered light of the Sun that is remaining at the beginning of the amateur-astronomical twilight phase is in general yet so faint that most astronomical observations can be made.

8

s

Calculates the approximate time when the center of the Sun passes a reference altitude of 15 degrees below a mathematical horizon in the evening; thus the ending of amateur-astronomical twilight.

8

u

Calculates the approximate period while the center of the Sun is above a reference altitude of 15 degrees below a mathematical horizon; thus the period, while the center of the Sun is always above -15 degrees.

8

z

Calculates the approximate period while the center of the Sun is below a reference altitude of 15 degrees below a mathematical horizon; thus the period, while the center of the Sun is always below -15 degrees.

8

o

Calculates the approximate time when the center of the Sun passes a reference altitude of 18 degrees below a mathematical horizon in the morning; thus the beginning of astronomical twilight. No appreciable scattered sunlight is remaining at the beginning of the astronomical twilight phase, the sky is completely dark yet. People of Islamic faith normally pray for the first time on the day during the period, which is between this clocktime and the time of standard sunrise. These people commonly use the term Fajr for this prayer time. See [Standard rise time of the Sun], page 168, for further details.

9

s

Calculates the approximate time when the center of the Sun passes a reference altitude of 18 degrees below a mathematical horizon in the evening; thus the ending of astronomical twilight. People of Islamic faith normally pray for the last time on the day at this clocktime, or some minutes later. These people commonly use the term Isha for this prayer time.

9

u

Calculates the approximate period while the center of the Sun is above a reference altitude of 18 degrees below a mathematical horizon; thus the period, while the center of the Sun is always above -18 degrees.

9

z

Calculates the approximate period while the center of the Sun is below a reference altitude of 18 degrees below a mathematical horizon; thus the period, while the center of the Sun is always below -18 degrees.

a

o, s

Calculates the approximate topocentric, apparent elevation of the Sun, thus the vertical angular distance between the Sun’s center and the horizon, in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Sun’s center is below the horizon at the moment, and results with a positive sign mean that the momentary center of the Sun is above the horizon. Observations of celestial objects that are done from the surface of the Earth yield in topocentrically based data. The locations of the celestial bodies are often at another place if the data is topocentrically determined instead of determine it geocentrically, i.e. at the fictitious center of the Earth. This is mainly caused by the refraction, which raises a celestial body to another location as it is been in reality. Because the terrestrial globe flattens towards the pole caps and therefore cannot be taken as an ideally shaped sphere, the individual Earth radius between the observer’s location and the center of the Earth also affects the computation of topocentrically based data.

b

o, s

Calculates the approximate topocentric, apparent azimuth of the Sun in degrees and arcminutes as it happen at civil midnight time.

c

o, s

Calculates the approximate topocentric, apparent declination of the Sun, thus the vertical angular distance between the Sun’s center and the celestial equator, in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Sun’s center is below the celestial equator at the moment, and results with a positive sign mean that the momentary center of the Sun is below the celestial equator.

d

o, s

Calculates the approximate topocentric, apparent ecliptic longitude of the Sun, thus the horizontal angular distance between the Sun’s center and the vernal equinox point on the ecliptic (the zodiacal line or Sun’s orbit), in degrees and arcminutes as it happen at civil midnight time.

e

o, s

Calculates the approximate topocentric, apparent right ascension of the Sun, thus the horizontal angular distance between the Sun’s center and the hour circle that passes through the vernal equinox point on the ecliptic, as time value in hours and minutes as it happen at civil midnight time.

f

o, s

Calculates the approximate topocentric, apparent distance of the Sun from the Earth in astronomical units as it happen at civil midnight time. An astronomical unit, abbreviated by ae, is equal to the mean distance of the Sun from the Earth, which is about 149,597,870.691 kilometers.

g

o, s

Calculates the approximate topocentric, apparent horizontal parallax of the Sun in degrees and arcminutes as it happen at civil midnight time. The horizontal parallax of the Sun specifies the diameter of the Earth as it is seen from the surface of the Sun.

h

o, s

Calculates the approximate topocentric, apparent semidiameter of the Sun in degrees and arcminutes as it happen at civil midnight time.

i

o, s

Calculates the approximate refraction of the Earth’s atmosphere in degrees and arcminutes as it happen at civil midnight time.

j

o, s

Calculates the approximate geocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Sun’s center is below the horizon at the moment, and results with a positive sign mean that the momentary center of the Sun is above the horizon.

k

o, s

Calculates the approximate geocentric, apparent azimuth of the Sun in degrees and arcminutes as it happen at civil midnight time.

l

o, s

Calculates the approximate geocentric, apparent declination of the Sun in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Sun’s center is below the celestial equator at the moment, and results with a positive sign mean that the momentary center of the Sun is above the celestial equator.

m

o, s

Calculates the approximate geocentric, apparent ecliptic longitude of the Sun in degrees and arcminutes as it happen at civil midnight time.

n

o, s

Calculates the approximate geocentric, apparent right ascension of the Sun as time value in hours and minutes as it happen at civil midnight time.

o

o, s

Calculates the approximate geocentric, apparent distance of the Sun from the Earth in astronomical units as it happen at civil midnight time.

p

o, s

Calculates the approximate geocentric, apparent horizontal parallax of the Sun in degrees and arcminutes as it happen at civil midnight time.

q

o, s

Calculates the approximate geocentric, apparent semidiameter of the Sun in degrees and arcminutes as it happen at civil midnight time.

r

o, s

Calculates the approximate delta-t in seconds as it happen at civil midnight time. Delta-t is the difference between the Terrestrial Dynamical time (abbreviated by TDT), that was formerly known as Ephemeris time (abbreviated by ET), and the Universal time (UT). Thus, ‘delta-t ≡ TDT - UT’.

s

o, s

Calculates the approximate, apparent location oriented sidereal time (local sidereal time (LAST), also known as local star time) in hours and minutes as it happen at civil midnight time. A star day is the period between two consecutive upper culminations of the vernal equinox point on the ecliptic in the meridian of the observer’s location. Therefore, the local star time is the momentary period, which is past between the last upper culmination of the vernal equinox point in the meridian of the observer’s location (the momentary hour angle of the vernal equinox point), thus the right ascension of the stars in the observer’s meridian at the moment.

t

o, s

Outputs the base time as time value in hours and minutes, for which the dynamical, i.e. depending on the respective clocktime, astronomical data and times of the Sun are calculated. Without a given --time-offset=argument option, the astronomical data and times of the Sun are always calculated for 0 o’clock Universal time (UTC/GMT). See [Calendar option --time-offset=argument], page 33, for further details.

u

o, s

Calculates the approximate Julian date in days as it happen at civil midnight time. See Section G.2.2.3 [Julian day number], page 162, for further information about the Julian date.

v

o, s

Calculates the approximate Julian Ephemeris date, thus a Julian date that is corrected by delta-t, in days as it happen at civil midnight time.

w

o, s

Calculates the approximate difference between true solar time and mean solar time as time value in hours and minutes as it happen at civil midnight time. This so-called equation of time is a correction to be added to the true solar time —as read on a sundial— to obtain the mean solar time. A true solar day is the period between two consecutive lower culminations of the Sun. This entity is taken as the base for deriving the true solar time (as it is also shown by a sundial during the day). A star day is also known as a mean solar day. Because the Sun apparently shifts with respect to the vernal equinox point on the ecliptic due to the Earth’s orbit around the Sun, the star day and the true solar day have a different length. As the true Sun namely moves irregularly through the ecliptic, a fictitious mean Sun with a symmetrical motion through the celestial equator is used for deriving the mean solar time. So, this difference in time is a consequence of the ellipticity and tilt of the Earth’s orbit, causing the irregular apparent movement of the Sun across the sky.

x

o, s

Calculates the difference of the approximate topocentric, apparent elevation of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the momentary center of the Sun is at an elevation that is below the momentary elevation of the Moon’s center; thus the Sun is lower than the Moon. Results with a positive sign signify that the momentary center of the Sun is at an elevation that is above the momentary elevation of the Moon’s center; thus the Sun is higher than the Moon.

y

o, s

Calculates the difference of the approximate topocentric, apparent azimuth of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at civil midnight time. The result specifies the horizontal angular distance, by which the momentary center of the Sun is distant from the momentary Moon’s center, and that measured at the vertical circles that pass the Sun and the North point and the Moon and the North point. Results with a negative sign signify that the Moon is to the right (clockwise) of the Sun if one looks to the Sun — or alternatively expressed, that the Sun is to the left (anti-clockwise) of the Moon. Results with a positive sign signify that the Moon is to the left (anti-clockwise) of the Sun if one looks to the Sun — or alternatively expressed, that the Sun is to the right (clockwise) of the Moon.

z

o, s

Calculates the difference of the approximate geocentric, apparent elevation of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at civil midnight time.

A

o, s

Calculates the difference of the approximate geocentric, apparent azimuth of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at civil midnight time.

B

o

Calculates the difference of the approximate topocentric, apparent elevation of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Sun. See [Standard rise time of the Sun], page 168, for further details.

B

s

Calculates the difference of the approximate topocentric, apparent elevation of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Sun. See [Standard set time of the Sun], page 168, for further details.

C

o

Calculates the difference of the approximate topocentric, apparent azimuth of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Sun. See [Standard rise time of the Sun], page 168, for further details.

C

s

Calculates the difference of the approximate topocentric, apparent azimuth of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Sun. See [Standard set time of the Sun], page 168, for further details.

D

o

Calculates the difference of the approximate geocentric, apparent elevation of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Sun. See [Standard rise time of the Sun], page 168, for further details.

D

s

Calculates the difference of the approximate geocentric, apparent elevation of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Sun. See [Standard set time of the Sun], page 168, for further details.

E

o

Calculates the difference of the approximate geocentric, apparent azimuth of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Sun. See [Standard rise time of the Sun], page 168, for further details.

E

s

Calculates the difference of the approximate geocentric, apparent azimuth of Sun and Moon (delta), at which the Sun is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Sun. See [Standard set time of the Sun], page 168, for further details.

F

o, s

Calculates the difference of the approximate astronomical midnight times of Sun and Moon (delta), at which the Sun is used as the reference point, as time value in hours and minutes as it happen at astronomical midnight time of the Sun. Results with a negative sign signify that the astronomical midnight time of the Sun is earlier than the astronomical midnight time of the Moon; thus the solar midnight is before the lunar midnight. Results with a positive sign signify that the astronomical midnight time of the Sun is later than the astronomical midnight time of the Moon; thus the solar midnight is after the lunar midnight. See [Astronomical midnight time of the Sun], page 165, and [Astronomical midnight time of the Moon], page 188, for further details.

G

o, s

Calculates the difference of the approximate astronomical noon times of Sun and Moon (delta), at which the Sun is used as the reference point, as time value in hours and minutes as it happen at astronomical noon time of the Sun. Results with a negative sign signify that the astronomical noon time of the Sun is earlier than the astronomical noon time of the Moon; thus the solar noon is before the lunar noon. Results with a positive sign signify that the astronomical noon time of the Sun is later than the astronomical noon time of the Moon; thus the solar noon is after the lunar noon. See [Astronomical noon time of the Sun], page 166, and [Astronomical noon time of the Moon], page 189, for further details.

H

o

Calculates the difference of the approximate standard rise times of Sun and Moon (delta), at which the Sun is used as the reference point, as time value in hours and minutes as it happen at standard rise time of the Sun. Results with a negative sign signify that the standard rise time of the Sun is earlier than the standard rise time of the Moon; thus the sunrise is before the moonrise. Results with a positive sign signify that the standard rise time of the Sun is later than the standard rise time of the Moon; thus the sunrise is after the moonrise. See [Standard rise time of the Sun], page 168, and [Standard rise time of the Moon], page 192, for further details.

H

s

Calculates the difference of the approximate standard set times of Sun and Moon (delta), at which the Sun is used as the reference point, as time value in hours and minutes as it happen at standard set time of the Sun. Results with a negative sign signify that the standard set time of the Sun is earlier than the standard set time of the Moon; thus the sunset is before the moonset. Results with a positive sign signify that the standard set time of the Sun is later than the standard set time of the Moon; thus the sunset is after the moonset. See [Standard set time of the Sun], page 168, and [Standard set time of the Moon], page 192, for further details.

I

o, s

Calculates the approximate topocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at astronomical midnight time of the Sun (topocentric midnight height). See [Astronomical midnight time of the Sun], page 165, for further details.

J

o, s

Calculates the approximate topocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at astronomical midnight time of the Sun (topocentric midnight height). See [Astronomical noon time of the Sun], page 166, for further details.

K

o

Calculates the approximate topocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at standard rise time of the Sun (topocentric rise height). See [Standard rise time of the Sun], page 168, for further details.

K

s

Calculates the approximate topocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at standard set time of the Sun (topocentric set height). See [Standard set time of the Sun], page 168, for further details.

L

o

Calculates the approximate topocentric, apparent azimuth of the Sun in degrees and arcminutes as it happen at standard rise time of the Sun (topocentric rise azimuth). The horizontal angular distance between the topocentric rise azimuth and the East direction is also known as the topocentric morning width of the Sun. See [Standard rise time of the Sun], page 168, for further details.

L

s

Calculates the approximate topocentric, apparent azimuth of the Sun in degrees and arcminutes as it happen at standard set time of the Sun (topocentric set azimuth). The horizontal angular distance between the topocentric set azimuth and the West direction is also known as the topocentric evening width of the Sun. See [Standard set time of the Sun], page 168, for further details.

M

o, s

Calculates the approximate geocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at astronomical midnight time of the Sun (geocentric midnight height). See [Astronomical midnight time of the Sun], page 165, for further details.

N

o, s

Calculates the approximate geocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at astronomical midnight time of the Sun (geocentric midnight height). See [Astronomical noon time of the Sun], page 166, for further details.

O

o

Calculates the approximate geocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at standard rise time of the Sun (geocentric rise height). See [Standard rise time of the Sun], page 168, for further details.

O

s

Calculates the approximate geocentric, apparent elevation of the Sun in degrees and arcminutes as it happen at standard set time of the Sun (geocentric set height). See [Standard set time of the Sun], page 168, for further details.

P

o

Calculates the approximate geocentric, apparent azimuth of the Sun in degrees and arcminutes as it happen at standard rise time of the Sun (geocentric rise azimuth). The horizontal angular distance between the geocentric rise azimuth and the East direction is also known as the geocentric morning width of the Sun. See [Standard rise time of the Sun], page 168, for further details.

P

s

Calculates the approximate geocentric, apparent azimuth of the Sun in degrees and arcminutes as it happen at standard set time of the Sun (geocentric set azimuth). The horizontal angular distance between the geocentric set azimuth and the West direction is also known as the geocentric evening width of the Sun. See [Standard set time of the Sun], page 168, for further details.

Q

o

Calculates the approximate time when the length of the shadow cast by a vertical pole in the forenoon is equal the length of the pole. Nevertheless, the minimum length of the shadow is subtracted from the length of the shadow before comparing it with the length of the pole. See [Fixed dates option --adjust-value=argument], page 39, how to change the shadow length factor.

Q

s

Calculates the approximate time when the length of the shadow cast by a vertical pole in the afternoon is equal the length of the pole. Nevertheless, the minimum length of the shadow is subtracted from the length of the shadow before comparing it with the length of the pole. People of Islamic faith, and that the people holding the Shafi school of jurisprudence, normally pray for the third time on the day at this clocktime, or some minutes later. These people commonly use the term Asr for this prayer time. See [Astronomical noon time of the Sun], page 166, for more information. And note [Fixed dates option --adjust-value=argument], page 39, how to change the shadow length factor.

Q

u

Calculates the approximate period while the center of the Sun is above a reference altitude, at which the length of the shadow cast by a vertical pole is of single or shorter length than the pole itself.

Q

z

Calculates the approximate period while the center of the Sun is below a reference altitude, at which a vertical pole either casts no shadow anymore, or casts a shadow that is longer than the single length of the pole itself.

R

o

Calculates the approximate time when the length of the shadow cast by a vertical pole at forenoon is twice the length of the pole. Nevertheless, the minimum length of the shadow is subtracted from the length of the shadow before comparing it with the length of the pole. See [Fixed dates option --adjust-value=argument], page 39, how to change the shadow length factor.

R

s

Calculates the approximate time when the length of the shadow cast by a vertical pole in the afternoon is twice the length of the pole. Nevertheless, the minimum length of the shadow is subtracted from the length of the shadow before comparing it with the length of the pole. People of Islamic faith, and that the people holding the Hanafi school of jurisprudence, normally pray for the third time on the day at this clocktime, or some minutes later. These people commonly use the term Asr for this prayer time. See [Astronomical noon time of the Sun], page 166, for more information. And note [Fixed dates option --adjust-value=argument], page 39, how to change the shadow length factor.

R

u

Calculates the approximate period while the center of the Sun is above a reference altitude, at which the length of the shadow cast by a vertical pole is of double or shorter length than the pole itself.

R

z

Calculates the approximate period while the center of the Sun is below a reference altitude, at which a vertical pole either casts no shadow anymore, or casts a shadow that is longer than twice the length of the pole itself.

If no mode is given, Gcal automatically uses that mode, which is enabled by the mode character ‘5’. If a mode character is given that is not according to one of the ‘0’…‘9’, ‘a’…‘z’ and ‘A’…‘R’ characters, Gcal also automatically uses that mode, which is enabled by the mode character ‘5’.

Gcal represents the Sun oriented special texts depending on the selected mode using the following types and styles:

1. Unsigned decimal based rational number value Unsigned decimal based rational number values are represented using the n.n... format by default. A ‘*’ character that is directly given before some mode characters causes Gcal to represent the value for another quantity. For the mode characters, which a. cause the calculation of Earth/Sun or Earth/Moon distances, the calculated dis-

tance is represented in kilometers. b. cause the calculation of phase angles of the Moon, the calculated phase angle is

represented as a phase value in percents.

If definite events happen, Gcal displays special event oriented texts instead of using the previously described representations. See [Event texts of the Sun oriented special texts], page 181, where you can find the event oriented texts that are created for clocktime values, which are schematically and analogously used for the type of representation as it is described here.

2. Signed decimal based rational number value Signed decimal based rational number values are represented using the +|-n.n... format by default. A ‘*’ character that is directly given before a mode character causes Gcal not to represent such values using another style. If definite events happen, Gcal displays special event oriented texts instead of using the previously described representations. See [Event texts of the Sun oriented special texts], page 181, where you can find the event oriented texts that are created for clocktime values, which are schematically and analogously used for the type of representation as it is described here.

3. Clocktime value Clocktime values are represented in hours and minutes, and that in the hh:mm 24-hour format by default. A ‘*’ character that is directly given before a mode character causes Gcal to represent the clocktime value using the 12-hour format, thus to provide it with a time suffix. See [Actual localtime in hh:mm format %t[argument] special text], page 159, for more details about the above mentioned time value template. If definite events happen, Gcal displays special event oriented texts instead of using the previously described representations:

• In case Gcal is unable to compute the approximate clocktime by reason of a misspecified argument, a ??:?? text will be created instead of the clocktime text.

• In case Gcal calculates an approximate clocktime that refers to a previous or next day, or Gcal is unable to calculate a clocktime because the event calculated does not occur during this day, a **:** text will be created instead of the clocktime text.

• In case Gcal is unable to calculate an approximate clocktime because the reference point of the astronomical object, thus either the center or the upper limb of the Sun’s or the Moon’s disk, is always above the respective reference altitude, a ++:++ text will be created instead of the clocktime text.

• In case Gcal is unable to calculate an approximate clocktime because the reference point of the astronomical object, thus either the center or the upper limb of the Sun’s or the Moon’s disk, is always below the respective reference altitude, a --:-- text will be created instead of the clocktime text.

4. Unsigned time value Unsigned time values, which mostly denote a period or interval of time, are represented in hours and minutes using the hhhmm' format by default. A ‘*’ character that is directly given before a mode character causes Gcal to represent the time value using another style, and that in decimal hours, i.e. in the hh.h... format. If definite events happen, Gcal displays special event oriented texts instead of using the previously described representations. See [Event texts of the Sun oriented special texts], page 181, where you can find the event oriented texts that are created for clocktime values, which are schematically and analogously used for the type of representation as it is described here.

5. Signed time value Signed time values, which mostly denote a period or interval of time, are represented in hours and minutes using the +|-hhhmm' format by default. A ‘*’ character that is directly given before a mode character causes Gcal to represent the time value using another style, and that in decimal hours, i.e. in the +|-hh.h... format. If definite events happen, Gcal displays special event oriented texts instead of using the previously described representations. See [Event texts of the Sun oriented special texts], page 181, where you can find the event oriented texts that are created for clocktime values, which are schematically and analogously used for the type of representation as it is described here.

6. Unsigned angular value Unsigned angular values are represented in degrees and arcminutes using the ddddmm' format by default. A ‘*’ character that is directly given before a mode character causes Gcal to represent the angular value using another style, and that in decimal degrees, i.e. in the ddd.d... format. If definite events happen, Gcal displays special event oriented texts instead of using the previously described representations. See [Event texts of the Sun oriented special texts], page 181, where you can find the event oriented texts that are created for clocktime values, which are schematically and analogously used for the type of representation as it is described here.

7. Signed angular value Signed angular values are represented in degrees and arcminutes using the +|-ddddmm' format by default. A ‘*’ character that is directly given before a mode character causes Gcal to represent the angular value using another style, and that in decimal degrees, i.e. in the +|-ddd.d... format. If definite events happen, Gcal displays special event oriented texts instead of using the previously described representations. See [Event texts of the Sun oriented special texts], page 181, where you can find the event oriented texts that are created for clocktime values, which are schematically and analogously used for the type of representation as it is described here.

After the optional style and mode characters, the latitude and longitude of the geographic co-ordinates follows, for which the calculations must be made. They must be conform the ISO-6709:1983 standard representation of latitude and longitude for geographic point locations, so that the co-ordinate has to be declared like this:

• Latitude (+ for North and on the equator, - for South of the equator.)

    +|-dd[.dd]

Degrees (2 digits), and optionally decimal degrees

    +|-ddmm[.mm]

Degrees (2 digits), arcminutes (2 digits), and optionally decimal minutes

    +|-ddmmss[.ss]

Degrees (2 digits), arcminutes (2 digits), arcseconds (2 digits), and optionally decimal seconds

• Longitude (+ for East and on the prime meridian (Greenwich), - for West of Greenwich and up to the 180th meridian.)

    +|-ddd[.dd]

Degrees (3 digits), and optionally decimal degrees

    +|-dddmm[.mm]

Degrees (3 digits), arcminutes (2 digits), and optionally decimal minutes

    +|-dddmmss[.ss]

Degrees (3 digits), arcminutes (2 digits), arcseconds (2 digits), and optionally decimal seconds

• Optionally, the relative height in meters above the sea level (+ for heights above and on the sea level, - for heights below the sea level.)

[+|-n[n[n[n]]]] Height in meters (integer number) in range -9999…+9999 (1…4 digits)

All components of the co-ordinates must have leadings zeroes in case they have less digits than the templates shown above. Declared decimal seconds are not respected by Gcal. Heights which have a negative sign remain unrespected if Gcal determinates Sun and Moon data and times, respectively. In such a case, Gcal always uses the height +0. Latitude and longitude co-ordinates, and the height of the observer’s location are connected without any separating characters, like ‘+40-075+61’, ‘+401213.1-0750015.1’ or ‘+40.20361-075.00417+0061’. See the pertinent literature for more details.

A time value [+|-]mmmm|hh:[mm], which is separated by a ‘,’ character, may trail the co-ordinate. Such a time value informs Gcal, about how many minutes mmmm respectively hours hh and minutes mm the geographic location is displaced from Universal time (UTC/GMT). This time displacement value defines the timezone, which is actually valid for this location. If summer- and wintertimes are respected for the location, you should include that change in time into the timezone value for the period in which the summertime is valid, by which the clock is put on during the summertime period — such a change is either subtracted from the timezone value for locations West of the prime meridian (Greenwich), or it is added for locations East of the prime meridian, because Gcal is actually unable to perform such operations automatically! See [Actual local time in hh:mm format %t[argument] special text], page 159, for more details about the above mentioned time value template. If no time displacement value is specified for a given co-ordinate, Gcal assumes a time displacement value of 0, which is equal to the actual Universal time (UTC/GMT).

The following table informs you about which type of representation is caused by a mode. The previously defined numbering scheme, as it has been used for the introduction of the types of representations, is used as key value in the column that holds the type of representation. The table also contains a column that shows whether a mode enables dynamical values, i.e. values that are depending on the respective clocktime (if you use the --time-offset=argument option, you can change the respective clocktime that is used for calculating such values). In a next table column, it is listed whether the given co-ordinate of the location influences the determination of a value, and the last column of the table gives you the information whether a given timezone value affects the values determination:

Mode  Representation Type  Dynamical  Co-ordinate  Timezone
0  3  No  Yes  Yes
1  3  No  Yes  Yes
2  3 or 4  No  Yes  Yes
3  3 or 4  No  Yes  Yes
4  3 or 4  No  Yes  Yes
5  3 or 4  No  Yes  Yes
6  3 or 4  No  Yes  Yes
7  3 or 4  No  Yes  Yes
8  3 or 4  No  Yes  Yes
9  3 or 4  No  Yes  Yes
a  7  Yes  Yes  Yes
b  6  Yes  Yes  Yes
c  7  Yes  Yes  Yes
d  6  Yes  Yes  Yes
e  4  Yes  Yes  Yes
f  1 or 1a  Yes  Yes  Yes
g  6  Yes  Yes  Yes
h  6  Yes  Yes  Yes
i  6  Yes  No  No
j  7  Yes  Yes  Yes
k  6  Yes  Yes  Yes
l  7  Yes  Yes  Yes
m  6  Yes  Yes  Yes
n  4  Yes  Yes  Yes
o  1 or 1a  Yes  Yes  Yes
p  6  Yes  Yes  Yes
q  6  Yes  Yes  Yes
r  2  Yes  No  No
s  3  Yes  Yes  Yes
t  3  Yes  No  No
u  1  Yes  No  No
v  1  Yes  No  No
w  5  Yes  No  Yes

x  7  Yes  Yes  Yes
y  7  Yes  Yes  Yes
z  7  Yes  Yes  Yes
A  7  Yes  Yes  Yes
B  7  No  Yes  Yes
C  7  No  Yes  Yes
D  7  No  Yes  Yes
E  7  No  Yes  Yes
F  5  No  Yes  Yes
G  5  No  Yes  Yes
H  5  No  Yes  Yes
I  7  No  Yes  Yes
J  7  No  Yes  Yes
K  7  No  Yes  Yes
L  6  No  Yes  Yes
M  7  No  Yes  Yes
N  7  No  Yes  Yes
O  7  No  Yes  Yes
P  6  No  Yes  Yes
Q  3 or 4  No  Yes  Yes
R  3 or 4  No  Yes  Yes

And now some examples to these special texts:

     The text ‘Sunrise at %o+5158+00738,120  in MS, BRD’ will be expanded to

↦ ‘Sunrise at 05:16 in MS, BRD’, in case the actual system date is the 1st June 1998.

     The text ‘Sunset at %s*5+5158+00738,120  in MS, BRD’ will be expanded to

↦ ‘Sunset at 09:39pm in MS, BRD’, in case the actual system date is the 1st June 1998.

     The text ‘Sun visible %u5+5158+00738,120  in MS, BRD’ will be expanded to

↦ ‘Sun visible 16h24' in MS, BRD’, in case the actual system date is the 1st June 1998.

     The text ‘Sun non-visible %z*+5158+00738,120  in MS, BRD’ will be

expanded to

↦ ‘Sun non-visible 7.607 in MS, BRD’, in case the actual system date is the 1st June 1998.

     The text ‘Sun azimuth 0 o'clock=%s*a+5158+00738,120  in MS, BRD’ will be

expanded to

     ↦ ‘Sun azimuth 0 o'clock=339d16' in MS, BRD’, in case the actual system

date is the 1st June 1998.

     The text ‘Equation of time %ot+00+000=%o*w+00+000,120  BRD’ will be ex-

panded to

     ↦ ‘Equation of time +16h00'=+00h02'13.201" BRD’, in case you call Gcal

with the --time-offset=16: and --precise options and the actual system date is the 1st June 1998.

 The text ‘Julian date at %ot+00+000 =%ou+00+000’ will be expanded to
 ↦ ‘Julian date at +10h15'=2450965.927’, in case you call Gcal with the

--time-offset=10:15 option and the actual system date is the 1st June 1998.

Here is a list that reports about the used reference systems in a short manner, describes other aspects that are unmentioned now, and informs about the lacks and limitations that are existing for the Sun oriented special texts:

• All astronomical calculations are based on methods that use —astronomically seen— low-precision elements, which of course is shown in the precision of the results that are caused by them. Nevertheless in most cases, these results can be treated as more as sufficiently precise if one measures their practical use.

• All azimuth and elevation values are always referring to the center of the respective astronomical object, except the case, it is stated otherwise.

• The base point of azimuth values, right ascension values and other ecliptic co-ordinate values is always measured clockwise relative to the geographical, true North (not the magnetic North as shown by a compass), where angular values for the North direction are both denoted as 0 degree and 360 degree.

• All topocentrically based azimuth and elevation values are always corrected by the amount of atmospheric refraction. Such a kind of correction is always done if the center of the respective astronomical object is above 2 degrees below the horizon.

• Generally, the rise and set times of the Sun cannot be precisely predicted, because depending on unpredictable atmospheric conditions, i.e. the local weather conditions as they appear during these times change the amount of atmospheric refraction at the horizon, so that the precalculated rise and set times may throughout be in error by a minute or more.

• The accuracy of rise and set computations decreases again at high latitudes, i.e. within in the solar Arctic Circle. There, small variations in atmospheric refraction can change the time of sunrise or sunset by many minutes, since the Sun intersects the horizon at a very shallow angle. For the same reason, at high latitudes, the effects of observer height and local topography are magnified and can substantially change the times of the phenomena actually observed, or even whether the phenomena are observed to occur at all.

• The short explanations that are given to the different twilight phases all assume the model of ideal atmospheric conditions by default. It is obvious that these explanations all become meaningless if other conditions occur in reality, like a cloudy sky or a badly polluted atmosphere.

Please also note the following references:

• See [Fixed dates option --precise], page 43, how to obtain a more precise representation of the values which are cause by these special texts.

• And note [Calendar option --time-offset=argument], page 33, how to change the clocktime, for which the dynamical, i.e. depending on the respective clocktime, Sun oriented special texts are calculated.

• Likewise, see [Fixed dates option --cycle-end=argument], page 41, how the dynamical, i.e. depending on the respective clocktime, Sun oriented special texts are calculated for a series of clocktimes for the current day.

• And also note [Fixed dates option --adjust-value=argument], page 39, how to change the reference altitude that is used to calculate the rise and set times, as well as the shadow length factor.

• Furthermore, see [Fixed dates option --atmosphere=air-pressure[,temperature]], page 40, how to change the base data of the atmosphere that influences the calculation of the Sun oriented special texts.

• And last, note [Fixed dates option --limit], page 40, how to limit the rise and set times of the Sun to the actual day.

All Sun oriented special texts must always be trailed by a whitespace character which is removed in output!

G.2.2.6 Moon data %[format]?argument special texts

%[format]([*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate time of moonrise by default,

%[format])[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate time of moonset by default,

%[format][[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate period of visibility of the Moon (lunar day length) by default,

%[format]][*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] references the

approximate period of non-visibility of the Moon (lunar night length) by default.

All these special texts can be used for at pleasure any geographic point location, i.e. it is possible to determine different astronomical values for any location on the globe, and that for at pleasure any clocktime with a resolution of a single minute within the period of the years AD 1 until AD 9999, that is respected by Gcal.

The selection which value has to be calculated by these special texts is done by specifying the mode part of the preceding argument. Actually, exactly 61 different modes can be used that are represented by the ‘0’…‘9’, ‘a’…‘z’ and ‘A’…‘Y’ characters, and which create different kind of results that are depending on the special text used. First of all, here is a table in which all usable modes are described and explained sufficiently. You can also see from this table, which Moon oriented special text or texts are corresponding to which mode, i.e. cause the determination of an astronomical value as it is described in the table:

Mode  Special text Description

0

          (, )

Calculates the approximate midnight time of the Moon. The astronomical midnight time of the Moon is at that clocktime, when the Moon holds an azimuth (horizontal angular distance between the vertical circle, that passes the Moon, and the North point) of either precisely 0 degrees of precisely 180 degrees, which depends on the season and the geographical location. At that clocktime, the Moon is close its lowest culmination point, i.e. close the lowest point below or above the horizontal plane the Moon transits during this day. Nevertheless, there is exactly one day during a synodic month (or lunation) —i.e. the mean time between two consecutive conjunctions (or New Moon phases)— at which no lunar midnight happens, because the Moon revolves the Earth within 24 hours and 50 minutes on the average — which also means, that the Moon rises on the average 50 minutes later each day.

1

(, )

Calculates the approximate noon time of the Moon. The astronomical noon time of the Moon is at that clocktime, when the Moon holds an azimuth of either precisely 180 degrees of precisely 0 degrees, which depends on the season and the geographical location. At that clocktime, the Moon is close its highest culmination point, i.e. close the highest point above or below the horizontal plane the Moon transits during this day. Nevertheless, there is exactly one day during a synodic month at which no lunar noon happens.

2

(

Calculates the approximate time when the center of the Moon passes a reference altitude which is between about 54 and 61 arcminutes above a mathematical-geocentric horizon before lunar noon time; thus rising. A mathematical horizon is a purely geometrically-built horizon which disregards the phenomenon of refraction as it arises in reality by the influence of the Earth’s atmosphere. A geocentrical horizon is the horizontal plane that passes through the Earth’s center, orthogonal to the observer’s local vertical. In the further context, the shorter term mathematical horizon is used which actually means the mathematicalgeocentric horizon. The above mentioned reference altitude is computed from the value of the Moon’s parallax as it appear at that clocktime. Nevertheless, there is exactly one day during a synodic month at which no such moonrise happens.

2

)

Calculates the approximate time when the center of the Moon passes a reference altitude which is between about 54 and 61 arcminutes above a mathematical horizon after lunar noon time; thus setting. The above mentioned reference altitude is computed from the value of the Moon’s parallax as it appear at that clocktime. Nevertheless, there is exactly one day during a synodic month at which no such moonset happens.

2

[

Calculates the approximate period while the center of the Moon is above a reference altitude which is between about 54 and 61 arcminutes above a mathematical horizon; thus is visible.

2

]

Calculates the approximate period while the center of the Moon is below a reference altitude which is between about 54 and 61 arcminutes above a mathematical horizon; thus is non-visible.

3

(

Calculates the approximate time when the upper limb of the Moon passes a reference altitude which is between about 54 and 61 arcminutes above a mathematical horizon before lunar noon time; thus rising. The above mentioned reference altitude is computed from the respective values of the Moon’s semidiameter and Moon’s parallax as they appear at that clocktime. If the reference altitude that is referring to the Moon’s upper limb is converted to a reference altitude that is referring to the Moon’s center, this results in a value which is between about 39 and 44 arcminutes above the geocentric horizon. Nevertheless, there is exactly one day during a synodic month at which no such moonrise happens.

3

)

Calculates the approximate time when the upper limb of the Moon passes a reference altitude which is between about 54 and 61 arcminutes above a mathematical horizon after lunar noon time; thus setting. The above mentioned reference altitude is computed from the respective values of the Moon’s semidiameter and Moon’s parallax as they appear at that clocktime. Nevertheless, there is exactly one day during a synodic month at which no such moonset happens.

3

[

Calculates the approximate period while the upper limb of the Moon is above a reference altitude which is between about 54 and 61 arcminutes above a mathematical horizon; thus is visible.

3

]

Calculates the approximate period while the upper limb of the Moon is below a reference altitude which is between about 54 and 61 arcminutes above a mathematical horizon; thus is nonvisible.

4

(

Calculates the approximate time when the center of the Moon passes a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon before lunar noon time; thus rising. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the

--atmosphere option. [Fixed dates option --atmosphere=air-

pressure[,temperature]], page 40, how to change the base data of the atmosphere, so that the atmospheric conditions as defined by it are used to calculate the amount of refraction. The above mentioned reference altitude is computed from the respective values of the Moon’s parallax and (standard) refraction as they appear at that clocktime. Nevertheless, there is exactly one day during a synodic month at which no such moonrise happens.

4

)

Calculates the approximate time when the center of the Moon passes a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon after lunar noon time; thus setting. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the --atmosphere option. The above mentioned reference altitude is computed from the respective values of the Moon’s parallax and (standard) refraction as they appear at that clocktime. Nevertheless, there is exactly one day during a synodic month at which no such moonset happens.

4

[

Calculates the approximate period while the center of the Moon is above a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon; thus is visible.

4

]

Calculates the approximate period while the center of the Moon is below a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon; thus is non-visible.

5

(

Calculates the approximate time when the upper limb of the Moon passes a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon before lunar noon time; thus rising. This kind of rise time calculation is done according to the standard calculation method as it is commonly used internationally. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the

--atmosphere option. [Fixed dates option --atmosphere=air-

pressure[,temperature]], page 40, how to change the base data of the atmosphere, so that the atmospheric conditions as defined by it are used to calculate the amount of refraction. The above mentioned reference altitude is computed from the respective values of the Moon’s semidiameter, Moon’s parallax and (standard) refraction as they appear at that clocktime. If the reference altitude that is referring to the Moon’s upper limb is converted to a reference altitude that is referring to the Moon’s center, this results in a value which is between about 5 and 10 arcminutes above the geocentric horizon. Nevertheless, there is exactly one day during a synodic month at which no such moonrise happens.

5

)

Calculates the approximate time at which the upper limb of the Moon passes a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon after lunar noon time; thus setting. This kind of set time calculation is done according to the standard calculation method as it is commonly used internationally. The phenomenon of refraction is already respected in this as it arises in reality by the influence of the Earth’s atmosphere, and that with the standard value of 34 arcminutes, which can indirectly be changed by using the --atmosphere option. The above mentioned reference altitude is computed from the respective values of the Moon’s semidiameter, Moon’s parallax and (standard) refraction as they appear at that clocktime. Nevertheless, there is exactly one day during a synodic month at which no such moonset happens.

5

[

Calculates the approximate period while the upper limb of the Moon is above a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon; thus is visible. This kind of visibility period calculation is done according to the standard calculation method as it is commonly used internationally.

5

]

Calculates the approximate period while the upper limb of the Moon is below a reference altitude which is between about 20 and 27 arcminutes above the geocentric horizon; thus is nonvisible. This kind of non-visibility period calculation is done according to the standard calculation method as it is commonly used internationally.

6

(, )

Calculates the approximate topocentric, apparent horizontal parallax of the Moon in degrees and arcminutes as it happen at civil midnight time. The Moon’s parallax states the diameter of the Earth as it is seen from the surface of the Moon. Observations of celestial objects that are done from the surface of the Earth yield in topocentrically based data. The locations of the celestial bodies are often at another place if the data is topocentrically determined instead of determine it geocentrically, i.e. at the fictitious center of the Earth. This is mainly caused by the refraction, which raises a celestial body to another location as it is been in reality. Because the terrestrial globe flattens towards the pole caps and therefore cannot be taken as an ideally shaped sphere, the individual Earth radius between the observer’s location and the center of the Earth also affects the computation of topocentrically based data.

7

(, )

Calculates the approximate topocentric, apparent semidiameter of the Moon in degrees and arcminutes as it happen at civil midnight time.

8

(, )

Calculates the approximate topocentric, apparent brightness of the Moon in magnitude units as it happen at civil midnight time. The magnitude (Latin term magnitudo, abbreviated m) is used to define the brightness of a star, and is a non-metrical value. The difference between two consecutive magnitudes is 1 to 2.512. Therefore, a star with the brightness of 1m is 2.512 times brighter than a star of 2m. A negative magnitude denotes a very bright star, for example almost -27m for the Sun, whereas the hardly visible planet Pluto has a magnitude of a bit more than +14m. The Full Moon has a visual brightness of about -12m.55.

9

(, )

Calculates the approximate topocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at civil midnight time.

a

(, )

Calculates the approximate topocentric, apparent elevation of the Moon, thus the vertical angular distance between the Moon’s center and the horizon, in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Moon’s center is below the horizon at the moment, and results with a positive sign mean that the momentary center of the Moon is above the horizon.

b

(, )

Calculates the approximate topocentric, apparent azimuth of the Moon in degrees and arcminutes as it happen at civil midnight time.

c

(, )

Calculates the approximate topocentric, apparent declination of the Moon, thus the vertical angular distance between the Moon’s center and the celestial equator, in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Moon’s center is below the celestial equator at the moment, and results with a positive sign mean that the momentary center of the Moon is above the celestial equator.

d

(, )

Calculates the approximate topocentric, apparent ecliptic longitude of the Moon, thus the horizontal angular distance between the Moon’s center and the vernal equinox point on the ecliptic (the zodiacal line or Sun’s orbit), in degrees and arcminutes as it happen at civil midnight time.

e

(, )

Calculates the approximate topocentric, apparent ecliptic latitude of the Moon, thus the vertical angular distance between the Moon’s center and the ecliptic (zodiacal line/Sun’s orbit), in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Moon’s center is North of the ecliptic at the moment, and results with a positive sign mean that the momentary center of the Moon is South of the ecliptic.

f

(, )

Calculates the approximate topocentric, apparent right ascension of the Moon, thus the horizontal angular distance between the Moon’s center and the hour circle that passes through the vernal equinox point on the ecliptic, as time value in hours and minutes as it happen at civil midnight time.

g

(, )

Calculates the approximate topocentric, apparent distance of the Moon from the Earth in mean Earth equator radii as it happen at civil midnight time. The mean radius of Earth at the equator is about 6,378.137 kilometer.

h

(, )

Calculates the approximate topocentric, apparent elongation of the Moon, thus the horizontal angular distance between the Moon’s center and the Sun’s center, in degrees and arcminutes as it happen at civil midnight time.

i

(, )

Calculates the approximate refraction of the Earth’s atmosphere in degrees and arcminutes as it happen at civil midnight time.

j

(, )

Calculates the approximate geocentric, apparent horizontal parallax of the Moon in degrees and arcminutes as it happen at civil midnight time.

k

(, )

Calculates the approximate geocentric, apparent semidiameter of the Moon in degrees and arcminutes as it happen at civil midnight time.

l

(, )

Calculates the approximate geocentric, apparent brightness of the Moon in magnitude units as it happen at civil midnight time.

m

(, )

Calculates the approximate geocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at civil midnight time.

n

(, )

Calculates the approximate geocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Moon’s center is below the horizon at the moment, and results with a positive sign mean that the momentary center of the Moon is above the horizon.

o

(, )

Calculates the approximate geocentric, apparent azimuth of the Moon in degrees and arcminutes as it happen at civil midnight time.

p

(, )

Calculates the approximate geocentric, apparent declination of the Moon in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Moon’s center is below the celestial equator at the moment, and results with a positive sign mean that the momentary center of the Moon is above the celestial equator.

q

(, )

Calculates the approximate geocentric, apparent ecliptic longitude of the Moon in degrees and arcminutes as it happen at civil midnight time.

r

(, )

Calculates the approximate geocentric, apparent ecliptic latitude of the Moon in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the Moon’s center is North of the ecliptic at the moment, and results with a positive sign mean that the momentary center of the Moon is South of the ecliptic.

s

(, )

Calculates the approximate geocentric, apparent right ascension of the Moon as time value in hours and minutes as it happen at civil midnight time.

t

(, )

Calculates the approximate geocentric, apparent distance of the Moon from the Earth in mean Earth equator radii as it happen at civil midnight time.

u

(, )

Calculates the approximate geocentric, apparent elongation of the Moon in degrees and arcminutes as it happen at civil midnight time.

v

(, )

Calculates the approximate delta-t in seconds as it happen at civil midnight time. Delta-t is the difference between the Terrestrial Dynamical time (abbreviated by TDT), that was formerly known as Ephemeris time (abbreviated by ET), and the Universal time (UT). Thus, ‘delta-t ≡ TDT - UT’.

w

(, )

Calculates the approximate, apparent location oriented sidereal time (local sidereal time (LAST), also known as local star time) in hours and minutes as it happen at civil midnight time. A star day is the period between two consecutive upper culminations of the vernal equinox point on the ecliptic in the meridian of the observer’s location. Therefore, the local star time is the momentary period, which is past between the last upper culmination of the vernal equinox point in the meridian of the observer’s location (the momentary hour angle of the vernal equinox point), thus the right ascension of the stars in the observer’s meridian at the moment.

x

(, )

Outputs the base time as time value in hours and minutes, for which the dynamical, i.e. depending on the respective clocktime, astronomical data and times of the Moon are calculated. Without a given --time-offset=argument option, the astronomical data and times of the Moon are always calculated for 0 o’clock Universal time (UTC/GMT). See [Calendar option --time-offset=argument], page 33, for further details.

y

(, )

Calculates the approximate Julian date in days as it happen at civil midnight time. See Section G.2.2.3 [Julian day number], page 162, for further information about the Julian date.

z

(, )

Calculates the approximate Julian Ephemeris date, thus a Julian date that is corrected by delta-t, in days as it happen at civil midnight time.

A

(, )

Calculates the difference of the approximate topocentric, apparent elevation of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at civil midnight time. Results with a negative sign signify that the momentary center of the Moon is at an elevation that is below the momentary elevation of the Sun’s center; thus the Moon is lower than the Sun. Results with a positive sign signify that the momentary center of the Moon is at an elevation that is above the momentary elevation of the Sun’s center; thus the Moon is higher than the Sun.

B

(, )

Calculates the difference of the approximate topocentric, apparent azimuth of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at civil midnight time. The result specifies the horizontal angular distance, by which the momentary center of the Moon is distant from the momentary Sun’s center, and that measured at the vertical circles that pass the Moon and the North point and the Sun and the North point. Results with a negative sign signify that the Sun is to the right (clockwise) of the Moon if one looks to the Moon — or alternatively expressed, that the Moon is to the left (anti-clockwise) of the Sun. Results with a positive sign signify that the Sun is to the left (anti-clockwise) of the Moon if one looks to the Moon — or alternatively expressed, that the Moon is to the right (clockwise) of the Sun.

C

(, )

Calculates the difference of the approximate geocentric, apparent elevation of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at civil midnight time.

D

(, )

Calculates the difference of the approximate geocentric, apparent azimuth of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at civil midnight time.

E

(

Calculates the difference of the approximate topocentric, apparent elevation of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Moon. See [Standard rise time of the Moon], page 192, for further details.

E

)

Calculates the difference of the approximate topocentric, apparent elevation of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Moon. See [Standard set time of the Moon], page 192, for further details.

F

(

Calculates the difference of the approximate topocentric, apparent azimuth of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Moon. See [Standard rise time of the Moon], page 192, for further details.

F

)

Calculates the difference of the approximate topocentric, apparent azimuth of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Moon. See [Standard set time of the Moon], page 192, for further details.

G

(

Calculates the difference of the approximate geocentric, apparent elevation of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Moon. See [Standard rise time of the Moon], page 192, for further details.

G

)

Calculates the difference of the approximate geocentric, apparent elevation of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Moon. See [Standard set time of the Moon], page 192, for further details.

H

(

Calculates the difference of the approximate geocentric, apparent azimuth of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard rise time of the Moon. See [Standard rise time of the Moon], page 192, for further details.

H

)

Calculates the difference of the approximate geocentric, apparent azimuth of Moon and Sun (delta), at which the Moon is used as the reference point, in degrees and arcminutes as it happen at standard set time of the Moon. See [Standard set time of the Moon], page 192, for further details.

I

(, )

Calculates the difference of the approximate astronomical midnight times of Moon and Sun (delta), at which the Moon is used as the reference point, as time value in hours and minutes as it happen at astronomical midnight time of the Moon. Results with a negative sign signify that the astronomical midnight time of the Moon is earlier than the astronomical midnight time of the Sun; thus the lunar midnight is before the solar midnight. Results with a positive sign signify that the astronomical midnight time of the Moon is later than the astronomical midnight time of the Sun; thus the lunar midnight is after the solar midnight. See [Astronomical midnight time of the Moon], page 188, and [Astronomical midnight time of the Sun], page 165, for further details.

J

(, )

Calculates the difference of the approximate astronomical noon times of Moon and Sun (delta), at which the Moon is used as the reference point, as time value in hours and minutes as it happen at astronomical noon time of the Moon. Results with a negative sign signify that the astronomical noon time of the Moon is earlier than the astronomical noon time of the Sun; thus the lunar noon is before the solar noon. Results with a positive sign signify that the astronomical noon time of the Moon is later than the astronomical noon time of the Sun; thus the lunar noon is after the solar noon. See [Astronomical noon time of the Moon], page 189, and [Astronomical noon time of the Sun], page 166, for further details.

K

(

Calculates the difference of the approximate standard rise times of Moon and Sun (delta), at which the Moon is used as the reference point, as time value in hours and minutes as it happen at standard rise time of the Moon. Results with a negative sign signify that the standard rise time of the Moon is earlier than the standard rise time of the Sun; thus the moonrise is before the sunrise. Results with a positive sign signify that the standard rise time of the Moon is later than the standard rise time of the Sun; thus the moonrise is after the sunrise. See [Standard rise time of the Moon], page 192, and [Standard rise time of the Sun], page 168, for further details.

K

)

Calculates the difference of the approximate standard set times of Moon and Sun (delta), at which the Moon is used as the reference point, as time value in hours and minutes as it happen at standard set time of the Moon. Results with a negative sign signify that the standard set time of the Moon is earlier than the standard set time of the Sun; thus the moonset is before the sunset. Results with a positive sign signify that the standard set time of the Moon is later than the standard set time of the Sun; thus the moonset is after the sunset. See [Standard set time of the Moon], page 192, and [Standard set time of the Sun], page 168, for further details.

L

(, )

Calculates the approximate topocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at astronomical midnight time of the Moon (topocentric midnight height). See [Astronomical midnight time of the Moon], page 188, for further details.

M

(, )

Calculates the approximate topocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at astronomical midnight time of the Moon (topocentric midnight phase angle). See [Astronomical midnight time of the Moon], page 188, for further details.

N

(, )

Calculates the approximate topocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at astronomical noon time of the Moon (topocentric noon height). See [Astronomical noon time of the Moon], page 189, for further details.

O

(, )

Calculates the approximate topocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at astronomical noon time of the Moon (topocentric noon phase angle). See [Astronomical noon time of the Moon], page 189, for further details.

P

(

Calculates the approximate topocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at standard rise time of the Moon (topocentric rise height). See [Standard rise time of the Moon], page 192, for further details.

P

)

Calculates the approximate topocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at standard set time of the Moon (topocentric set height). See [Standard set time of the Moon], page 192, for further details.

Q

(

Calculates the approximate topocentric, apparent azimuth of the Moon in degrees and arcminutes as it happen at standard rise time of the Moon (topocentric rise azimuth). The horizontal angular distance between the topocentric rise azimuth and the East direction is also known as the topocentric rise width of the Moon. See [Standard rise time of the Moon], page 192, for further details.

Q

)

Calculates the approximate topocentric, apparent azimuth of the Moon in degrees and arcminutes as it happen at standard set time of the Moon (topocentric set azimuth). The horizontal angular distance between the topocentric set azimuth and the West direction is also known as the topocentric set width of the Moon. See [Standard set time of the Moon], page 192, for further details.

R

(

Calculates the approximate topocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at standard rise time of the Moon (topocentric rise phase angle). See [Standard rise time of the Moon], page 192, for further details.

R

)

Calculates the approximate topocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at standard set time of the Moon (topocentric set phase angle). See [Standard set time of the Moon], page 192, for further details.

S

(, )

Calculates the approximate geocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at astronomical midnight time of the Moon (geocentric midnight height). See [Astronomical midnight time of the Moon], page 188, for further details.

T

(, )

Calculates the approximate geocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at astronomical midnight time of the Moon (geocentric midnight phase angle). See [Astronomical midnight time of the Moon], page 188, for further details.

U

(, )

Calculates the approximate geocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at astronomical noon time of the Moon (geocentric noon height). See [Astronomical noon time of the Moon], page 189, for further details.

V

(, )

Calculates the approximate geocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at astronomical noon time of the Moon (geocentric noon phase angle). See [Astronomical noon time of the Moon], page 189, for further details.

W

(

Calculates the approximate geocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at standard rise time of the Moon (geocentric rise height). See [Standard rise time of the Moon], page 192, for further details.

W

)

Calculates the approximate geocentric, apparent elevation of the Moon in degrees and arcminutes as it happen at standard set time of the Moon (geocentric set height). See [Standard set time of the Moon], page 192, for further details.

X

(

Calculates the approximate geocentric, apparent azimuth of the Moon in degrees and arcminutes as it happen at standard rise time of the Moon (geocentric rise azimuth). The horizontal angular distance between the geocentric rise azimuth and the East direction is also known as the geocentric rise width of the Moon. See [Standard rise time of the Moon], page 192, for further details.

X

)

Calculates the approximate geocentric, apparent azimuth of the Moon in degrees and arcminutes as it happen at standard set time of the Moon (geocentric set azimuth). The horizontal angular distance between the geocentric set azimuth and the West direction is also known as the geocentric set width of the Moon. See [Standard set time of the Moon], page 192, for further details.

Y

(

Calculates the approximate geocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at standard rise time of the Moon (geocentric rise phase angle). See [Standard rise time of the Moon], page 192, for further details.

Y

)

Calculates the approximate geocentric, apparent phase angle of the Moon in range 0.0…1.0 as it happen at standard set time of the Moon (geocentric set phase angle). See [Standard set time of the Moon], page 192, for further details.

If no mode is given, Gcal automatically uses that mode, which is enabled by the mode character ‘5’. If a mode character is given that is not according to one of the ‘0’…‘9’, ‘a’…‘z’ and ‘A’…‘Y’ characters, Gcal also automatically uses that mode, which is enabled by the mode character ‘5’.

Depending on the selected mode, Gcal represents the Moon oriented special texts using the same types and styles as they are used by the Sun oriented special texts, these are analogously valid! See [x-Sun-representation], page 180, for the detailed description of the different types of representation used by the Sun oriented special texts, which are likewise valid for the Moon oriented special texts.

The argument the Moon oriented special texts must have is exactly equivalent the argument the Sun oriented special texts must have! See [Arguments of the Sun oriented special texts], page 182, for the detailed description of the components of the argument which also has to be given to the Moon oriented special texts.

The following table informs you about which type of representation is caused by a mode. The previously defined numbering scheme, as it has been used for the introduction of the types of representation, is used as key value in the column that holds the type of representation. The table also contains a column that shows whether a mode enables dynamical values, i.e. values that are depending on the respective clocktime (if you use the --time-offset=argument option, you can change the respective clocktime that is used for calculating such values). In a next table column, it is listed whether the given co-ordinate of the location influences the determination of a value, and the last column of the table gives you the information whether a given timezone value affects the values determination:

Mode  Representation Type  Dynamical  Co-ordinate  Timezone
0  3  No  Yes  Yes
1  3  No  Yes  Yes
2  3 or 4  No  Yes  Yes
3  3 or 4  No  Yes  Yes
4  3 or 4  No  Yes  Yes
5  3 or 4  No  Yes  Yes
6  6  Yes  Yes  Yes
7  6  Yes  Yes  Yes
8  2  Yes  Yes  Yes
9  1 or 1b  Yes  Yes  Yes
a  7  Yes  Yes  Yes
b  6  Yes  Yes  Yes
c  7  Yes  Yes  Yes
d  6  Yes  Yes  Yes
e  7  Yes  Yes  Yes
f  4  Yes  Yes  Yes
g  1 or 1a  Yes  Yes  Yes
h  6  Yes  Yes  Yes
i  6  Yes  No  No
j  6  Yes  Yes  Yes
k  6  Yes  Yes  Yes
l  2  Yes  Yes  Yes
m  1 or 1b  Yes  Yes  Yes
n  7  Yes  Yes  Yes

o  6  Yes  Yes  Yes
p  7  Yes  Yes  Yes
q  6  Yes  Yes  Yes
r  7  Yes  Yes  Yes
s  4  Yes  Yes  Yes
t  1 or 1a  Yes  Yes  Yes
u  6  Yes  Yes  Yes
v  2  Yes  No  No
w  3  Yes  Yes  Yes
x  3  Yes  No  No
y  1  Yes  No  No
z  1  Yes  No  No
A  7  Yes  Yes  Yes
B  7  Yes  Yes  Yes
C  7  Yes  Yes  Yes
D  7  Yes  Yes  Yes
E  7  No  Yes  Yes
F  7  No  Yes  Yes
G  7  No  Yes  Yes
H  7  No  Yes  Yes
I  5  No  Yes  Yes
J  5  No  Yes  Yes
K  5  No  Yes  Yes
L  7  No  Yes  Yes
M  1 or 1b  No  Yes  Yes
N  7  No  Yes  Yes
O  1 or 1b  No  Yes  Yes
P  7  No  Yes  Yes
Q  6  No  Yes  Yes
R  1 or 1b  No  Yes  Yes
S  7  No  Yes  Yes
T  1 or 1b  No  Yes  Yes
U  7  No  Yes  Yes
V  1 or 1b  No  Yes  Yes
W  7  No  Yes  Yes
X  6  No  Yes  Yes
Y  1 or 1b  No  Yes  Yes

And now some examples to these special texts:

     The text ‘Moonrise at %(+5158+00738,120  in MS, BRD’ will be expanded to

↦ ‘Moonrise at 12:21 in MS, BRD’, in case the actual system date is the 1st June 1998.

     The text ‘Moonset at %)*5+5158+00738,120  in MS, BRD’ will be expanded to

↦ ‘Moonset at 01:53am in MS, BRD’, in case the actual system date is the 1st June 1998.

The text ‘Moon visible %[5+5158+00738,120  in MS, BRD’ will be expanded

to ↦ ‘Moon visible 13h32' in MS, BRD’, in case the actual system date is the 1st June 1998.

The text ‘Moon non-visible %]*+5158+00738,120  in MS, BRD’ will be ex-

panded to ↦ ‘Moon non-visible 10.469 in MS, BRD’, in case the actual system date is the 1st June 1998.

The text ‘Moon azimuth 0 o'clock=%(*a+5158+00738,120  in MS, BRD’ will

be expanded to

↦ ‘Moon azimuth 0 o'clock=267d37' in MS, BRD’, in case the actual system

date is the 1st June 1998.

The text ‘Moonphase %(x+00+000 =%(*m+5158+00738,120 % in MS, BRD’ will

be expanded to

↦ ‘Moonphase +16h00'=45.248% in MS, BRD’, in case you call Gcal with the

--time-offset=16: option and the actual system date is the 1st June 1998.

The text ‘Julian date at %(x+00+000 =%(y+00+000’ will be expanded to
↦ ‘Julian date at +10h15'=2450965.927’, in case you call Gcal with the

--time-offset=10:15 option and the actual system date is the 1st June 1998.

Here is a list that reports about the used reference systems in a short manner, describes other aspects that are unmentioned now, and informs about the lacks and limitations that are existing for the Moon oriented special texts:

• All astronomical calculations are based on methods that use —astronomically seen— low-precision elements, which of course is shown in the precision of the results that are caused by them. Nevertheless in most cases, these results can be treated as more as sufficiently precise if one measures their practical use.

• All azimuth and elevation values are always referring to the center of the respective astronomical object, except the case, it is stated otherwise.

• The base point of azimuth values, right ascension values and other ecliptic co-ordinate values is always measured clockwise relative to the geographical, true North (not the magnetic North as shown by a compass), where angular values for the North direction are both denoted as 0 degree and 360 degree.

• All topocentrically based azimuth and elevation values are always corrected by the amount of atmospheric refraction. Such a kind of correction is always done if the center of the respective astronomical object is above 2 degrees below the horizon.

• Generally, the rise and set times of the Moon cannot be precisely predicted, because depending on unpredictable atmospheric conditions, i.e. the local weather conditions as they appear during these times change the amount of atmospheric refraction at the horizon, so that the precalculated rise and set times may throughout be in error by a minute or more.

• The accuracy of rise and set computations decreases again at high latitudes, i.e. within in the lunar Arctic Circle. There, small variations in atmospheric refraction can change the time of moonrise or moonset by many minutes, since the Moon intersects the horizon at a very shallow angle. For the same reason, at high latitudes, the effects of observer height and local topography are magnified and can substantially change the times of the phenomena actually observed, or even whether the phenomena are observed to occur at all.

• It is possible that calculations, which are done for locations within the lunar Arctic Circle, sometimes result in the wrongly **.** text instead of the more precisely ++:++ or –:– texts that are used for marking special events. Unfortunately, therefore, then the reason why the Moon does not rise or set is no longer clearly evident, respectively, it is possible to get the false suggestion that the reason for non-rising or non-setting of the Moon is caused by just the one day within a synodic month, at which the Moon does not rise or set anyhow. Unfortunately, it is just only obvious that the Moon does not rise or set during this day.

Please also note the following references:

• See [Fixed dates option --precise], page 43, how to obtain a more precise representation of the values which are cause by these special texts.

• And note [Calendar option --time-offset=argument], page 33, how to change the clocktime, for which the dynamical, i.e. depending on the respective clocktime, Moon oriented special texts are calculated.

• Likewise, see [Fixed dates option --cycle-end=argument], page 41, how the dynamical, i.e. depending on the respective clocktime, Moon oriented special texts are calculated for a series of clocktimes for the current day.

• Furthermore, see [Fixed dates option --adjust-value=argument], page 39, how to change the reference altitude that is used to calculate the rise and set times.

• And last, note [Fixed dates option --atmosphere=air-pressure[,temperature]],

page 40, how to change the base data of the atmosphere that affects the calculation of the Moon oriented special texts.

All Moon oriented special texts must always be trailed by a whitespace character which is removed in output!

G.2.2.7 Environment variable %[format]-[argument] special text

%[format]-[argument] references the contents of an environment variable, e.g.: The text ‘I am `%-USER '-user’ will be expanded to ↦ ‘I am `guest'-user’, in case you logged-in as guest on your system. This special text must always be trailed by a whitespace character which is removed in

output!

G.2.3 Replacements without any argument %? special texts

Apart from further useful difference values, the text of a fixed date can be provided with different texts used for highlighting.

G.2.3.1 Difference value %[format]? special texts

The use of the other difference values as listed here is to calculate distance values between two dates. In case these special texts are directly lead by a ‘-’ character, e.g. ‘-%d’, Gcal switches the sign of the computed value.

You may depreciate the special meaning of the ‘-’ character —in case this character itself is needed— by placing a ‘\’ (backslash) character before it, e.g. ‘\-’. If you need the ‘\-’ characters themselves, you have to protect the ‘\’ (backslash) character by another ‘\’ (backslash) character, e.g. ‘\\-’.

The following other difference values %[format]? special texts are respected:

%d Specifies the current respectively queried day relative to the actual system date (≡today), e.g.:

The resource file line ‘0 %d days gone’ will be expanded to ↦ ‘-10 days gone’, in case you call Gcal with the -c10- option and no command.

%w Specifies the current respectively queried week relative to the actual system date (≡today). Started weeks are counted as complete weeks.

%m Specifies the current respectively queried month relative to the actual system date (≡today). Started months are counted as complete months.

%y Specifies the current respectively queried year relative to the actual system date (≡today). Started years are counted as complete years. This special text must always be trailed by a whitespace character which is removed in output!

An example:

Supposing, the actual system date is the 4th September 1999. To calculate, how many days, weeks, months and years are between the actual system date and the birthday of a person, who is born on 21st September 1962, Gcal can be called as follows:

      $ gcal -f/dev/null -Ux -u -#'19620921 %y %m %w %d~' 1962

⊣

      ⊣ -37 -444 -1928 -13497

⊣

      $ gcal -f/dev/null -Ux -u -#'19620921 \\-%y \-%m %w -%d~' 1962

⊣

      ⊣ \--37 --444 -1928 13497

⊣

      $ gcal -f/dev/null -Ux -u -#'00000904 %y %m %w %d~' %19620921 1999

⊣

      ⊣ 37 444 1928 13497

⊣

      $ gcal -f/dev/null -Ux -u -#'00000904 %y %B19620921' %19620921 1962

⊣

      ⊣ 37 36

Please do not confound the relative year number %[format]y special text with the age value %B special text (see [Age value %[format]B[date] special text], page 158). The %y special text counts started years as complete years, while the %B special text does not.

G.2.3.2 Highlighting %? special texts

The following highlighting %? special texts are respected:

%1 %1 is replaced by the starting highlighting sequence respectively the starting marking character that is used for highlighting the actual day (see [Global option --highlighting=text], page 12). This attains, that all succeeding text of the line after this special text is displayed in the same way as the highlighted respectively marked actual day. %1 is used together with the %2 special text, which turns off this enabled highlighting sequence respectively produces the ending marking character. If a %1 text is not succeeded by a %2 text on the line, Gcal automatically inserts such a %2 text at the end of the line. For example:

     Only %1THIS%2 word is highlighted in this line.
     %1This text is highlighted up to%2 here.
     All from %1here up to the end of the line is highlighted.

%2 %2 is replaced by the ending highlighting sequence respectively the ending marking character that is used for highlighting the actual day. This attains, that a possibly active highlighting according to the actual day is turned off respectively an ending marking character is produced. %2 is used together with the %1 special text. The %2 text has no affect if no preceding %1 text was found on the line.

%3 %3 is replaced by the starting highlighting sequence respectively the starting marking character that is used for highlighting a holiday, (see [Global option --highlighting=text], page 12). This attains, that all succeeding text of the line after this special text is displayed in the same way as the highlighted respectively marked holiday. %3 is used together with the %4 special text, which turns off this enabled highlighting sequence respectively produces the ending marking character. If a %3 text is not succeeded by a %4 text on the line, Gcal automatically inserts such a %4 text at the end of the line. For example:

     Only %3THIS%4 word is highlighted in this line.
     %3This text is highlighted up to%4 here.
     All from %3here up to the end of the line is highlighted.

%4 %4 is replaced by the ending highlighting sequence respectively the ending marking character that is used for highlighting a holiday. This attains, that a possibly active highlighting according to a holiday is turned off respectively an ending marking character is produced. %4 is used together with the %3 special text. The %4 text has no affect if no preceding %3 text was found on the line.

%5 %5 is replaced by the starting highlighting sequence respectively the starting marking character that is used for highlighting the actual day if a fixed date is on today’s date, (see [Global option --highlighting=text], page 12). This attains, that all succeeding text of the line after this special text is displayed in the same way as the highlighted respectively marked actual day. %5 is used together with the %6 special text, which turns off this enabled highlighting sequence respectively produces the ending marking character. If a %5 text is not succeeded by a %6 text on the line, Gcal automatically inserts such a %6 text at the end of the line.

%6 %6 is replaced by the ending highlighting sequence respectively the ending marking character that is used for highlighting the actual day if a fixed date is on today’s date. This attains, that a possibly active highlighting according to the actual day is turned off respectively an ending marking character is produced. %6 is used together with the %5 special text. The %6 text has no affect if no preceding %5 text was found on the line.

%7 %7 is replaced by the starting highlighting sequence respectively the starting marking character that is used for highlighting a holiday if a fixed date is on a legal holiday date, (see [Global option --highlighting=text], page 12). This attains, that all succeeding text of the line after this special text is displayed in the same way as the highlighted respectively marked holiday. %7 is used together with the %8 special text, which turns off this enabled highlighting sequence respectively produces the ending marking character. If a %7 text is not succeeded by a %8 text on the line, Gcal automatically inserts such a %8 text at the end of the line.

%8 %8 is replaced by the ending highlighting sequence respectively the ending marking character that is used for highlighting a holiday if a fixed date is on a legal holiday date. This attains, that a possibly active highlighting according to a holiday is turned off respectively an ending marking character is produced. %8 is used together with the %7 special text. The %8 text has no affect if no preceding %7 text was found on the line.

%9 %9 is replaced by the starting highlighting sequence respectively the starting marking character that is used for highlighting the actual day if a fixed date is on today’s date; otherwise %9 is replaced by the starting highlighting sequence respectively the starting marking character that is used for highlighting a holiday if a fixed date is on a legal holiday date, (see [Global option --highlighting=text], page 12). This attains, that all succeeding text of the line after this special text is displayed in the same way as the highlighted respectively marked actual day or holiday. %9 is used together with the %0 special text, which turns off this enabled highlighting sequence respectively produces the ending marking character. If a %9 text is not succeeded by a %0 text on the line, Gcal automatically inserts such a %0 text at the end of the line.

%0 %0 is replaced by the ending highlighting sequence respectively the ending marking character that is used for highlighting the actual day if a fixed date is on today’s date; otherwise %0 is replaced by the ending highlighting sequence respectively the ending marking character that is used for highlighting a holiday if a fixed date is on a legal holiday date. This attains, that a possibly active highlighting according to today’s date or a holiday is turned off respectively an ending marking character is produced. %0 is used together with the %9 special text. The %0 text has no affect if no preceding %9 text was found on the line.

G.3 Shell command %![argument] special text

The argument which trails %! until the end of the line, is passed to the shell as a command for further processing. But before, all text variable references and %?… special texts for replacing particular texts which are given in argument are interpreted by Gcal. See Section 5.1.9 [Text variables], page 106, and Section G.2 [%?… special texts for text replacement], page 150. All leading text before the %! special text is processed by Gcal as usual. If the simple --debug respectively --debug=internal option is given at program start-up, informational messages about the executed command and its exit code will be shown on the standard error channel (see [Global option --debug=internal], page 11).

If the --debug=abort option is given, the Gcal program will be aborted with an error code in case an exit code not equal zero has occurred during the execution of the command. See [Error Code 2], page 257, and [Global option --debug=abort], page 11, for further information.

Use the --execute-command option if you want to execute all %![argument] special texts by the shell instead of seeing them textually only. See [Fixed date option --execute-command], page 43, for more details.

Appendix H Summary of all %?… Special Texts

Here is a short-list of all %?… special texts which can be used in the text part of Gcal resource file line. Of course, not all possibilities concerning the representation —which are offered by the format instruction— are listed here for those special texts which expanded representation may be modified by an optional format, but only some few exemplary. See Appendix J [Format Instruction], page 249, for further details.

%i[date][#[date]]

Exclusion of all non-inclusive date periods

%e[date][#[date]]

Exclusion of all exclusive date periods

%? Inclusive day period %v Exclusion of all non-legal holidays %x Exclusion of all non-holidays %a Exclusion of all non-Mondays %c Exclusion of all non-Tuesdays %f Exclusion of all non-Wednesdays %g Exclusion of all non-Thursdays %h Exclusion of all non-Fridays %l Exclusion of all non-Saturdays %p Exclusion of all non-Sundays %q Exclusion of all non-Mondays…Thursdays %r Exclusion of all non-Mondays…Fridays

%? Exclusive day period %V Exclusion of all legal holidays %X Exclusion of all holidays %A Exclusion of all Mondays %C Exclusion of all Tuesdays %F Exclusion of all Wednesdays %G Exclusion of all Thursdays %H Exclusion of all Fridays %L Exclusion of all Saturdays %P Exclusion of all Sundays %Q Exclusion of all Mondays…Thursdays %R Exclusion of all Mondays…Fridays %[format]K[date] weekday name

%K[date] Complete weekday name

%>3#K[date]

3-Letter weekday name

%>2#K[date]

2-Letter weekday name

%[format]?[date] weekday number %W[date] Weekday number, Mon≡1…Sun≡7

%>1&*W[date]

Weekday number, Mon≡1…Sun≡7, Ordinal Number suffix

%E[date] Weekday number, Mon≡0…Sun≡6

%>1&*E[date]

Weekday number, Mon≡0…Sun≡6, Ordinal Number suffix

%I[date] Weekday number, Sun≡1…Sat≡7

%>1&*I[date]

Weekday number, Sun≡1…Sat≡7, Ordinal Number suffix

%J[date] Weekday number, Sun≡0…Sat≡6

%>1&*J[date]

Weekday number, Sun≡0…Sat≡6, Ordinal Number suffix

%S[date] Weekday number, starting day of week ≡1

%>1&*S[date]

Weekday number, starting day of week ≡1, Ordinal Number suffix

%T[date] Weekday number, starting day of week ≡0

%>1&*T[date]

Weekday number, starting day of week ≡0, Ordinal Number suffix

%[format]N[date] day number %N[date] Day-of-year number

%>03*N[date]

Day-of-year number, 3-digits with leading zeroes

%>1&*N[date]

Day-of-year number, Ordinal Number suffix

%>03&*N[date]

Day-of-year number, 3-digits with leading zeroes, Ordinal Number suffix

%[format]D[date] day number %D[date] Day number

%>02*D[date]

Day number, 2-digits with leading zero

%>1&*D[date]

Day number, Ordinal Number suffix

%>02&*D[date]

Day number, 2-digits with leading zero, Ordinal Number suffix

%k[date] ISO week number or standard week number, either 2-digits with leading zero,

or 4-alphanumeric characters text

%[format]U[date] month name %U[date] Complete month name

%>3#U[date]

3-Letter month name

%[format]M[date] month

%M[date] Month number

%>02*M[date]

Month number, 2-digits with leading zero

%>1&*M[date]

Month number, Ordinal Number suffix

%>02&*M[date]

Month number, 2-digits with leading zero, Ordinal Number suffix

%[format]Y[date] year number %Y[date] Year number

%>04*Y[date]

Complete year number, 4-digits with leading zeroes

%[format]?[date] Moon phase %O[date] Moon phase text

%>03*O[date]

Moon phase text, 3-digits with leading zeroes

%Z[date] Moon phase text graphics

%?[date] biorhythm %,[date] Biorhythm text (series of values) %;[date] Biorhythm text graphics (bar)

%[format]?[date] difference value %ydate Relative year number %b[date] Age value (birthday) %B[date] Age value (birthday), Ordinal Number suffix

%[format]?[argument] actual clocktime

%t[*][[+|-]mmmm|hh:[mm]]

Actual local time in hh:mm format

%[format]'[[+|-]mmmm|hh:[mm]]

Actual local time in m[m…] format

%[format]_[*][[+|-]mmmm|hh:[mm]]

Actual local time hour in h[h…] format

%[format]?[[+|-]mmmm|hh:[mm]]

Actual local time minute in m[m…] format

%[format]{[[+|-]mmmm|hh:[mm]]

Actual local time 12-hour format suffix

%@[*][[+|-]mmmm|hh:[mm]]

Actual Universal time in hh:mm/date format

%[format]`[[+|-]mmmm|hh:[mm]]

Actual Universal time in m[m…]/date format

%[format].[*][[+|-]mmmm|hh:[mm]]

Actual Universal time hour in h[h…] format

%[format]/[[+|-]mmmm|hh:[mm]]

Actual Universal time minute in m[m…] format

%[format]}[[+|-]mmmm|hh:[mm]]

Actual Universal time 12-hour format suffix

%"[[+|-]mmmm|hh:[mm]]

Numerical Universal time timezone offset value

%=[[+|-]mmmm|hh:[mm]]

Actual numerical local time timezone offset value

%n[[+|-]n] Textual date in fixed ‘%>02*D-%>3#U-%>04*Y’ format

%[format]j[[+|-]n]

Julian day number

%[format]b[*][mode]ISO-6709:1983-co-ordinate-1/ISO-6709:1983-co-ordinate-2

geographical distance and course angle

%b0ISO-6709:1983-co-ordinate-1/ISO-6709:1983-co-ordinate-2

Geographical distance (air line distance) in kilometers

%b*0ISO-6709:1983-co-ordinate-1/ISO-6709:1983-co-ordinate-2

Geographical distance (air line distance) in statute miles

%b1ISO-6709:1983-co-ordinate-1/ISO-6709:1983-coordinate-2

Course angle (true track) between ISO-6709:1983-co-ordinate-1 and ISO-6709:1983-co-ordinate-2 in degrees and arcminutes

%b*1ISO-6709:1983-co-ordinate-1/ISO-6709:1983-co-ordinate-2

Course angle (true track) between ISO-6709:1983-co-ordinate-1 and ISO-6709:1983-co-ordinate-2 in decimal degrees

%b2ISO-6709:1983-co-ordinate-1/ISO-6709:1983-co-ordinate-2

Course angle (true track) between ISO-6709:1983-co-ordinate-2 and ISO-6709:1983-co-ordinate-1 in degrees and arcminutes

%b*2ISO-6709:1983-co-ordinate-1/ISO-6709:1983-co-ordinate-2

Course angle (true track) between ISO-6709:1983-co-ordinate-2 and ISO-6709:1983-co-ordinate-1 in decimal degrees

%[format]?[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] Sun data
%o0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Sun’s astronomical midnight time in hh:mm 24-hours format

%o*0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Sun’s astronomical midnight time in hh:mm 12-hours format with time suffix

%o1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Sun’s astronomical noon time in hh:mm 24-hours format

%o*1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Sun’s astronomical noon time in hh:mm 12-hours format with time suffix

%o2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Sun’s center in hh:mm 24-hours format

%o*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Sun’s center in hh:mm 12-hours format with time suffix

%s2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Sun’s center in hh:mm 24-hours format

%o*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Sun’s center in hh:mm 12-hours format with time suffix

%u2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Sun’s center above reference altitude, in hours and minutes in hhhmm' format

%u*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Sun’s center above reference altitude, in decimal hours in hh.h... format

%z2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Sun’s center below reference altitude, in hours and minutes in hhhmm' format

%z*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Sun’s center below reference altitude, in decimal hours in hh.h... format

%o3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Sun’s upper limb in hh:mm 24-hours format

%o*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Sun’s upper limb in hh:mm 12-hours format with time suffix

%s3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Sun’s upper limb in hh:mm 24-hours format

%s*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Sun’s upper limb in hh:mm 12-hours format with time suffix

%u3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Sun’s upper limb above reference altitude, in hours and minutes in hhhmm' format

%u*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Sun’s upper limb above reference altitude, in decimal hours in hh.h... format

%z3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Sun’s upper limb below reference altitude, in hours and minutes in hhhmm' format

%z*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Sun’s upper limb below reference altitude, in decimal hours in hh.h... format

%o4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Rise time of Sun’s center in hh:mm 24-hours format

%o*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Rise time of Sun’s center in hh:mm 12-hours format with time suffix

%s4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Set time of Sun’s center in hh:mm 24-hours format

%s*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Set time of Sun’s center in hh:mm 12-hours format with time suffix

%u4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of visibility, Sun’s center above reference altitude, in hours and minutes in hhhmm' format

%u*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of visibility, Sun’s center above reference altitude, in decimal hours in hh.h... format

%z4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of non-visibility, Sun’s center below reference altitude, in hours and minutes in hhhmm' format

%z*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of non-visibility, Sun’s center below reference altitude, in decimal hours in hh.h... format

%o5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard rise time of Sun’s upper limb in hh:mm 24-hours format

%o*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard rise time of Sun’s upper limb in hh:mm 12-hours format with time suffix

%s5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard set time of Sun’s upper limb in hh:mm 24-hours format

%s*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard set time of Sun’s upper limb in hh:mm 12-hours format with time suffix

%u5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard visibility, Sun’s upper limb above reference altitude, in hours and minutes in hhhmm' format

%u*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard visibility, Sun’s upper limb above reference altitude, in decimal hours in hh.h... format

%z5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard non-visibility, Sun’s upper limb below reference altitude, in hours and minutes in hhhmm' format

%z*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard non-visibility, Sun’s upper limb below reference altitude, in decimal hours in hh.h... format

%o6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of civil twilight in the morning in hh:mm 24-hours format

%o*6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of civil twilight in the morning in hh:mm 12-hours format with time suffix

%s6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of civil twilight in the evening in hh:mm 24-hours format

%s*6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of civil twilight in the evening in hh:mm 12-hours format with time suffix

%u6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of civil twilight in hours and minutes in hhhmm' format

%u*6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of civil twilight in decimal hours in hh.h... format

%z6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of civil twilight in hours and minutes in hhhmm' format

%z*6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of civil twilight in decimal hours in hh.h... format

%o7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of nautical twilight in the morning in hh:mm 24-hours format

%o*7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of nautical twilight in the morning in hh:mm 12-hours format with time suffix

%s7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of nautical twilight in the evening in hh:mm 24-hours format

%s*7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of nautical twilight in the evening in hh:mm 12-hours format with time suffix

%u7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of nautical twilight in hours and minutes in hhhmm' format

%u*7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of nautical twilight in decimal hours in hh.h... format

%z7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of nautical twilight in hours and minutes in hhhmm' format

%z*7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of nautical twilight in decimal hours in hh.h... format

%o8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of amateur-astronomical twilight in the morning in hh:mm 24-hours format

%o*8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of amateur-astronomical twilight in the morning in hh:mm 12-hours format with time suffix

%s8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of amateur-astronomical twilight in the evening in hh:mm 24-hours format

%s*8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of amateur-astronomical twilight in the evening in hh:mm 12-hours format with time suffix

%u8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of amateurastronomical twilight in hours and minutes in hhhmm' format

%u*8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of amateurastronomical twilight in decimal hours in hh.h... format

%z8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of amateurastronomical twilight in hours and minutes in hhhmm' format

%z*8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of amateurastronomical twilight in decimal hours in hh.h... format

%o9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of astronomical twilight in the morning in hh:mm 24-hours format

%o*9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Begin time of astronomical twilight in the morning in hh:mm 12-hours format with time suffix

%s9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of astronomical twilight in the evening in hh:mm 24-hours format

%s*9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

End time of astronomical twilight in the evening in hh:mm 12-hours format with time suffix

%u9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of astronomical twilight in hours and minutes in hhhmm' format

%u*9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center above reference altitude of astronomical twilight in decimal hours in hh.h... format

%z9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of astronomical twilight in hours and minutes in hhhmm' format

%z*9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of Sun’s center below reference altitude of astronomical twilight in decimal hours in hh.h... format

%oaISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%saISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun in degrees and arcminutes in

+|-ddddmm' format
%o*aISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*aISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun in decimal degrees in +|-ddd.d... format

%obISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sbISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Sun in degrees and arcminutes in ddddmm' format

%o*bISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*bISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Sun in decimal degrees in ddd.d... format

%ocISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%scISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric declination of Sun in degrees and arcminutes in

          +|-ddddmm' format
%o*cISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*cISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric declination of Sun in decimal degrees in +|-ddd.d... format

%odISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sdISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric ecliptic longitude of Sun in degrees and arcminutes in

          ddddmm' format
%o*dISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*dISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric ecliptic longitude of Sun in decimal degrees in

          ddd.d... format
%oeISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%seISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric right ascension of Sun as time value in hours and minutes in hhhmm' format

%o*eISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*eISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric right ascension of Sun as time value in decimal hours in hh.h... format

%ofISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sfISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric distance of Sun in astronomical units in n.n... format

%o*fISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*fISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric distance of Sun in kilometers in n.n... format

%ogISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sgISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric horizontal parallax of Sun in degrees and arcminutes in ddddmm' format

%o*gISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*gISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric horizontal parallax of Sun in decimal degrees in

          ddd.d... format
%ohISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%shISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric semidiameter of Sun in degrees and arcminutes in

          ddddmm' format
%o*hISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*hISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric semidiameter of Sun in decimal degrees in ddd.d... format

%oiISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%siISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Atmospheric refraction in degrees and arcminutes in ddddmm' format

%o*iISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*iISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Atmospheric refraction in decimal degrees in ddd.d... format

%ojISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sjISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun in degrees and arcminutes in

          +|-ddddmm' format
%o*jISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*jISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun in decimal degrees in +|-ddd.d... format

%okISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%skISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Sun in degrees and arcminutes in ddddmm' format

%o*kISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*kISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Sun in decimal degrees in ddd.d... format

%olISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%slISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric declination of Sun in degrees and arcminutes in

          +|-ddddmm' format
%o*lISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*lISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric declination of Sun in decimal degrees in +|-ddd.d... format

%omISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%smISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric ecliptic longitude of Sun in degrees and arcminutes in

          ddddmm' format
%o*mISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*mISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric ecliptic longitude of Sun in decimal degrees in ddd.d... format

%onISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%snISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric right ascension of Sun as time value in hours and minutes in hhhmm' format

%o*nISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*nISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric right ascension of Sun as time value in decimal hours in

          hh.h... format
%ooISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%soISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric distance of Sun in astronomical units in n.n... format

%o*oISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*oISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric distance of Sun in kilometers in n.n... format

%opISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%spISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric horizontal parallax of Sun in degrees and arcminutes in

          ddddmm' format
%o*pISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*pISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric horizontal parallax of Sun in decimal degrees in

          ddd.d... format
%oqISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sqISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric semidiameter of Sun in degrees and arcminutes in

          ddddmm' format
%o*qISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*qISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric semidiameter of Sun in decimal degrees in ddd.d... format

%orISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%srISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%o*rISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*rISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta-t in seconds in +|-n.n... format

%osISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%ssISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Local sidereal time in hh:mm 24-hours format

%o*sISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*sISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Local sidereal time in hh:mm 12-hours format with time suffix

%otISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%stISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Base time as time value in hours and minutes in hhhmm' format

%o*tISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*tISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Base time as time value in decimal hours in hh.h... format

%ouISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%suISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%o*uISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*uISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Julian date in days in n.n... format

%ovISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%svISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%o*vISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*vISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Julian Ephemeris date in days in n.n... format

%owISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%swISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Equation of time as time value in hours and minutes in +|-hhhmm' format

%o*wISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*wISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Equation of time as time value in decimal hours in +|-hh.h... format

%oxISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sxISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Sun/Moon in degrees and arcminutes

          in +|-ddddmm' format
%o*xISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*xISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Sun/Moon in decimal degrees in

          +|-ddd.d... format
%oyISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%syISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Sun/Moon in degrees and arcminutes

          in +|-ddddmm' format

%o*yISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*yISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Sun/Moon in decimal degrees in

+|-ddd.d... format
%ozISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%szISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Sun/Moon in degrees and arcminutes

in +|-ddddmm' format
%o*zISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*zISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Sun/Moon in decimal degrees in

+|-ddd.d... format
%oAISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sAISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Sun/Moon in degrees and arcminutes

in +|-ddddmm' format
%o*AISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*AISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Sun/Moon in decimal degrees in

+|-ddd.d... format
%oBISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Sun/Moon at standard sunrise time in degrees and arcminutes in +|-ddddmm' format

%o*BISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Sun/Moon at standard sunrise time in decimal degrees in +|-ddd.d... format

%sBISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Sun/Moon at standard sunset time in degrees and arcminutes in +|-ddddmm' format

%s*BISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Sun/Moon at standard sunset time in decimal degrees in +|-ddd.d... format

%oCISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Sun/Moon at standard sunrise time in degrees and arcminutes in +|-ddddmm' format

%o*CISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Sun/Moon at standard sunrise time in decimal degrees in +|-ddd.d... format

%sCISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Sun/Moon at standard sunset time in degrees and arcminutes in +|-ddddmm' format

%s*CISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Sun/Moon at standard sunset time in decimal degrees in +|-ddd.d... format

%oDISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Sun/Moon at standard sunrise time in degrees and arcminutes in +|-ddddmm' format

%o*DISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Sun/Moon at standard sunrise time in decimal degrees in +|-ddd.d... format

%sDISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Sun/Moon at standard sunset time in degrees and arcminutes in +|-ddddmm' format

%s*DISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Sun/Moon at standard sunset time in decimal degrees in +|-ddd.d... format

%oEISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Sun/Moon at standard sunrise time in degrees and arcminutes in +|-ddddmm' format

%o*EISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Sun/Moon at standard sunrise time in decimal degrees in +|-ddd.d... format

%sEISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Sun/Moon at standard sunset time in degrees and arcminutes in +|-ddddmm' format

%s*EISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Sun/Moon at standard sunset time in decimal degrees in +|-ddd.d... format

%oFISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sFISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical midnight time of Sun/Moon as time value in hours and minutes in +|-hhhmm' format

%o*FISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*FISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical midnight time of Sun/Moon as time value in decimal hours in +|-hh.h... format

%oGISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sGISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical noon of Sun/Moon as time value in hours and minutes in +|-hhhmm' format

%o*GISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*GISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical noon of Sun/Moon as time value in decimal hours in +|-hh.h... format

%oHISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time of standard rise of Sun/Moon at standard sunrise time as time value in hours and minutes in +|-hhhmm' format

%o*HISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time of standard rise of Sun/Moon at standard sunrise time as time value in decimal hours in +|-hh.h... format

%sHISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time of standard set of Sun/Moon at standard sunset time as time value in hours and minutes in +|-hhhmm' format

%s*HISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time of standard set of Sun/Moon at standard sunset time as time value in decimal hours in +|-hh.h... format

%oIISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sIISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at Sun’s astronomical midnight time in degrees and arcminutes in +|-ddddmm' format

%o*IISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*IISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at Sun’s astronomical midnight time in decimal degrees in +|-ddd.d... format

%oJISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sJISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at Sun’s astronomical noon time in degrees and arcminutes in +|-ddddmm' format

%o*JISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*JISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at Sun’s astronomical noon time in decimal degrees in +|-ddd.d... format

%oKISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at standard sunrise time in degrees and arcminutes in +|-ddddmm' format

%o*KISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at standard sunrise time in decimal degrees in +|-ddd.d... format

%sKISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at standard sunset time in degrees and arcminutes in +|-ddddmm' format

%s*KISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Sun at standard sunset time in decimal degrees in +|-ddd.d... format

%oLISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Sun at standard sunrise time in degrees and arcminutes in ddddmm' format

%o*LISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Sun at standard sunrise time in decimal degrees in ddd.d... format

%sLISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Sun at standard sunset time in degrees and arcminutes in ddddmm' format

%s*LISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Sun at standard sunset time in decimal degrees in ddd.d... format

%oMISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sMISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at Sun’s astronomical midnight time in degrees and arcminutes in +|-ddddmm' format

%o*MISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*MISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at Sun’s astronomical midnight time in decimal degrees in +|-ddd.d... format

%oNISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%sNISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at Sun’s astronomical noon time in degrees and arcminutes in +|-ddddmm' format

%o*NISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%s*NISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at Sun’s astronomical noon time in decimal degrees in +|-ddd.d... format

%oOISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at standard sunrise time in degrees and arcminutes in +|-ddddmm' format

%o*OISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at standard sunrise time in decimal degrees in +|-ddd.d... format

%sOISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at standard sunset time in degrees and arcminutes in +|-ddddmm' format

%s*OISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Sun at standard sunset time in decimal degrees in +|-ddd.d... format

%oPISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Sun at standard sunrise time in degrees and arcminutes in ddddmm' format

%o*PISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Sun at standard sunrise time in decimal degrees in ddd.d... format

%sPISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Sun at standard sunset time in degrees and arcminutes in ddddmm' format

%s*PISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Sun at standard sunset time in decimal degrees in ddd.d... format

%oQISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Single shadow length time in the forenoon in hh:mm 24-hours format

%o*QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Single shadow length time in the forenoon in hh:mm 12-hours format with time suffix

%sQISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Single shadow length time in the afternoon in hh:mm 24-hours format

%s*QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Single shadow length time in the afternoon in hh:mm 12-hours format with time suffix

%uQISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of single or shorter shadow length in hours and minutes in

hhhmm' format
%u*QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of single or shorter shadow length in decimal hours in

hh.h... format
%zQISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of longer than single shadow length in hours and minutes in

hhhmm' format
%z*QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of longer than single shadow length in decimal hours in

hh.h... format
%oRISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Double shadow length time in the forenoon in hh:mm 24-hours format

%o*RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Double shadow length time in the forenoon in hh:mm 12-hours format with time suffix

%sRISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Double shadow length time in the afternoon in hh:mm 24-hours format

%s*RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Double shadow length time in the afternoon in hh:mm 12-hours format with time suffix

%uRISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of double or shorter shadow length in hours and minutes in

hhhmm' format
%u*RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of double or shorter shadow length in decimal hours in

hh.h... format
%zRISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of longer than double shadow length in hours and minutes in hhhmm' format

%z*RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of longer than double shadow length in decimal hours in

hh.h... format
%[format]?[*][mode]ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]] Moon data
%(0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Moon’s astronomical midnight time in hh:mm 24-hours format

%(*0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*0ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Moon’s astronomical midnight time in hh:mm 12-hours format with time suffix

%(1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Moon’s astronomical noon time in hh:mm 24-hours format

%(*1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*1ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Moon’s astronomical noon time in hh:mm 12-hours format with time suffix

%(2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Moon’s center in hh:mm 24-hours format

%(*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Moon’s center in hh:mm 12-hours format with time suffix

%)2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Moon’s center in hh:mm 24-hours format

%(*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Moon’s center in hh:mm 12-hours format with time suffix

%[2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Moon’s center above reference altitude, in hours and minutes in hhhmm' format

%[*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Moon’s center above reference altitude, in decimal hours in hh.h... format

%]2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Moon’s center below reference altitude, in hours and minutes in hhhmm' format

%]*2ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Moon’s center below reference altitude, in decimal hours in hh.h... format

%(3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Moon’s upper limb in hh:mm 24-hours format

%(*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical rise time of Moon’s upper limb in hh:mm 12-hours format with time suffix

%)3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Moon’s upper limb in hh:mm 24-hours format

%)*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geometrical set time of Moon’s upper limb in hh:mm 12-hours format with time suffix

%[3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Moon’s upper limb above reference altitude, in hours and minutes in hhhmm' format

%[*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical visibility, Moon’s upper limb above reference altitude, in decimal hours in hh.h... format

%]3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Moon’s upper limb below reference altitude, in hours and minutes in hhhmm' format

%]*3ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of geometrical non-visibility, Moon’s upper limb below reference altitude, in decimal hours in hh.h... format

%(4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Rise time of Moon’s center in hh:mm 24-hours format

%(*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Rise time of Moon’s center in hh:mm 12-hours format with time suffix

%)4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Set time of Moon’s center in hh:mm 24-hours format

%)*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Set time of Moon’s center in hh:mm 12-hours format with time suffix

%[4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of visibility, Moon’s center above reference altitude, in hours and minutes in hhhmm' format

%[*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of visibility, Moon’s center above reference altitude, in decimal hours in hh.h... format

%]4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of non-visibility, Moon’s center below reference altitude, in hours and minutes in hhhmm' format

%]*4ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of non-visibility, Moon’s center below reference altitude, in decimal hours in hh.h... format

%(5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard rise time of Moon’s upper limb in hh:mm 24-hours format

%(*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard rise time of Moon’s upper limb in hh:mm 12-hours format with time suffix

%)5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard set time of Moon’s upper limb in hh:mm 24-hours format

%)*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Standard set time of Moon’s upper limb in hh:mm 12-hours format with time suffix

%[5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard visibility, Moon’s upper limb above reference altitude, in hours and minutes in hhhmm' format

%[*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard visibility, Moon’s upper limb above reference altitude, in decimal hours in hh.h... format

%]5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard non-visibility, Moon’s upper limb below reference altitude, in hours and minutes in hhhmm' format

%]*5ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Period of standard non-visibility, Moon’s upper limb below reference altitude, in decimal hours in hh.h... format

%(6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric horizontal parallax of Moon in degrees and arcminutes in ddddmm' format

%(*6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*6ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric horizontal parallax of Moon in decimal degrees in

          ddd.d... format
%(7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric semidiameter of Moon in degrees and arcminutes in

          ddddmm' format
%(*7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*7ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric semidiameter of Moon in decimal degrees in ddd.d... format

%(8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%(*8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*8ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric brightness of Moon in magnitude units in +|-n.n... format

%(9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon in range 0.0…1.0 in n.n... format

%(*9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*9ISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon as percent value in range 0.0…100.0 in n.n... format

%(aISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)aISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon in degrees and arcminutes in

          +|-ddddmm' format
%(*aISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*aISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon in decimal degrees in +|-ddd.d... format

%(bISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)bISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Moon in degrees and arcminutes in ddddmm' format

%(*bISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*bISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Moon in decimal degrees in ddd.d... format

%(cISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)cISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric declination of Moon in degrees and arcminutes in

          +|-ddddmm' format
%(*cISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*cISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric declination of Moon in decimal degrees in +|-ddd.d... format

%(dISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)dISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric ecliptic longitude of Moon in degrees and arcminutes in ddddmm' format

%(*dISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*dISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric ecliptic longitude of Moon in decimal degrees in

          ddd.d... format
%(eISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)eISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric ecliptic latitude of Moon in degrees and arcminutes in

          +|-ddddmm' format
%(*eISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*eISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric ecliptic latitude of Moon in decimal degrees in

          +|-ddd.d... format
%(fISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)fISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric right ascension of Moon as time value in hours and minutes in hhhmm' format

%(*fISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*fISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric right ascension of Moon as time value in decimal hours in hh.h... format

%(gISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)gISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric distance of Moon in mean Earth equator radii in

          n.n... format
%(*gISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*gISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric distance of Moon in kilometers in n.n... format

%(hISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)hISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elongation of Moon in degrees and arcminutes in

          ddddmm' format

%(*hISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*hISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elongation of Moon in decimal degrees in ddd.d... format

%(iISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)iISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Atmospheric refraction in degrees and arcminutes in ddddmm' format

%(*iISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*iISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Atmospheric refraction in decimal degrees in ddd.d... format

%(jISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)jISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric horizontal parallax of Moon in degrees and arcminutes in ddddmm' format

%(*jISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*jISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric horizontal parallax of Moon in decimal degrees in

          ddd.d... format
%(kISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)kISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric semidiameter of Moon in degrees and arcminutes in

          ddddmm' format
%(*kISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*kISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric semidiameter of Moon in decimal degrees in ddd.d... format

%(lISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)lISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%(*lISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*lISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric brightness of Moon in magnitude units in +|-n.n... format

%(mISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)mISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon in range 0.0…1.0 in n.n... format

%(*mISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*mISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon as percent value in range 0.0…100.0 in n.n... format

%(nISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)nISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon in degrees and arcminutes in

          +|-ddddmm' format
%(*nISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*nISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon in decimal degrees in +|-ddd.d... format

%(oISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)oISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Moon in degrees and arcminutes in ddddmm' format

%(*oISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*oISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Moon in decimal degrees in ddd.d... format

%(pISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)pISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric declination of Moon in degrees and arcminutes in

          +|-ddddmm' format
%(*pISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*pISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric declination of Moon in decimal degrees in +|-ddd.d... format

%(qISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)qISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric ecliptic longitude of Moon in degrees and arcminutes in

          ddddmm' format
%(*qISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*qISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric ecliptic longitude of Moon in decimal degrees in

          ddd.d... format
%(rISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)rISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric ecliptic latitude of Moon in degrees and arcminutes in

          +|-ddddmm' format
%(*rISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*rISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric ecliptic latitude of Moon in decimal degrees in

          +|-ddd.d... format
%(sISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)sISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric right ascension of Moon as time value in hours and minutes in hhhmm' format

%(*sISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*sISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric right ascension of Moon as time value in decimal hours in hh.h... format

%(tISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)tISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric distance of Moon in mean Earth equator radii in n.n... format

%(*tISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*tISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric distance of Moon in kilometers in n.n... format

%(uISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)uISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elongation of Moon in degrees and arcminutes in

          ddddmm' format
%(*uISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*uISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elongation of Moon in decimal degrees in ddd.d... format

%(vISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)vISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%(*vISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*vISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta-t in seconds in +|-n.n... format

%(wISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)wISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Local sidereal time in hh:mm 24-hours format

%(*wISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*wISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Local sidereal time in hh:mm 12-hours format with time suffix

%(xISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)xISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Base time as time value in hours and minutes in hhhmm' format

%(*xISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*xISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Base time as time value in decimal hours in hh.h... format

%(yISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)yISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%(*yISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*yISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Julian date in days in n.n... format

%(zISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)zISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%(*zISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*zISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Julian Ephemeris date in days in n.n... format

%(AISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)AISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Moon/Sun in degrees and arcminutes

          in +|-ddddmm' format
%(*AISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*AISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Moon/Sun in decimal degrees in

          +|-ddd.d... format
%(BISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)BISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Moon/Sun in degrees and arcminutes

          in +|-ddddmm' format
%(*BISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*BISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Moon/Sun in decimal degrees in

          +|-ddd.d... format
%(CISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)CISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Moon/Sun in degrees and arcminutes

          in +|-ddddmm' format
%(*CISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*CISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Moon/Sun in decimal degrees in

          +|-ddd.d... format
%(DISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)DISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Moon/Sun in degrees and arcminutes

          in +|-ddddmm' format
%(*DISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*DISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Moon/Sun in decimal degrees in

          +|-ddd.d... format
%(EISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Moon/Sun at standard moonrise time in degrees and arcminutes in +|-ddddmm' format

%(*EISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Moon/Sun at standard moonrise time in decimal degrees in +|-ddd.d... format

%)EISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Moon/Sun at standard moonset time in degrees and arcminutes in +|-ddddmm' format

%)*EISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric elevation of Moon/Sun at standard moonset time in decimal degrees in +|-ddd.d... format

%(FISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Moon/Sun at standard moonrise time in degrees and arcminutes in +|-ddddmm' format

%(*FISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Moon/Sun at standard moonrise time in decimal degrees in +|-ddd.d... format

%)FISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Moon/Sun at standard moonset time in degrees and arcminutes in +|-ddddmm' format

%)*FISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta topocentric azimuth of Moon/Sun at standard moonset time in decimal degrees in +|-ddd.d... format

%(GISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Moon/Sun at standard moonrise time in degrees and arcminutes in +|-ddddmm' format

%(*GISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Moon/Sun at standard moonrise time in decimal degrees in +|-ddd.d... format

%)GISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Moon/Sun at standard moonset time in degrees and arcminutes in +|-ddddmm' format

%)*GISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric elevation of Moon/Sun at standard moonset time in decimal degrees in +|-ddd.d... format

%(HISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Moon/Sun at standard moonrise time in degrees and arcminutes in +|-ddddmm' format

%(*HISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Moon/Sun at standard moonrise time in decimal degrees in +|-ddd.d... format

%)HISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Moon/Sun at standard moonset time in degrees and arcminutes in +|-ddddmm' format

%)*HISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta geocentric azimuth of Moon/Sun at standard moonset time in decimal degrees in +|-ddd.d... format

%(IISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)IISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical midnight time of Moon/Sun as time value in hours and minutes in +|-hhhmm' format

%(*IISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*IISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical midnight time of Moon/Sun as time value in decimal hours in +|-hh.h... format

%(JISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)JISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical noon time of Moon/Sun as time value in hours and minutes in +|-hhhmm' format

%(*JISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*JISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta astronomical noon time of Moon/Sun as time value in decimal hours in +|-hh.h... format

%(KISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time standard rise of Moon/Sun at standard moonrise time as time value in hours and minutes in +|-hhhmm' format

%(*KISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time standard rise of Moon/Sun at standard moonrise time as time value in decimal hours in +|-hh.h... format

%)KISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time standard set of of Moon/Sun at standard moonset time as time value in hours and minutes in +|-hhhmm' format

%)*KISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Delta time standard set of of Moon/Sun at standard moonset time as time value in decimal hours in +|-hh.h... format

%(LISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)LISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at Moon’s astronomical midnight time in degrees and arcminutes in +|-ddddmm' format

%(*LISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*LISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at Moon’s astronomical midnight time in decimal degrees in +|-ddd.d... format

%(MISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)MISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at Moon’s astronomical midnight time in range 0.0…1.0 in n.n... format

%(*MISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*MISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at Moon’s astronomical midnight time as percent value in range 0.0…100.0 in n.n... format

%(NISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)NISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at Moon’s astronomical noon time in degrees and arcminutes in +|-ddddmm' format

%(*NISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*NISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at Moon’s astronomical noon time in decimal degrees in +|-ddd.d... format

%(OISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)OISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at Moon’s astronomical noon time in range 0.0…1.0 in n.n... format

%(*OISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*OISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at Moon’s astronomical noon time as percent value in range 0.0…100.0 in n.n... format

%(KISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at standard moonrise time in degrees and arcminutes in +|-ddddmm' format

%(*PISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at standard moonrise time in decimal degrees in +|-ddd.d... format

%)PISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at standard moonset time in degrees and arcminutes in +|-ddddmm' format

%)*PISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric elevation of Moon at standard moonset time in decimal degrees in +|-ddd.d... format

%(QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Moon at standard moonrise time in degrees and arcminutes in ddddmm' format

%(*QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Moon at standard moonrise time in decimal degrees in ddd.d... format

%)QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Moon at standard moonset time in degrees and arcminutes in ddddmm' format

%)*QISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric azimuth of Moon at standard moonset time in decimal degrees in ddd.d... format

%(RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at standard moonrise time in range 0.0…1.0 in n.n... format

%(*RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at standard moonrise time as percent value in range 0.0…100.0 in n.n... format

%)RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at standard moonset time in range 0.0…1.0 in n.n... format

%)*RISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Topocentric phase angle of Moon at standard moonset time as percent value in range 0.0…100.0 in n.n... format

%(SISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)SISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at Moon’s astronomical midnight time in degrees and arcminutes in +|-ddddmm' format

%(*SISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*SISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at Moon’s astronomical midnight time in decimal degrees in +|-ddd.d... format

%(TISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)TISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at Moon’s astronomical midnight time in range 0.0…1.0 in n.n... format

%(*TISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*TISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at Moon’s astronomical midnight time as percent value in range 0.0…100.0 in n.n... format

%(UISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)UISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at Moon’s astronomical noon time in degrees and arcminutes in +|-ddddmm' format

%(*UISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*UISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at Moon’s astronomical noon time in decimal degrees in +|-ddd.d... format

%(VISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)VISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at Moon’s astronomical noon time in range 0.0…1.0 in n.n... format

%(*VISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]
%)*VISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at Moon’s astronomical noon time as percent value in range 0.0…100.0 in n.n... format

%(WISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at standard moonrise time in degrees and arcminutes in +|-ddddmm' format

%(*WISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at standard moonrise time in decimal degrees in +|-ddd.d... format

%)WISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at standard moonset time in degrees and arcminutes in +|-ddddmm' format

%)*WISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric elevation of Moon at standard moonset time in decimal degrees in +|-ddd.d... format

%(XISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Moon at standard moonrise time in degrees and arcminutes in ddddmm' format

%(*XISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Moon at standard moonrise time in decimal degrees in ddd.d... format

%)XISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Moon at standard moonset time in degrees and arcminutes in ddddmm' format

%)*XISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric azimuth of Moon at standard moonset time in decimal degrees in ddd.d... format

%(YISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at standard moonrise time in range 0.0…1.0 in n.n... format

%(*YISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at standard moonrise time as percent value in range 0.0…100.0 in n.n... format

%)YISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at standard moonset time in range 0.0…1.0 in n.n... format

%)*YISO-6709:1983-co-ordinate[,[+|-]mmmm|hh:[mm]]

Geocentric phase angle of Moon at standard moonset time as percent value in range 0.0…100.0 in n.n... format

%[format]-[argument]

Contents of environment variable %[format]? difference value

%d Relative day number %w Relative week number %m Relative month number %y Relative year number

%? highlighting %1 Start of highlighting sequence 1 %2 End of highlighting sequence 1 %3 Start of highlighting sequence 2 %4 End of highlighting sequence 2 %5 Start of highlighting sequence 1 if a fixed date is on today’s date %6 End of highlighting sequence 1 if a fixed date is on today’s date %7 Start of highlighting sequence 2 if a fixed date is on a legal holiday

date

%8 End of highlighting sequence 2 if a fixed date is on a legal holiday

date

%9 Start of highlighting sequence 1 if a fixed date is on today’s date, otherwise start of highlighting sequence 2 if a fixed date is on a legal holiday date

%0 End of highlighting sequence 1 if a fixed date is on today’s date, otherwise end of highlighting sequence 2 if a fixed date is on a legal holiday date

%![argument]

Shell executes argument as a command

Appendix I Obsolete Special Texts and Format Elements

Because Gcal 4 is able to represent most special texts which may occur in the text part of a resource file line (see Appendix G [Description of all %?… Special Texts], page 145), and most of the format elements as they may be used in the date format text of the --date-format See [Calendar option --date-format=de|us|gb|text], page 35, by using a format instruction, some of them —mainly introduced in the Gcal versions 2.20 and 2.40— are obsolete now. The following complete tables of all these obsolete special texts respectively format elements with their new representation, as used by Gcal 4 now, show the changes made.

I.1 Table of Obsolete Special Texts

This table opposes all obsolete special texts as they may be used in the text part of a resource file line (see Appendix G [Description of all %?… Special Texts], page 145), and their new representation, as used by Gcal 4 now:

Symbolic name  Old-style  New-style
%complete weekday name  %o  ≡ %K
%3-letter weekday name  %O  ≡ %>3#K
%2-letter weekday name  %K  ≡ %>2#K
%weekday number (Mon=1…Sun=7)  %S  ≡ %W
%weekday number (Mon=1…Sun=7) ONS  %I  ≡ %>1&*W
%weekday number (Mon=0…Sun=6) ONS  %J  ≡ %>1&*E
%weekday number (Sun=1…Sat=7)  %(  ≡ %I
%weekday number (Sun=1…Sat=7) ONS  %<  ≡ %>1&*I
%weekday number (Sun=0…Sat=6)  %)  ≡ %J
%weekday number (Sun=0…Sat=6) ONS  %>  ≡ %>1&*J
%weekday number (Mon=1…Sun=7) FLX  %[  ≡ %S
%weekday number (Mon=1…Sun=7) FLX ONS  %{  ≡ %>1&*S
%weekday number (Mon=0…Sun=6) FLX  %]  ≡ %T
%weekday number (Mon=0…Sun=6) FLX ONS  %}  ≡ %>1&*T
%day-of-year number  %+  ≡ %N
%day-of-year number +LZ  %*  ≡ %>03*N
%day-of-year number ONS  %&  ≡ %>1&*N
%day-of-year number +LZ ONS  %#  ≡ %>03&*N
%day number +LZ  %N  ≡ %>02*D
%day number ONS  %s  ≡ %>1&*D
%day number +LZ ONS  %u  ≡ %>02&*D
%complete month name  %M  ≡ %U
%3-letter month name  %T  ≡ %>3#U
%month number  %U  ≡ %M
%month number +LZ  %W  ≡ %>02*M
%month number ONS  %z  ≡ %>1&*M
%month number +LZ ONS  %Z  ≡ %>02&*M
%complete year number +LZ  %=  ≡ %>04*Y
%age value (birthday) number  %b  ≡ %B
%age value (birthday) number ONS  %B  ≡ %>1&*B
%moon phase  %-  ≡ %O
%moon phase +LZ  %_≡ %>03*O
%moon phase text graphics image %: ≡ %Z

Legend:

+LZ ≡ With leading zero(es)

FLX

≡

Flexible starting day of week (see [Calendar option --starting-day=argument], page 33)

ONS ≡ With Ordinal Number suffix

I.2 Table of Obsolete Date Format Elements

This table opposes all obsolete date format elements as they may be used in the date format

text of the --date-format option (see [Calendar option --date-format=de|us|gb|text],

page 35), and their new representation, as used by Gcal 4 now:

Symbolic name  Old-style  New-style
%day number  %d  ≡ %>02*D
%day number +LS  %D  ≡ %>2*D
%day number +LZ ONS  %u  ≡ %02&*D
%day number +LS ONS  %U  ≡ %>2&*D
%complete month name  %B  ≡ %U
%3-letter month name  %b  ≡ %<3#U
%month number +LZ  %m  ≡ %>02*M
%month number +LS  %M  ≡ %>2*M
%complete year number +LZ  %y  ≡ %>04*Y
%complete year number +LZ  %Y  ≡ %>4*Y
%last-2-digits year number +LZ  %z  ≡ %>02#Y
%last-2-digits year number +LS  %Z  ≡ %>2#Y
%complete weekday name  %A  ≡ %K
%3-letter weekday name  %W  ≡ %<3#K
%2-letter weekday name  %w  ≡ %<2#K

Legend:

+LZ ≡ With leading zero(es)
+LS ≡ With leading space(s)
ONS ≡ With Ordinal Number suffix

I.3 Table of Obsolete Date Formats

This table opposes all obsolete date format texts as they may be used in the fixed defaults of the country specific date formats of the --date-format option for German users, U.S. american users and users in Great Britain (see [Calendar option --date-format=de|us|gb|text], page 35), and their new representation, as used by Gcal 4 now:

CC  Old-style  New-style

DE  %w, %1%D%2 %b %y  ≡ %<2#K, %1%>2*D%2 %<3#U %>04*Y
US  %W, %b %1%U%2 %y  ≡ %<3#K, %<3#U %1%>2&*D%2 %>04*Y
GB  %W, %1%U%2 %b %y  ≡ %<3#K, %1%>2&*D%2 %<3#U %>04*Y

Appendix J Format Instruction

Gcal 4 is able to represent text variables, special texts and date formats especially in a very comfortable and uniform manner by using a format instruction.

Here comes the description of the format instruction components, now called format in the further context. In general, the template of a format instruction is:

alignment [sign] [zero] width [style] [suffix] format

If a format of above mentioned template is used, its components must be specified in the above mentioned order, at which the definition of the components embraced by ‘[ ]’ is of course optional. The single components of the format instruction are assembled as follows:

Character  Description of the components

Alignment group (exactly one member must be defined):

<

Field contents is placed at the left margin of the field by using the width width

:

Field contents is placed in centered manner in the field by using the width width

>

Field contents is placed at the right margin of the field by using the width width

Sign (may be defined):

+  Numerical value is provided with leading sign

Zero (may be defined):

0  Numerical value is filled with leading zero(es)

Width (must be defined):

n  Field has the width n (actual range 1…256).

Style group (exactly one member may be defined):

u  Field contents is converted to upper-case letters
l  Field contents is converted to lower-case letters
w  Field contents is converted to capitalized words

Suffix (may be defined):

&  Numerical value is provided with an Ordinal Number suffix

Format group (exactly one member must be defined):

*  Field contents is not cut after position width

#  Field contents is cut after position width

If a given width component is greater 256 but less than 999, a field width of 256 is assumed. But if a given width component is zero, this interrupts the interpretation phase of Gcal so that the complete format is not interpreted anymore and is therefore displayed respectively assigned as is. The same happens if a component is misspelled, given in wrong order, or one of the required components is unspecified.

Furthermore, a style component may be specified case-independently. See [Calendar option --translate-string=text], page 37, for further annotations how the style component is used correctly.

Appendix K Coding Scheme

The following coding scheme tables explain the relations between the date part (yyyy …) of a resource file line and the modifiers, the short-style option -c|C[] —which is used for enabling the fixed date feature— respectively the --period-of-fixed-dates=argument long-style option may have, i.e. they point out all possible correspondences. This is necessary because Gcal does not respect all entries found in a resource file if a fixed date argument is given, it respects only those entries which have a definite relation to the fixed date argument to avoid the output of redundant information! See [Fixed dates option

--period-of-fixed-dates=argument], page 55, for further details.

In the date part yyyy …:

yyyy represents any year number (0001…9999).

mm represents any month number (01…12 or 99) respectively any textual short month name with length of 3 characters (Jan, Feb…).

dd represents any day number (01…28, 29, 30, 31 or 99).

www represents any textual short weekday name with length of 2…3 characters

          (Mon. . .Sun).

? represents any case-insensitive single letter.

n represents any positive integer number (one or more digits).

K.1 Coding Scheme Table 1

This table opposes all possible encode variations of a date given in the date part of resource file with all fitting modifiers of the short-style option -c|C[].

Please note:

The -c, -ct, -cnd, -c*d|wn[www], -c@?[[+|-]n[www]], -cmmdd and -cmmwwwn options respect

all encode variations of the date part!

Date part  Short-style option -c|C[]

00000000                    -cn+|-
000000dd                    -cnw|+|-, -cw[+|-], -cm[+|-]
0000mm00                    -cn+|-

0000mmdd

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn
yyyy0000                    -cn+|-
yyyy00dd                    -cnw|+|-, -cw[+|-], -cm[+|-]
yyyymm00                    -cn+|-

yyyymmdd

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn
000000wwwn                  -cnw|+|-, -cw[+|-], -cm[+|-]

0000mmwwwn

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn
yyyy00wwwn                  -cnw|+|-, -cw[+|-], -cm[+|-]

yyyymmwwwn

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn

0000*dn[www]

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn

yyyy*dn[www]

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn

0000*wn[www]

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn

yyyy*wn[www]

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn

0000@?[+|-]n[www]

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn

yyyy@?[+|-]n[www]

                            -cnw|+|-, -cw[+|-], -cm[+|-], -cy[+|-], -clnd|w|+|-,
                            -cl*d|wn[www], -cl@?[[+|-]n[www]], -clmmdd, -clmmwwwn

K.2 Coding Scheme Table 2

This table opposes all fitting modifiers of the short-style option -c|C[] with all possible encode variations of a date given in the date part of resource file.

Short-style option -c|C[]  Date part

-c
-ct
-cnd|+|-
-cmmdd
-cmmwwwn
-c*d|wn[www]
-c@?[[+|-]n[www]]

                            00000000, 0000mm00, yyyy0000, yyyymm00, 000000dd,
                            0000mmdd, yyyy00dd, yyyymmdd, 000000wwwn,
                            0000mmwwwn, yyyy00wwwn, yyyymmwwwn, 0000*dn,
                            0000*dnwww, yyyy*dn, yyyy*dnwww, 0000*wn,
                            0000*wnwww, yyyy*wn, yyyy*wnwww, 0000@?[+|-]n,
                            0000@?[+|-]nwww, yyyy@?[+|-]n, yyyy@?[+|-]nwww

-cnw
-cw[+|-]

                            000000dd, 0000mmdd, yyyy00dd, yyyymmdd, 000000wwwn,
                            0000mmwwwn, yyyy00wwwn, yyyymmwwwn, 0000*dn,
                            0000*dnwww, yyyy*dn, yyyy*dnwww, 0000*wn,
                            0000*wnwww, yyyy*wn, yyyy*wnwww, 0000@?[+|-]n,
                            0000@?[+|-]nwww, yyyy@?[+|-]n, yyyy@?[+|-]nwww

-cm[+|-]

                            000000dd, 0000mmdd, yyyy00dd, yyyymmdd, 000000wwwn,
                            0000mmwwwn, yyyy00wwwn, yyyymmwwwn, 0000*dn,
                            0000*dnwww, yyyy*dn, yyyy*dnwww, 0000*wn,
                            0000*wnwww, yyyy*wn, yyyy*wnwww, 0000@?[+|-]n,
                            0000@?[+|-]nwww, yyyy@?[+|-]n, yyyy@?[+|-]nwww

-cy[+|-]
-clnd|w|+|-
-clmmdd
-clmmwwwn
-cl*d|wn[www]
-cl@?[[+|-]n[www]]

                            0000mmdd, yyyymmdd, 0000mmwwwn, yyyymmwwwn,
                            0000*dn, 0000*dnwww, yyyy*dn, yyyy*dnwww, 0000*wn,
                            0000*wnwww, yyyy*wn, yyyy*wnwww, 0000@?[+|-]n,
                            0000@?[+|-]nwww, yyyy@?[+|-]n, yyyy@?[+|-]nwww

Appendix L Environment Variables

Here is a table of all environment variables which are used by Gcal:

GCAL

GCAL_OPT Flags which are passed to the Gcal program automatically.

GCAL_COLUMNS

COLUMNS Sets the number of columns on the screen. Takes precedence over the number of columns specified by the TERM environment variable or the system itself.

GCAL_DATE_FORMAT

Contains the date format text which defines the ordering of a displayed date. Takes precedence over the default date format text.

GCAL_LINES

LINES Sets the number of lines on the screen. Takes precedence over the number of lines specified by the TERM environment variable or the system itself.

GCAL_SYS_DATADIR

Name of a file directory absolute to the root/$prefix file directory, where the system-administrator stores common, public access resource files (dates files).

GCAL_USR_DATADIR

Name of a file directory relative to the HOME file directory, where the user stores personal, respectively private access resource files (dates files). Takes prece-

dence over GCAL_SYS_DATADIR.

GCALANSI If defined or set, ANSI escape highlighting sequences are used respectively substituted by default if no other highlighting sequences are defined by command line option or found in the Termcap library.

GCALPATH (‘:’) colon-separated list of file directories which are searched to find the .gcalrc file. Takes precedence over HOME, GCAL_USR_DATADIR and

GCAL_SYS_DATADIR.

GCALPROG Contains the file name of the executable Gcal program that is used by the Tcal program to call the Gcal program. Takes precedence over the file name gcal which is burned-in during the compilation step of the Tcal program. The GCALPROG environment variable must always be set if the Gcal program is installed under another name than the standard name gcal, otherwise Tcal is unable to execute the Gcal program automatically!

HOME Name of the user’s home file directory. This is used to find the .gcalrc file.

Takes precedence over GCAL_USR_DATADIR and GCAL_SYS_DATADIR.

LANGUAGE The national language and other cultural-specific conventions used by Gcal at run-time. Takes precedence over LC_ALL, LC_MESSAGES and LANG.

LC_ALL The national language and other cultural-specific conventions used by Gcal at run-time. Takes precedence over LC_MESSAGES and LANG.

LC_MESSAGES

The national language and other cultural-specific conventions used by Gcal at run-time. Takes precedence over LANG.

LANG The national language and other cultural-specific conventions used by Gcal at run-time.

LOGNAME The eMail address which is used if the mailing option is wanted. See [Global option --mail[=address]], page 14, for more details.

MAILPROG Name of the mailer program which is used if the mailing option is wanted. Takes precedence over the file name mail which is burned-in during the compilation step of the Gcal program. See [Global option --mail[=address]], page 14, for more details.

MAILTO The eMail address which is used if the mailing option is wanted. Takes precedence over USER and LOGNAME. See [Global option --mail[=address]], page 14, for more details.

PAGER Name of the external pager program which is used if the paging option is wanted. See [Global option --pager], page 11, for more details.

PATH (‘:’) colon-separated list of file directories which are searched to find the external pager program. See [Global option --pager], page 11, for more details.

TERM The type of terminal on which the Gcal program is being run.

TXT2GCALPROG

Contains the file name of the executable Txt2gcal program that is used by the Gcal program to call the Txt2gcal program. Takes precedence over the file name txt2gcal which is burned-in during the compilation step of the Gcal program. The TXT2GCALPROG environment variable must always be set if the Txt2gcal program is installed under another name than the standard name txt2gcal, otherwise Gcal is unable to execute the Txt2gcal program automatically! See Section 5.1.9 [Text variables], page 106, for further details.

TZ The timezone information which is added to the date text of sent eMails, created response files and created shell script files.

USER The eMail address which is used if the mailing option is wanted. Takes precedence over LOGNAME. See [Global option --mail[=address]], page 14, for more details.

Appendix M Error Codes

Gcal produces the following exit codes:

127 Request for help screen, software license or program version (adjustable). 126 Invalid command line argument / option. 125 Year out of range for computing the Easter Sunday’s date. 124 Virtual memory exhausted. 123 Invalid date part in resource file or in include file. 122 Invalid month field entry in resource file or in include file. 121 Invalid day field entry in resource file or in include file. 120 Malformed #include statement in resource file or in include file. 119 Illegal recursive/cyclic #include statement in resource file or in include file. 118 Resource file or include file not found. 117 Invalid n’th weekday field (only 1…5 or 9 valid). 116 Missing whitespace separator character between date part and text part. 115 Response file respectively shell script cannot be written. 114 Invalid date variable operation. 113 Referenced date variable undefined. 112 Invalid date value assigned to date variable. 111 Sending output to eMail address failed. 110 Internal error, C function pipe(), fork(), exec*() or unlink() failed. 109 Error reading resource, include respectively response file. 108 Illegal character found in response file. 107 Internal error, invalid value for the size of an internal table occurred. 106 Invalid date format text specified. 105 Invalid text variable definition. 104 Invalid regular expression search pattern specified. 3 Program aborted by signal. 2 Exit code not equal zero returned by external command. 1 Activated -c|C[], -f|F name[+…] or -# line option caused no output of re-

source file entries.

0 Normal program termination.

Appendix N Glossary

Here is an alphabetical list of most terms with short description that are introduced and used in this manual:

Actual date modifier Gcal uses that date as today’s date which is declared by the actual date modifier, instead of using the date that is given by computer’s clock as system date.

Alternation operator The ‘|’ metacharacter is used to specify alternatives in a regular expression.

Amateur-astronomical twilight The time when the center of the Sun passes a reference altitude of 15 degrees below a mathematical horizon.

Anchor The ‘^’ and ‘$’ metacharacters in a regular expression.

ANSI The American National Standards Institute. This organization produces many standards, among them the standards for the C and C++ programming languages.

Asr prayer time The third Islamic prayer time. Begins in the afternoon when the length of the shadow cast by a vertical pole has a definite length, and ends just before sunset.

Astronomical twilight The time when the center of the Sun passes a reference altitude of 18 degrees below a mathematical horizon.

Astronomical unit An astronomical unit, abbreviated by ae, is equal to the mean distance of the Sun from the Earth, which is about 149,597,870.691 kilometers.

Autumnal equinox point on the ecliptic The point where the descending Sun’s orbit intersects the celestial equator.

Azimuth The horizontal angular distance between the vertical circle, that passes a celestial body, and the North point.

Biorhythm Ascending and descending curves which show three coherent biological cycles, the physical, the emotional and the intellectual cycle.

BSD Short term for Berkeley Software Distribution, a Unix operating system variant.

C The system programming language that most GNU software is written in.

Calendar option class The option which control the calendar layout.

Celestial axis The connecting line which runs between the celestial North and South Pole.

Celestial body elevation The vertical angular distance between the celestial body and the horizon.

Celestial equator A celestial great circle, which is vertical to the celestial axis and simultaneously represents the Earth’s equator that is projected onto the celestial sphere.

Celestial Pole The northern and southern lengthening of the Earth’s rotation axis, respectively, projected on the celestial sphere.

Character class A character class in a regular expression is a special notation for describing lists of characters that have a specific attribute, but where the actual characters themselves can vary from country to country and/or from character set to character set.

Character list Matches any one of the characters that are enclosed in the square brackets in a text, which is processed by a regular expression.

Character set The set of numeric codes used by a computer system to represent the characters (letters, numbers, punctuation, etc.) of a particular country or place. The most common character set in use today is ASCII (American Standard Code for Information Interchange). Many European countries use an extension of ASCII known as ISO-8859-1 (ISO Latin-1).

Civil Islamic calendar The arithmetically determined Islamic lunar calendar.

Civil midnight time In terms of Gcal, this is 0 o’clock, the day change just took part. This clocktime is alternatively also denoted as 24 o’clock, or 12pm.

Civil twilight The time when the center of the Sun passes a reference altitude of 6 degrees below a mathematical horizon.

Collating element A special sequence that can appear in a character lists of a regular expression. This apply to non-ascii character sets, which can have single symbols that are represented with more than one character, as well as several characters that are equivalent for collating, or sorting, purposes.

Collating symbol A multi-character collating element of a regular expression that is enclosed in ‘[.’ and ‘.]’ metacharacters.

Command line argument A special text that is mostly given on the command line to run Gcal, and which serves for controlling the program.

Command A special type of command line argument which controls the periods Gcal respects.

Common option class The standard options all GNU software should implement at least partially.

Conjunction A celestial body is either behind the Sun or is between Sun and Earth (seen from the Earth), and is therefore non-visible from the Earth. In terms of the Moon, this is the New Moon phase.

Critical day Those days in the biorhythm in which one or more of the biological cycles crosses the zero point. At that time, one’s system is said to be in a state of flux and it may be desirable to exhibit caution.

Cycle mode The dynamical, i.e. depending on the respective clocktime, Sun and Moon oriented special texts are calculated for a series of clocktimes for the current day, i.e. determination is done for any time interval that can be individually defined by the user.

Date part That part of a Gcal resource file line that contains the information at which date a fixed date happens.

Decimal system Base-ten notation, where the digits are 0…9.

Declination The vertical angular distance between the center of a celestial body and the celestial equator.

Delta-t Delta-t is the difference between the Terrestrial Dynamical time (abbreviated by TDT), that was formerly known as Ephemeris time (abbreviated by ET), and the Universal time (UT). Thus, ‘delta-t ≡ TDT - UT’.

Early years In terms of eternal holidays that are determined by Gcal, these are the years within the period AD 30 until the midth of the 20th century.

Ecliptic latitude The vertical angular distance between the center of a celestial body and the ecliptic.

Ecliptic longitude The vertical angular distance between the center of a celestial body and the vernal equinox point on the ecliptic.

Ecliptic The zodiacal line or Sun’s orbit.

Elongation The horizontal angular distance between the center of a celestial body and the Sun’s center, as seen from the Earth.

Ending date A part of a Gcal special text that defines the last or latest date at which the fixed date may occur.

Environment A collection of strings, of the form name=value, that each program has available to it. Users generally place values into the environment in order to provide information to various programs. Typical examples are the environment variables

HOME and PATH.

Equation of time A correction to be added to the true solar time —as read on a sundial— to obtain the mean solar time.

Equivalence class A list of equivalent characters of a regular expression that is enclosed in ‘[=’ and ‘=]’ metacharacters.

Escape sequence A special sequence of characters used for describing non-printing characters, such as \E or \033 or \x1B for the ASCII ESC (escape) character itself.

Evening width Set width The horizontal angular distance between the set azimuth of a celestial body

and the West direction.

Exclusive day period Exclusive date period

A Gcal special text that can be specified with the effect, that only those fixed dates are respected and displayed which are not part of a definite period.

Extended file name A name that contains one ore more ‘/’ characters, and which denotes a file, whose fixed access path either starts from the root file directory, like ‘/foo/bar/file’, or from the actual file directory, like ‘./bar/file’ or

‘../bar/file’.

Extended language code information A simple language code information, which additionally has a territory information that is separated by a ‘_’, like en_us for american English.

Fajr prayer time The first Islamic prayer time. Begins with the beginning of astronomical morning twilight, and ends just before sunrise.

Fixed date option class The options which control the fixed date layout and intensity.

Format instruction Text variables, special texts and date formats can be provided with such an instruction text that results the special representation of the contents of text variables, and the evaluated special texts and date formats.

FSF Free Software Foundation

A non-profit organization dedicated to the production and distribution of freely distributable software. It was founded by Richard M. Stallman, the author of the original Emacs editor.

Geocentric data Observations of celestial bodies that are done from the fictitious center of the Earth yield in geocentrically based values.

Geocentrical horizon That fictitious horizontal plane that passes through the Earth’s center, orthogonal to the observer’s local vertical.

Geometrical horizon Mathematical horizon Mathematical-geocentric horizon

A purely geometrically-built horizon which disregards the phenomenon of refraction as it arises in reality by the influence of the Earth’s atmosphere.

Global date variable A Gcal date variable, which is defined in the command line. Is visible in all referenced resource files.

Global option class The options which modify the program output.

Global text variable A Gcal text variable, which is defined in the command line. Is visible in all referenced resource files.

Gnomon A vertical pole which is used to ascertain the length of a shadow.

GNU GNU’s not Unix. An on-going project of the Free Software Foundation to create a complete, freely distributable, POSIX-compliant computing environment.

GPL General Public License

The terms under which Gcal and its source code may be distributed.

Gregorian calendar The calendar which was introduced in AD 1582 by Pope Gregory XIII, that has an average year length of 365.2425 days. This calendar is most respected around the world today.

Hexadecimal system Base 16 notation, where the digits are 0…9 and A…F, with ‘A’ representing 10, ‘B’ representing 11, and so on up to ‘F’ for 15.

Highest culmination point The highest point above or below the horizontal plane which transits a celestial body during a day.

Highlighting sequence A control character sequence which cause a color or intensity switch in output text. Typical control character sequences are the ANSI escape sequences which have a leading escape character, and trailing more characters that define the type of the ANSI escape sequence.

Horizontal parallax The horizontal parallax of a celestial body specifies the diameter of the Earth as it is seen from the surface of the celestial body.

Hybrid calendar mode The automatical change from one calendar system to another if there are relations to a date for which the other calendar system is valid. Thus, Gcal automatically changes from the Julian calendar system to the Gregorian calendar system if output is related to dates after the Gregorian Reformation has happened.

Inclusive day period Inclusive date period

A Gcal special text that can be specified with the effect, that only those fixed dates are respected and displayed which are part of a definite period.

Integer number A whole number, i.e. a number that does not have a fractional part.

Interval expression One or two numbers inside braces in a regular expression.

Isha prayer time The fifth Islamic prayer time. Begins at the end of the astronomical evening twilight, and ends just before the astronomical midnight time of the Sun.

ISO ISO is the abbreviation of International Organization for Standardization.

ISO week number The week starts on a Monday, and the first week of a year is the one which includes the first Thursday; equivalently, the one which includes the 4th January.

Julian Ephemeris date A Julian date that is corrected by delta-t.

Julian calendar The calendar which was introduced in 46 BC by Julius Caesar and in the later years corrected by emperor Augustus, that has an average year length of 365.25 years. This calendar was replaced by the Gregorian calendar in AD 1582.

Julian date Julian day number

A day number that is based on the date 1st January 4713 BCE, which is the starting day zero of a consecutive day counting used in astronomical computations. The real zero of this date is at 12 o’clock Universal time (UTC/GMT); the day does not change at midnight, but at noon Universal time.

Leap day The 29th February in the Gregorian calendar.

Leap year In the Gregorian calendar, a leap year is any year which number can be divided by 4 without a remainder, and years ending in hundreds are no leap years unless they are divisible by 400. But the Eastern Orthodox churches compute leap years within the Gregorian calendar by using another rule; and that, a leap year is any year which number can be divided by 4 without a remainder, and years ending in hundreds are leap years, if a remainder of 2 or 6 occurs when such a year is divided by 9.

Legal holiday A day people don’t have to work — as in most companies, and businesses are closed; like ‘New Year's Day’ on January 1st.

Lists of commands Ranges of commands

Produce more than one calendar sheet per program run.

Local date variable A Gcal date variable, which is defined in a resource file. Is visible only in this specific resource file, and that file which is included by it.

Local star time Local sidereal time

The momentary period which is past between the last upper culmination of the vernal equinox point in the meridian of the observer’s location (the momentary hour angle of the vernal equinox point), thus the right ascension of the stars in the observer’s meridian at the moment.

Local text variable A Gcal text variable, which is defined in a resource file. Is visible only in this specific resource file, and that file which is included by it.

Long-style option A special —non-standard— style to write an option, e.g. --foo.

Lowest culmination point The lowest point below or above the horizontal plane which transits a celestial body during a day.

Lunar Arctic Circle The geographical latitude circle that is between about 61.5 and 71.9 decimal degrees, and the geographical pole.

Lunation Synodic month

The mean time between two consecutive New Moon phases.

Maghrib prayer time The forth Islamic prayer time. Begins at sunset, and ends at the beginning of the Isha prayer time.

Magnitude unit Is used to define the brightness of a star, and is a non-metrical value (Latin term magnitudo, abbreviated m). The difference between two consecutive magnitudes is 1 to 2.512. Therefore, a star with the brightness of 1m is 2.512 times brighter than a star of 2m. A negative magnitude denotes a very bright star, for example almost -27m for the Sun, whereas the hardly visible planet Pluto has a magnitude of a bit more than +14m. The Full Moon has a visual brightness of about -12m.55.

Marking character Single, printable characters which lead and trail the output text for marking purposes.

Mean Earth equator radius The mean radius of Earth at the equator is about 6,378.137 kilometer.

Mean solar time Mean solar day Mean sun A star day is also known as a mean solar day. Because the Sun apparently shifts

with respect to the vernal equinox point on the ecliptic due to the Earth’s orbit around the Sun, the star day and the true solar day have a different length. As the true Sun namely moves irregularly through the ecliptic, a fictitious mean Sun with a symmetrical motion through the celestial equator is used for deriving the mean solar time.

Meridian Either a geographical longitude or an astronomical vertical circle.

Metacharacters Regular expression operators

Special characters in a regular expression.

Midnight height The elevation of a celestial body as it happen at astronomical midnight time of the celestial body.

Midnight phase angle The phase angle of a celestial body as it happen at astronomical midnight time of the celestial body.

Modifier One or more characters which directly trail a short-style option, and which define a special mode of operation for this option.

Morning width Rise width

The horizontal angular distance between the rise azimuth of a celestial body and the East direction.

Nautical twilight The time when the center of the Sun passes a reference altitude of 12 degrees below a mathematical horizon.

Negated character list Complemented character list

Matches any one of the characters, except those in the square brackets, in a text, that is processed by a regular expression.

Negative day Those days in the biorhythm in which one or more of the biological cycles have a negative maximum value. At that time, one’s system is said to be in a lessen state.

Noon height The elevation of a celestial body as it happen at astronomical noon time of the celestial body.

Noon line The straight line which connects the North and the South point.

Noon phase angle The phase angle of a celestial body as it happen at astronomical noon time of the celestial body.

Octal system Base-eight notation, where the digits are 0…7.

Opposition The Earth is in one line between the celestial body and the Sun, and therefore, the celestial body is completely visible from the Earth. In terms of the Moon, this is the Full Moon phase.

Option A special type of command line argument which controls how Gcal behaves.

Orthodrome A great circles arc that is casted between two points on a surface of a sphere. Is the shortest geodetic connecting line between two points on a sphere.

Pattern A pattern is an arbitrary conditional expression against which input is tested. If the condition is satisfied, the pattern is said to match the input. A typical pattern might compare the input against a regular expression.

Phase angle The phase angle states the phase of a celestial body, i.e. the illuminated fraction of it as it can be seen from the Earth.

Pipelining The output of a program is used as an input data stream for another program.

Positive day Those days in the biorhythm in which one or more of the biological cycles have a positive maximum value. At that time, one’s system is said to be in a raised state.

POSIX The name for a series of standards being developed by the IEEE that specify a Portable Operating System interface. The IX denotes the Unix heritage of these standards.

Proleptic calendar mode This means, Gcal only uses a definite calendar system during a definite period, although there was a change to another calendar system in the historic reality during this definite period.

Rational number Floating Point number

This is a number that can have a fractional part.

RFC RFC is the abbreviation of Request for Comments.

Redirection The output of a program is sent to another device instead of sending it to the standard output channel.

Refraction Refraction is an optical phenomenon caused by the Earth’s atmosphere, which leads to an apparent raising of the location of a celestial body. The amount of refraction increases with growing air pressure and sinking temperature and vice-versa.

Regexp Regular expression

A way of describing a set of strings.

Relative time offset value A time offset value which is related to a base time.

Religious Islamic calendar The astronomically determined Islamic lunar calendar.

Resource file A file that essentially contains the fixed date entries.

Response file A file used by Gcal to preload options and commands at program start-up. A response file may contain options and commands, but no references to further response files.

Right ascension The horizontal angular distance between the center of a celestial body and the hour circle that passes through the vernal equinox point on the ecliptic.

Rise azimuth The azimuth of a celestial body as it happen at rise time of the celestial body.

Rise height The elevation of a celestial body as it happen at rise time of the celestial body.

Rise phase angle The phase angle of a celestial body as it happen at rise time of the celestial body.

Set azimuth The azimuth of a celestial body as it happen at set time of the celestial body.

Set height The elevation of a celestial body as it happen at set time of the celestial body.

Set phase angle The apparent phase angle of a celestial body as it happen at set time of the celestial body.

Shell The command processor of the operating system. This operating system program interprets and runs given command lines.

Short-style option The traditional —standard— style to write an option, e.g. -x.

Simple file name A name that denotes a file, whose access-path either does not start from the root file directory or from the actual file directory, like ‘file’ or ‘bar/file’.

Simple language code information A language code that consists of two-letters like de for the German language, as it is defined in the ISO-639:1988.

Single command 3-Month mode command

Produce only one calendar sheet per program run.

Solar Arctic Circle The geographical latitude circle that is between about 66.7 and 65.9 (corrected for atmospheric refraction) decimal degrees, and the geographical pole. For computing the different twilight phases, this latitude circle decreases to about:

• 60.7 decimal degrees for civil twilight.

• 54.7 decimal degrees for nautical twilight.

• 51.7 decimal degrees for amateur-astronomers twilight.

• 48.7 decimal degrees for astronomical twilight.

Special text A text in a Gcal resource file which is started by a ‘%’ character and whose purpose is to suppress output of fixed dates in definite cases, either, or to provide them with particular texts respectively to start external commands.

Summer point on the ecliptic Summer solstice Solstice in summer

The Sun’s orbit crosses the highest point above the celestial equator that is passed during a revolution of the Earth around the Sun.

Standard week number The week starts on the respective starting day of the week, and the days in a new year that are preceding the first starting day of the week are in the last week of the previous year, respectively in week zero of the new year.

Star day The period between two consecutive upper culminations of the vernal equinox point on the ecliptic in the meridian of the observer’s location.

Starting date A part of a Gcal special text that defines the first or earliest date at which the fixed date may occur.

Statute mile A linear measure, one statute mile is equal to 1.609344 kilometer.

String A chain of characters.

Switch character An option is defined by a leading switch character, either the ‘-’ or the ‘/’ character for traditional short-style options, or ‘--’ for mnemonic long-style options; a command may not have a leading switch character.

Text part That part of a Gcal resource file line that contains the fixed date text.

Topocentric data Observations of celestial bodies that are done from the surface of the Earth yield in topocentrically based values.

True solar time True solar day

A true solar day is the period between two consecutive lower culminations of the Sun. This entity is taken as the base for deriving the true solar time (as it is also shown by a sundial during the day).

True track Course angle

The course angle gives the direction, which is the vertex of a great circles arc (Orthodrome) that is casted between two points on a surface of a sphere, at a meridian (Gcal uses a geoid (rotation ellipsoid) that is calculated taking pattern from the World Geodetic System, short W.G.S, that was designed in 1961). Such an arc is the shortest geodetic connecting line between two points on a sphere. Indeed, the course angles on such an arc are changing continually, because the circular line that is casted intersects each meridian with another angle — except the case, that both co-ordinates are referring to exactly the same longitude or to exactly the same latitude.

Twilight The interval of time before sunrise and again after sunset, during which natural light is provided by the upper atmosphere that does receive direct sunlight and reflects parts of it towards the Earth’s surface.

Unix A computer operating system originally developed in the early 1970’s at AT&T Bell Laboratories. It initially became popular in universities around the world, and later moved into commercial environments as a software development system and network server system. There are many commercial versions of Unix, as well as several work-alike systems whose source code is freely available (such as Linux, NetBSD, and FreeBSD).

Vernal equinox point on the ecliptic The point where the ascending Sun’s orbit intersects the celestial equator.

Whitespace character For example Tab, Space, Form-feed and similar characters.

Winter point on the ecliptic Winter solstice Solstice in winter

The Sun’s orbit crosses the lowest point below the celestial equator that is passed during a revolution of the Earth around the Sun.

Word A sequence of one or more letters, digits, or underscores (‘_’) in a regular expression.

Zuhr prayer time The second Islamic prayer time. Begins after the astronomical noon time of the Sun, and ends at the beginning of the Asr prayer time.

Command Line Argument Index 271

Command Line Argument Index

Here is an alphabetical list of the command line arguments processed by Gcal:

%

%date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

-

-# line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 -*dn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -*dnwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 -*wn[www] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

--adjust-value=argument. . . . . . . . . . . . . . . . . . . . . 39

--alternative-format . . . . . . . . . . . . . . . . . . . . . . . . 50

--astronomical-holidays. . . . . . . . . . . . . . . . . . . . . 15
--atmosphere=air-pressure[,temperature] . . . . 40

--bahai-holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 --bahai-months . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

--biorhythm-axis=number. . . . . . . . . . . . . . . . . . . . . 49

--blocks=1. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --blocks=12 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --blocks=2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --blocks=3. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --blocks=4. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --blocks=6. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --blocks=number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --calendar-dates=both . . . . . . . . . . . . . . . . . . . . . . . 32

--calendar-dates=special . . . . . . . . . . . . . . . . . . . 32
--calendar-dates=special|both . . . . . . . . . . . . . . 32

--cc-holidays=cc[_tt][+…] . . . . . . . . . . . . . . . . . . 17 --celtic-holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

--chinese-flexible-holidays . . . . . . . . . . . . . . . . 15
--chinese-flexible-months . . . . . . . . . . . . . . . . . . 28

--chinese-holidays . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --chinese-months. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 --christian-holidays . . . . . . . . . . . . . . . . . . . . . . . . 16 --coptic-months . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 --copyleft. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --cycle-end=argument . . . . . . . . . . . . . . . . . . . . . . . . 41 --cycle-step=argument . . . . . . . . . . . . . . . . . . . . . . . 42

--date-format=de|us|gb|text . . . . . . . . . . . . . . . . 35
--date-variable=argument . . . . . . . . . . . . . . . . . . . 37

--debug=abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --debug=all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --debug=handled . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --debug=internal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --debug=unhandled. . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

--debug[=internal|handled|unhandled|all|abort]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

--descending-fixed-dates=long . . . . . . . . . . . . . . 45
--descending-fixed-dates=short. . . . . . . . . . . . . 45
--descending-fixed-dates[=short|long]. . . . . . 45
--descending-holiday-list=long. . . . . . . . . . . . . 15
--descending-holiday-list=short . . . . . . . . . . . 15
--descending-holiday-list[=long|short] . . . . 14

--disable-highlighting . . . . . . . . . . . . . . . . . . . . . . 12 --end-of-month . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --end-of-week . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --end-of-year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 --ethiopic-months. . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

--exclude-fixed-dates-list-title . . . . . . . . . . 53
--exclude-holiday-list-title . . . . . . . . . . . . . . . 15

--execute-command. . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

--exit-status-help-non-zero . . . . . . . . . . . . . . . . 10
--export-date-variables . . . . . . . . . . . . . . . . 38, 105
--export-text-variables . . . . . . . . . . . . . . . . 39, 111

--filter-day=argument . . . . . . . . . . . . . . . . . . . . . . . 44

--filter-period=argument . . . . . . . . . . . . . . . . . . . 44

--filter-text=pattern . . . . . . . . . . . . . . . . . . . . . . . 44 --fixed-dates=both . . . . . . . . . . . . . . . . . . . . . . . . . . 32 --fixed-dates=special . . . . . . . . . . . . . . . . . . . . . . . 32

--fixed-dates=special|both . . . . . . . . . . . . . . . . . 32

--force-highlighting . . . . . . . . . . . . . . . . . . . . . . . . 12

--french-revolutionary-months . . . . . . . . . . . . . . 28
--gregorian-reform=1582|1700|1752|1753|argument

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 --grouping-text[=text] . . . . . . . . . . . . . . . . . . . . . . . 51 --heading-text=text . . . . . . . . . . . . . . . . . . . . . . . . . 53 --hebrew-holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --hebrew-months . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 --help . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 127, 128, 129 --here=line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 --highlighting=no. . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --highlighting=text . . . . . . . . . . . . . . . . . . . . . . . . . 12 --highlighting=yes . . . . . . . . . . . . . . . . . . . . . . . . . . 12 --holiday-dates=both . . . . . . . . . . . . . . . . . . . . . . . . 32

--holiday-dates=special. . . . . . . . . . . . . . . . . . . . . 32
--holiday-dates=special|both . . . . . . . . . . . . . . . 32

--holiday-list=long . . . . . . . . . . . . . . . . . . . . . . . . . 15 --holiday-list=short . . . . . . . . . . . . . . . . . . . . . . . . 15

--holiday-list[=long|short] . . . . . . . . . . . . . . . . . 14

--ignore-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

--include-consecutive-number . . . . . . . . . . . . . . . 54
--include-holidays=long. . . . . . . . . . . . . . . . . . . . . 50
--include-holidays=short . . . . . . . . . . . . . . . . . . . 50
--include-resource-file-name . . . . . . . . . . . . . . . 50

--include-today . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 --include-week-number . . . . . . . . . . . . . . . . . . . . . . . 52 --indian-civil-months . . . . . . . . . . . . . . . . . . . . . . . 28

--islamic-civil-holidays . . . . . . . . . . . . . . . . . . . 16

--islamic-civil-months . . . . . . . . . . . . . . . . . . . . . . 28 --iso-week-number=no . . . . . . . . . . . . . . . . . . . . . . . . 31 --iso-week-number=yes . . . . . . . . . . . . . . . . . . . . . . . 31

--iso-week-number=yes|no . . . . . . . . . . . . . . . . . . . 30
--japanese-flexible-holidays . . . . . . . . . . . . . . . 16
--japanese-flexible-months . . . . . . . . . . . . . . . . . 28

--japanese-holidays . . . . . . . . . . . . . . . . . . . . . . . . . 16 --japanese-months. . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 --leap-day=february . . . . . . . . . . . . . . . . . . . . . . . . . 55

--leap-day=february|march . . . . . . . . . . . . . . . . . . 55

--leap-day=march. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 --license . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --limit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 --list-mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

--list-of-fixed-dates=long . . . . . . . . . . . . . . . . . 45
--list-of-fixed-dates=short . . . . . . . . . . . . . . . . 45
--list-of-fixed-dates[=short|long] . . . . . . . . . 45

--long-help=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --long-help[=argument] . . . . . . . . . . . . . . . . . . . . . . . 10 --mail[=address] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --month . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

--moonimage-lines=number . . . . . . . . . . . . . . . . . . . 49
--multicultural-new-year-holidays . . . . . . . . . 16

--old-armenic-months . . . . . . . . . . . . . . . . . . . . . . . . 29 --old-egyptic-months . . . . . . . . . . . . . . . . . . . . . . . . 29

--omit-multiple-date-part . . . . . . . . . . . . . . . . . . 52

--orthodox-calendar . . . . . . . . . . . . . . . . . . . . . . . . . 30

--orthodox-new-holidays. . . . . . . . . . . . . . . . . . . . . 16
--orthodox-old-holidays. . . . . . . . . . . . . . . . . . . . . 16

--pager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

--period-of-fixed-dates=argument . . . . . . . . . . 55
--persian-jalaali-holidays . . . . . . . . . . . . . . . . . 16
--persian-jalaali-months . . . . . . . . . . . . . . . . . . . 29

--precise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 --resource-file=name[+…] . . . . . . . . . . . . . . . . . . 46 --response-file=name . . . . . . . . . . . . . . . . . . . . . . . . 10 --revert-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 --shell-script=name . . . . . . . . . . . . . . . . . . . . . . . . . 10 --shift=[+|-]number . . . . . . . . . . . . . . . . . . . . . . . . . 127 --start-of-month. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 --start-of-week . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --start-of-year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

--starting-day=argument. . . . . . . . . . . . . . . . . . . . . 33

--suppress-calendar . . . . . . . . . . . . . . . . . . . . . . . . . 31 --suppress-date-part . . . . . . . . . . . . . . . . . . . . . . . . 52

--suppress-fixed-dates-list-separator. . . . . 52
--suppress-holiday-list-separator . . . . . . . . . 15

--suppress-text-part . . . . . . . . . . . . . . . . . . . . . . . . 53

--text-variable=argument . . . . . . . . . . . . . . . . . . . 38

--time-offset=argument . . . . . . . . . . . . . . . . . . . . . . 33 --today . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 --tomorrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

--transform-year=argument . . . . . . . . . . . . . . . . . . 34
--translate-string=text. . . . . . . . . . . . . . . . . . . . . 37
--type=special|standard. . . . . . . . . . . . . . . . . . . . . 29

--usage=? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --usage[=argument] . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --version . . . . . . . . . . . . . . . . . . . . . . 10, 127, 128, 129 --week . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --with-week-number . . . . . . . . . . . . . . . . . . . . . . . . . . 30 --year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 --zero-dates-only. . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

--zodical-marker-holidays . . . . . . . . . . . . . . . . . . 16

-? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -??. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -@dvar[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

-@dvar[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
-@e[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
-@e[[+|-]n]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59
-@t[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
-@t[[+|-]n]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
-a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
-A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
-b 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
-b 12 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
-b 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
-b 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
-b 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
-b 6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
-b number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
-C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
-c*dn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
-c*dnwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57
-c*wn[www]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58
-c-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
-C-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
-c@dvar[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
-c@dvar[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
-c@e[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
-c@e[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59
-c@t[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
-c@t[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
-c[m|M[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63
-c[t|T] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
-c[w|W[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63
-c[y|Y[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64
-c|C[-] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45
-ca. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
-cA. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
-cd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
-ce. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
-cE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
-cg[text] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51
-cJ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53
-ck. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
-cl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
-cl*dn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
-cl*dnwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58
-cl*wn[www] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58
-cl@dvar[+|-]nwww. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
-cl@dvar[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
-cl@e[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
-cl@e[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60
-cl@t[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
-cl@t[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
-clmmdd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57
-clmmwwwn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57
-cln+|- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
-clnd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
-clnw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
-cm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63
-cm+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63
-cm- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

Command Line Argument Index 273

-cmmdd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 -cmmwwwn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 -cn+|- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 -cnd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -cnw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 -co. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -cQ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -ct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 -cT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 -cU. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -cw. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -cw+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -cw- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -cx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 -cy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -cy+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -cy- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -cz. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 -cZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 -d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -D argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 -e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 -E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 -f|F name[+…] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 -G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -g[text] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 -h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -H no . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -H text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -H yes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 -hh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 -I pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 -i[-] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 -j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -J . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 -j[b] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -jb. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -jc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -jc[b] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -jcb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -jn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -jn[b] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -jnb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 -k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -K . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 -l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -l*dn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -l*dnwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 -l*wn[www]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 -l@dvar[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 -l@dvar[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 -l@e[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -l@e[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 -l@t[+|-]nwww . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 -l@t[[+|-]n] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 -lmmdd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 -lmmwwwn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 -ln+|- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 -lnd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 -lnw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 -m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -m+. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -m-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -m[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -M[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -m|M[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -n-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -N-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -n|N[-] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 -p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -P argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 -Q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -q cc[_tt][+…] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 -r argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 -R name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -s argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 -S name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 -T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 -u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 -U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 -V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 -v argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -w+. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -w-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -w[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -W[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -w|W[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 -X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -Y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -y+. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -y-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -y[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -Y[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -y|Y[+|-]] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 -z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 -Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

@

@file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

C

Command, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, . yyyy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, .+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .+ yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .- yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .. yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, : . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, :yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, mm yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, mm yyyy-yyyy . . . . . . . . . . . . . . . . . . . . . . . . 75 Command, mm,…,mm . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, mm,…,mm yyyy . . . . . . . . . . . . . . . . . . . . . 74 Command, mm,mm/yyyy,. . . . . . . . . . . . . . . . . . . . . . . . 74 Command, mm-mm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Command, mm-mm yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . 75

Command, mm-mm yyyy+yyyy . . . . . . . . . . . . . . . . . . . . 75

Command, mm/yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

                                                Command, mm/yyyy,. . .,mm/yyyy . . . . . . . . . . . . . . . 74
                                                Command, mm/yyyy-mm/yyyy . . . . . . . . . . . . . . . . . . . . 75

Command, mm/yyyy-yyyy . . . . . . . . . . . . . . . . . . . . . . . 75
Command, mm:. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73
Command, mm:;mm:;. . . . . . . . . . . . . . . . . . . . . . . . . . . . 75
Command, mm:;mm:yyyy;. . . . . . . . . . . . . . . . . . . . . . . 75
Command, mm:yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

                                                Command, mm:yyyy+mm:yyyy . . . . . . . . . . . . . . . . . . . . 75
                                                Command, mm:yyyy;. . .;mm:yyyy . . . . . . . . . . . . . . . 74

Command, month name . . . . . . . . . . . . . . . . . . . . . . . . . . 72
Command, NOTHING . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72
Command, yyyy. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73
Command, yyyy+yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . 75
Command, yyyy;…;yyyy . . . . . . . . . . . . . . . . . . . . . . 74

Environment Variable Index 275

Environment Variable Index

Here is an alphabetical list of the environment variables checked by Gcal:

C

COLUMNS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 255

G

GCAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 76, 255 GCAL_COLUMNS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 255 GCAL_DATE_FORMAT . . . . . . . . . . . . . . . . . . . . . . . . 35, 255 GCAL_LINES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 255 GCAL_OPT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76, 255 GCAL_SYS_DATADIR . . . . . . . . . . . . . . . . . . . . . . . . 82, 255 GCAL_USR_DATADIR . . . . . . . . . . . . . . . . . . . . . . . . 82, 255 GCALANSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13, 255 GCALPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82, 255 GCALPROG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127, 255

H

HOME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46, 82, 255

L

LANG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 133, 256 LANGUAGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 255 LC_ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 255 LC_MESSAGES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131, 256 LINES. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 255 LOGNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 256

M

MAILPROG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 256 MAILTO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 256

P

PAGER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 256 PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 256

T

TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 256 TXT2GCALPROG . . . . . . . . . . . . . . . . . . . . . . . . . . . 107, 256 TZ. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256

U

USER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14, 256

Table Index 277

Table Index

Here is an alphabetical list of the most import tables, summaries and other general views that can be found in this manual:

A

Actual date modifier, date formats. . . . . . . . . . 67

C

Co-ordinate, ISO-6709:1983 . . . . . . . . . . . . . . . . . 183
Coding scheme 1, date part. . . . . . . . . . . . . . . . . . 252
Coding scheme 2, date part. . . . . . . . . . . . . . . . . . 253
Country specific holidays . . . . . . . . . . . . . . . . . . . 17

D

Date format elements, obsolete individual

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247

Date format text, format elements and
    character replacement instructions . . . . 36
Date formats of actual date modifier. . . . . . . . 67
Date formats, obsolete fixed . . . . . . . . . . . . . . . 247
Date part, coding scheme 1. . . . . . . . . . . . . . . . . . 252
Date part, coding scheme 2. . . . . . . . . . . . . . . . . . 253
Date variables, operations . . . . . . . . . . . . . . . . . 105

E

Entries in resource file. . . . . . . . . . . . . . . . . . . . . 85 Environment variables . . . . . . . . . . . . . . . . . . . . . . 255 Error codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

Eternal holiday list, marking characters used

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

Event oriented texts of Moon data and times,

special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

Event oriented texts of Sun data and times,

special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Exit codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

F

Fixed date, line structure . . . . . . . . . . . . . . . . . . 84
Format elements and character replacement
    instructions in date format text . . . . . . . 36

Format Instruction. . . . . . . . . . . . . . . . . . . . . . . . . . 249

G

Geographical distance and course angle, modes

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 GNU long-style options. . . . . . . . . . . . . . . . . . . . . . . . 9

H

                                                Holidays, country specific . . . . . . . . . . . . . . . . . . 17

I

                                                ISO-6709:1983 co-ordinate . . . . . . . . . . . . . . . . . . 183

L

                                                Line structure of fixed date . . . . . . . . . . . . . . . . 84

Long-style options, GNU. . . . . . . . . . . . . . . . . . . . . . . 9

M

                                                Marking characters used in eternal holiday

list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Metasymbols, used . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

                                                Modes, geographical distance and course angle

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163

                                                Modes, Moon oriented special texts . . . . . . . . 188
                                                Modes, Sun oriented special texts. . . . . . . . . . 165
                                                Moon data and times, special event oriented

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

                                                Moon oriented special texts, modes . . . . . . . . 188
                                                Moon oriented special texts, representation

types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203

N

                                                Native language, selection . . . . . . . . . . . . . . . . . 131
                                                Native languages, supported. . . . . . . . . . . . . . . . 131

O

                                                Obsolete fixed date formats . . . . . . . . . . . . . . . . 247
                                                Obsolete individual date format elements

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 Obsolete special texts . . . . . . . . . . . . . . . . . . . . . 246

                                                Operations on date variables . . . . . . . . . . . . . . . 105
                                                Operations on text variables . . . . . . . . . . . . . . . 111

R

                                                Regular expression programming libraries,

supported . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Regular expressions . . . . . . . . . . . . . . . . . . . . . . . . 143

                                                Representation types of the Moon oriented

special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . 203

                                                Representation types of the Sun oriented

special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 Resource file, entries. . . . . . . . . . . . . . . . . . . . . . . 85 Run-time error codes . . . . . . . . . . . . . . . . . . . . . . . . 257

S

Selection of native language . . . . . . . . . . . . . . . 131
Short-style options, traditional . . . . . . . . . . . . 9
Special event oriented texts, Moon data and

times. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

Special event oriented texts, Sun data and

times. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 Special texts, obsolete . . . . . . . . . . . . . . . . . . . . 246

Sun data and times, special event oriented

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181

                                                Sun oriented special texts, modes. . . . . . . . . . 165
                                                Sun oriented special texts, representation

types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184

                                                Supported native languages . . . . . . . . . . . . . . . . . 131
                                                Supported regular expression programming

libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

T

                                                Text variables, operations . . . . . . . . . . . . . . . . . 111
                                                Traditional short-style options . . . . . . . . . . . . . 9

U

Used metasymbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

Appendix O GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008, 2010, 2011, 2013 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING

You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects.

If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public.

It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS

You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version:

A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 287

Concept Index

#

‘#’ separator character of ranges of days. . . . . . . . 90

$

$ regexp, anchor metacharacter . . . . . . . . . . . . . . . 138

%

‘%’ special text starting character . . . . . . . . . . . . . 100 %![argument] special text, shell command . . . . . 210 %"[argument] special text, numerical Universal

time timezone offset value . . . . . . . . . . . . . . . . 161 %,[date] special text, biorhythm text . . . . . . . . . 157 %;[date] special text, biorhythm text graphics

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 %=[argument] special text, numerical local time

timezone offset value . . . . . . . . . . . . . . . . . . . . . 161 %>02&*D[date] special text, day number . . . . . . . 154 %>02&*M[date] special text, month number …. 155 %>02*D[date] special text, day number . . . . . . . . 154 %>02*M[date] special text, month number . . . . . 155 %>03&*N[date] special text, day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 %>03*N[date] special text, day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 %>03*O[date] special text, Moon phase text … 156 %>04*Y[date] special text, year number . . . . . . . 155 %>1&*B[date] special text, age value . . . . . . . . . . . 158 %>1&*D[date] special text, day number . . . . . . . . 154 %>1&*E[date] special text, weekday number … 151 %>1&*I[date] special text, weekday number … 151 %>1&*J[date] special text, weekday number … 152 %>1&*M[date] special text, month number . . . . . 155 %>1&*N[date] special text, day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 %>1&*S[date] special text, weekday number … 152 %>1&*T[date] special text, weekday number … 153 %>1&*W[date] special text, weekday number … 151 %>2#K[date] special text, 2-letter weekday name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 %>3#K[date] special text, 3-letter weekday name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 %>3#U[date] special text, 3-letter month name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 %? special texts, exclusions without any argument

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 %? special texts, exclusive day period . . . . . . . . . 149 %? special texts, highlighting. . . . . . . . . . . . . . . . . . 208 %? special texts, inclusive day period . . . . . . . . . . 149 %? special texts, replacements without any

argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 %?… special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 %?… special texts, date exclusion . . . . . . . . . . . . 146 %?… special texts, description . . . . . . . . . . . . . . . 145 %?… special texts, short-list . . . . . . . . . . . . . . . . . 211 %?… special texts, summary . . . . . . . . . . . . . . . . . 211 %?… special texts, text replacement . . . . . . . . . 150 %?[date] special texts, biorhythm . . . . . . . . . . . . . 156 %?[date] special texts, exclusions with date

argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 %?[date] special texts, replacements with date

argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 %@[argument] special text, actual Universal time in hh:mm/date format. . . . . . . . . . . . . . . . . . . . . . . 160 %[format]'[argument] special text, actual local time in m[m…] format . . . . . . . . . . . . . . . . . . . . . . . 159

%[format]-[argument] special text, environment

variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 %[format].[argument] special text, actual Universal time hour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 %[format]/[argument] special text, actual Universal time minute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 %[format]? special texts, difference value . . . . . . 206 %[format]?[argument] special text, actual local time minute. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160

%[format]?[argument] special texts, actual

clocktime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 %[format]?[argument] special texts, replacements

with other argument . . . . . . . . . . . . . . . . . . . . . 158 %[format]?[date] special texts, difference value

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 %[format]?[date] special texts, Moon phase … 155 %[format]?[date] special texts, weekday number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 %[format]?argument special texts, Moon data

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 %[format]?argument special texts, Sun data … 165 %[format]_[argument] special text, actual local time hour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 %[format]`[argument] special text, actual Universal time in m[m…]/date format. . . . . . . . . . . . . . 160 %[format]{[argument] special text, actual local time 12-hour format suffix. . . . . . . . . . . . . . . . . . . . . 160 %[format]}[argument] special text, actual Universal time 12-hour format suffix. . . . . . . . . . . . . . . . 161

%[format]b[argument] special text, geographical

distance and course angle . . . . . . . . . . . . . . . . 162 %[format]D[date] special text, day number. . . . . 153 %[format]j[argument] special text, Julian day

number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 %[format]K[date] special text, weekday name . . 150 %[format]M[date] special text, month number . . 154 %[format]N[date] special text, day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 %[format]U[date] special text, month name …. 154 %[format]Y[date] special text, year number …. 155 %0 special text, end of highlighting sequence 1 or 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 %1 special text, start of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 %2 special text, end of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 %3 special text, start of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 %4 special text, end of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 %5 special text, start of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 %6 special text, end of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 %7 special text, start of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 %8 special text, end of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 %9 special text, start of highlighting sequence 1 or

2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 %A special text, exclusive Monday . . . . . . . . . . . . . 149 %a special text, inclusive Monday . . . . . . . . . . . . . 149 %B[date] special text, age value . . . . . . . . . . . . . . . 158 %C special text, exclusive Tuesday . . . . . . . . . . . . . 149 %c special text, inclusive Tuesday . . . . . . . . . . . . . 149 %d special text, relative day number. . . . . . . . . . . 207 %D[date] special text, day number. . . . . . . . . . . . . 153 %date actual date modifier . . . . . . . . . . . . . . . . . . . . . 67 %date, command line argument . . . . . . . . . . . . . . . . 67 %date, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 %E[date] special text, weekday number . . . . . . . . 151 %e[date][#[date]] special text, exclusive date period

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 %F special text, exclusive Wednesday . . . . . . . . . . 149 %f special text, inclusive Wednesday . . . . . . . . . . 149 %G special text, exclusive Thursday . . . . . . . . . . . 149 %g special text, inclusive Thursday . . . . . . . . . . . . 149 %H special text, exclusive Friday . . . . . . . . . . . . . . 149 %h special text, inclusive Friday . . . . . . . . . . . . . . . 149 %I[date] special text, weekday number . . . . . . . . 151 %i[date][#[date]] special text, inclusive date period

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 %J[date] special text, weekday number . . . . . . . . 152 %K[date] special text, complete weekday name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 %k[date] special text, week number . . . . . . . . . . . 154 %L special text, exclusive Saturday . . . . . . . . . . . . 150 %l special text, inclusive Saturday . . . . . . . . . . . . 149 %m special text, relative month number . . . . . . . . 207 %M[date] special text, month number . . . . . . . . . . 154 %n[argument] special text, textual date . . . . . . . . 162 %N[date] special text, day-of-year number . . . . . 153 %O[date] special text, Moon phase text . . . . . . . . 155 %P special text, exclusive Sunday. . . . . . . . . . . . . . 150 %p special text, inclusive Sunday . . . . . . . . . . . . . . 149 %Q special text, exclusive Monday…Thursday

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 %q special text, inclusive Monday…Thursday

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 %R special text, exclusive Monday…Friday … 150 %r special text, inclusive Monday…Friday … 149 %S[date] special text, weekday number . . . . . . . . 152 %t* special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 %t[argument] special text, actual local time in

hh:mm format . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 %T[date] special text, weekday number . . . . . . . . 152 %U[date] special text, complete month name… 154 %V special text, exclusive legal holidays. . . . . . . . 149 %v special text, inclusive legal holidays . . . . . . . . 149 %w special text, relative week number . . . . . . . . . 207 %W[date] special text, weekday number . . . . . . . . 151 %X special text, exclusive all holidays. . . . . . . . . . 149 %x special text, inclusive all holidays . . . . . . . . . . 149 %y special text, relative year number . . . . . . . . . . 207 %Y[date] special text, year number . . . . . . . . . . . . 155 %ydate special text, relative year number . . . . . 158 %Z[date] special text, Moon phase text graphics

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156

(

(…) regexp, grouping . . . . . . . . . . . . . . . . . . . . . . . 140

*

* regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

                              +

+ regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 ‘+’ separator character . . . . . . . . . . . . . . . . . . . . . 17, 46

,

‘,’ separator character of lists of days . . . . . . . . . . 89

-

‘-’ switch character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ‘--’ switch character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

.

. regexp, metacharacter . . . . . . . . . . . . . . . . . . . . . . 138 .gcalrc resource file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 ‘.n’ appearance factor of days. . . . . . . . . . . . . . . . . . 94

/

‘/’ co-ordinate termination character. . . . . . . . . . 163 ‘/’ switch character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

:

‘:’ separator character . . . . . . . . . . . . . . 12, 13, 37, 38 ‘:n’ repetition factor of days . . . . . . . . . . . . . . . . . . . 92

;

‘;’ comment character . . . . . . . . . . . . . . . . . . . . 65, 101 ‘;’ remark character . . . . . . . . . . . . . . . . . . . . . . 65, 101

?

? regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

@

@file response file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 @file, command line argument . . . . . . . . . . . . . . . . 65 @file, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

[

[…] regexp, character list . . . . . . . . . . . . . . . . . . . 138 [.element.] regexp, collating symbols . . . . . . . . 140 [:alnum:] regexp, character class. . . . . . . . . . . . . 139 [:alpha:] regexp, character class. . . . . . . . . . . . . 139 [:blank:] regexp, character class. . . . . . . . . . . . . 139 [:cntrl:] regexp, character class. . . . . . . . . . . . . 139 [:digit:] regexp, character class. . . . . . . . . . . . . 139 [:graph:] regexp, character class. . . . . . . . . . . . . 139 [:keyword:] regexp, character classes . . . . . . . . 138 [:lower:] regexp, character class. . . . . . . . . . . . . 139 [:print:] regexp, character class. . . . . . . . . . . . . 139 [:punct:] regexp, character class. . . . . . . . . . . . . 139 [:space:] regexp, character class. . . . . . . . . . . . . 139 [:upper:] regexp, character class. . . . . . . . . . . . . 139 [:xdigit:] regexp, character class. . . . . . . . . . . . 139 [=list=] regexp, equivalence classes . . . . . . . . . . 140 [^…] regexp, complemented character list . . 140

^

‘^’ line break-up character . . . . . . . . . . 44, 51, 53, 97 ^regexp, anchor metacharacter . . . . . . . . . . . . . . . 138

‘_’ pseudo space character … 39, 44, 46, 47, 51, 53

\

‘\’ depreciation character … 39, 44, 46, 47, 51, 53,

97, 98, 100, 107, 206 ‘\’ line continuation character. . . . . . . . . . . . . . . . . . 97 ‘\’ protection character . . 39, 44, 46, 47, 51, 53, 97,

98, 100, 107, 206 \ regexp, metacharacter . . . . . . . . . . . . . . . . . . . . . . 138 \' regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 \< regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 \> regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 \` regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 \b regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 \B regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 \w regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 \W regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 141

                              |

| regexp, operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 {expression} regexp, interval expressions …. 141

                              ~

‘~’ line break-up character . . . . . . . . . . 44, 51, 53, 97

1

12-hours display format of time . . . . . . . . . . . . . . . 159

2

2-letter weekday name %>2#K[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151, 212 24-hours display format of time . . . . . . . . . . . . . . . 159

3

3-letter month name %>3#U[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154, 213 3-letter weekday name %>3#K[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150, 212 3-Month mode command . . . . . . . . . . . . . . . . . . . . . 269 3-Month mode command . yyyy, special . . . . . . . 73 3-Month mode command .+ yyyy, special . . . . . . 74 3-Month mode command .+, special . . . . . . . . . . . 74 3-Month mode command ., special. . . . . . . . . . . . . 73 3-Month mode command .- yyyy, special . . . . . . 74 3-Month mode command .-, special . . . . . . . . . . . 74 3-Month mode command .. yyyy, special . . . . . . 74 3-Month mode command .., special . . . . . . . . . . . 74 3-Month mode commands . . . . . . . . . . . . . . . . . . 72, 73 3-Month mode, limitations . . . . . . . . . . . . . . . . 89, 103

A

Accuracy of astronomical calculations . . . . . . . . . . . 1 Activate fixed date function . . . . . . . . . . . . 45, 46, 48 Actual clocktime %[format]?[argument] special

texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158, 213 Actual date modifier. . . . . . . . . . . . . . . . . . . . . . . . . . 259 Actual date modifier, %date. . . . . . . . . . . . . . . . . . . . 67 Actual file directory . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Actual local time 12-hour format suffix

                                                    %[format]{[[+|-]mmmm|hh:[mm]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 Actual local time 12-hour format suffix

%[format]{[argument] special text. . . . . . . . . 160 Actual local time hour

%[format]_[*][[+|-]mmmm|hh:[mm]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 Actual local time hour %[format]_[argument]

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Actual local time in hh:mm format %t[argument]

special text . . . . . . . . . . . . . . . . . . . . . . . . . 159, 213 Actual local time in m[m…] format

%[format]'[argument] special text …. 159, 214 Actual local time minute

%[format]?[[+|-]mmmm|hh:[mm]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 Actual local time minute %[format]?[argument]

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 Actual Universal time 12-hour format suffix

%[format]}[[+|-]mmmm|hh:[mm]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 Actual Universal time 12-hour format suffix

%[format]}[argument] special text. . . . . . . . . 161 Actual Universal time hour

%[format].[*][[+|-]mmmm|hh:[mm]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 Actual Universal time hour %[format].[argument]

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 Actual Universal time in hh:mm/date format

%@[argument] special text . . . . . . . . . . . . 160, 214 Actual Universal time in m[m…]/date format

%[format]`[argument] special text …. 160, 214 Actual Universal time minute

%[format]/[[+|-]mmmm|hh:[mm]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 Actual Universal time minute

%[format]/[argument] special text. . . . . . . . . 161 Addition by days of date variables . . . . . . . . . . . . 105 Addition by weekdays of date variables . . . . . . . 105 Addition of text variables. . . . . . . . . . . . . . . . . . . . . 111 Additional resource file . . . . . . . . . . . . . . . . . . . . . . . . 46 Additional resource file line . . . . . . . . . . . . . . . . . . . . 46 Age value %>1&*B[date] special text . . . . . . . . . . . 158 Age value %B[date] special text. . . . . . . . . . . 158, 213 Age value (birthday) %b[date] special text. . . . . 213 Air pressure of Earth’s atmosphere, set . . . . . . . . 40 Alternation operator. . . . . . . . . . . . . . . . . . . . . . . . . . 259 Alternative date format in calendar sheet . . . . . . 32 Alternative date format in eternal holiday list . . 32 Alternative date format in fixed date list . . . . . . . 32 Amateur-astronomical twilight . . . . . . . . . . . . . . . . 259 Amateur-astronomical twilight, Sun. . . . . . . . . . . 170 Anchor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Anchor metacharacter, $ regexp . . . . . . . . . . . . . . 138 Anchor metacharacter, ^regexp . . . . . . . . . . . . . . 138 Anchors in regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138

ANSI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259

Appearance factor of days, ‘.n’ . . . . . . . . . . . . . . . . 94 Appearance factors of days, examples . . . . . . . . . . 95 Appearance factors of days, valid date parts. … 94 Arguments of command line . . . . . . . . . . . . . . . . . . . . 8 Aspects in Internationalization. . . . . . . . . . . . . . . . 131 Asr prayer time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Asr, Islamic prayer time. . . . . . . . . . . . . 166, 179, 180 Assignment of special texts to text variables,

limitations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 Assignments to date variables . . . . . . . . . 37, 85, 103 Assignments to text variables. . . . . . . . . . 38, 85, 106 Astronomical calculations, accuracy . . . . . . . . . . . . . 1 Astronomical calculations, precision . . . . . . . . . . . . . 1 Astronomical data and times, more precise

representation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Astronomical data in eternal holiday list . . . . . . . 15 Astronomical midnight time, Moon . . . . . . . . . . . 188 Astronomical midnight time, Sun . . . . . . . . . . . . . 165 Astronomical noon time, Moon . . . . . . . . . . . . . . . 189 Astronomical noon time, Sun . . . . . . . . . . . . . . . . . 166 Astronomical twilight. . . . . . . . . . . . . . . . . . . . . . . . . 259 Astronomical twilight, Sun . . . . . . . . . . . . . . . . . . . 170 Astronomical unit . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Astronomical unit, Sun . . . . . . . . . . . . . . . . . . . . . . . 172 Atmospheric refraction . . . . . . . . . . . . . . . . . . . . . . . . 40 Autumnal equinox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Autumnal equinox point on the ecliptic . . . . . . . 259 Azimuth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Azimuth, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Azimuth, Sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165

B

Bahá’ì holidays in eternal holiday list . . . . . . . . . . 15 Base data of Earth’s atmosphere to calculate

refraction, set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Base time of astronomical functions, output . . 173,

196 Base time of the astronomical functions, change

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Base year of calendar, change . . . . . . . . . . . . . . . . . . 34 Biorhythm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Biorhythm %?[date] special texts . . . . . . . . . 156, 213 Biorhythm text %,[date] special text . . . . . 157, 213 Biorhythm text graphics %;[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157, 213 Biorhythm, critical day . . . . . . . . . . . . . . . . . . . . . . . 157 Biorhythm, negative day. . . . . . . . . . . . . . . . . . . . . . 157 Biorhythm, positive day . . . . . . . . . . . . . . . . . . . . . . 157 Birthday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 BSD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Buffer matching operators, regexp . . . . . . . . . . . . 142 Buffer, regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142

C

C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 C Preprocessor symbol USE PAGER . . . . . . . . . . 13 Calendar format, special . . . . . . . . . . . . . . . . . . . . . . . 29 Calendar format, standard . . . . . . . . . . . . . . . . . . . . . 29 Calendar mode, hybrid . . . . . . . . . . . . . . . . . . . . . . . . 34 Calendar mode, proleptic . . . . . . . . . . . . . . . . . . . . . . 34 Calendar option class . . . . . . . . . . . . . . . . . . . . . . 9, 259 Calendar options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Calendar sheet, alternative date format . . . . . . . . 32 Calendar sheet, provide with week numbers …. 30 Calendar sheet, special date format . . . . . . . . . . . . 32 Calendar sheet, standard and special date format

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Calendar sheet, suppress . . . . . . . . . . . . . . . . . . . . . . . 31 Calendar, Chinese . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Calendar, Gregorian . . . . . . . . . . . . . . . . . . . . . . . . . . 125 Calendar, Japanese . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Calendar, Julian. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 Calendar, leap year rule as used by Eastern

churches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Case sensitivity in matching, regexp . . . . . . . . . . 142 Celestial axis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Celestial body elevation . . . . . . . . . . . . . . . . . . . . . . 260 Celestial equator . . . . . . . . . . . . . . . . . . . . 171, 194, 260 Celestial Pole. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 Celtic holidays in eternal holiday list . . . . . . . . . . . 15 Change base time of the astronomical functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Change base year of calendar . . . . . . . . . . . . . . . . . . 34 Character class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 Character class, [:alnum:] regexp . . . . . . . . . . . . 139 Character class, [:alpha:] regexp . . . . . . . . . . . . 139 Character class, [:blank:] regexp . . . . . . . . . . . . 139 Character class, [:cntrl:] regexp . . . . . . . . . . . . 139 Character class, [:digit:] regexp . . . . . . . . . . . . 139 Character class, [:graph:] regexp . . . . . . . . . . . . 139 Character class, [:lower:] regexp . . . . . . . . . . . . 139 Character class, [:print:] regexp . . . . . . . . . . . . 139 Character class, [:punct:] regexp . . . . . . . . . . . . 139 Character class, [:space:] regexp . . . . . . . . . . . . 139 Character class, [:upper:] regexp . . . . . . . . . . . . 139 Character class, [:xdigit:] regexp . . . . . . . . . . . 139 Character classes, [:keyword:] regexp . . . . . . . . 138 Character classes, regexp . . . . . . . . . . . . . . . . . . . . . 138 Character list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 Character list in regexp, complemented . . . . . . . 140 Character list in regexp, negated. . . . . . . . . . . . . . 140 Character list, […] regexp . . . . . . . . . . . . . . . . . . 138 Character list, regexp. . . . . . . . . . . . . . . . . . . . . . . . . 138 Character replacement instruction . . . . . . . . . . . . . 36 Character set. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 Characters, command separator. . . . . . . . . . . . . . . . 72 Chinese calendar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Chinese holidays in eternal holiday list . . . . . . . . . 16 Christian holidays in eternal holiday list . . . . . . . 16 Civil Islamic calendar . . . . . . . . . . . . . . . . . . . . . . 3, 260 Civil midnight time . . . . . . . . . . . . . . . . . . . . . . . 33, 260 Civil twilight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 Civil twilight, Sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Clocktime value, Sun and Moon oriented special

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Co-ordinate termination character, ‘/’ . . . . . . . . 163 Co-ordinate, ISO-6709:1983. . . . . . . . . . 163, 182, 203 Coding scheme of resource file . . . . . . . . . . . . . . . . 251 Coding Scheme Table 1 . . . . . . . . . . . . . . . . . . . . . . . 252 Coding Scheme Table 2 . . . . . . . . . . . . . . . . . . . . . . . 253 Collating element. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 Collating elements, regexp . . . . . . . . . . . . . . . . . . . . 139 Collating symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 260 Collating symbols, [.element.] regexp . . . . . . . 140 Collating symbols, regexp. . . . . . . . . . . . . . . . . . . . . 140 COLUMNS environment variable . . . . . . . . . . . . . . . . 255 Command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Command execution . . . . . . . . . . . . . . 38, 43, 106, 210 Command line argument . . . . . . . . . . . . . . . . . . . . . 260 Command line argument %date . . . . . . . . . . . . . . . . 67 Command line argument @file . . . . . . . . . . . . . . . . 65 Command line interpreter . . . . . . . . . 7, 43, 107, 210 Command line options, gcal . . . . . . . . . . . . . . . . . . . . 9 Command line options, gcal2txt . . . . . . . . . . . . . 128 Command line options, tcal. . . . . . . . . . . . . . . . . . 127 Command line options, txt2gcal . . . . . . . . . . . . . 127 Command line word, single . . . . . . . . . . . . . . . . . . . . 48 Command line, arguments . . . . . . . . . . . . . . . . . . . . . . 8 Command separator characters . . . . . . . . . . . . . . . . 72 Command separator characters, table . . . . . . . . . . 72 Command, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, . yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, .+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .+ yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .- yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, .. yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, : . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, :yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Command, mm yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, mm yyyy-yyyy . . . . . . . . . . . . . . . . . . . . . . 75 Command, mm,…,mm . . . . . . . . . . . . . . . . . . . . . . . . . 74 Command, mm,…,mm yyyy . . . . . . . . . . . . . . . . . . . . 74 Command, mm,mm/yyyy,. . . . . . . . . . . . . . . . . . . . . . . 74 Command, mm-mm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Command, mm-mm yyyy . . . . . . . . . . . . . . . . . . . . . . . . 75 Command, mm-mm yyyy+yyyy . . . . . . . . . . . . . . . . . . . 75 Command, mm/yyyy. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, mm/yyyy,…,mm/yyyy . . . . . . . . . . . . . . 74 Command, mm/yyyy-mm/yyyy . . . . . . . . . . . . . . . . . . 75 Command, mm/yyyy-yyyy . . . . . . . . . . . . . . . . . . . . . . 75 Command, mm: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, mm:;mm:;. . . . . . . . . . . . . . . . . . . . . . . . . . 75 Command, mm:;mm:yyyy;. . . . . . . . . . . . . . . . . . . . . 75 Command, mm:yyyy. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, mm:yyyy+mm:yyyy . . . . . . . . . . . . . . . . . . 75 Command, mm:yyyy;…;mm:yyyy . . . . . . . . . . . . . . 74 Command, month name . . . . . . . . . . . . . . . . . . . . . . . . 72 Command, NOTHING. . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Command, ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Command, yyyy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Command, yyyy+yyyy . . . . . . . . . . . . . . . . . . . . . . . . . 75 Command, yyyy;…;yyyy. . . . . . . . . . . . . . . . . . . . . 74 Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Commands, general view . . . . . . . . . . . . . . . . . . . . . . 72 Commands, lists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Commands, ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Commands, single . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Comment character, ‘;’ . . . . . . . . . . . . . . . . . . . 65, 101 Comment line in resource file . . . . . . . . . . . . . 85, 101 Common option class . . . . . . . . . . . . . . . . . . . . . . 9, 261 Common options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Complemented character list. . . . . . . . . . . . . . . . . . 266 Complemented character list, [^…] regexp . . 140 Complemented character list, regexp . . . . . . . . . . 140 Complemented regular expression . . . . . . . . . . . . . . 45 Complete month name %U[date] special text . . 154,

213 Complete weekday name %K[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150, 212 Conjunction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Conjunction, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Continue line in resource file . . . . . . . . . . . . . . . . . . . 97 Control character sequence . . . . . . . . . . . . . . . . . . . . 12 Copyleft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Country code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 133 Country specific holidays in eternal holiday list

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Country specific special characters . . . . . . . . . . . . . 37 Course angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Course angle (true track) %b*1argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 214 Course angle (true track) %b*2argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 215 Course angle (true track) %b1argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 214 Course angle (true track) %b2argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 214 Critical day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Critical day, biorhythm . . . . . . . . . . . . . . . . . . . . . . . 157 Current month of current year . . . . . . . . . . . . . . . . . 72 Current months, range. . . . . . . . . . . . . . . . . . . . . . . . . 75 Cycle displacement value for Sun and Moon data

and times, set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Cycle ending time value for Sun and Moon data

and times, set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Cycle mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41, 261 Cycle timestep value for Sun and Moon data and

times, set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

D

Date exclusion %?… special texts . . . . . . . . . . . . 146 Date format elements . . . . . . . . . . . . . . . . . . . . . 36, 247 Date format texts, fixed default . . . . . . . . . . . 35, 247 Date format texts, individual . . . . . . . . . . . . . 35, 247 Date formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Date formats, fixed default . . . . . . . . . . . . . . . . . . . . 35 Date formats, individual . . . . . . . . . . . . . . . . . . . . . . . 35 Date part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Date part of a resource file line . . . . . . . . . . . . . . . . 86 Date part, appearance factor of days . . . . . . . . . . . 94 Date part, further attributes . . . . . . . . . . . . . . . . . . . 89 Date part, lists of days . . . . . . . . . . . . . . . . . . . . . . . . 89 Date part, ranges of days . . . . . . . . . . . . . . . . . . . . . . 90 Date part, repetition factor of days . . . . . . . . . . . . 92 Date part, structure . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 Date part, validity of omission values . . . . . . . . . . 88 Date variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Date variables, addition by days . . . . . . . . . . . . . . 105 Date variables, addition by weekdays . . . . . . . . . 105 Date variables, assignments. . . . . . . . . . . . 37, 85, 103 Date variables, definition. . . . . . . . . . . . . . . . . . 37, 103 Date variables, export of local. . . . . . . . . . . . . 38, 105 Date variables, global . . . . . . . . . . . . . . . . . . . . . 37, 103 Date variables, local . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Date variables, management . . . . . . . . . . . . . . 37, 103 Date variables, name space. . . . . . . . . . . . . . . . 37, 103 Date variables, operations . . . . . . . . . . . . . 37, 85, 105 Date variables, scope. . . . . . . . . . . . . . . . . . . . . . 37, 105 Date variables, simple decrement . . . . . . . . . . . . . 105 Date variables, simple increment . . . . . . . . . . . . . . 105 Date variables, subtraction by days . . . . . . . . . . . 105 Date variables, subtraction by weekdays . . . . . . 105 Date variables, total number. . . . . . . . . . . . . . . . . . 103 Date variables, visibility . . . . . . . . . . . . . . . . . . . . . . 103 Date variables, visibility of local. . . . . . . . . . . 38, 105 Day number %>02&*D[date] special text . . 154, 213 Day number %>02*D[date] special text. … 154, 212 Day number %>1&*D[date] special text. … 154, 213 Day number %[format]D[date] special text …. 153,

212 Day number %[format]N[date] special text . . . . . 212 Day number %D[date] special text . . . . . . . . 153, 212 Day-of-year number %>03&*N[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153, 212 Day-of-year number %>03*N[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153, 212 Day-of-year number %>1&*N[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153, 212 Day-of-year number %[format]N[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Day-of-year number %N[date] special text . . . . . 153,

212 Daylight Savings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 Decimal system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Declination . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Declination, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 Declination, Sun. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Default mode, geographical distance and course

angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Default mode, Moon data and times . . . . . . . . . . 202 Default mode, Sun data and times . . . . . . . . . . . . 180 Defaults of native language . . . . . . . . . . . . . . . . . . . 132 Definition of date variables. . . . . . . . . . . . . . . . 37, 103 Definition of global date variables . . . . . . . . . . . . . . 37 Definition of global text variables . . . . . . . . . . . . . . 38 Definition of local date variables . . . . . . . . . . . . . . 103 Definition of local text variables . . . . . . . . . . . . . . 106 Definition of text variables . . . . . . . . . . . . . . . . 38, 106 Deletion of local date variables . . . . . . . . . . . . . . . 104 Delta-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Delta-t, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 Delta-t, Sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 Depreciation character, ‘\’ . . 39, 44, 46, 47, 51, 53,

97, 98, 100, 107, 206 Depreciation of special characters. … 8, 39, 44, 46,

47, 51, 53, 65, 97, 98, 100, 107, 206 Description of all %?… special texts. . . . . . . . . . 145 Description of format instruction . . . . . . . . . . . . . 249 Determine type of week number. . . . . . . . . . . . . . . . 30 Difference value %[format]? special texts . . 206, 243 Difference value %[format]?[date] special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158, 213 Direction to Makkah, Islamic prayer . . . . . . . . . . 164 Directory, actual. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Disable highlighting sequences . . . . . . . . . . . . . . . . . 12 Disable marking characters . . . . . . . . . . . . . . . . . . . . 12 Double or shorter shadow length, period . . . . . . 180 Double shadow length at afternoon . . . . . . . . . . . 180 Double shadow length at forenoon . . . . . . . . . . . . 180

E

Early years . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Earth equator radius, unit . . . . . . . . . . . . . . . . . . . . 194 Earth’s atmosphere, set air pressure. . . . . . . . . . . . 40 Earth’s atmosphere, set base data to calculate

refraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Earth’s atmosphere, set temperature . . . . . . . . . . . 40 Eastern churches’ leap year rule, use . . . . . . . . . . . 30 Eclipse, lunar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Eclipse, solar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Ecliptic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 194, 261 Ecliptic latitude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Ecliptic latitude, Moon . . . . . . . . . . . . . . . . . . . . . . . 194 Ecliptic longitude. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Ecliptic longitude, Moon. . . . . . . . . . . . . . . . . . . . . . 194 Ecliptic longitude, Sun . . . . . . . . . . . . . . . . . . . . . . . 171 Electronic Mail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Elevation, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 Elevation, Sun. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Elongation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 261 Elongation, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 EMail, empty message body . . . . . . . . . . . . . . . . . . . 14 EMail, error code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 EMail, methods of sending. . . . . . . . . . . . . . . . . . . . . 14 EMail, option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Empty local text variables . . . . . . . . . . . . . . . 106, 110 Empty message body in Mail . . . . . . . . . . . . . . . . . . 14 End of highlighting sequence 1 %2 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208, 243 End of highlighting sequence 1 %6 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209, 243 End of highlighting sequence 1 or 2 %0 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209, 243 End of highlighting sequence 2 %4 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208, 243 End of highlighting sequence 2 %8 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209, 243 Ending date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Ending date of fixed date . . . . . . . . . . . . . . . . 146, 148 Environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Environment variable %[format]-[argument] special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206, 242 Environment variable, COLUMNS . . . . . . . . . . . . . . . 255 Environment variable, GCAL. . . . . . . . . . . . 10, 76, 255 Environment variable, GCAL_COLUMNS. . . . . . . . . . 255 Environment variable, GCAL_DATE_FORMAT . . . . . 255 Environment variable, GCAL_LINES . . . . . . . . . . . . 255 Environment variable, GCAL_OPT. . . . . . . . . . . 76, 255 Environment variable, GCAL_SYS_DATADIR . . . . . 82,

255 Environment variable, GCAL_USR_DATADIR . . . . . 82,

255 Environment variable, GCALANSI. . . . . . . . . . . 13, 255 Environment variable, GCALPATH. . . . . . . . . . . 82, 255 Environment variable, GCALPROG . . . . . . . . . 127, 255 Environment variable, HOME . . . . . . . . . . . . . . . 82, 255 Environment variable, LANG . . . . . . . . . . . . . . 131, 256 Environment variable, LANGUAGE . . . . . . . . . 131, 255 Environment variable, LC_ALL . . . . . . . . . . . . 131, 255 Environment variable, LC_MESSAGES . . . . . . 131, 256 Environment variable, LINES. . . . . . . . . . . . . . . . . . 255 Environment variable, LOGNAME . . . . . . . . . . . . 14, 256 Environment variable, MAILPROG. . . . . . . . . . . 14, 256 Environment variable, MAILTO . . . . . . . . . . . . . 14, 256 Environment variable, PAGER . . . . . . . . . . . . . . 11, 256 Environment variable, PATH . . . . . . . . . . . . . . . 11, 256 Environment variable, TERM . . . . . . . . . . . . . . . . . . . 256 Environment variable, TXT2GCALPROG . . . . . 107, 256 Environment variable, TZ . . . . . . . . . . . . . . . . . . . . . 256 Environment variable, USER . . . . . . . . . . . . . . . 14, 256 Environment variables, table. . . . . . . . . . . . . . . . . . 255 Ephemeris date, Julian . . . . . . . . . . . . . . . . . . 174, 197 Ephemeris time (ET) . . . . . . . . . . . . . . . . . . . . 173, 196 Equation of time. . . . . . . . . . . . . . . . . . . . . . . . . 174, 262 Equinox, autumnal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Equinox, vernal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Equivalence class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Equivalence classes, [=list=] regexp . . . . . . . . . 140 Equivalence classes, regexp . . . . . . . . . . . . . . . . . . . 140 Error code 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 46 Error code 113 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Error code 118 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Error code 119 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 Error code 126. . . . . . . . . . . . . . . . . . . . . . 56, 57, 58, 59 Error code 127 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Error code 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108, 210 Error code of executed commands . . . . . . . . . . . . 210 Error codes, table . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 Escape sequence . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 262 Eternal holiday list, alternative date format …. 32 Eternal holiday list, astronomical data . . . . . . . . . 15 Eternal holiday list, Bahá’ì holidays. . . . . . . . . . . . 15 Eternal holiday list, Celtic holidays . . . . . . . . . . . . 15 Eternal holiday list, Chinese holidays . . . . . . . . . . 16 Eternal holiday list, Christian holidays . . . . . . . . . 16 Eternal holiday list, country specific holidays… 17 Eternal holiday list, Hebrew holidays. . . . . . . . . . . 16 Eternal holiday list, holidays from the flexible

Chinese calendar . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Eternal holiday list, holidays from the flexible

Japanese calendar . . . . . . . . . . . . . . . . . . . . . . . . . 16 Eternal holiday list, Islamic holidays . . . . . . . . . . . 16 Eternal holiday list, Japanese holidays . . . . . . . . . 16 Eternal holiday list, months from Bahá’ì calendar

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from flexible Chinese

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from flexible Japanese

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the Chinese

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the civil Indian

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the civil Islamic

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the Coptic

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the Ethiopic

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the French

Revolutionary calendar. . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the Hebrew

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Eternal holiday list, months from the Japanese

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Eternal holiday list, months from the Old-Armenic

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Eternal holiday list, months from the Old-Egyptic

calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Eternal holiday list, months from the Persian

Jalaali calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Eternal holiday list, multicultural New Year

holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Eternal holiday list, option . . . . . . . . . . . . . . . . . . . . 14 Eternal holiday list, Orthodox new calendar

holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Eternal holiday list, Orthodox old calendar

holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Eternal holiday list, Persian Jalaali holidays. … 16 Eternal holiday list, special date format . . . . . . . . 32 Eternal holiday list, standard and special date

format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Eternal holiday list, suppression of leading blank

line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Eternal holiday list, suppression of title text … 15 Eternal holiday list, zodiacal marker holidays . . 16 Eternal Holidays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 Evening width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Evening width geocentric, Sun . . . . . . . . . . . . . . . . 179 Evening width topocentric, Sun . . . . . . . . . . . . . . . 178 Event oriented texts of Moon data and times,

special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Event oriented texts of Sun data and times, special

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Examples of resource files. . . . . . . . . . . . . . . . . . . . . 114 Examples to appearance factors of days . . . . . . . . 95 Examples to lists of days . . . . . . . . . . . . . . . . . . . . . . 89 Examples to ranges of days . . . . . . . . . . . . . . . . . . . . 91 Examples to repetition factors of days . . . . . . . . . 93 Exclusions with date argument %?[date] special

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 Exclusions without any argument %? special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 Exclusive all holidays %X special text . . . . . 149, 211 Exclusive date period . . . . . . . . . . . . . . . . . . . . 146, 262 Exclusive date period %e[date][#[date]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147, 211 Exclusive day period. . . . . . . . . . . . . . . . . . . . . 148, 262 Exclusive day period %? special texts . . . . . 149, 211 Exclusive Friday %H special text . . . . . . . . . . 149, 211 Exclusive legal holidays %V special text … 149, 211 Exclusive Monday %A special text . . . . . . . . 149, 211 Exclusive Monday…Friday %R special text… 150,

211 Exclusive Monday…Thursday %Q special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150, 211 Exclusive Saturday %L special text . . . . . . . 150, 211 Exclusive Sunday %P special text . . . . . . . . . 150, 211 Exclusive Thursday %G special text . . . . . . . 149, 211 Exclusive Tuesday %C special text . . . . . . . . 149, 211 Exclusive Wednesday %F special text . . . . . 149, 211 Execute command . . . . . . . . . . . . . . . . 38, 43, 106, 210 Executed commands, error code . . . . . . . . . . . . . . 210 Executed commands, exit code . . . . . . . . . . . . . . . 210 Exit code 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 46 Exit code 113 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Exit code 118. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Exit code 119 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 Exit code 126. . . . . . . . . . . . . . . . . . . . . . . 56, 57, 58, 59 Exit code 127. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Exit code 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108, 210 Exit code of executed commands. . . . . . . . . . . . . . 210 Exit codes, table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 Explicit fiscal year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Export of local date variables . . . . . . . . . . . . . 38, 105 Export of local text variables . . . . . . . . . . . . . 39, 111 Extended file name . . . . . . . . . . . . . . . . . . . . . . . 83, 262 Extended help text . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Extended language code information . . . . . 133, 262 External pager program . . . . . . . . . . . . . . . . . . . . . . . 11

F

Fajr prayer time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Fajr, Islamic prayer time . . . . . . . . . . . . . . . . . . . . . 170 File directory, actual. . . . . . . . . . . . . . . . . . . . . . . . . . . 82 File name, extended . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 File name, simple. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 File searching mechanism . . . . . . . . . . . . . . . . . . . . . . 82 Filtration, days of fixed dates . . . . . . . . . . . . . . . . . . 44 Filtration, fixed date texts . . . . . . . . . . . . . . . . . 44, 45 Filtration, periods of fixed dates . . . . . . . . . . . . . . . 44 Fiscal year, explicit . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Fiscal year, implicit. . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Fiscal year, limitations . . . . . . . . . . . . . . . . . . . . 89, 103 Fiscal year, single. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Fiscal years of the actual year, list . . . . . . . . . . . . . 75 Fiscal years, list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Fiscal years, mixed list . . . . . . . . . . . . . . . . . . . . . . . . 75 Fiscal years, range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Fixed date function, activate . . . . . . . . . . . 45, 46, 48 Fixed date list, alternative date format. . . . . . . . . 32 Fixed date list, special date format . . . . . . . . . . . . 32 Fixed date list, standard and special date format

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Fixed date list, suppression of leading blank line

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 Fixed date list, suppression of title text . . . . . . . . 53 Fixed date option class. . . . . . . . . . . . . . . . . . . . . 9, 262 Fixed date option, --heading-text=text . . . . . . 53 Fixed date options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 Fixed date period long-style option,

--end-of-month . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Fixed date period long-style option,

--end-of-week. . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Fixed date period long-style option,

--end-of-year. . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 Fixed date period long-style option, --month … 63 Fixed date period long-style option,

--start-of-month . . . . . . . . . . . . . . . . . . . . . . . . 63 Fixed date period long-style option,

--start-of-week . . . . . . . . . . . . . . . . . . . . . . . . . 63 Fixed date period long-style option,

--start-of-year . . . . . . . . . . . . . . . . . . . . . . . . . 64 Fixed date period long-style option, --today … 45 Fixed date period long-style option, --tomorrow

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 Fixed date period long-style option, --week …. 63 Fixed date period long-style option, --year …. 64 Fixed date period modifier, *dn . . . . . . . . . . . . . . . . 55 Fixed date period modifier, *dnwww . . . . . . . . . . . . 57 Fixed date period modifier, *wn[www] . . . . . . . . . . . 58 Fixed date period modifier, @dvar[+|-]nwww …. 62 Fixed date period modifier, @dvar[[+|-]n] . . . . . . 61 Fixed date period modifier, @e[+|-]nwww . . . . . . . . 60 Fixed date period modifier, @e[[+|-]n] . . . . . . . . . . 59 Fixed date period modifier, @t[+|-]nwww . . . . . . . . 61 Fixed date period modifier, @t[[+|-]n] . . . . . . . . . . 60 Fixed date period modifier, 0mmdd. . . . . . . . . . . . . . 57 Fixed date period modifier, d. . . . . . . . . . . . . . . 48, 55 Fixed date period modifier, l. . . . . . . . . . . . . . . 48, 55 Fixed date period modifier, l*dn . . . . . . . . . . . . . . . 55 Fixed date period modifier, l*dnwww . . . . . . . . . . . 58 Fixed date period modifier, l*wn[www] . . . . . . . . . . 58 Fixed date period modifier, l@dvar[+|-]nwww … 62 Fixed date period modifier, l@dvar[[+|-]n] . . . . . 62 Fixed date period modifier, l@e[+|-]nwww. . . . . . . 60 Fixed date period modifier, l@e[[+|-]n] . . . . . . . . . 60 Fixed date period modifier, l@t[+|-]nwww. . . . . . . 61 Fixed date period modifier, l@t[[+|-]n] . . . . . . . . . 61 Fixed date period modifier, lmmwwwn . . . . . . . . . . . 57 Fixed date period modifier, ln+|-. . . . . . . . . . . . . . 56 Fixed date period modifier, lnd . . . . . . . . . . . . . . . . 55 Fixed date period modifier, lnw . . . . . . . . . . . . . . . . 56 Fixed date period modifier, m|M[+|-] . . . . . . . . . . . 63 Fixed date period modifier, mmdd . . . . . . . . . . . . . . . 57 Fixed date period modifier, mmwwwn . . . . . . . . . . . . 57 Fixed date period modifier, n+|- . . . . . . . . . . . . . . . 56 Fixed date period modifier, nd . . . . . . . . . . . . . . . . . 55 Fixed date period modifier, nw . . . . . . . . . . . . . . . . . 56 Fixed date period modifier, t|T . . . . . . . . . . . . . . . . 62 Fixed date period modifier, w|W[+|-] . . . . . . . . . . . 63 Fixed date period modifier, y|Y[+|-] . . . . . . . . . . . 64 Fixed date period modifiers and options. . . . . . . . 55 Fixed date period option, --leap-day=february

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 Fixed date period option,

    --leap-day=february|march . . . . . . . . . . 48, 55

Fixed date period option, --leap-day=march … 55 Fixed date text modifier, a . . . . . . . . . . . . . . . . . . . . 50 Fixed date text modifier, A . . . . . . . . . . . . . . . . . . . . 50 Fixed date text modifier, alternative list format

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 Fixed date text modifier, e . . . . . . . . . . . . . . . . . . . . 50 Fixed date text modifier, E . . . . . . . . . . . . . . . . . . . . 50 Fixed date text modifier, g[text]. . . . . . . . . . . . . . . 51 Fixed date text modifier, J . . . . . . . . . . . . . . . . . . . . 53 Fixed date text modifier, k . . . . . . . . . . . . . . . . . . . . 52 Fixed date text modifier, o . . . . . . . . . . . . . . . . . . . . 52 Fixed date text modifier, Q . . . . . . . . . . . . . . . . . . . . 52 Fixed date text modifier, U . . . . . . . . . . . . . . . . . . . . 52 Fixed date text modifier, x . . . . . . . . . . . . . . . . . . . . 53 Fixed date text modifier, z . . . . . . . . . . . . . . . . . . . . 54 Fixed date text modifier, Z . . . . . . . . . . . . . . . . . . . . 54 Fixed date text modifiers . . . . . . . . . . . . . . . . . . . . . . 49

Fixed date text option, --biorhythm-axis=number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 Fixed date text option,

    --moonimage-lines=number . . . . . . . . . . . . . . . 49

Fixed date texts, filtration . . . . . . . . . . . . . . . . . 44, 45 Fixed date, ending date . . . . . . . . . . . . . . . . . . 146, 148 Fixed date, line structure . . . . . . . . . . . . . . . . . . . . . . 84 Fixed date, option . . . . . . . . . . . . . . . . . . . . . . . . . 45, 48 Fixed date, starting date. . . . . . . . . . . . . . . . . 146, 147 Fixed dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 48, 79 Fixed dates for the month . . . . . . . . . . . . . . . . . . . . . 63 Fixed dates for the week . . . . . . . . . . . . . . . . . . . . . . . 63 Fixed dates for the year . . . . . . . . . . . . . . . . . . . . . . . 64 Fixed dates for today . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Fixed dates for tomorrow . . . . . . . . . . . . . . . . . . . . . . 62 Fixed dates, filtration of days . . . . . . . . . . . . . . . . . . 44 Fixed dates, filtration of periods . . . . . . . . . . . . . . . 44 Fixed default date format texts . . . . . . . . . . . 35, 247 Fixed default date formats . . . . . . . . . . . . . . . . . . . . . 35 Fixed format date text . . . . . . . . . . . . . . . . . . . . . . . 162 Floating Point number . . . . . . . . . . . . . . . . . . . . . . . 267 Force line break-up . . . . . . . . . . . . . . . . . . . . . . . . 51, 53 Force line break-up in resource file . . . . . . . . . . . . . 97 Force year calendar sheet . . . . . . . . . . . . . . . . . . . . . . 31 Forcing highlighting sequences . . . . . . . . . . . . . . . . . 12 Format elements . . . . . . . . . . . . . . . . . . . . . . . . . . 36, 247 Format elements, obsolete . . . . . . . . . . . . . . . . . . . . 245 Format instruction . . . . . . . . . 36, 100, 106, 145, 262 Format instruction, description . . . . . . . . . . . . . . . 249 Free Software Foundation. . . . . . . . . . . . . . . . . . . . . 263 FSF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Full Moon phase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Further date part attributes . . . . . . . . . . . . . . . . . . . 89 Further resource file line . . . . . . . . . . . . . . . . . . . . . . . 46

G

GCAL environment variable. . . . . . . . . . . . . 10, 76, 255 Gcal Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 gcal options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Gcal Utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 gcal, options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Gcal, preliminary remarks . . . . . . . . . . . . . . . . . . . . . . 1 GCAL_COLUMNS environment variable . . . . . . . . . . . 255 GCAL_DATE_FORMAT environment variable . . . . . . 255 GCAL_LINES environment variable . . . . . . . . . . . . . 255 GCAL_OPT environment variable . . . . . . . . . . . . 76, 255 GCAL_SYS_DATADIR environment variable… 82, 255 GCAL_USR_DATADIR environment variable… 82, 255 gcal2txt, options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 gcal2txt, utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 GCALANSI environment variable . . . . . . . . . . . . 13, 255 GCALPATH environment variable . . . . . . . . . . . . 82, 255 GCALPROG environment variable. . . . . . . . . . . 127, 255 General Public License . . . . . . . . . . . . . . . . . . . . . . . 263 Geocentric data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Geocentric data, Moon . . . . . . . . . . . . . . . . . . . . . . . 193 Geocentric data, Sun . . . . . . . . . . . . . . . . . . . . . . . . . 171 Geocentrical horizon. . . . . . . . . . . . . . . . . . . . . . . . . . 263 Geographical data, more precise representation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Geographical distance %b*0argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 214 Geographical distance %b0argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 214 Geographical distance and course angle

%[format]b[argument] special text. . . . . . . . . 162 Geographical distance and course angle

%[format]bargument special texts . . . . . . . . . 214 Geographical distance and course angle, default

mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Geographical distance and course angle, mode

characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Geographical distance and course angle, style of

representation . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Geometric horizon, Moon . . . . . . . . . . . . . . . . . . . . . 189 Geometrical horizon . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Global date variable . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Global date variables, definition. . . . . . . . . . . . . . . . 37 Global option class. . . . . . . . . . . . . . . . . . . . . . . . . 9, 263 Global options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Global text variable . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Global text variables, definition . . . . . . . . . . . . . . . . 38 Glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 259 Gnomon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166, 263 GNU . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 GNU specific regexp operators . . . . . . . . . . . . . . . . 141 GNU style options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 GPL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Greenwich Mean time (GMT)… 33, 155, 158, 160,

161, 162, 173, 183, 196 Gregorian calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Gregorian calendar, genesis . . . . . . . . . . . . . . . . . . . 125 Gregorian Reformation, set period . . . . . . . . . . . . . 34 Grouping, (…) regexp . . . . . . . . . . . . . . . . . . . . . . 140

H

Handling of non-printable characters in the text

part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Handling of NUL characters in the text part. . . . . 97 Handling of special character sequences in the text part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Hebrew holidays in eternal holiday list . . . . . . . . . 16 Help text, extended. . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Help text, long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Help text, short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Help, online . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Hexadecimal system . . . . . . . . . . . . . . . . . . . . . . . . . . 263 Highest culmination point . . . . . . . . . . . . . . . . . . . . 263 Highest culmination point, Moon . . . . . . . . . . . . . 189 Highest culmination point, Sun . . . . . . . . . . . . . . . 166 Highlighting %? special texts . . . . . . . . . . . . . 208, 243 Highlighting sequence . . . . . . . . . . . . . . . . . . . . . . . . 264 Highlighting sequences, disabling. . . . . . . . . . . . . . . 12 Highlighting sequences, forcing. . . . . . . . . . . . . . . . . 12 Highlighting sequences, setting . . . . . . . . . . . . . . . . . 12 Holiday list, eternal. . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 Holiday, legal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Holidays from the flexible Chinese calendar in

eternal holiday list . . . . . . . . . . . . . . . . . . . . . . . . 15 Holidays from the flexible Japanese calendar in

eternal holiday list . . . . . . . . . . . . . . . . . . . . . . . . 16 HOME environment variable . . . . . . . . . . . . . . . . 82, 255 Horizontal parallax . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 Horizontal parallax, moon . . . . . . . . . . . . . . . . . . . . 193 Horizontal parallax, Sun . . . . . . . . . . . . . . . . . . . . . . 172 How to use regular expressions. . . . . . . . . . . . . . . . 137 Hybrid calendar mode . . . . . . . . . . . . . . . . . . . . 34, 264

I

Ignoring case distinctions in regular expression

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45, 142 Ignoring case, regexp. . . . . . . . . . . . . . . . . . . . . . 45, 142 Implicit fiscal year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Include directives . . . . . . . . . . . . . . . . . . . . . . . . . 85, 102 Inclusive all holidays %x special text. . . . . . 149, 211 Inclusive date period. . . . . . . . . . . . . . . . . . . . . 146, 264 Inclusive date period %i[date][#[date]] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146, 211 Inclusive day period . . . . . . . . . . . . . . . . . . . . . 148, 264 Inclusive day period %? special texts . . . . . 149, 211 Inclusive Friday %h special text . . . . . . . . . . 149, 211 Inclusive legal holidays %v special text. … 149, 211 Inclusive Monday %a special text . . . . . . . . . 149, 211 Inclusive Monday…Friday %r special text … 149,

211 Inclusive Monday…Thursday %q special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149, 211 Inclusive Saturday %l special text . . . . . . . . 149, 211 Inclusive Sunday %p special text. . . . . . . . . . 149, 211 Inclusive Thursday %g special text . . . . . . . 149, 211 Inclusive Tuesday %c special text . . . . . . . . . 149, 211 Inclusive Wednesday %f special text . . . . . . 149, 211 Individual date format texts . . . . . . . . . . . . . . 35, 247 Individual date formats . . . . . . . . . . . . . . . . . . . . . . . . 35 Integer number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 Internal pager, simple. . . . . . . . . . . . . . . . . . . . . . . . . . 11 Internationalization aspects . . . . . . . . . . . . . . . . . . . 131 Interval expression. . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 Interval expressions, {expression} regexp …. 141 Interval expressions, regexp . . . . . . . . . . . . . . . . . . . 141 Introduction, Gcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Invoking gcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Invoking gcal2txt. . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 Invoking tcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Invoking txt2gcal. . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Isha prayer time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 Isha, Islamic prayer time . . . . . . . . . . . . . . . . . . . . . 170 Islamic calendar, civil . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Islamic calendar, religious . . . . . . . . . . . . . . . . . . . . . . . 3 Islamic holidays in eternal holiday list. . . . . . . . . . 16 Islamic prayer time, Asr. . . . . . . . . . . . . 166, 179, 180 Islamic prayer time, Fajr . . . . . . . . . . . . . . . . . . . . . 170 Islamic prayer time, Isha . . . . . . . . . . . . . . . . . . . . . 170 Islamic prayer time, Maghrib . . . . . . . . . . . . . . . . . 168 Islamic prayer time, Zuhr . . . . . . . . . . . . . . . . . . . . . 166 Islamic prayer, direction to Makkah . . . . . . . . . . . 164 ISO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 ISO week number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 ISO-3166. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 133 ISO-639:1988 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 ISO-6709:1983 co-ordinate . . . . . . . . . . . 163, 182, 203 ISO-8601:1988 week number …. 31, 52, 56, 58, 86,

154

J

Japanese calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Japanese holidays in eternal holiday list . . . . . . . . 16 Julian calendar . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 264 Julian date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 Julian day number . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 Julian day number %[format]j[argument] special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162, 214 Julian Ephemeris date . . . . . . . . . . . . . . 174, 197, 264

L

LANG environment variable . . . . . . . . . . . . . . . 131, 256 Language code information, extended . . . . . . . . . 133 Language code information, simple . . . . . . . . . . . 131 Language codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 LANGUAGE environment variable. . . . . . . . . . . 131, 255 LC_ALL environment variable . . . . . . . . . . . . . 131, 255 LC_MESSAGES environment variable . . . . . . . 131, 256 Leap day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 264 Leap year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Leap year rule as used by Eastern Orthodox

churches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Legal holiday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 265 Length preserving behavior of numerical text

variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Limit sunrise and sunset to actual day . . . . . . . . . 40 Limitations concerning the assignment of special

texts to text variables . . . . . . . . . . . . . . . . . . . . 145 Limitations in a resource file . . . . . . . 14, 84, 97, 103 Limitations in the text part of a line . . . . . . . . . . . 97 Limitations of use of 3-Month modes . . . . . . 89, 103 Limitations of use of fiscal years. . . . . . . . . . . 89, 103 Limitations, Moon oriented special texts . . . . . . 205 Limitations, Sun oriented special texts . . . . . . . . 186 Line break-up character, ‘^’ . . . . . . . . . 44, 51, 53, 97 Line break-up character, ‘~’ . . . . . . . . . 44, 51, 53, 97 Line continuation character, ‘\’ . . . . . . . . . . . . . . . . 97 Line structure of fixed date . . . . . . . . . . . . . . . . . . . . 84 Line, date part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 LINES environment variable . . . . . . . . . . . . . . . . . . . 255 List of all long-style option names. . . . . . . . . . . . . . 10 List of fiscal years . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 List of fiscal years of the actual year . . . . . . . . . . . 75 List of months in specified year . . . . . . . . . . . . . . . . 74 List of specified fiscal years . . . . . . . . . . . . . . . . . . . . 74 List of specified months of selected year . . . . . . . . 74 List of specified months of the current year. . . . . 74 List of specified years . . . . . . . . . . . . . . . . . . . . . . . . . . 74 List of whitespace character . . . . . . . . . . . . . . . . . . . 84 Lists of commands . . . . . . . . . . . . . . . . . . . . 72, 74, 265 Lists of days. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 Lists of days, examples . . . . . . . . . . . . . . . . . . . . . . . . 89 Lists of days, valid date parts . . . . . . . . . . . . . . . . . . 89 Local date variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Local date variables . . . . . . . . . . . . . . . . . . . . . . . . . . 103 Local date variables, definition. . . . . . . . . . . . . . . . 103 Local date variables, deletion . . . . . . . . . . . . . . . . . 104 Local sidereal time . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Local star time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Local text variable . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Local text variables. . . . . . . . . . . . . . . . . . . . . . . . . . . 106 Local text variables, definition . . . . . . . . . . . . . . . . 106 Local text variables, empty . . . . . . . . . . . . . . 106, 110 Local time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 Local time 12-hour format suffix, display. . . . . . 160 Local time hour, display . . . . . . . . . . . . . . . . . . . . . . 159 Local time in hh:mm format, display . . . . . . . . . . . 159 Local time in m[m…] format, display . . . . . . . . 159 Local time minute, display. . . . . . . . . . . . . . . . . . . . 160 LOGNAME environment variable . . . . . . . . . . . . . 14, 256 Long help text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Long text part in resource file line . . . . . . . . . . . . . 97 Long-style option . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 265 Long-style option names, list. . . . . . . . . . . . . . . . . . . 10 Longer than double shadow length, period …. 180 Longer than single shadow length, period . . . . . 180 Lowest culmination point . . . . . . . . . . . . . . . . . . . . . 265 Lowest culmination point, Moon . . . . . . . . . . . . . . 188 Lowest culmination point, Sun . . . . . . . . . . . . . . . . 165 Lunar Arctic Circle . . . . . . . . . . . . . . . . . . . . . . 205, 265 Lunar day length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Lunar eclipse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Lunar night length . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Lunation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Lunation, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188

M

Maghrib prayer time. . . . . . . . . . . . . . . . . . . . . . . . . . 265 Maghrib, Islamic prayer time . . . . . . . . . . . . . . . . . 168 Magnitude unit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Magnitude units, Moon . . . . . . . . . . . . . . . . . . . . . . . 193 Mail, empty message body . . . . . . . . . . . . . . . . . . . . . 14 Mail, methods of sending . . . . . . . . . . . . . . . . . . . . . . 14 Mail, sending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 MAILPROG environment variable . . . . . . . . . . . . 14, 256 MAILTO environment variable . . . . . . . . . . . . . . 14, 256 Management of date variables . . . . . . . . . . . . . 37, 103 Management of text variables . . . . . . . . . . . . . 38, 106 Marking character . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Marking characters, disabling . . . . . . . . . . . . . . . . . . 12 Marking characters, setting . . . . . . . . . . . . . . . . . . . . 12 Matching word boundaries in regexp . . . . . . . . . . 141 Mathematical horizon . . . . . . . . . . . . . . . . . . . . . . . . 263 Mathematical-geocentric horizon . . . . . . . . . . . . . . 263 Mathematical-geocentric horizon, Moon . . . . . . . 189 Mean Earth equator radius . . . . . . . . . . . . . . . . . . . 266 Mean solar day . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 266 Mean solar time . . . . . . . . . . . . . . . . . . . . . . . . . 174, 266 Mean sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Mean Sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 Meridian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162, 266 Message catalog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 Metacharacter, . regexp . . . . . . . . . . . . . . . . . . . . . . 138 Metacharacter, \ regexp . . . . . . . . . . . . . . . . . . . . . . 138 Metacharacters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Metacharacters, regexp . . . . . . . . . . . . . . . . . . . . . . . 137 Metasymbols within this document, table . . . . . 135 Methods of sending Mail. . . . . . . . . . . . . . . . . . . . . . . 14 Midnight height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Midnight height geocentric, Moon . . . . . . . . . . . . 201 Midnight height geocentric, Sun . . . . . . . . . . . . . . 178 Midnight height topocentric, Moon . . . . . . . . . . . 200 Midnight height topocentric, Sun . . . . . . . . . . . . . 177 Midnight phase angle . . . . . . . . . . . . . . . . . . . . . . . . . 266 Midnight phase angle geocentric, Moon . . . . . . . 201 Midnight phase angle topocentric, Moon . . . . . . 200 Midnight time of Moon, astronomical . . . . . . . . . 188 Midnight time of Sun, astronomical . . . . . . . . . . . 165 Mixed list of fiscal years . . . . . . . . . . . . . . . . . . . . . . . 75 Mixed list of months. . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Mode characters, geographical distance and course angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Mode characters, Moon data and times . . . . . . . 188 Mode characters, Sun data and times . . . . . . . . . 165 Modifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 266 Modifiers, fixed date period . . . . . . . . . . . . . . . . . . . . 55 Modifiers, fixed date text . . . . . . . . . . . . . . . . . . . . . . 49 Month %[format]M[date] special text . . . . . . . . . . 213 Month mode, selected . . . . . . . . . . . . . . . . . . . . . . . . . 72 Month mode, single. . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Month name %[format]U[date] special text …. 154,

213 Month number %>02&*M[date] special text …. 155,

213 Month number %>02*M[date] special text . . . . . 155,

213 Month number %>1&*M[date] special text . . . . . 155,

213 Month number %[format]M[date] special text . . 154 Month number %M[date] special text. . . . . . 154, 213 Months from the Bahá’ì calendar in eternal holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Months from the Chinese calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Months from the civil Indian calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Months from the civil Islamic calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Months from the Coptic calendar in eternal holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Months from the Ethiopic calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Months from the flexible Chinese calendar in

eternal holiday list . . . . . . . . . . . . . . . . . . . . . . . . 28 Months from the flexible Japanese calendar in

eternal holiday list . . . . . . . . . . . . . . . . . . . . . . . . 28 Months from the French Revolutionary calendar in eternal holiday list . . . . . . . . . . . . . . . . . . . . . . . . 28

Months from the Hebrew calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 Months from the Japanese calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Months from the Old-Armenic calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Months from the Old-Egyptic calendar in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Months from the Persian Jalaali calendar in

eternal holiday list . . . . . . . . . . . . . . . . . . . . . . . . 29 Months in specified year, list. . . . . . . . . . . . . . . . . . . 74 Months of year, range . . . . . . . . . . . . . . . . . . . . . . . . . 75 Months’ fixed dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Months, mixed list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Moon data %[format]?argument special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188, 229 Moon data and times, default mode. . . . . . . . . . . 202 Moon data and times, mode characters . . . . . . . 188 Moon data and times, special event oriented texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Moon data and times, style of representation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180, 202 Moon data and times, type of representation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180, 202 Moon oriented special texts, limitations . . . . . . . 205 Moon oriented special texts, reference systems used

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 Moon oriented special texts, table of modes … 188 Moon oriented special texts, table of representation

types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 Moon phase %[format]?[date] special texts … 155,

213 Moon phase text %>03*O[date] special text … 156,

213 Moon phase text %O[date] special text. … 155, 213 Moon phase text graphics %Z[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156, 213 Moon phase, Full . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Moon phase, New . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Moon phase, waning Half . . . . . . . . . . . . . . . . . . . . . . 15 Moon phase, waxing Half . . . . . . . . . . . . . . . . . . . . . . 15 Moon’s center, geocentric brightness %(*largument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric brightness %(largument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric brightness %)*largument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric brightness %)largument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric horizontal parallax

%(*jargument special text. . . . . . . . . . . . . . . . 195 Moon’s center, geocentric horizontal parallax

%(jargument special text . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric horizontal parallax

%)*jargument special text. . . . . . . . . . . . . . . . 195 Moon’s center, geocentric horizontal parallax

%)jargument special text . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric phase angle

%(*margument special text. . . . . . . . . . . . . . . . 195 Moon’s center, geocentric phase angle %(margument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric phase angle

%)*margument special text. . . . . . . . . . . . . . . . 195 Moon’s center, geocentric phase angle %)margument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric semidiameter

%(*kargument special text. . . . . . . . . . . . . . . . 195 Moon’s center, geocentric semidiameter

%(kargument special text . . . . . . . . . . . . . . . . . 195 Moon’s center, geocentric semidiameter

%)*kargument special text. . . . . . . . . . . . . . . . 195 Moon’s center, geocentric semidiameter

%)kargument special text . . . . . . . . . . . . . . . . . 195 Moon’s center, geometrical non-visibility

%]*2argument special text . . . . . . . . . . . 189, 230 Moon’s center, geometrical non-visibility

%]2argument special text . . . . . . . . . . . . 189, 230 Moon’s center, geometrical rise time %(*2argument special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229

Moon’s center, geometrical rise time %(2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon’s center, geometrical set time %)*2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon’s center, geometrical set time %)2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon’s center, geometrical visibility %[*2argument special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 230

Moon’s center, geometrical visibility %[2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon’s center, non-visibility %]*4argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 231

Moon’s center, non-visibility %]4argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 231 Moon’s center, rise time %(*4argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 230

Moon’s center, rise time %(4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 230 Moon’s center, set time %)*4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 231 Moon’s center, set time %)4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 230 Moon’s center, topocentric brightness

%(*8argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric brightness %(8argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

Moon’s center, topocentric brightness

%)*8argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric brightness %)8argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

Moon’s center, topocentric horizontal parallax

%(*6argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric horizontal parallax

%(6argument special text . . . . . . . . . . . . . . . . . 193 Moon’s center, topocentric horizontal parallax

%)*6argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric horizontal parallax

%)6argument special text . . . . . . . . . . . . . . . . . 193 Moon’s center, topocentric phase angle

%(*9argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric phase angle

%(9argument special text . . . . . . . . . . . . . . . . . 193 Moon’s center, topocentric phase angle

%)*9argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric phase angle

%)9argument special text . . . . . . . . . . . . . . . . . 193 Moon’s center, topocentric semidiameter

%(*7argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric semidiameter

%(7argument special text . . . . . . . . . . . . . . . . . 193 Moon’s center, topocentric semidiameter

%)*7argument special text. . . . . . . . . . . . . . . . 193 Moon’s center, topocentric semidiameter

%)7argument special text . . . . . . . . . . . . . . . . . 193 Moon’s center, visibility %[*4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 231 Moon’s center, visibility %[4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191, 231 Moon’s upper limb, geometrical non-visibility

%]*3argument special text . . . . . . . . . . . 190, 230 Moon’s upper limb, geometrical non-visibility

%]3argument special text . . . . . . . . . . . . 190, 230 Moon’s upper limb, geometrical rise time

%(*3argument special text . . . . . . . . . . . 190, 230 Moon’s upper limb, geometrical rise time

%(3argument special text . . . . . . . . . . . . 190, 230 Moon’s upper limb, geometrical set time

%)*3argument special text . . . . . . . . . . . 190, 230 Moon’s upper limb, geometrical set time

%)3argument special text . . . . . . . . . . . . 190, 230 Moon’s upper limb, geometrical visibility

%[*3argument special text . . . . . . . . . . . 190, 230 Moon’s upper limb, geometrical visibility

%[3argument special text . . . . . . . . . . . . 190, 230 Moon’s upper limb, non-visibility standard

%]*5argument special text. . . . . . . . . . . . . . . . 231 Moon’s upper limb, non-visibility standard

%]5argument special text . . . . . . . . . . . . . . . . . 231 Moon’s upper limb, standard non-visibility

%]*5argument special text. . . . . . . . . . . . . . . . 193 Moon’s upper limb, standard non-visibility

%]5argument special text . . . . . . . . . . . . . . . . . 193 Moon’s upper limb, standard rise time

%(*5argument special text . . . . . . . . . . . 192, 231 Moon’s upper limb, standard rise time

%(5argument special text . . . . . . . . . . . . 192, 231 Moon’s upper limb, standard set time

%)*5argument special text . . . . . . . . . . . 192, 231 Moon’s upper limb, standard set time %)5argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 192, 231 Moon’s upper limb, standard visibility

%[*5argument special text. . . . . . . . . . . . . . . . 192 Moon’s upper limb, standard visibility

%[5argument special text . . . . . . . . . . . . . . . . . 192 Moon’s upper limb, visibility standard

%[*5argument special text. . . . . . . . . . . . . . . . 231 Moon’s upper limb, visibility standard

%[5argument special text . . . . . . . . . . . . . . . . . 231 Moon, astronomical midnight time %(*0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 188, 229 Moon, astronomical midnight time %(0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 188, 229 Moon, astronomical midnight time %)*0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 188, 229 Moon, astronomical midnight time %)0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 188, 229 Moon, astronomical noon time %(*1argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon, astronomical noon time %(1argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon, astronomical noon time %)*1argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon, astronomical noon time %)1argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 189, 229 Moon, azimuth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Moon, base time %(*xargument special text . . 196,

236 Moon, base time %(xargument special text. … 196,

236 Moon, base time %)*xargument special text . . 196,

236 Moon, base time %)xargument special text. … 196,

236 Moon, conjunction . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Moon, delta-t %(*vargument special text . . . . . 196,

236 Moon, delta-t %(vargument special text . . 196, 236 Moon, delta-t %)*vargument special text . . . . . 196,

236 Moon, delta-t %)vargument special text . . 196, 236 Moon, geocentric azimuth %(*oargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric azimuth %(oargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric azimuth %)*oargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric azimuth %)oargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric azimuth at standard rise time

%(*Xargument special text . . . . . . . . . . . 202, 242 Moon, geocentric azimuth at standard rise time

%(Xargument special text . . . . . . . . . . . . 202, 242 Moon, geocentric azimuth at standard set time

%)*Xargument special text . . . . . . . . . . . 202, 242 Moon, geocentric azimuth at standard set time

%)Xargument special text . . . . . . . . . . . . 202, 242 Moon, geocentric brightness %(*largument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric brightness %(largument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric brightness %)*largument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric brightness %)largument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric data . . . . . . . . . . . . . . . . . . . . . . . . 193 Moon, geocentric declination %(*pargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric declination %(pargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric declination %)*pargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric declination %)pargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric distance %(*targument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric distance %(targument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric distance %)*targument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric distance %)targument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric ecliptic latitude %(*rargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 235 Moon, geocentric ecliptic latitude %(rargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 235 Moon, geocentric ecliptic latitude %)*rargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 235 Moon, geocentric ecliptic latitude %)rargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 235 Moon, geocentric ecliptic longitude %(*qargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric ecliptic longitude %(qargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric ecliptic longitude %)*qargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric ecliptic longitude %)qargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric elevation %(*nargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric elevation %(nargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric elevation %)*nargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric elevation %)nargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 235 Moon, geocentric elevation at Moon’s astronomical

midnight time %(*Sargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric elevation at Moon’s astronomical midnight time %(Sargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric elevation at Moon’s astronomical midnight time %)*Sargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric elevation at Moon’s astronomical midnight time %)Sargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric elevation at Moon’s astronomical noon time %(*Uargument special text …. 201, 241

Moon, geocentric elevation at Moon’s astronomical

noon time %(Uargument special text . . 201, 241 Moon, geocentric elevation at Moon’s astronomical

noon time %)*Uargument special text …. 201, 241

Moon, geocentric elevation at Moon’s astronomical noon time %)Uargument special text . . 201, 241 Moon, geocentric elevation at standard rise time

%(*Wargument special text . . . . . . . . . . . 202, 242 Moon, geocentric elevation at standard rise time

%(Wargument special text . . . . . . . . . . . . 202, 242 Moon, geocentric elevation at standard set time

%)*Wargument special text . . . . . . . . . . . 202, 242 Moon, geocentric elevation at standard set time

%)Wargument special text . . . . . . . . . . . . 202, 242 Moon, geocentric elongation %(*uargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric elongation %(uargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric elongation %)*uargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric elongation %)uargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric horizontal parallax %(*jargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric horizontal parallax %(jargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric horizontal parallax %)*jargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric horizontal parallax %)jargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric phase angle %(*margument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric phase angle %(margument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric phase angle %)*margument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric phase angle %)margument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric phase angle at Moon’s

astronomical midnight time %(*Targument special text . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric phase angle at Moon’s astronomical midnight time %(Targument special text . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric phase angle at Moon’s astronomical midnight time %)*Targument special text . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric phase angle at Moon’s astronomical midnight time %)Targument special text . . . . . . . . . . . . . . . . . . . . . . . . . 201, 241

Moon, geocentric phase angle at Moon’s astronomical noon time %(*Vargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202, 242

Moon, geocentric phase angle at Moon’s astronomical noon time %(Vargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202, 241

Moon, geocentric phase angle at Moon’s astronomical noon time %)*Vargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202, 242

Moon, geocentric phase angle at Moon’s astronomical noon time %)Vargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202, 241

Moon, geocentric phase angle at standard rise time %(*Yargument special text . . . . . . . . . . . 202, 242 Moon, geocentric phase angle at standard rise time

%(Yargument special text . . . . . . . . . . . . 202, 242 Moon, geocentric phase angle at standard set time

%)*Yargument special text . . . . . . . . . . . 202, 242 Moon, geocentric phase angle at standard set time

%)Yargument special text . . . . . . . . . . . . 202, 242 Moon, geocentric right ascension %(*sargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric right ascension %(sargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 235 Moon, geocentric right ascension %)*sargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, geocentric right ascension %)sargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 196, 235 Moon, geocentric semidiameter %(*kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric semidiameter %(kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric semidiameter %)*kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geocentric semidiameter %)kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 234 Moon, geometric horizon . . . . . . . . . . . . . . . . . . . . . 189 Moon, highest culmination point. . . . . . . . . . . . . . 189 Moon, Julian date %(*yargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197, 236 Moon, Julian date %(yargument special text . . 197,

236 Moon, Julian date %)*yargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197, 236 Moon, Julian date %)yargument special text . . 197,

236 Moon, Julian Ephemeris date %(*zargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon, Julian Ephemeris date %(zargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon, Julian Ephemeris date %)*zargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon, Julian Ephemeris date %)zargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon, local sidereal time %(*wargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, local sidereal time %(wargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, local sidereal time %)*wargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, local sidereal time %)wargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196, 236 Moon, lowest culmination point. . . . . . . . . . . . . . . 188 Moon, lunation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Moon, magnitude units . . . . . . . . . . . . . . . . . . . . . . . 193 Moon, mathematical-geocentric horizon . . . . . . . 189 Moon, midnight height geocentric. . . . . . . . . . . . . 201 Moon, midnight height topocentric . . . . . . . . . . . 200 Moon, midnight phase angle geocentric . . . . . . . 201 Moon, midnight phase angle topocentric . . . . . . 200 Moon, noon height geocentric. . . . . . . . . . . . . . . . . 201 Moon, noon height topocentric . . . . . . . . . . . . . . . 200 Moon, noon phase angle geocentric . . . . . . . . . . . 202 Moon, noon phase angle topocentric . . . . . . . . . . 200 Moon, refraction %(*iargument special text . . 195,

234 Moon, refraction %(iargument special text … 195,

234 Moon, refraction %)*iargument special text . . 195,

234 Moon, refraction %)iargument special text … 195,

234 Moon, rise azimuth geocentric . . . . . . . . . . . . . . . . 202 Moon, rise azimuth topocentric . . . . . . . . . . . . . . . 201 Moon, rise height geocentric . . . . . . . . . . . . . . . . . . 202 Moon, rise height topocentric . . . . . . . . . . . . . . . . . 200 Moon, rise phase angle geocentric . . . . . . . . . . . . . 202 Moon, rise phase angle topocentric. . . . . . . . . . . . 201 Moon, rise width geocentric. . . . . . . . . . . . . . . . . . . 202 Moon, rise width topocentric . . . . . . . . . . . . . . . . . 201 Moon, set azimuth geocentric . . . . . . . . . . . . . . . . . 202 Moon, set azimuth topocentric. . . . . . . . . . . . . . . . 201 Moon, set height geocentric . . . . . . . . . . . . . . . . . . . 202 Moon, set height topocentric. . . . . . . . . . . . . . . . . . 200 Moon, set phase angle geocentric . . . . . . . . . . . . . 202 Moon, set phase angle topocentric . . . . . . . . . . . . 201 Moon, set width geocentric . . . . . . . . . . . . . . . . . . . 202 Moon, set width topocentric . . . . . . . . . . . . . . . . . . 201 Moon, standard non-visibility period . . . . . . . . . . 193 Moon, standard rise time . . . . . . . . . . . . . . . . . . . . . 192 Moon, standard set time. . . . . . . . . . . . . . . . . . . . . . 192 Moon, standard visibility period . . . . . . . . . . . . . . 192 Moon, synodic month . . . . . . . . . . . . . . . . . . . . . . . . 188 Moon, topocentric azimuth %(*bargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric azimuth %(bargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric azimuth %)*bargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric azimuth %)bargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric azimuth at standard rise time

%(*Qargument special text . . . . . . . . . . . 201, 240 Moon, topocentric azimuth at standard rise time

%(Qargument special text . . . . . . . . . . . . 201, 240 Moon, topocentric azimuth at standard set time

%)*Qargument special text . . . . . . . . . . . 201, 241 Moon, topocentric azimuth at standard set time

%)Qargument special text . . . . . . . . . . . . 201, 240 Moon, topocentric brightness %(*8argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric brightness %(8argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric brightness %)*8argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric brightness %)8argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric data . . . . . . . . . . . . . . . . . . . . . . . 193 Moon, topocentric declination %(*cargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric declination %(cargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric declination %)*cargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric declination %)cargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric distance %(*gargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric distance %(gargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric distance %)*gargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric distance %)gargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic latitude %(*eargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic latitude %(eargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic latitude %)*eargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic latitude %)eargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic longitude %(*dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic longitude %(dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic longitude %)*dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric ecliptic longitude %)dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric elevation %(*aargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric elevation %(aargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric elevation %)*aargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric elevation %)aargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194, 232 Moon, topocentric elevation at Moon’s

astronomical midnight time %(*Largument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 239

Moon, topocentric elevation at Moon’s astronomical midnight time %(Largument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 239

Moon, topocentric elevation at Moon’s astronomical midnight time %)*Largument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 239

Moon, topocentric elevation at Moon’s astronomical midnight time %)Largument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 239

Moon, topocentric elevation at Moon’s astronomical noon time %(*Nargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric elevation at Moon’s astronomical noon time %(Nargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240 Moon, topocentric elevation at Moon’s

astronomical noon time %)*Nargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric elevation at Moon’s astronomical noon time %)Nargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric elevation at standard rise time

%(*Pargument special text . . . . . . . . . . . 200, 240 Moon, topocentric elevation at standard rise time

%(Pargument special text . . . . . . . . . . . . 200, 240 Moon, topocentric elevation at standard set time

%)*Pargument special text . . . . . . . . . . . 200, 240 Moon, topocentric elevation at standard set time

%)Pargument special text . . . . . . . . . . . . 200, 240 Moon, topocentric elongation %(*hargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 195, 234 Moon, topocentric elongation %(hargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 233 Moon, topocentric elongation %)*hargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 195, 234 Moon, topocentric elongation %)hargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195, 233 Moon, topocentric horizontal parallax

%(*6argument special text. . . . . . . . . . . . . . . . 232 Moon, topocentric horizontal parallax

%(6argument special text . . . . . . . . . . . . . . . . . 231 Moon, topocentric horizontal parallax

%)*6argument special text. . . . . . . . . . . . . . . . 232 Moon, topocentric horizontal parallax

%)6argument special text . . . . . . . . . . . . . . . . . 231 Moon, topocentric phase angle %(*9argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric phase angle %(9argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric phase angle %)*9argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric phase angle %)9argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric phase angle at Moon’s

astronomical midnight time %(*Margument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric phase angle at Moon’s astronomical midnight time %(Margument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 239

Moon, topocentric phase angle at Moon’s astronomical midnight time %)*Margument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric phase angle at Moon’s astronomical midnight time %)Margument special text . . . . . . . . . . . . . . . . . . . . . . . . . 200, 239

Moon, topocentric phase angle at Moon’s astronomical noon time %(*Oargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric phase angle at Moon’s astronomical noon time %(Oargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240 Moon, topocentric phase angle at Moon’s

astronomical noon time %)*Oargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric phase angle at Moon’s astronomical noon time %)Oargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200, 240

Moon, topocentric phase angle at standard rise

time %(*Rargument special text . . . . . . 201, 241 Moon, topocentric phase angle at standard rise

time %(Rargument special text . . . . . . . 201, 241 Moon, topocentric phase angle at standard set

time %)*Rargument special text . . . . . . 201, 241 Moon, topocentric phase angle at standard set

time %)Rargument special text . . . . . . . 201, 241 Moon, topocentric right ascension %(*fargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric right ascension %(fargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric right ascension %)*fargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric right ascension %)fargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 194, 233 Moon, topocentric semidiameter %(*7argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric semidiameter %(7argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric semidiameter %)*7argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon, topocentric semidiameter %)7argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 Moon/Sun, delta astronomical midnight time

%(*Iargument special text . . . . . . . . . . . 199, 239 Moon/Sun, delta astronomical midnight time

%(Iargument special text . . . . . . . . . . . . 199, 239 Moon/Sun, delta astronomical midnight time

%)*Iargument special text . . . . . . . . . . . 199, 239 Moon/Sun, delta astronomical midnight time

%)Iargument special text . . . . . . . . . . . . 199, 239 Moon/Sun, delta astronomical noon time

%(*Jargument special text . . . . . . . . . . . 199, 239 Moon/Sun, delta astronomical noon time

%(Jargument special text . . . . . . . . . . . . 199, 239 Moon/Sun, delta astronomical noon time

%)*Jargument special text . . . . . . . . . . . 199, 239 Moon/Sun, delta astronomical noon time

%)Jargument special text . . . . . . . . . . . . 199, 239 Moon/Sun, delta geocentric azimuth %(*Dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon/Sun, delta geocentric azimuth %(Dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon/Sun, delta geocentric azimuth %)*Dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon/Sun, delta geocentric azimuth %)Dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237 Moon/Sun, delta geocentric azimuth at standard

moonrise time %(*Hargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238 Moon/Sun, delta geocentric azimuth at standard moonrise time %(Hargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta geocentric azimuth at standard moonset time %)*Hargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199, 238

Moon/Sun, delta geocentric azimuth at standard moonset time %)Hargument special text . . 199, 238

Moon/Sun, delta geocentric elevation

%(*Cargument special text . . . . . . . . . . . 197, 237 Moon/Sun, delta geocentric elevation %(Cargument special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237

Moon/Sun, delta geocentric elevation

%)*Cargument special text . . . . . . . . . . . 197, 237 Moon/Sun, delta geocentric elevation %)Cargument special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237

Moon/Sun, delta geocentric elevation at standard moonrise time %(*Gargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta geocentric elevation at standard moonrise time %(Gargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta geocentric elevation at standard moonset time %)*Gargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta geocentric elevation at standard moonset time %)Gargument special text . . 198, 238

Moon/Sun, delta standard rise time %(*Kargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239

Moon/Sun, delta standard rise time %(Kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 Moon/Sun, delta standard set time %)*Kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 Moon/Sun, delta standard set time %)Kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 Moon/Sun, delta topocentric azimuth

%(*Bargument special text . . . . . . . . . . . 197, 237 Moon/Sun, delta topocentric azimuth %(Bargument special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237

Moon/Sun, delta topocentric azimuth

%)*Bargument special text . . . . . . . . . . . 197, 237 Moon/Sun, delta topocentric azimuth %)Bargument special text . . . . . . . . . . . . . . . . . . . . . . . . . 197, 237

Moon/Sun, delta topocentric azimuth at standard moonrise time %(*Fargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta topocentric azimuth at standard moonrise time %(Fargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta topocentric azimuth at standard moonset time %)*Fargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta topocentric azimuth at standard moonset time %)Fargument special text . . 198, 238 Moon/Sun, delta topocentric elevation

%(*Aargument special text . . . . . . . . . . . 197, 237 Moon/Sun, delta topocentric elevation

%(Aargument special text . . . . . . . . . . . . 197, 237 Moon/Sun, delta topocentric elevation

%)*Aargument special text . . . . . . . . . . . 197, 237 Moon/Sun, delta topocentric elevation

%)Aargument special text . . . . . . . . . . . . 197, 237 Moon/Sun, delta topocentric elevation at standard

moonrise time %(*Eargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 237

Moon/Sun, delta topocentric elevation at standard moonrise time %(Eargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 237

Moon/Sun, delta topocentric elevation at standard moonset time %)*Eargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198, 238

Moon/Sun, delta topocentric elevation at standard moonset time %)Eargument special text . . 198, 238

Moonrise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Moonrise and moonset, set reference altitude … 39 Moonset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 More precise representation of astronomical data

and times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 More precise representation of geographical data

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 Morning width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Morning width geocentric, Sun . . . . . . . . . . . . . . . 179 Morning width topocentric, Sun . . . . . . . . . . . . . . 178 Move leap day to 1st March . . . . . . . . . . . . . . . . . . . 55 Move leap day to 28th February . . . . . . . . . . . . . . . 55 Multicultural New Year holidays in eternal holiday

list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

N

Name space of date variables . . . . . . . . . . . . . . 37, 103 Name space of text variables . . . . . . . . . . . . . . 38, 110 Native language defaults. . . . . . . . . . . . . . . . . . . . . . 132 Native language, selection . . . . . . . . . . . . . . . . . . . . 131 Native language, used . . . . . . . . . . . . . . . . . . . . . . . . 131 Nautical twilight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Nautical twilight, Sun . . . . . . . . . . . . . . . . . . . . . . . . 169 Negated character list . . . . . . . . . . . . . . . . . . . . . . . . 266 Negated character list, regexp . . . . . . . . . . . . . . . . 140 Negated regular expression. . . . . . . . . . . . . . . . . . . . . 45 Negative day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Negative day, biorhythm. . . . . . . . . . . . . . . . . . . . . . 157 New Moon phase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Non-printable characters in the text part, handling

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Noon height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Noon height geocentric, Moon . . . . . . . . . . . . . . . . 201 Noon height geocentric, Sun . . . . . . . . . . . . . . . . . . 178 Noon height topocentric, Moon . . . . . . . . . . . . . . . 200 Noon height topocentric, Sun . . . . . . . . . . . . . . . . . 178 Noon line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Noon line, Sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Noon phase angle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Noon phase angle geocentric, Moon . . . . . . . . . . . 202 Noon phase angle topocentric, Moon . . . . . . . . . . 200 Noon time of Moon, astronomical . . . . . . . . . . . . . 189 Noon time of Sun, astronomical. . . . . . . . . . . . . . . 166 NUL characters in the text part, handling . . . . . . . 97 Number of year calendar sheet blocks . . . . . . . . . . 31 Numerical local time timezone offset value

%=[argument] special text . . . . . . . . . . . . 161, 214 Numerical local time timezone offset value using

RFC-822 style, display . . . . . . . . . . . . . . . . . . . 161 Numerical text variables . . . . . . . . . . . . . . . . . . . . . . 111 Numerical text variables, length preserving

behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 Numerical Universal time timezone offset value

%"[argument] special text . . . . . . . . . . . . 161, 214 Numerical Universal time timezone offset value

using RFC-822 style, display. . . . . . . . . . . . . . 161

O

Obsolete date format elements, table. . . . . . . . . . 247 Obsolete date format texts, fixed default . . . . . . 247 Obsolete date format texts, individual . . . . . . . . 247 Obsolete date formats, table . . . . . . . . . . . . . . . . . . 247 Obsolete fixed default date format texts. . . . . . . 247 Obsolete fixed default date format texts, table

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 Obsolete individual date format texts . . . . . . . . . 247 Obsolete special texts and format elements. … 245 Obsolete special texts, table . . . . . . . . . . . . . . . . . . 246 Octal system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Omission values in date part, validity . . . . . . . . . . 88 Operations on date variables . . . . . . . . . . 37, 85, 105 Operations on text variables . . . . . . . . . . . . . . 85, 111 Operator, * regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 Operator, + regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, ? regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, \' regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Operator, \< regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, \> regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, \` regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Operator, \b regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, \B regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, \w regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, \W regexp . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Operator, | regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 Operators, regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Opposition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Option %date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Option @file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Option class, calendar . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Option class, common . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Option class, fixed date . . . . . . . . . . . . . . . . . . . . . . . . . 9 Option class, global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Option modifier . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 9 Options of gcal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Options of gcal2txt. . . . . . . . . . . . . . . . . . . . . . . . . . 128 Options of tcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Options of txt2gcal. . . . . . . . . . . . . . . . . . . . . . . . . . 127 Options, fixed date period . . . . . . . . . . . . . . . . . . . . . 55 Options, fixed date text. . . . . . . . . . . . . . . . . . . . . . . . 49 Options, long-style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Options, short-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Options, single character. . . . . . . . . . . . . . . . . . . . . . . 48 Orbit, Sun’s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 194 Orthodox new calendar holidays in eternal holiday

list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Orthodox old calendar holidays in eternal holiday

list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Orthodrome . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162, 267 Other resource file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Output of base time of astronomical functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 196 Output, pipelining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Output, redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

P

PAGER environment variable . . . . . . . . . . . . . . . 11, 256 Pager program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Pager, external . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Pager, internal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Pager, using . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 PATH environment variable . . . . . . . . . . . . . . . . 11, 256 Pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44, 267 Pattern, regular expressions . . . . . . . . . . . . . . . . . . 137 Period of double or shorter shadow length …. 180 Period of Gregorian Reformation, set . . . . . . . . . . 34 Period of longer than double shadow length … 180 Period of longer than single shadow length …. 180 Period of single or shorter shadow length . . . . . 180 Period, Exclusive date . . . . . . . . . . . . . . . . . . . . . . . . 146 Period, exclusive day . . . . . . . . . . . . . . . . . . . . . . . . . 148 Period, inclusive date . . . . . . . . . . . . . . . . . . . . . . . . . 146 Period, inclusive day . . . . . . . . . . . . . . . . . . . . . . . . . 148 Period-of-fixed-dates-modifier . . . . . . . . . . . . . . . . . . 47 Persian Jalaali holidays in eternal holiday list . . 16 Phase angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Pipelining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 267 Pipelining of output . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Positive day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Positive day, biorhythm . . . . . . . . . . . . . . . . . . . . . . 157

POSIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267

POSIX regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . 138, 141 Prayer time Asr, Islamic . . . . . . . . . . . . 166, 179, 180 Prayer time Fajr, Islamic . . . . . . . . . . . . . . . . . . . . . 170 Prayer time Isha, Islamic . . . . . . . . . . . . . . . . . . . . . 170 Prayer time Maghrib, Islamic . . . . . . . . . . . . . . . . . 168 Prayer time Zuhr, Islamic . . . . . . . . . . . . . . . . . . . . 166 Precedence, regexp operators . . . . . . . . . . . . . . . . . 142 Precision of astronomical calculations . . . . . . . . . . . 1 Preliminary remarks to Gcal . . . . . . . . . . . . . . . . . . . . 1 Private-access resource files . . . . . . . . . . . . . . . . . . . 255 Program exit state 127 . . . . . . . . . . . . . . . . . . . . . . . . 10 Program exit states, table . . . . . . . . . . . . . . . . . . . . 257 Proleptic calendar mode . . . . . . . . . . . . . . . . . . 34, 267 Proleptic Gregorian calendar. . . . . . . . . . . . . . . . . . . 35 Proleptic Julian calendar . . . . . . . . . . . . . . . . . . . . . . 35 Protection character, ‘\’ …. 39, 44, 46, 47, 51, 53,

97, 98, 100, 107, 206 Protection of special characters . . 8, 39, 44, 46, 47,

51, 53, 65, 97, 98, 100, 107, 206 Provide calendar sheet with week numbers . . . . . 30 Pseudo space character, ‘_’ . . 39, 44, 46, 47, 51, 53 Public-access resource files . . . . . . . . . . . . . . . . . . . . 255

Q

Quotation of special characters . . . . . . . . . . . . . . 8, 65

R

Range of current months. . . . . . . . . . . . . . . . . . . . . . . 75 Range of fiscal years . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Range of months of year . . . . . . . . . . . . . . . . . . . . . . . 75 Range of specified fiscal years . . . . . . . . . . . . . . . . . . 75 Range of specified month of selected years . . . . . 75 Range of specified months of selected year . . . . . 75 Range of specified months of selected years. . . . . 75 Range of specified years . . . . . . . . . . . . . . . . . . . . . . . 75 Ranges of commands . . . . . . . . . . . . . . . . . . 72, 75, 265 Ranges of days . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 Ranges of days, examples . . . . . . . . . . . . . . . . . . . . . . 91 Ranges of days, valid date parts . . . . . . . . . . . . . . . 91 Rational number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 267 Redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12, 268 Redirection of output . . . . . . . . . . . . . . . . . . . . . . . . . . 13 Redirection of standard output . . . . . . . . . . . . . . . . 84 Redirection of standard output, appending to file

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 Reference altitude used for moonrise and moonset, set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

Reference altitude used for sunrise and sunset, set

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 Reference systems used, Moon oriented special

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 Reference systems used, Sun oriented special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 References of text variables . . . . . . . . . . . . . . . . . . . . 85 Refraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Refraction, atmospheric. . . . . . . . . . . . . . . . . . . . . . . . 40 Regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137, 268 Regexp anchor metacharacter, $ . . . . . . . . . . . . . . 138 Regexp anchor metacharacter, ^. . . . . . . . . . . . . . 138 Regexp character class, [:alnum:] . . . . . . . . . . . . 139 Regexp character class, [:alpha:] . . . . . . . . . . . . 139 Regexp character class, [:blank:] . . . . . . . . . . . . 139 Regexp character class, [:cntrl:] . . . . . . . . . . . . 139 Regexp character class, [:digit:] . . . . . . . . . . . . 139 Regexp character class, [:graph:] . . . . . . . . . . . . 139 Regexp character class, [:lower:] . . . . . . . . . . . . 139 Regexp character class, [:print:] . . . . . . . . . . . . 139 Regexp character class, [:punct:] . . . . . . . . . . . . 139 Regexp character class, [:space:] . . . . . . . . . . . . 139 Regexp character class, [:upper:] . . . . . . . . . . . . 139 Regexp character class, [:xdigit:] . . . . . . . . . . . 139 Regexp character classes, [:keyword:] . . . . . . . . 138 Regexp character list, […] . . . . . . . . . . . . . . . . . . 138 Regexp collating symbols, [.element.] . . . . . . . 140 Regexp complemented character list, [^…] . . 140 Regexp equivalence classes, [=list=] . . . . . . . . . 140 Regexp grouping, (…) . . . . . . . . . . . . . . . . . . . . . . 140 Regexp interval expressions, {expression}…. 141 Regexp metacharacter, . . . . . . . . . . . . . . . . . . . . . . 138 Regexp metacharacter, \ . . . . . . . . . . . . . . . . . . . . . 138 Regexp metacharacters . . . . . . . . . . . . . . . . . . . . . . . 137 Regexp operator, * . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 Regexp operator, + . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, ? . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, \' . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Regexp operator, \< . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, \> . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, \` . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Regexp operator, \b . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, \B . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, \w . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, \W . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regexp operator, | . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 Regexp operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 Regexp operators, GNU specific . . . . . . . . . . . . . . 141 Regexp operators, precedence . . . . . . . . . . . . . . . . . 142 Regexp, anchors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 Regexp, buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 Regexp, buffer matching operators . . . . . . . . . . . . 142 Regexp, case sensitivity in matching . . . . . . . . . . 142 Regexp, character classes . . . . . . . . . . . . . . . . . . . . . 138 Regexp, character list . . . . . . . . . . . . . . . . . . . . . . . . 138 Regexp, collating elements . . . . . . . . . . . . . . . . . . . . 139 Regexp, collating symbols . . . . . . . . . . . . . . . . . . . . 140 Regexp, complemented character list. . . . . . . . . . 140 Regexp, definition of a word . . . . . . . . . . . . . . . . . . 141 Regexp, equivalence classes . . . . . . . . . . . . . . . . . . . 140 Regexp, ignoring case . . . . . . . . . . . . . . . . . . . . . 45, 142 Regexp, interval expressions . . . . . . . . . . . . . . . . . . 141 Regexp, matching word boundaries . . . . . . . . . . . 141 Regexp, negated character list . . . . . . . . . . . . . . . . 140 Regexp, POSIX. . . . . . . . . . . . . . . . . . . . . . . . . . 138, 141 Regular expression . . . . . . . . . . . . . . . . . . . 44, 137, 268 Regular expression anchor metacharacter, $ … 138 Regular expression anchor metacharacter, ^… 138 Regular expression character class, [:alnum:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:alpha:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:blank:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:cntrl:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:digit:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:graph:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:lower:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:print:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:punct:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:space:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:upper:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character class, [:xdigit:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 Regular expression character classes, [:keyword:]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 Regular expression character list, […] . . . . . . . 138 Regular expression collating symbols, [.element.]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 Regular expression complemented character list, [^ …] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

Regular expression equivalence class, [=list=]

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 Regular expression grouping, (…) . . . . . . . . . . . 140 Regular expression interval expressions,

{expression}. . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 Regular expression metacharacter, . . . . . . . . . . . 138 Regular expression metacharacter, \ . . . . . . . . . . 138 Regular expression metacharacters . . . . . . . . . . . . 137 Regular expression operator, * . . . . . . . . . . . . . . . . 140 Regular expression operator, + . . . . . . . . . . . . . . . . 141 Regular expression operator, ? . . . . . . . . . . . . . . . . 141 Regular expression operator, \'. . . . . . . . . . . . . . . 142 Regular expression operator, \<. . . . . . . . . . . . . . . 141 Regular expression operator, \>. . . . . . . . . . . . . . . 141 Regular expression operator, \`. . . . . . . . . . . . . . . 142 Regular expression operator, \b. . . . . . . . . . . . . . . 141 Regular expression operator, \B. . . . . . . . . . . . . . . 141 Regular expression operator, \w. . . . . . . . . . . . . . . 141 Regular expression operator, \W. . . . . . . . . . . . . . . 141 Regular expression operator, | . . . . . . . . . . . . . . . . 140 Regular expression operators . . . . . . . . . . . . . 137, 266 Regular expression, complemented . . . . . . . . . . . . . 45 Regular expression, ignoring case distinctions . . 45,

142 Regular expression, negated. . . . . . . . . . . . . . . . . . . . 45 Regular expression, reverted . . . . . . . . . . . . . . . . . . . 45 Regular expressions as patterns . . . . . . . . . . . . . . . 137 Regular expressions, how to use. . . . . . . . . . . . . . . 137 Regular expressions, short-list . . . . . . . . . . . . . . . . 143 Regular expressions, summary . . . . . . . . . . . . . . . . 143 Relative day number %d special text. . . . . . 207, 243 Relative month number %m special text … 207, 243 Relative time offset value . . . . . . . . . . . . . . . . . 33, 268 Relative week number %w special text …. 207, 243 Relative year number %y special text . . . . . 207, 243 Relative year number %ydate special text. . . . . 158,

213 Religious Islamic calendar . . . . . . . . . . . . . . . . . . 3, 268 Remark character, ‘;’ . . . . . . . . . . . . . . . . . . . . . 65, 101 Remarked line in resource file . . . . . . . . . . . . . 85, 101 Repetition factor of days, ‘:n’. . . . . . . . . . . . . . . . . . 92 Repetition factors of days, examples . . . . . . . . . . . 93 Repetition factors of days, valid date parts . . . . . 93 Replacements with date argument %?[date] special

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Replacements with other argument

%[format]?[argument] special texts. . . . . . . . 158 Replacements without any argument %? special

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 Representation-of-text-modifier . . . . . . . . . . . . . . . . 47 Resource file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 268 Resource file examples . . . . . . . . . . . . . . . . . . . . . . . . 114 Resource file line, additional . . . . . . . . . . . . . . . . . . . 46 Resource file line, commented . . . . . . . . . . . . . 85, 101 Resource file line, further . . . . . . . . . . . . . . . . . . . . . . 46 Resource file line, long text part . . . . . . . . . . . . . . . 97 Resource file line, remarked . . . . . . . . . . . . . . . 85, 101 Resource file, .gcalrc . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Resource file, additional . . . . . . . . . . . . . . . . . . . . 46, 82 Resource file, coding scheme . . . . . . . . . . . . . . . . . . 251 Resource file, commented line . . . . . . . . . . . . . 85, 101 Resource file, date part . . . . . . . . . . . . . . . . . . . . . . . 251 Resource file, date part of a line . . . . . . . . . . . . . . . 86 Resource file, force line break-up . . . . . . . . . . . . . . . 97 Resource file, include directives . . . . . . . . . . . . . . . 102 Resource file, limitations . . . . . . . . . . . 14, 84, 97, 103 Resource file, line continuation . . . . . . . . . . . . . . . . . 97 Resource file, loading . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Resource file, name . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Resource file, other . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 Resource file, remarked line . . . . . . . . . . . . . . . 85, 101 Resource file, remarking of text part . . . . . . . . . . . 97 Resource file, searching mechanism. . . . . . . . . . . . . 82 Resource file, standard. . . . . . . . . . . . . . . . . . . . . . . . . 82 Resource file, structure . . . . . . . . . . . . . . . . . . . . . . . . 84 Resource file, text part . . . . . . . . . . . . . . . . . . . . . . . . 97 Resource file, two parts of a line . . . . . . . . . . . . . . . 84 Resource file, use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 Respected environment variables, table . . . . . . . 255 Response file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Response file, @file . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Response file, creation . . . . . . . . . . . . . . . . . . . . . . . . . 10 Reverted regular expression . . . . . . . . . . . . . . . . . . . . 45 RFC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 RFC-822, numerical local time timezone offset

value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 RFC-822, numerical Universal time timezone offset

value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 Right ascension. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Right ascension, Moon. . . . . . . . . . . . . . . . . . . . . . . . 194 Right ascension, Sun . . . . . . . . . . . . . . . . . . . . . . . . . 172 Rise azimuth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Rise azimuth geocentric, Moon . . . . . . . . . . . . . . . 202 Rise azimuth geocentric, Sun . . . . . . . . . . . . . . . . . 179 Rise azimuth topocentric, Moon . . . . . . . . . . . . . . 201 Rise azimuth topocentric, Sun . . . . . . . . . . . . . . . . 178 Rise height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Rise height geocentric, Moon . . . . . . . . . . . . . . . . . 202 Rise height geocentric, Sun . . . . . . . . . . . . . . . . . . . 179 Rise height topocentric, Moon . . . . . . . . . . . . . . . . 200 Rise height topocentric, Sun . . . . . . . . . . . . . . . . . . 178 Rise phase angle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Rise phase angle geocentric, Moon . . . . . . . . . . . . 202 Rise phase angle topocentric, Moon . . . . . . . . . . . 201 Rise width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 Rise width geocentric, Moon . . . . . . . . . . . . . . . . . . 202 Rise width topocentric, Moon. . . . . . . . . . . . . . . . . 201 Run-time error codes, table . . . . . . . . . . . . . . . . . . . 257

S

Scope of date variables . . . . . . . . . . . . . . . . . . . . 37, 105 Scope of text variables . . . . . . . . . . . . . . . . . . . . 38, 112 Searching resource file . . . . . . . . . . . . . . . . . . . . . . . . . 82 Selected month mode . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Selected month of current year . . . . . . . . . . . . . . . . . 72 Selected year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Selection of native language . . . . . . . . . . . . . . . . . . 131 Sending eMail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Separator character of lists of days, ‘,’ . . . . . . . . . 89 Separator character of ranges of days, ‘#’. . . . . . . 90 Separator character, ‘+’ . . . . . . . . . . . . . . . . . . . . 17, 46 Separator character, ‘:’ . . . . . . . . . . . . . 12, 13, 37, 38 Separator character, whitespace . . . . . . . . 9, 84, 102 Sequence of control characters . . . . . . . . . . . . . . . . . 12 Set azimuth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Set azimuth geocentric, Moon . . . . . . . . . . . . . . . . 202 Set azimuth geocentric, Sun . . . . . . . . . . . . . . . . . . 179 Set azimuth topocentric, Moon . . . . . . . . . . . . . . . 201 Set azimuth topocentric, Sun . . . . . . . . . . . . . . . . . 178 Set cycle displacement value for Sun and Moon

data and times . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Set cycle ending time value for Sun and Moon data and times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

Set cycle timestep value for Sun and Moon data

and times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Set height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Set height geocentric, Moon . . . . . . . . . . . . . . . . . . 202 Set height geocentric, Sun . . . . . . . . . . . . . . . . . . . . 179 Set height topocentric, Moon . . . . . . . . . . . . . . . . . 200 Set height topocentric, Sun . . . . . . . . . . . . . . . . . . . 178 Set highlighting sequences . . . . . . . . . . . . . . . . . . . . . 12 Set marking characters . . . . . . . . . . . . . . . . . . . . . . . . 12 Set phase angle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 Set phase angle geocentric, Moon . . . . . . . . . . . . . 202 Set phase angle topocentric, Moon . . . . . . . . . . . . 201 Set shadow length factor. . . . . . . . . . . . . . . . . . . . . . . 39 Set starting day of week . . . . . . . . . . . . . . . . . . . . . . . 33 Set width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Set width geocentric, Moon . . . . . . . . . . . . . . . . . . . 202 Set width topocentric, Moon. . . . . . . . . . . . . . . . . . 201 Shadow length at afternoon, double . . . . . . . . . . . 180 Shadow length at afternoon, single . . . . . . . . . . . . 179 Shadow length at forenoon, double. . . . . . . . . . . . 180 Shadow length at forenoon, single. . . . . . . . . . . . . 179 Shadow length factor, set . . . . . . . . . . . . . . . . . . . . . . 39 Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 8, 43, 107, 268 Shell command %![argument] special text . . . . . 210,

243 Shell command, execution . . . . . . . . . . . . . . . . . . . . 210 Shell script, creation . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Short help text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Short-list of all %?… special texts . . . . . . . . . . . . 211 Short-list of all regular expressions . . . . . . . . . . . . 143 Short-style option . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 268 Sidereal time, Moon . . . . . . . . . . . . . . . . . . . . . . . . . . 196 Sidereal time, Sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 Signed angular value, Sun and Moon oriented

special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Signed decimal based rational number value, Sun

and Moon oriented special texts . . . . . . . . . . 181 Signed time value, Sun and Moon oriented special

texts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Simple decrement of date variables. . . . . . . . . . . . 105 Simple decrement of text variables . . . . . . . . . . . . 111 Simple file name . . . . . . . . . . . . . . . . . . . . . . . . . . 83, 269 Simple increment of date variables . . . . . . . . . . . . 105 Simple increment of text variables . . . . . . . . . . . . 111 Simple language code information . . . . . . . . 131, 269 Single character options . . . . . . . . . . . . . . . . . . . . . 9, 48 Single command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 Single command line word. . . . . . . . . . . . . . . . . . . 9, 48 Single commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Single Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Single fiscal year. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Single month mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 Single month of selected year . . . . . . . . . . . . . . . . . . 73 Single or shorter shadow length, period . . . . . . . 180 Single shadow length at afternoon . . . . . . . . . . . . 179 Single shadow length at forenoon . . . . . . . . . . . . . 179 Single year mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Software license, finding . . . . . . . . . . . . . . . . . . . . . . . 10 Solar Arctic Circle . . . . . . . . . . . . . . . . . . . . . . . 186, 269 Solar day length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Solar day, mean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 Solar day, true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 Solar eclipse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Solar night length . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Solar time, mean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 Solar time, true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 Solstice in summer . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 Solstice in winter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Solstice, summer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Solstice, winter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Special 3-Month mode commands . . . . . . . . . . 72, 73 Special 3-Month mode, . . . . . . . . . . . . . . . . . . . . . . . 73 Special 3-Month mode, . yyyy . . . . . . . . . . . . . . . . . 73 Special 3-Month mode, .+ . . . . . . . . . . . . . . . . . . . . . 74 Special 3-Month mode, .+ yyyy . . . . . . . . . . . . . . . . 74 Special 3-Month mode, .- . . . . . . . . . . . . . . . . . . . . . 74 Special 3-Month mode, .- yyyy . . . . . . . . . . . . . . . . 74 Special 3-Month mode, .. . . . . . . . . . . . . . . . . . . . . . 74 Special 3-Month mode, .. yyyy . . . . . . . . . . . . . . . . 74 Special calendar format . . . . . . . . . . . . . . . . . . . . . . . . 29 Special character sequences in the text part,

handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Special characters, country specific. . . . . . . . . . . . . 37 Special characters, depreciation . . 8, 39, 44, 46, 47,

51, 53, 65, 97, 98, 100, 107, 206 Special characters, protection …. 8, 39, 44, 46, 47,

51, 53, 65, 97, 98, 100, 107, 206 Special characters, quotation . . . . . . . . . . . . . . . . 8, 65 Special date format in calendar sheet . . . . . . . . . . 32 Special date format in eternal holiday list . . . . . . 32 Special date format in fixed date list . . . . . . . . . . . 32 Special event oriented texts, Moon data and times

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Special event oriented texts, Sun data and times

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 Special text %![argument], shell command . . . . . 210 Special text %"[argument], numerical Universal

time timezone offset value . . . . . . . . . . . . . . . . 161 Special text %,[date], biorhythm text . . . . . . . . . 157 Special text %;[date], biorhythm text graphics

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 Special text %=[argument], numerical local time

timezone offset value . . . . . . . . . . . . . . . . . . . . . 161 Special text %>02&*D[date], day number. . . . . . . 154 Special text %>02&*M[date], month number …. 155 Special text %>02*D[date], day number . . . . . . . . 154 Special text %>02*M[date], month number . . . . . 155 Special text %>03&*N[date], day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Special text %>03*N[date], day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Special text %>03*O[date], Moon phase text … 156 Special text %>04*Y[date], year number . . . . . . . 155 Special text %>1&*B[date], age value . . . . . . . . . . 158 Special text %>1&*D[date], day number . . . . . . . . 154 Special text %>1&*E[date], weekday number … 151 Special text %>1&*I[date], weekday number … 151 Special text %>1&*J[date], weekday number … 152 Special text %>1&*M[date], month number . . . . . 155 Special text %>1&*N[date], day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Special text %>1&*S[date], weekday number … 152 Special text %>1&*T[date], weekday number … 153 Special text %>1&*W[date], weekday number … 151 Special text %>2#K[date], 2-letter weekday name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 Special text %>3#K[date], 3-letter weekday name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Special text %>3#U[date], 3-letter month name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 Special text %@[argument], actual Universal time in hh:mm/date format. . . . . . . . . . . . . . . . . . . . . . . 160 Special text %[format]'[argument], actual local

time in m[m…] format . . . . . . . . . . . . . . . . . . 159 Special text %[format]-[argument], environment

varibale . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 Special text %[format].[argument], actual Universal

time hour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 Special text %[format]/[argument], actual Universal

time minute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 Special text %[format]?[argument], actual local

time minute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 Special text %[format]_[argument], actual local

time hour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Special text %[format]`[argument], actual Universal

time in m[m…]/date format. . . . . . . . . . . . . . 160 Special text %[format]{[argument], actual local

time 12-hour format suffix. . . . . . . . . . . . . . . . 160 Special text %[format]}[argument], actual Universal

time 12-hour format suffix. . . . . . . . . . . . . . . . 161 Special text %[format]b[argument], geographical

distance and course angle . . . . . . . . . . . . . . . . 162 Special text %[format]D[date], day number …. 153 Special text %[format]j[argument], Julian day

number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 Special text %[format]K[date], weekday name . . 150 Special text %[format]M[date], month number . . 154 Special text %[format]N[date], day-of-year number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 Special text %[format]U[date], month name. … 154 Special text %[format]Y[date], year number. … 155 Special text %0, end of highlighting sequence 1 or 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 Special text %1, start of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 Special text %2, end of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 Special text %3, start of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 Special text %4, end of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 Special text %5, start of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 Special text %6, end of highlighting sequence 1

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 Special text %7, start of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 Special text %8, end of highlighting sequence 2

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 Special text %9, start of highlighting sequence 1 or

2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 Special text %A, exclusive Monday . . . . . . . . . . . . 149 Special text %a, inclusive Monday . . . . . . . . . . . . . 149 Special text %B[date], age value . . . . . . . . . . . . . . . 158 Special text %C, exclusive Tuesday . . . . . . . . . . . . 149 Special text %c, inclusive Tuesday. . . . . . . . . . . . . 149 Special text %d, relative day number . . . . . . . . . . 207 Special text %D[date], day number . . . . . . . . . . . . 153 Special text %E[date], weekday number. . . . . . . . 151 Special text %e[date][#[date]], exclusive date

period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 Special text %F, exclusive Wednesday . . . . . . . . . 149 Special text %f, inclusive Wednesday . . . . . . . . . . 149 Special text %G, exclusive Thursday . . . . . . . . . . . 149 Special text %g, inclusive Thursday . . . . . . . . . . . 149 Special text %H, exclusive Friday . . . . . . . . . . . . . . 149 Special text %h, inclusive Friday . . . . . . . . . . . . . . 149 Special text %I[date], weekday number. . . . . . . . 151 Special text %i[date][#[date]], inclusive date period

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 Special text %J[date], weekday number. . . . . . . . 152 Special text %K[date], complete weekday name

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Special text %k[date], week number . . . . . . . . . . . 154 Special text %L, exclusive Saturday. . . . . . . . . . . . 150 Special text %l, inclusive Saturday . . . . . . . . . . . . 149 Special text %m, relative month number . . . . . . . 207 Special text %M[date], month number. . . . . . . . . . 154 Special text %n[argument], textual date . . . . . . . 162 Special text %N[date], day-of-year number . . . . . 153 Special text %O[date], Moon phase text . . . . . . . 155 Special text %P, exclusive Sunday . . . . . . . . . . . . . 150 Special text %p, inclusive Sunday . . . . . . . . . . . . . 149 Special text %Q, exclusive Monday…Thursday

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Special text %q, inclusive Monday…Thursday

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 Special text %R, exclusive Monday…Friday … 150 Special text %r, inclusive Monday…Friday … 149 Special text %S[date], weekday number. . . . . . . . 152 Special text %t[argument], actual local time in

hh:mm format . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 Special text %T[date], weekday number. . . . . . . . 152 Special text %U[date], complete month name . . 154 Special text %V, exclusive legal holidays . . . . . . . 149 Special text %v, inclusive legal holidays . . . . . . . 149 Special text %w, relative week number . . . . . . . . . 207 Special text %W[date], weekday number. . . . . . . . 151 Special text %X, exclusive all holidays . . . . . . . . . 149 Special text %x, inclusive all holidays. . . . . . . . . . 149 Special text %y, relative year number. . . . . . . . . . 207 Special text %Y[date], year number . . . . . . . . . . . . 155 Special text %ydate, relative year number . . . . . 158 Special text %Z[date], Moon phase text graphics

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 Special text starting character, ‘%’ . . . . . . . . . . . . 100 Special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 Special texts %?, exclusions without any argument

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 Special texts %?, exclusive day period . . . . . . . . . 149 Special texts %?, highlighting . . . . . . . . . . . . . . . . . 208 Special texts %?, inclusive day period . . . . . . . . . 149 Special texts %?, replacements without any

argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 Special texts %?…, date exclusion. . . . . . . . . . . . 146 Special texts %?…, text replacement . . . . . . . . . 150 Special texts %?[date], biorhythm . . . . . . . . . . . . . 156 Special texts %?[date], exclusions with date

argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 Special texts %?[date], replacements with date

argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 Special texts %[format]?, difference value . . . . . . 206 Special texts %[format]?[argument], actual

clocktime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 Special texts %[format]?[argument], replacements

with other argument . . . . . . . . . . . . . . . . . . . . . 158 Special texts %[format]?[date], difference value

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 Special texts %[format]?[date], Moon phase … 155 Special texts %[format]?[date], weekday number

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 Special texts %[format]?argument, Moon data

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Special texts %[format]?argument, Sun data … 165 Special texts, obsolete . . . . . . . . . . . . . . . . . . . . . . . . 245 Specified fiscal years, list. . . . . . . . . . . . . . . . . . . . . . . 74 Specified fiscal years, range . . . . . . . . . . . . . . . . . . . . 75 Specified month of selected years, range . . . . . . . . 75 Specified months of selected year, list . . . . . . . . . . 74 Specified months of selected year, range . . . . . . . . 75 Specified months of selected years, range . . . . . . . 75 Specified months of the current year, list . . . . . . . 74 Specified years, list . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 Specified years, range . . . . . . . . . . . . . . . . . . . . . . . . . . 75 Standard and special date format in calendar sheet

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Standard and special date format in eternal

holiday list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Standard and special date format in fixed date list

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Standard calendar format . . . . . . . . . . . . . . . . . . . . . . 29 Standard non-visibility period, Moon . . . . . . . . . 193 Standard non-visibility period, Sun . . . . . . . . . . . 169 Standard rise time, Moon. . . . . . . . . . . . . . . . . . . . . 192 Standard rise time, Sun. . . . . . . . . . . . . . . . . . . . . . . 168 Standard set time, Moon . . . . . . . . . . . . . . . . . . . . . 192 Standard set time, Sun . . . . . . . . . . . . . . . . . . . . . . . 168 Standard visibility period, Moon . . . . . . . . . . . . . . 192 Standard visibility period, Sun. . . . . . . . . . . . . . . . 168 Standard week number …. 31, 52, 56, 58, 86, 154,

269 Star day . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 196, 269 Star time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 196 Start of highlighting sequence 1 %1 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208, 243 Start of highlighting sequence 1 %5 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208, 243 Start of highlighting sequence 1 or 2 %9 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209, 243 Start of highlighting sequence 2 %3 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208, 243 Start of highlighting sequence 2 %7 special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209, 243 Starting date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 Starting date of fixed date . . . . . . . . . . . . . . . 146, 147 Starting day of week, setting . . . . . . . . . . . . . . . . . . . 33 Statute mile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163, 269 String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 Structure of date part . . . . . . . . . . . . . . . . . . . . . . . . . 86 Structure of resource file . . . . . . . . . . . . . . . . . . . . . . . 84 Style of representation of Moon data and times

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180, 202 Style of representation of Sun data and times . . 180 Style of representation, geographical distance and

course angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 Subtraction by days of date variables . . . . . . . . . 105 Subtraction by weekdays of date variables. . . . . 105 Subtraction of text variables . . . . . . . . . . . . . . . . . . 111 Summary of all %?… special texts . . . . . . . . . . . 211 Summary of all regular expressions . . . . . . . . . . . 143 Summer point on the ecliptic . . . . . . . . . . . . . . . . . 269 Summer solstice . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 269 Summertime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 Sun and Moon data and times, set cycle

displacement value . . . . . . . . . . . . . . . . . . . . . . . . 42 Sun and Moon data and times, set cycle ending

time value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Sun and Moon data and times, set cycle timestep

value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 Sun and Moon oriented special texts, Clocktime

value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Sun and Moon oriented special texts, signed

angular value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Sun and Moon oriented special texts, signed

decimal based rational number value . . . . . 181 Sun and Moon oriented special texts, signed time

value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Sun and Moon oriented special texts, unsigned

angular value . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Sun and Moon oriented special texts, unsigned

decimal based rational number value . . . . . 181 Sun and Moon oriented special texts, unsigned

time value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Sun data %[format]?argument special texts … 165,

215 Sun data and times, default mode . . . . . . . . . . . . 180 Sun data and times, mode characters . . . . . . . . . 165 Sun data and times, special event oriented texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 Sun data and times, style of representation … 180 Sun data and times, type of representation …. 180 Sun oriented special texts, limitations. . . . . . . . . 186 Sun oriented special texts, reference systems used

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 Sun oriented special texts, table of modes . . . . . 165 Sun oriented special texts, table of representation

types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 Sun’s center, begin of amateur-astronomical

%o*8argument special text. . . . . . . . . . . . . . . . 218 Sun’s center, begin of amateur-astronomical

%o8argument special text . . . . . . . . . . . . . . . . . 218 Sun’s center, begin of amateur-astronomical

twilight %o*8argument special text . . . . . . . 170 Sun’s center, begin of amateur-astronomical

twilight %o8argument special text. . . . . . . . . 170 Sun’s center, begin of astronomical twilight

%o*9argument special text . . . . . . . . . . . 170, 219 Sun’s center, begin of astronomical twilight

%o9argument special text . . . . . . . . . . . . 170, 219 Sun’s center, begin of civil twilight %o*6argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 169, 217 Sun’s center, begin of civil twilight %o6argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 169, 217 Sun’s center, begin of nautical twilight

%o*7argument special text . . . . . . . . . . . 169, 218 Sun’s center, begin of nautical twilight

%o7argument special text . . . . . . . . . . . . 169, 218 Sun’s center, end of amateur-astronomical

%s*8argument special text. . . . . . . . . . . . . . . . 218 Sun’s center, end of amateur-astronomical

%s8argument special text . . . . . . . . . . . . . . . . . 218 Sun’s center, end of amateur-astronomical twilight

%s*8argument special text. . . . . . . . . . . . . . . . 170 Sun’s center, end of amateur-astronomical twilight

%s8argument special text . . . . . . . . . . . . . . . . . 170 Sun’s center, end of astronomical twilight

%s*9argument special text . . . . . . . . . . . 170, 219 Sun’s center, end of astronomical twilight

%s9argument special text . . . . . . . . . . . . 170, 219 Sun’s center, end of civil twilight %s*6argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 169, 217 Sun’s center, end of civil twilight %s6argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 169, 217 Sun’s center, end of nautical twilight

%s*7argument special text . . . . . . . . . . . 169, 218 Sun’s center, end of nautical twilight %s7argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 169, 218 Sun’s center, geometrical non-visibility

%z*2argument special text . . . . . . . . . . . 166, 215 Sun’s center, geometrical non-visibility

%z2argument special text . . . . . . . . . . . . 166, 215 Sun’s center, geometrical rise time %o*2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun’s center, geometrical rise time %o2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun’s center, geometrical set time %s*2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun’s center, geometrical set time %s2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun’s center, geometrical visibility %u*2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun’s center, geometrical visibility %u2argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun’s center, non-visibility %z*4argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168, 216 Sun’s center, non-visibility %z4argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168, 216 Sun’s center, period above amateur-astronomical

twilight reference altitude %u*8argument special text . . . . . . . . . . . . . . . . . . . . . . . . . 170, 219 Sun’s center, period above amateur-astronomical

twilight reference altitude %u8argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170, 218

Sun’s center, period above astronomical twilight reference altitude %u*9argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219

Sun’s center, period above astronomical twilight reference altitude %u9argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219

Sun’s center, period above civil twilight reference

altitude %u*6argument special text . . 169, 217 Sun’s center, period above civil twilight reference

altitude %u6argument special text. … 169, 217 Sun’s center, period above nautical twilight

reference altitude %u*7argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169, 218

Sun’s center, period above nautical twilight reference altitude %u7argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169, 218

Sun’s center, period below amateur-astronomical twilight reference altitude %z*8argument special text . . . . . . . . . . . . . . . . . . . . . . . . . 170, 219

Sun’s center, period below amateur-astronomical twilight reference altitude %z8argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170, 219

Sun’s center, period below astronomical twilight reference altitude %z*9argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219

Sun’s center, period below astronomical twilight reference altitude %z9argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219

Sun’s center, period below civil twilight reference

altitude %z*6argument special text . . 169, 218 Sun’s center, period below civil twilight reference

altitude %z6argument special text. … 169, 217 Sun’s center, period below nautical twilight

reference altitude %z*7argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170, 218

Sun’s center, period below nautical twilight reference altitude %z7argument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170, 218

Sun’s center, rise time %o*4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 216 Sun’s center, rise time %o4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 216 Sun’s center, set time %s*4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 216 Sun’s center, set time %s4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 216 Sun’s center, visibility %u*4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 216 Sun’s center, visibility %u4argument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167, 216 Sun’s orbit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 194 Sun’s upper limb, geometrical non-visibility

%z*3argument special text . . . . . . . . . . . 167, 216 Sun’s upper limb, geometrical non-visibility

%z3argument special text . . . . . . . . . . . . 167, 216 Sun’s upper limb, geometrical rise time

%o*3argument special text . . . . . . . . . . . 167, 216 Sun’s upper limb, geometrical rise time

%o3argument special text . . . . . . . . . . . . 167, 215 Sun’s upper limb, geometrical set time

%s*3argument special text . . . . . . . . . . . 167, 216 Sun’s upper limb, geometrical set time

%s3argument special text . . . . . . . . . . . . 167, 216 Sun’s upper limb, geometrical visibility

%u*3argument special text . . . . . . . . . . . 167, 216 Sun’s upper limb, geometrical visibility

%u3argument special text . . . . . . . . . . . . 167, 216 Sun’s upper limb, standard non-visibility

%z*5argument special text . . . . . . . . . . . 169, 217 Sun’s upper limb, standard non-visibility

%z5argument special text . . . . . . . . . . . . 169, 217 Sun’s upper limb, standard rise time %o*5argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 168, 217 Sun’s upper limb, standard rise time %o5argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 168, 217 Sun’s upper limb, standard set time %s*5argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 168, 217 Sun’s upper limb, standard set time %s5argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 168, 217 Sun’s upper limb, standard visibility %u*5argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 168, 217 Sun’s upper limb, standard visibility %u5argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 168, 217 Sun, amateur-astronomical twilight . . . . . . . . . . . 170 Sun, astronomical midnight time %o*0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 165, 215 Sun, astronomical midnight time %o0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 165, 215 Sun, astronomical midnight time %s*0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 165, 215 Sun, astronomical midnight time %s0argument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 165, 215 Sun, astronomical noon time %o*1argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun, astronomical noon time %o1argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun, astronomical noon time %s*1argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun, astronomical noon time %s1argument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166, 215 Sun, astronomical twilight . . . . . . . . . . . . . . . . . . . . 170 Sun, astronomical unit. . . . . . . . . . . . . . . . . . . . . . . . 172 Sun, azimuth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Sun, base time %o*targument special text …. 173,

223 Sun, base time %otargument special text . . 173, 223 Sun, base time %s*targument special text …. 173,

223 Sun, base time %stargument special text . . 173, 223 Sun, civil twilight . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Sun, delta-t %o*rargument special text … 173, 222 Sun, delta-t %orargument special text …. 173, 222 Sun, delta-t %s*rargument special text … 173, 222 Sun, delta-t %srargument special text …. 173, 222 Sun, double shadow length time at afternoon

%s*Rargument special text . . . . . . . . . . . 180, 228 Sun, double shadow length time at afternoon

%sRargument special text . . . . . . . . . . . . 180, 228 Sun, double shadow length time at forenoon

%o*Rargument special text . . . . . . . . . . . 180, 228 Sun, double shadow length time at forenoon

%oRargument special text . . . . . . . . . . . . 180, 228 Sun, equation of time %o*wargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, equation of time %owargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, equation of time %s*wargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, equation of time %swargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, evening width geocentric . . . . . . . . . . . . . . . . 179 Sun, evening width topocentric . . . . . . . . . . . . . . . 178 Sun, geocentric azimuth %o*kargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221

Sun, geocentric azimuth %okargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric azimuth %s*kargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221

Sun, geocentric azimuth %skargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric azimuth at standard rise time

%o*Pargument special text . . . . . . . . . . . 179, 227 Sun, geocentric azimuth at standard rise time

%oPargument special text . . . . . . . . . . . . 179, 227 Sun, geocentric azimuth at standard set time

%s*Pargument special text . . . . . . . . . . . 179, 228 Sun, geocentric azimuth at standard set time

%sPargument special text . . . . . . . . . . . . 179, 228 Sun, geocentric data . . . . . . . . . . . . . . . . . . . . . . . . . . 171 Sun, geocentric declination %o*largument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric declination %olargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric declination %s*largument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric declination %slargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric distance %o*oargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222

Sun, geocentric distance %ooargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric distance %s*oargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222

Sun, geocentric distance %soargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric ecliptic longitude %o*margument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric ecliptic longitude %omargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric ecliptic longitude %s*margument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric ecliptic longitude %smargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric elevation %o*jargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric elevation %ojargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric elevation %s*jargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric elevation %sjargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, geocentric elevation at standard rise time

%o*Oargument special text . . . . . . . . . . . 179, 227 Sun, geocentric elevation at standard rise time

%oOargument special text . . . . . . . . . . . . 179, 227 Sun, geocentric elevation at standard set time

%s*Oargument special text . . . . . . . . . . . 179, 227 Sun, geocentric elevation at standard set time

%sOargument special text . . . . . . . . . . . . 179, 227 Sun, geocentric elevation at Sun’s astronomical

midnight time %o*Margument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 227

Sun, geocentric elevation at Sun’s astronomical midnight time %oMargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 227

Sun, geocentric elevation at Sun’s astronomical midnight time %s*Margument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 227

Sun, geocentric elevation at Sun’s astronomical midnight time %sMargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178, 227

Sun, geocentric elevation at Sun’s astronomical noon time %o*Nargument special text …. 178, 227

Sun, geocentric elevation at Sun’s astronomical

noon time %oNargument special text . . 178, 227 Sun, geocentric elevation at Sun’s astronomical

noon time %s*Nargument special text …. 178, 227

Sun, geocentric elevation at Sun’s astronomical

noon time %sNargument special text . . 178, 227 Sun, geocentric horizontal parallax %o*pargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric horizontal parallax %opargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric horizontal parallax %s*pargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric horizontal parallax %spargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric right ascension %o*nargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric right ascension %onargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric right ascension %s*nargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric right ascension %snargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric semidiameter %o*qargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric semidiameter %oqargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric semidiameter %s*qargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, geocentric semidiameter %sqargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 222 Sun, highest culmination point. . . . . . . . . . . . . . . . 166 Sun, Julian date %o*uargument special text… 174,

223 Sun, Julian date %ouargument special text …. 174,

223 Sun, Julian date %s*uargument special text… 174,

223 Sun, Julian date %suargument special text …. 174,

223 Sun, Julian Ephemeris date %o*vargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, Julian Ephemeris date %ovargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, Julian Ephemeris date %s*vargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, Julian Ephemeris date %svargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 223 Sun, local sidereal time %o*sargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 223 Sun, local sidereal time %osargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 223 Sun, local sidereal time %s*sargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 223 Sun, local sidereal time %ssargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173, 223 Sun, lowest culmination point. . . . . . . . . . . . . . . . . 165 Sun, mean. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 Sun, midnight height geocentric. . . . . . . . . . . . . . . 178 Sun, midnight height topocentric . . . . . . . . . . . . . 177 Sun, morning width geocentric. . . . . . . . . . . . . . . . 179 Sun, morning width topocentric . . . . . . . . . . . . . . 178 Sun, nautical twilight. . . . . . . . . . . . . . . . . . . . . . . . . 169 Sun, noon height geocentric. . . . . . . . . . . . . . . . . . . 178 Sun, noon height topocentric . . . . . . . . . . . . . . . . . 178 Sun, noon line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Sun, period of double or shorter shadow length

%u*Rargument special text . . . . . . . . . . . 180, 229 Sun, period of double or shorter shadow length

%uRargument special text . . . . . . . . . . . . 180, 229 Sun, period of longer than double shadow length

%z*Rargument special text . . . . . . . . . . . 180, 229 Sun, period of longer than double shadow length

%zRargument special text . . . . . . . . . . . . 180, 229 Sun, period of longer than single shadow length

%z*Qargument special text . . . . . . . . . . . 180, 228 Sun, period of longer than single shadow length

%zQargument special text . . . . . . . . . . . . 180, 228 Sun, period of single or shorter shadow length

%u*Qargument special text . . . . . . . . . . . 180, 228 Sun, period of single or shorter shadow length

%uQargument special text . . . . . . . . . . . . 180, 228 Sun, refraction %o*iargument special text …. 172,

221 Sun, refraction %oiargument special text . . . . . 172,

221 Sun, refraction %s*iargument special text …. 172,

221 Sun, refraction %siargument special text . . . . . 172,

221 Sun, rise azimuth geocentric . . . . . . . . . . . . . . . . . . 179 Sun, rise azimuth topocentric . . . . . . . . . . . . . . . . . 178 Sun, rise height geocentric . . . . . . . . . . . . . . . . . . . . 179 Sun, rise height topocentric . . . . . . . . . . . . . . . . . . . 178 Sun, set azimuth geocentric . . . . . . . . . . . . . . . . . . . 179 Sun, set azimuth topocentric. . . . . . . . . . . . . . . . . . 178 Sun, set height geocentric. . . . . . . . . . . . . . . . . . . . . 179 Sun, set height topocentric . . . . . . . . . . . . . . . . . . . 178 Sun, single shadow length time at afternoon

%s*Qargument special text . . . . . . . . . . . 179, 228 Sun, single shadow length time at afternoon

%sQargument special text . . . . . . . . . . . . 179, 228 Sun, single shadow length time at forenoon

%o*Qargument special text . . . . . . . . . . . 179, 228 Sun, single shadow length time at forenoon

%oQargument special text . . . . . . . . . . . . 179, 228 Sun, standard non-visibility period. . . . . . . . . . . . 169 Sun, standard rise time . . . . . . . . . . . . . . . . . . . . . . . 168 Sun, standard set time. . . . . . . . . . . . . . . . . . . . . . . . 168 Sun, standard visibility period . . . . . . . . . . . . . . . . 168 Sun, topocentric azimuth %o*bargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric azimuth %obargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric azimuth %s*bargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric azimuth %sbargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric azimuth at standard rise time

%o*Largument special text . . . . . . . . . . . 178, 227 Sun, topocentric azimuth at standard rise time

%oLargument special text . . . . . . . . . . . . 178, 226 Sun, topocentric azimuth at standard set time

%s*Largument special text . . . . . . . . . . . 178, 227 Sun, topocentric azimuth at standard set time

%sLargument special text . . . . . . . . . . . . 178, 227 Sun, topocentric data. . . . . . . . . . . . . . . . . . . . . . . . . 171 Sun, topocentric declination %o*cargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric declination %ocargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric declination %s*cargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric declination %scargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric distance %o*fargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric distance %ofargument special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric distance %s*fargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric distance %sfargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220

Sun, topocentric ecliptic longitude %o*dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric ecliptic longitude %odargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric ecliptic longitude %s*dargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric ecliptic longitude %sdargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 171, 220 Sun, topocentric elevation %o*aargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219 Sun, topocentric elevation %oaargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219 Sun, topocentric elevation %s*aargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219 Sun, topocentric elevation %saargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 219 Sun, topocentric elevation at standard rise time

%o*Kargument special text . . . . . . . . . . . 178, 226 Sun, topocentric elevation at standard rise time

%oKargument special text . . . . . . . . . . . . 178, 226 Sun, topocentric elevation at standard set time

%s*Kargument special text . . . . . . . . . . . 178, 226 Sun, topocentric elevation at standard set time

%sKargument special text . . . . . . . . . . . . 178, 226 Sun, topocentric elevation at Sun’s astronomical midnight time %o*Iargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226

Sun, topocentric elevation at Sun’s astronomical midnight time %oIargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226

Sun, topocentric elevation at Sun’s astronomical midnight time %s*Iargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226

Sun, topocentric elevation at Sun’s astronomical midnight time %sIargument special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226

Sun, topocentric elevation at Sun’s astronomical noon time %o*Jargument special text …. 178, 226

Sun, topocentric elevation at Sun’s astronomical

noon time %oJargument special text . . 178, 226 Sun, topocentric elevation at Sun’s astronomical

noon time %s*Jargument special text …. 178, 226

Sun, topocentric elevation at Sun’s astronomical

noon time %sJargument special text . . 178, 226 Sun, topocentric horizontal parallax %o*gargument special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221

Sun, topocentric horizontal parallax %ogargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric horizontal parallax %s*gargument special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221

Sun, topocentric horizontal parallax %sgargument special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric right ascension %o*eargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric right ascension %oeargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric right ascension %s*eargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric right ascension %seargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 220 Sun, topocentric semidiameter %o*hargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, topocentric semidiameter %ohargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, topocentric semidiameter %s*hargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun, topocentric semidiameter %shargument special

text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172, 221 Sun/Moon, delta astronomical midnight time

%o*Fargument special text . . . . . . . . . . . 176, 225 Sun/Moon, delta astronomical midnight time

%oFargument special text . . . . . . . . . . . . 176, 225 Sun/Moon, delta astronomical midnight time

%s*Fargument special text . . . . . . . . . . . 176, 225 Sun/Moon, delta astronomical midnight time

%sFargument special text . . . . . . . . . . . . 176, 225 Sun/Moon, delta astronomical noon time

%o*Gargument special text . . . . . . . . . . . 177, 225 Sun/Moon, delta astronomical noon time

%oGargument special text . . . . . . . . . . . . 177, 225 Sun/Moon, delta astronomical noon time

%s*Gargument special text . . . . . . . . . . . 177, 225 Sun/Moon, delta astronomical noon time

%sGargument special text . . . . . . . . . . . . 177, 225 Sun/Moon, delta geocentric azimuth %o*Aargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric azimuth %oAargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric azimuth %s*Aargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric azimuth %sAargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric azimuth at standard

sunrise time %o*Eargument special text . . 176, 225

Sun/Moon, delta geocentric azimuth at standard sunrise time %oEargument special text … 176, 225

Sun/Moon, delta geocentric azimuth at standard sunset time %s*Eargument special text … 176, 225

Sun/Moon, delta geocentric azimuth at standard sunset time %sEargument special text …. 176, 225

Sun/Moon, delta geocentric elevation

%o*zargument special text . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric elevation %ozargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric elevation

%s*zargument special text . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric elevation %szargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 224 Sun/Moon, delta geocentric elevation at standard

sunrise time %o*Dargument special text . . 176, 225

Sun/Moon, delta geocentric elevation at standard sunrise time %oDargument special text … 176, 225

Sun/Moon, delta geocentric elevation at standard sunset time %s*Dargument special text … 176, 225

Sun/Moon, delta geocentric elevation at standard sunset time %sDargument special text …. 176, 225

Sun/Moon, delta standard rise time %(*Kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 Sun/Moon, delta standard rise time %(Kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 Sun/Moon, delta standard rise time %o*Hargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226 Sun/Moon, delta standard rise time %oHargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226 Sun/Moon, delta standard set time %)*Kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Sun/Moon, delta standard set time %)Kargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 Sun/Moon, delta standard set time %s*Hargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226 Sun/Moon, delta standard set time %sHargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 177, 226 Sun/Moon, delta topocentric azimuth

%o*yargument special text . . . . . . . . . . . 175, 224 Sun/Moon, delta topocentric azimuth %oyargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 223 Sun/Moon, delta topocentric azimuth

%s*yargument special text . . . . . . . . . . . 175, 224 Sun/Moon, delta topocentric azimuth %syargument

special text . . . . . . . . . . . . . . . . . . . . . . . . . 175, 223 Sun/Moon, delta topocentric azimuth at standard

sunrise time %o*Cargument special text . . 175, 224

Sun/Moon, delta topocentric azimuth at standard sunrise time %oCargument special text … 175, 224

Sun/Moon, delta topocentric azimuth at standard sunset time %s*Cargument special text … 176, 225

Sun/Moon, delta topocentric azimuth at standard sunset time %sCargument special text …. 176, 224

Sun/Moon, delta topocentric elevation

%o*xargument special text . . . . . . . . . . . 174, 223 Sun/Moon, delta topocentric elevation

%oxargument special text . . . . . . . . . . . . 174, 223 Sun/Moon, delta topocentric elevation

%s*xargument special text . . . . . . . . . . . 174, 223 Sun/Moon, delta topocentric elevation

%sxargument special text . . . . . . . . . . . . 174, 223 Sun/Moon, delta topocentric elevation at standard

sunrise time %o*Bargument special text . . 175, 224

Sun/Moon, delta topocentric elevation at standard sunrise time %oBargument special text … 175, 224

Sun/Moon, delta topocentric elevation at standard sunset time %s*Bargument special text … 175, 224

Sun/Moon, delta topocentric elevation at standard sunset time %sBargument special text …. 175, 224

Sundial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 Sunrise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Sunrise and sunset, limit to actual day . . . . . . . . . 40 Sunrise and sunset, set reference altitude. . . . . . . 39 Sunset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Suppress calendar sheet. . . . . . . . . . . . . . . . . . . . . . . . 31 Suppress leading blank line of eternal holiday list

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Suppress leading blank line of fixed date list … 52 Suppress title text of eternal holiday list . . . . . . . 15 Suppress title text of fixed date list . . . . . . . . . . . . 53 Switch character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 269 Switch character, ‘-’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Switch character, ‘--’ . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Switch character, ‘/’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Synodic month . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 Synodic month, Moon . . . . . . . . . . . . . . . . . . . . . . . . 188

T

Table of all environment variables. . . . . . . . . . . . . 255 Table of all error codes . . . . . . . . . . . . . . . . . . . . . . . 257 Table of all exit codes . . . . . . . . . . . . . . . . . . . . . . . . 257 Table of all Metasymbols within this document

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 Table of command separator characters . . . . . . . . 72 Table of modes of the Moon oriented special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 Table of modes of the Sun oriented special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 Table of obsolete date format elements . . . . . . . . 247 Table of obsolete date formats . . . . . . . . . . . . . . . . 247 Table of obsolete fixed default date format texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 Table of obsolete special texts . . . . . . . . . . . . . . . . 246 Table of representation types of the Moon oriented

special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 Table of representation types of the Sun oriented

special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 tcal, options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 tcal, utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Temperature of Earth’s atmosphere, set . . . . . . . . 40 TERM environment variable . . . . . . . . . . . . . . . . . . . . 256 Terrestrial Dynamical time (TDT) . . . . . . . 173, 196 Territory code . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 133 Text part. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Text part of a line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 Text part of a line, limitations . . . . . . . . . . . . . . . . . 97 Text part of a resource file line. . . . . . . . . . . . . . . . . 97 Text replacement %?… special texts. . . . . . . . . . 150 Text variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 Text variables, addition. . . . . . . . . . . . . . . . . . . . . . . 111 Text variables, assignments . . . . . . . . . . . . 38, 85, 106 Text variables, definition . . . . . . . . . . . . . . . . . . 38, 106 Text variables, export of local . . . . . . . . . . . . . 39, 111 Text variables, global . . . . . . . . . . . . . . . . . . . . . 38, 106 Text variables, local . . . . . . . . . . . . . . . . . . . . . . . . . . 106 Text variables, management. . . . . . . . . . . . . . . 38, 106 Text variables, name space . . . . . . . . . . . . . . . . 38, 110 Text variables, numerical . . . . . . . . . . . . . . . . . . . . . 111 Text variables, operations . . . . . . . . . . . . . . . . . 85, 111 Text variables, references . . . . . . . . . . . . . . . . . . . . . . 85 Text variables, scope . . . . . . . . . . . . . . . . . . . . . . 38, 112 Text variables, simple decrement. . . . . . . . . . . . . . 111 Text variables, simple increment . . . . . . . . . . . . . . 111 Text variables, subtraction. . . . . . . . . . . . . . . . . . . . 111 Text variables, total number . . . . . . . . . . . . . . . . . . 106 Text variables, visibility . . . . . . . . . . . . . . . . . . . . . . 110 Text variables, visibility of local . . . . . . . . . . . 39, 111 Textual date %n[argument] special text … 162, 214 The gcal2txt utility . . . . . . . . . . . . . . . . . . . . . . . . . 128 The tcal utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 The txt2gcal utility . . . . . . . . . . . . . . . . . . . . . . . . . 127 Time offset value, relative. . . . . . . . . . . . . . . . . . . . . . 33 Time, 12-hours display format . . . . . . . . . . . . . . . . 159 Time, 24-hours display format . . . . . . . . . . . . . . . . 159 Timezone. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183, 203 Today’s fixed dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 Tomorrows’ fixed dates . . . . . . . . . . . . . . . . . . . . . . . . 62 Topocentric data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Topocentric data, Moon . . . . . . . . . . . . . . . . . . . . . . 193 Topocentric data, Sun . . . . . . . . . . . . . . . . . . . . . . . . 171 Total number of date variables. . . . . . . . . . . . . . . . 103 Total number of text variables . . . . . . . . . . . . . . . . 106 True solar day . . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 270 True solar time . . . . . . . . . . . . . . . . . . . . . . . . . . 174, 270 True track. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Twilight. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Twilight, amateur-astronomical . . . . . . . . . . . . . . . 170 Twilight, astronomical . . . . . . . . . . . . . . . . . . . . . . . . 170 Twilight, civil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 Twilight, nautical . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 txt2gcal, options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 txt2gcal, utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 TXT2GCALPROG environment variable . . . . . . 107, 256 Type of representation of Moon data and times

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180, 202 Type of representation of Sun data and times

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 TZ environment variable . . . . . . . . . . . . . . . . . . . . . . 256

U

Unit, Earth equator radius . . . . . . . . . . . . . . . . . . . 194 Universal time (UTC). . . . . . 33, 155, 158, 160, 161,

162, 173, 183, 196 Universal time 12-hour format suffix, display . . 161 Universal time hour, display . . . . . . . . . . . . . . . . . . 160 Universal time in hh:mm/date format, display

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 Universal time in m[m…]/date format, display

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 Universal time minute, display . . . . . . . . . . . . . . . . 161 Unix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Unsigned angular value, Sun and Moon oriented

special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Unsigned decimal based rational number value,

Sun and Moon oriented special texts . . . . . 181 Unsigned time value, Sun and Moon oriented

special texts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 Use Eastern Orthodox churches’ leap year rule . . 30 USER environment variable . . . . . . . . . . . . . . . . 14, 256 Utilities for Gcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Utility gcal2txt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 Utility tcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 Utility txt2gcal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

V

Valid date parts of appearance factors of days . . 94 Valid date parts of lists of days . . . . . . . . . . . . . . . . 89 Valid date parts of ranges of days . . . . . . . . . . . . . . 91 Valid date parts of repetition factors of days … 93 Validity of omission values in date part . . . . . . . . 88 Vernal equinox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Vernal equinox point on the ecliptic . . . . . 171, 172,

173, 194, 196, 270 Version number, finding . . . . . . . . . . . . . . . . . . . . . . . 10 Visibility of date variables . . . . . . . . . . . . . . . . . . . . 103 Visibility of local date variables . . . . . . . . . . . 38, 105 Visibility of local text variables . . . . . . . . . . . 39, 111 Visibility of text variables . . . . . . . . . . . . . . . . . . . . 110

W

Waning Half Moon phase . . . . . . . . . . . . . . . . . . . . . . 15 Waxing Half Moon phase . . . . . . . . . . . . . . . . . . . . . . 15 Week number %?[date] special text. . . . . . . . . . . . 213 Week number %k[date] special text. . . . . . . . . . . . 154 Week number, determine type . . . . . . . . . . . . . . . . . 30 Week number, ISO-8601:1988 … 31, 52, 56, 58, 86,

154 Week number, standard …. 31, 52, 56, 58, 86, 154 Week numbers in calendar sheet . . . . . . . . . . . . . . . 30 Weekday name %[format]K[date] special text

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150, 212 Weekday number %>1&*E[date] special text … 151,

212 Weekday number %>1&*I[date] special text … 151,

212 Weekday number %>1&*J[date] special text … 152,

212 Weekday number %>1&*S[date] special text … 152,

212 Weekday number %>1&*T[date] special text … 153,

212 Weekday number %>1&*W[date] special text … 151,

212 Weekday number %[format]?[date] special texts

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151, 212 Weekday number %E[date] special text … 151, 212 Weekday number %I[date] special text … 151, 212 Weekday number %J[date] special text … 152, 212 Weekday number %S[date] special text … 152, 212 Weekday number %T[date] special text … 152, 212 Weekday number %W[date] special text … 151, 212 Weeks’ fixed dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Whitespace character. . . . . . . . . . . . 98, 101, 257, 270 Whitespace separator character . . . . . . . . . 9, 84, 102 Winter point on the ecliptic . . . . . . . . . . . . . . . . . . 270 Winter solstice. . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 270 Wintertime. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 Word. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Word boundaries, matching in regexp . . . . . . . . . 141 Word, regexp definition . . . . . . . . . . . . . . . . . . . . . . . 141

Y

Year calendar sheet blocks, number . . . . . . . . . . . . 31 Year calendar sheet, force output . . . . . . . . . . . . . . 31 Year difference value. . . . . . . . . . . . . . . . . . . . . . . . . . 158 Year mode, single. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 Year number %>04*Y[date] special text … 155, 213 Year number %[format]Y[date] special text. … 155,

213 Year number %Y[date] special text. . . . . . . . 155, 213 Years’ fixed dates . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

Z

Zodiacal line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171, 194 Zodiacal marker holidays in eternal holiday list

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Zuhr prayer time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 Zuhr, Islamic prayer time . . . . . . . . . . . . . . . . . . . . . 166